A Complete Bibliography of Publications by, and about, Louis de Broglie

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

16 November 2022 Version 1.28

Title word cross-reference

(2 + 1) [KMSU02]. 1 [KdB51, Mum65, Vrk51, dB39h, dBT50, dB50e, dB50f]. 1/2 [Mum65, dB48g, dB51b, dB52c, dB48g, dB51k]. $120.00 [Lyn02]. 19.2 [SNG65]. 2 [dB41f]. $22.95 [Kar05]. $24.95 [Eck09]. $27.95 [Kar05]. $29.95 [Kle14]. $32.00 [Sch07b]. $4.50 [Kae53]. $4.75 [Mor54]. $49.50 [Sch85a]. $7.50 [Tis65]. = mc^{2} [For80]. ^{e} [Cos85]. B^{(3)} [Eva95]. c [DK93]. E^{∞} [ACIB02, ABN01]. δn · δφ ≥ 2π [dB65d]. ΔW = ΔQ + ΔL [dB67g]. \frac{1}{2} [dB51k]. γ [dBdB28]. H [Gom35c]. H_{0p} [dBA66]. K [dB20b]. L [dB20b]. M [DdB22]. M_{op} [dBA66]. N [SKFS15, SOG06]. Ψ [dB27b]. q [DK93]. Q = Q_{0}\sqrt{1 − β^{2}} [dB67h, dB67f]. \sqrt{N} [SKFS15]. SU(2) [Hof17]. t [Vrk51]. u [Des55a, dB58d]. |N, 0 > +|0, N > [KRK10]. |Ψ|^{2} [dB56c].

-atom [SKFS15]. -dimensional [KMSU02]. -il [Vrk51]. -New [Whi62]. -Numbers [DK93]. -photon [SOG06].

/Herder [RFS^{+}91]. /van [Sch85a]. /Vigier [Sch85a]. /Weisbuch [RFS^{+}91].

0 [Eck09, Kar05, Lyn02, Sch07b]. 0-19-850027-0 [Lyn02]. 0-309-08998-0 [Kar05]. 0-393-05165-X [Kar05]. 0-415-06588-7 [Dic96]. 0-521-31911-0 [Ble86, Sch85a]. 0-521-35404-8 [Dic96]. 0-521-48543-6 [Dic96].

1 [dBA71]. 1911 [LdB12, Som11, Bri54]. 1927 [Jäh10]. 1932 [For80]. 1933 [For80]. 1945 [dB45c]. 1948 [Aug49]. 1949 [Aug49]. 1954/55 [dB55e]. 1955/56 [dB56a]. 1957/58 [dB58a]. 1960 [Wil61]. 1965 [Gar66, Str66]. 1968 [Lin69]. 1987 [Abr88]. 1re [Ano71b, dBA71].

20th [FL09]. 23 [Som11]. 24e [dB55e]. 24th [dB55e]. 25 [Ble86]. 25e [dB56a]. 25th [dB56a]. 27e [dB58a]. 27th [dB58a]. ’29 [Enn87]. 2nd [Str65].

30 [LdB12]. 31 [dB45c].

5 [Kle14]. 50-picometer [Day09]. 50th [PCR73]. 5ème [dB28c]. 5th [dB28c].

649 [Gru54].

7.80 [Deh36]. 75-letiju [Suv67]. 75th [Suv67].

90th [Cos85, Gud85]. 978 [Eck09, Kle14, Sch07b]. 978-0-19-536909-0 [Eck09]. 978-0-691-13968-5 [Kle14]. 978-0-7432-6473-0 [Sch07b].

A. [Gar66, WW97, Wer65]. Abb [Deh36]. Abel [B.36]. abermals [Haa34]. Abragam [Wer65]. Abraham [Kie06]. Absolute [For80, MS67]. absorption [dB20a, dB21b]. Abstract [Ber88]. Académie [Ano58b, Ano75, dB45c]. Academy [Ano75, Ano45, Ano58b]. accord [dB27k]. accordera [Ano73]. according [Cos01, Geo32]. Account [Dic96, Gol94b, Hol93, Hol95]. Achieve [Lev96]. acids [dB25e]. Acoustics [MIS69]. action [Muc74, Siv11, dB48d]. active [Bil01]. Activity [CEC72, dB46d]. Actualités [Car35]. actuel [dB32e, dB37a]. actuelle [dB27h, dB63a]. Additional [dB50e, dB50f]. Adiabatic [dB76a]. adiabatique [dB76a]. Advances [BVH69, KS62, FS77, dB40c]. adventures [Fed08]. adverses [dB27d]. Aerodynamics [SCC^{+}65]. affect [dB50c]. After [Kae31, RY13, Kae48]. again [Pik77]. age [DT74, Van64]. agitation [Met38]. Ago [Bla02, Geh78]. agreement [Mei77]. Aharonov [BT09, BT10]. aid [GP72]. aide [dB53f, dB53g, dB53d]. aimants [dB49a]. Aimé [Ano53, dB53c]. AIP [Ano93, Lev93]. air [Géh33b]. airwaves [dB30e]. Akademische [Deh36]. al [Sch85a]. Alamos [Sch85c]. Alan [WW97]. Albert [For80, Kar05, L.58, BB71a, BB71b, Elb13, For80, Sch49a, Sch49b, dB49b, dB49c, dB55b, dB55c, dB67d]. Albin [Lup49]. Alfred [You76]. algebra [Bau54]. Algebraic [Var02]. algébre [Bau54]. Almanach [Aug49, Aug49]. along [Kie10]. Alternatives [GMG^{+}79]. Alto [Lev93]. Alwyn [Ble86, Tor87]. Ambition [Eck09, Jon08]. America [Sla68]. American [Van64]. amplitude [Feo09, Fie06]. Amsterdam [Gar66]. Analiz [Poi74]. analogie [dB32l]. Analogies [Buh37, dBD22b]. analogous [Cos01, Vrk51]. analogue [Cos01, Vrk51]. Analogy [DK93, dB32l]. Analyses [Aug49, Bou49, Bru48, Cos53b, Del52, Hum51, Spe68]. Analysis [MLG67, RT75, dB55a, Poi74]. analytic [KKTY01]. Analytical [Uts02]. anciennes [dB34l]. Anderson [Cas12, LvTW09]. André [Ano52b, Gil66, Hum51, Lup49, dB44a, dB71c, dB71d, dB42c]. angle [Kar29]. Angles [IB93]. angulaire [Cos01, dBA66]. angular [Cos01, dB36b, dBA66]. angular-momentum [dBA66]. anisotropes [dB42e]. anisotropic [dB42e]. Anlaß [dB35b]. annealing [Mar79]. année [dB55e, dB56a, dB58a]. années [dB46d]. annihilation [dBT44b]. anniversaire [Cos85]. Anniversaries [Kha20]. anniversary [PCR73]. Annual [Ano14]. anonymity [WG67]. Anri [Poi74]. Antidote [Dic96]. antiparticles [dB67b]. antiparticules [dB67b]. Antiprotons [Sch86b]. Antoinette [Spe68]. Antoon [Ano51b, dB51c]. Anwendungen [Nic54]. Anxiété [dB70a, dBxx]. anxiety [dB70a, dBxx]. any [Géh33b, dB77b]. apparent [Squ92]. Application [Dav98, dBV63a, Ano51a, dB46f, dB68f, Dav98, TP43, dB49a, dBV63a]. Applications [DRC68, RFS^{+}91, RZB97, Wer65, Ano51a, Nic54]. Applied [Mor54, BJK03]. Apply [Lev01b]. Appraisal [CFG96]. Approach [Esp99, Joh02, Nor22, BJK03, Mar05]. Approaches [Ber17]. Approximation [Nor22]. April [Ano53]. APS [Mar64]. arbitrary [SV09, dB39g, dB39f]. arbitrary-spin [dB39f]. Archaeology [LH89]. Archaeometry [LH89]. Archives [For80]. argument [Cos92]. arising [BH82a]. Aristocratic [Nye97]. Arrays [HK93]. Arrows [SSL98]. Art [LH89]. Arthur [Deh36, For80]. article [Kle27]. articles [Suv67]. Asher [TBMW05]. Asim [Ble86, Tor87]. Aspects [Čap91b, Hut54b, ME76, dB49d, dB49d]. associated [Des55a, Vig91, dBV63b]. Associates [Lev93]. association [AD95, Cos82]. associé [Des55a]. associées [dBV63b]. assumption [AD95, GV82]. Astronomers [Cla67]. astronomical [dB66c]. astronomique [dB66c]. Astronomy [HÖ66, KS62, WCÖ67]. Astrophysics [FCP85, KS62]. Atlas [Kar05]. Atmosphere [LD93]. Atom [BK97, DGR^{+}99, DAJZ96, Kle97, LPM98, Lev95, Lev97a, Lev04, Lub99, PSZM02, Rod19, GS12, HTH^{+}14, SKFS15, Sch07a, dBD24, dB42b]. atome [dBD24, dB42b]. atomes [DdB21, dBD21]. Atomic [BL89, BBSR69, DGR^{+}99, EUG98, FR13b, HBR13, Lee96, Lev95, PDEK06, Sch85b, SSDD95, Ano51a, DEW93, GM96, SKFS15, TTS05, Wit98, ZW92, dBdB21, dB27e, dB27g]. atomique [Ano51a, dB27e, dB27g]. Atoms [Lev91, Lev04, Lub99, Sch86a, Tso06, DdB21, dBD21]. attains [Day09]. attempt [dB56e, dB58d, dB77c]. Audretsch [RFS^{+}91]. Aufl [Deh36, Haa34]. August [Abr88]. Auguste [Ano53]. auspices [LdB12]. author [WG67]. Autobiographical [For80]. autowave [Bil01]. autres [dB40c]. aux [TP43, dB46e]. avant [Koj04]. avenir [dB41b]. Award [K.52, Ano73]. awarded [Sch13, Sch87b]. awards [Fed09, Ano93]. Axiomatic [Sum88, Des34]. Axiomatique [Des34].

B [Ano52b, Sch85a]. Bachynski [Ano74]. bad [Mer09]. Balance [dB55a]. Ballistic [Khu88]. Baltimore [Ano01]. Ban [Dre97]. Banesh [For80]. Bang [GA98]. Bargmann [Stu03, Stu04]. barriéres [dB33a, dB32d]. barriers [dB32d, dB33a]. Barrillon [Ano39b]. Barut [Ble86, Tor87, Sch85a]. base [dB73]. based [Haa34]. bases [ABdBC33, dB63a, dB67f]. Basic [Fri62, LB66, dB73]. Basil [Dic96]. Basis [Sum88]. Batchelor [Mor54]. Batzel [Tar00]. Bauer [B.36]. Be [Ano45, Lev01b]. Beam [KT66, Sch86a]. Beams [BL89, Lub96b, Sch05, Sch85b, Ben12, EFS31, EFS32, dB68a, dBV72a]. Beauregard [Ber69]. Become [Kae31]. Becoming [Kas13]. Becquerel [dB64d]. Bedeutung [Sch55a]. Been [Kae31, Bon94]. before [Koj04]. Began [Kae48]. Beginner [Kle97]. Beginning [Mil85]. beginnings [dB42d]. Behavior [Ros15]. behind [Fer96, Wal06]. Belgium [dB28c]. Bell [Vin22, Whi95, dB74b]. Below [Lub96a]. Benefit [BB79]. Bergson [Čap71a, Čap71b, Čap91a, dB41c]. Berkeley [Bro96]. Berlin [For80, Str65, Str66]. Bernard [dB69a]. Berr [Ano52b]. Besprechungen [WBBE30]. Best [K.52]. Beta [RP53]. Beta-Decay [RP53]. Bethe [Whi95]. Between [AGG^{+}93, Kas13, BB90, BB71a, BB71b, Cos50, DC48, Loc87a, Sla68, Squ92, dBD22b, dB26f, dB32l, dB35d, dB38a, dB38b, dB42a, dB42f, dB49s, dB49a, dB49r, dB49j, dB49i, dB52g, dB72b]. Bewegungen [Gar53]. Beyond [CL66, KJB^{+}10, MH79, Nor22, Rod19, SM69, Smo19b, dB47a]. bibliographiques [Gag53]. Big [GA98]. bilocal [dBHV59]. bilocale [dBHV59]. binding [Str65]. Bindung [Str65]. Binoux [Ano39c]. Biographical [Ano29a]. Biology [Gat68, dB66d, LS74, Gat68, Lin67, Mil67]. Bird [Pei85]. birth [BH82b]. birthday [Cos85, Suv67, Gud85]. Birzeit [HM79]. bis [For80]. Black [Gid13, KKTY98, KKTY00b, KKTY00a, Wan09]. Blackbody [dB22a]. Blair [Kar05]. bleiben [dB53j, dB53k]. Bloch [Hof84]. Blondel [dB42c, dB44a]. boîte [GH05]. bodhisattva [Yan12]. bodies [dB67e]. Body [Nor22, dB36b]. Bohm [Dic96, Gol94b, Vin22, Acu16, AW17, Anoxx, BT09, BT10, Bel80, Ben10, Ber04, Bri17, BSB99, BW05, Col12, CDE12, DH96, DH02, ECT17, Esp99, Fab22, FF03, Gha06, Gha14, Gol94a, Hol92, Hol93, Hol95, HV13, Joh02, JM13, KKTY98, KKTY00b, KKTY00a, KKTY01, KMSU02, Kie06, Kie10, Kob07, Mar05, MM01, NMEV12, PNSS12, PNF13, PNSS14, Ros15, SS00, SS01, SR13, SC97, Squ92, SV09, Sut97, TBMC07, Van94, Vig94, Vig95, Wan09]. Bohmian [CFG96]. Bohr [Bou49, Len48, Ser58, Shi85, Ano62, Pel47b, Pel47a, Wei14, Whe85, dBdB21, dB24c, dB63d]. Boileua [Ano39a]. Bolles [Kar05]. Boltzmann [dB46e]. Bondi [Mor54]. Bonding [SS66]. boojum [PPF^{+}81]. Book [Bal65, Bal72, Ber69, Bis57, BB73, Ble86, Bri54, Car35, Coh40, Coh51, Dic96, Eck09, For80, Gar66, Gud85, Kae53, Kar05, Kle14, Len48, Lin64, Lin65c, Lin65b, Lin67, Lin69, Lyn02, Mal39, Mor54, Pes88, Sch85a, Sch07b, Shi85, Str65, Str66, Sum88, Tis65, Tor87, Wer65, Bru48, WBBE30]. Books [Ano64a, Ano64b, Ano65a, Ano65b, Ano66a, Ano71a, Ano76b, Ano84b, Ano94a, Ano94b, Ano95, For80, Kar05, Ano13]. Border [AGG^{+}93]. Borel [Ano39a, Ano39b, Ano39c, dB57d]. Born [ABdB^{+}53, BB71a, BB71b, Dre17]. Bose [Lev01b, BEC99, Col95, Ket99, Lev97a, Lev98, Lev01a, Sha74, Wal06, Wil11]. Boson [DH96]. Bouligand [Car35]. bouncing [PNSS14]. Bourgeois [Ano39b]. bourses [Ano73]. Bouvier [Ano39c]. box [GH05]. Bragg [TP43]. Brans [Gha14, Gha06]. breaking [ABN01]. breather [Enz85]. Bridge [Kas13]. Brillouin [Fow29]. bringing [dB27k]. Broglie [Abr88, Acu16, AW17, And30, Ano29b, Ano30b, Ano35a, Ano39a, Ano39b, Ano39c, Ano52b, Ano54, Ano71b, Ano73, Anoxx, Aug49, B.36, Bel80, Ben10, Ber04, Ble86, Bou01, Bou49, Bri17, BSB99, BW05, Cas75, Coh40, Col12, CDE12, CGLM92, Dan37, Deh36, DH02, Dic96, Din41, ECT17, Esp99, EFS31, EFS32, Fab22, FF03, Fes56, For80, Fow29, Gam54, Gar53, Gha06, Gha14, Gol94b, Gru54, Gud85, Haa34, Hil28, HV91, Hol92, Hol93, Hol95, Hum51, HV13, Joh02, JM13, Kae53, KKTY98, KKTY00b, KKTY00a, KKTY01, KMSU02, Kie06, Kie10, Kob07, Kor30, Kub70, L.58, Lin64, Lin65c, Lin65b, Lin67, Lin69, Lup49, Mal39, Mar05, MM01, Mor54, NMEV12, Pes88, PNSS12, PNF13, PNSS14, PCR73, RFS^{+}91, Ros15, Ros48, Sch85a, SS00]. Broglie [Sie97, SR13, SC97, Spe68, Squ92, SV09, Stu03, Stu04, Sut97, TBMC07, Tis65, Tor87, Van94, Var02, Vig70, Vig94, Vig95, Vin22, Wan09, Whi62, WBBE30, Wil61, vdM82, Bou01, Abe57, AD95, ABN01, ACIB02, And30, Ano29a, Ano32, Ano34, Ano35a, Ano35b, Ano37a, Ano37b, Ano45, Ano48, Ano51b, Ano51a, Ano52a, Ano53, Ano55b, Ano55a, Ano58b, Ano58a, Ano59, Ano62, Ano66b, Ano73, Ano75, Ano76a, Ano82, Ano87c, Ano87b, Ano87a, Ano87d, Ano92a, Ano92b, App99, AKC^{+}89, Arn14, Aug49, Bac54, BH79, BJK03, BVV84, BB90, Bar90b, BB92, Bau54, Bau02, Bay07, Bil01, Bla56, Bla63, Bod97, Bod03, BH82a, Boi57, Bol50, Bon94, Bri54, Bri30a, Bri30b, Bru48, CKS^{+}97, Cap52, Čap71a, Čap91a, Car16, Car27, CCG^{+}08, CDW17, Cos53b]. Broglie [Cos50, Cos53a, Cos59, Cos82, Cos85, Cos92, Cos01, Cro87, CGS88, CGLM90, CFGL97, Cro04, Cro01b, Dar11, Dar13, Dav98, Ano30a, Del52, DF48, Des34, DC48, Des55b, Des55a, DEW93, DH96, DGR^{+}99, DAJZ96, Dre17, Dur17, ESI02, ERS^{+}17, EU26, EN52, Elb13, EK51, Enn87, Enz85, EFS31, EFS32, Eva95, Far17, FS97, FS98, Feo09, Fer96, FR13a, Fie06, Fil47, Fil02, F^{+}76, FMP99, FPN^{+}01, FL09, Gam26, Gar53, Gat68, Gau16, Géh33a, Géh33b, Geh37, Geo53, Ger87, Gil66, Gom35c, Gom35a, Gom35b, GH05, GM96, GV82, Haa34, Ham07, Han77, Hil28, Hof17, Hol87, Hol29, HB50, HDN02, Hou16, HTH^{+}14, Hut54a, IS31a, IS31b, JBCY95, Joh77, JNA^{+}12, K.52, Kar29, Kas13, Kir99, Kle27, Koj96a, Koj96b, Koj01]. Broglie [Koj04, Kor30, KJB^{+}10, Kub70, Kul69, Kul70, Kul71, KRK10, Lab35, Lee96, Len48, Lep92, Ley05, LSG^{+}08, LO10, Loc82, Loc87a, Loc87b, Loc93, Loc95, LKF96, Loc04a, Loc04b, Loc07, Mac78, May46, MS54, Mil67, Mil58b, Mil58a, Mol95, Muc74, Mum65, ODHP00, Osc01, Osc11, PW95, Pat45, PDEK06, Pel47b, Pel47a, Pro38, PSZM02, RF69, Rei44a, Rei98a, Rob17, RP53, Ros58, Rub50, S.52, SKFS15, Sch03, Sel93, SS01, Sie97, Siv11, Smo19a, Smy15, Som11, SSDD95, Ste35, Ste02, Stu04, SLH^{+}06, SOG06, Sur82, Suv67, Swa39, Syn54, Syn10, SI31, SI32, Tak27, Tap38, TP43, Ton56, TTS05, Uts02, Val90, Van05, VV03, Vig82, Vig91, Vig93, VV14, Vis51, Vrk51, Wai95, WZM91, Wei06b, Whi95, Wig85, Wil48, Win31, Win34, Wit98, YY99]. Broglie [ZW92, ZGWM92, dB24a, dB56a, dB57a, dB58a, dB75, dBDW^{+}84, dH04, dX33, And82, Coh51, Ser58]. Broglie-Bohm [Gol94b]. Broglie-Feldes [Cap52]. Broglie-type [Bod97, Bod03]. Broglie-Wellen [EFS31, EFS32, Hil28, Kor30]. Broglie... [Car35]. Broglie’a [SI32]. Broglies [Nye97]. Broglie’schen [HB50, EU26]. Brojlja [Suv67, Suv67]. Brownian [Met38, dB67c]. brownien [Met38, dB67c]. Brückner [RFS^{+}91]. Brunschvicg [B.36, dB45g]. Brussels [dB28c, LdB12, Pel47a, Som11, dB49l]. Bruxelles [Far01, dB28c, LdB12, Pel47a, Som11, dB49l, Bou49]. Buchbesprechung [Deh36, RFS^{+}91]. Builds [Lub96b].

C [Ano39c, For80, Hil28]. C. [Hil28]. cachée [Lin65b, Tis65, dB63g, dB64b, dB64c, dB66b, dB68c, dB76a]. cachés [Pau53]. calcul [dB20b, dB49o, dB49p]. calculation [dB20b, dB49o, dB49p]. Caldirola [TBMC07]. Calendar [Ano80, Ano84a]. Cambridge [Ble86, Dic96, Mor54, Sch85a, HJS70, HJS70]. Camille [dB65c]. Can [GP72, Lev01b, ZGWM92, Cro87]. canonical [SR13]. Cantorian [ABN01, ACIB02]. Caquot [Ano39b, Ano39a]. caractéristique [DC48]. carbon [GM07]. Carlhian [Ano30b]. carré [dBA66]. Cartan [For80, ACIB02]. cas [Mum65, dB39h, dB48h, dB48b]. Case [Kul69, Mum65, dB39h]. cases [dB48h, dB48b]. Cassidy [Ano93]. Cat [Kle96]. Catalog [For80]. catalysis [SJ74]. catalyst [Hof79]. Caullery [Ano39c]. Causal [Dic96, Gol94b, Kir98, dB57e, Hol93, Hol95, Vig94, Vig95, dB52e, dB53e, dB56c, dB56e, dBA57, dB57b, dB60b]. causale [Abe57, Mil58a, Ros58, dB52e, dB53e, dB56c, dB56e, dBA57, dB57b]. causalité [Car35, Geo32, dB29a, dB77a, dB77b]. Causality [Boh57a, Boh57b, Boh61, Boh71, Boh84a, Boh84b, Boh99, dB77a, F^{+}76, Geo32, dB29a, dB77b]. Causally [Kir99]. Cavity [DAJZ96]. Cécile [Fed08]. celebrated [F^{+}76]. celebrates [Ano51b]. célèbre [Ano51b]. celle [Cos50]. celles [Pro38]. Centenary [For80]. Center [For80, FL09, Hof17, dB46d]. Centimetric [de 49]. centimétriques [de 49]. Centre [dB46d, FL09]. Century [FL09, Bra09, dB76b]. CERN [Kha05, Lub84]. certain [Des55b, dB53b]. certaines [Des55b, dB41f]. certains [Hut54a, dB53b]. Certainties [dB66a]. Certainty [CL66]. Certitudes [dB66a]. Chadwick [Bro96]. Chair [ABdB^{+}53]. chaleur [dB66b, dB67e]. challenge [Bru07, Smo19a]. chamber [Lin65a]. champ [Cos50, EK51, Géh33b, Vis51, dB32j, dB35d, dB38c, dB49s, dB49r, dB49i, dB49n, dBR50, dB50h, dB51m, dB51i, dB51h, dB52g, dB52f, dB98b]. Champs [Wer65, dB40a, dB37d, dB43g, dB49q, dB49h, dBR50, dB57c]. chance [Boh57a, Boh57b, Boh61, Boh71, Boh84a, Boh84b, Boh99, dB45e]. Channel [Khu88]. Channeling [CCG^{+}08, Osc01, Osc11]. Chaos [ECT17, LM95, Sto05]. chaotic [Met38]. chaotique [Met38]. characteristic [DC48, dB48g]. Charge [IS31b, dB47h, dB49i, dB52g]. charged [AW17, Gau16, dB49s]. chargées [dB49s]. charges [Gou74, dB49r, dB49j]. Charles [B.36, dB46c]. Chemical [SS66, Str65, Gat68]. Chemie [Str66]. chemische [Str65]. Chemistry [GAL66, Les79, NM12, Ben12, SJ74]. Chern [Yan12]. Chevalier [Ano30b]. children [GP72]. China [Lub80]. chmury [SI32]. choc [dBA69]. choice [Bel80]. Chromosphere [AB76]. cinétique [dB23a, dB36b]. cinétiques [dB63f]. cinq [Kle27, dB27i]. Clarified [Dre97]. Clark [Bal72]. classement [Bri30a, Bri30b]. classes [Yan12]. classica [Lab35]. Classical [AGG^{+}93, DK93, KMSU02, Ros15, Val92, Zur91, Bar90b, Cos59, GS12, Lab35, Mer52, PNSS12, PNSS14, dB49o, dB54a, dB95, dBHV59]. classification [Bri30a, Bri30b]. classique [Cos59, Mer52, dB49o, dB54a]. classiques [LKF96, dBHV59, dB95]. CLEO [Ano01]. CLEO/QELS [Ano01]. Clock [CCG^{+}08, Bay07, Osc01, Osc11]. clocks [HBR13]. Cloud [IS31a, LD93, SI31, SI32]. Co [Mal39, dB65a]. co-workers [dB65a]. coefficients [dB52g]. Coexistence [SH69, dB65b]. coherence [ZGWM92]. Coherent [JM13, KMC05, Lev97a]. Cold [TJLB89]. collaborateurs [dB61a]. collaborators [dB61a]. Collana [Vig70]. Collected [Bal65, Sel93, Fow29]. Collection [Bru48, Dan37, Hum51, Boi57, F^{+}76, dBUC^{+}29, dBUC^{+}30, dB30e]. Collections [For80]. collective [SKFS15]. collective-state [SKFS15]. College [Tha93, Koj96a, Koj96b, Koj96a, Koj96b]. collider [Lub84]. Collision [SNG65]. Colloquium [SLF69]. combination [dH04]. comes [Van64]. Comets [HÖ66]. Comment [CGLM92, HV91]. compatibilité [Des55b]. compatibility [Des55b]. Competition [Gru74]. compiled [For80]. complémentaires [dB50e, dB50f]. complémentarité [DF48, DC48, dB48e]. Complementarity [DF48, DC48, dB48e, Shi85]. complementary [Loc07]. Complements [Ang49, Ang52, Ang49, Ang52]. Complete [Kir98, Kir99, dB77b, dB77b]. Completion [Bri17]. Complex [RFS^{+}91, Bod03, JM13, dB40a, dBT50, dB50e, dB50f, dB51k]. complexe [dBT50, dB50e, dB50f, dB51k]. complexes [dB40a]. Complexity [Kir98]. composées [dB46f]. composite [dB46f]. Compton [Lab35, SKFS15, dBdB24]. computation [Les79]. Computer [MW79, MLG67, Lin65a]. Computers [MLG67]. Computing [RG66]. Concept [Kir98, Pel47b]. Conception [dB34e, Ano35b, Gom35a, Gom35b, Loc93, Loc04a, dB34j, dB34k, dB49s, dB49j]. Conceptions [dL66a, dB27d, dB30h, dB37e, dB41c, dB41c]. Concepts [AS70, dB37e].

Conceptual [HC98, JM67]. concernant [DdB22, Som11, dBA57, dB97b, dB97a, dB98a]. concerning [dBA57, dB97b, dB97a, dB98a]. concordance [Car27]. concordanza [Car27]. Condensate [Col95, Lev98, Wil11]. Condensates [Lev97a, Lev97b, Lev01a, Lev01b, SL10]. Condensation [BEC99, Ket99]. Condensed [Gar66]. condition [Cos92]. Conference [Far01, Jäh10]. Conferences [ME76]. configuration [dB52h, dB54b]. configurations [BB92]. conflict [Squ92]. congrès [Bri54, Som11]. Congress [dB28c]. connaissance [dB49l, Lup49]. connaissances [dB37a]. connecting [dB58d]. Connection [Cos50]. Connexion [Cos50]. Conseil [dB28c]. consensus [Lev17]. Conséquences [dB32a, dB62c, dB62c, dB32a]. Considérations [dB54a, Gom35a, Gom35b, KdB51, dBdB24, dBdB34, Gom35a, Gom35b, KdB51, dBdB24, dBdB34, dB54a]. considered [dBBH^{+}63]. constant [dB49k]. constante [dB49k]. constantes [dBT44a]. constants [dBT44a]. construct [dB77c]. construction [Vig91]. Contact [Tha93]. Contemporaine [dL66a, Ano62, dB29a, dB35a, dB41c, dB48c, dB49c, dB76c]. Contemporanei [Vig70]. Contemporary [CL66, Ano62, dB29a, dB35a, dB41c, dB48c, dB76c, dB49c]. Contextuality [Vin22]. contingence [dB45e]. contingency [dB45e]. Continu [dBUC^{+}29, dBUC^{+}30, dB41a, Ano30b]. Continue [BSM^{+}99]. continued [dB97b]. Continues [Ble86, Sch85a, Tor87, BVV84, dB27l, dBDW^{+}84]. Continuité [dB27a]. Continuity [Čap91b, dB27a]. continuous [dB27l, dBUC^{+}29, dBUC^{+}30, dB41a]. contribution [Elb13, Hut54a, Suv67, dB53b, And82]. Contributions [Dre88, DFF13, F^{+}76]. controlled [Vig91]. controlling [HBR13]. Convenes [Ano01]. Convergence [Loc87a, dB50g]. Conversion [ESI02]. Cool [Lub96a]. coordinates [TdB71]. coordonnées [TdB71]. Copenhagen [BW00]. Cornell [Lev01a]. Corona [AB76]. Coronas [LB66]. Corporate [Lev93]. corps [dB36b, dB67e]. corpuscles [dB30c, dB30e, dB32d, dB49g, dB55b, dB27c]. corpusculaire [Bol50, dB46g, dB47f, dB50d]. corpusculaires [Buh37, dBdB21, dBdB22]. Corpuscular [dB26a, Bol50, dB46g, dB50d, Buh37]. Corpuscule [dB27b, Bor17, KdB51, dB33e, dBdB34, dB61c]. Corpuscules [dB27c, dB43a, dB45a, DF48, Mum65, dB30d, dB30c, dB30e, dB32d, dB33a, dB39a, dB39e, dB47c, dB49g, dB50b, dB52h, dB55b, dBA57, dB62d, dB97b, dB97a]. correction [Cou67]. Corrélation [DC48, DC48]. correspondance [dB21b, dB38a, dB38b]. correspondantes [Bri30a, Bri30b]. correspondence [BB71a, BB71b, Lib84, dB38a, dB38b, dB21b]. corresponding [Bri30a, Bri30b, Vig91]. Cosmic [Bos88, Kor63, RFS^{+}91, Sch88b, DFF13]. cosmic-ray [DFF13]. cosmological [PNSS12, PNSS14]. cosmology [MM01, PNF13, SR13, Squ92]. Cosmos [Kar05]. Costa [Ber69]. costs [Mal77]. Cotton [Ano53, Ano53, dB53c]. Coulomb [dBT44a]. counter [Lin65a]. Coup [dB34a, dB51a]. Coupole [dB45c]. courants [dB49a]. cours [Koj96a, Koj96b, dB97a]. Course [dB97a]. created [dB49q]. Creation [GA98]. Creativity [DT74, YF97]. crées [dB49q]. crise [dB29c]. crisis [dB29c]. cristaux [TP43]. Critical [dB63a]. criticality [SK11]. critique [dB63a]. Crossroads [Jäh10, ET07]. Crystal [BO67, TP43]. Crystalline [Nar84]. Crystallography [RG66]. Culture [Nye97]. Cultures [SA65]. Cupola [dB45c]. cuprates [Mil13]. Curie [Ano58a, Ano59]. Curies [Kae48]. curiosities [Kim15]. Current [Lin65c, dB27h, dB32e, dB63a, dB64a, dB37a]. currents [dB49a]. curvature [HTH^{+}14, Kul70]. cut [Rub50]. cut-off [Rub50]. CW [Lub99]. Cybernetics [dB51j, Val90]. cybernétique [dB51j]. cyclical [Muc74].

D [Gud85, Pat45]. d’absorption [dB20b]. d’acides [dB25e]. d’Aimé [Ano53, dB53c]. d’Albert [Elb13, dB49c, dB55b, dB67d]. d’Alembert [dB51l]. damped [Van94]. Dancing [GA98]. d’André [dB42c, dB44a, dB71c]. Danjon [dB71c, dB71d]. d’annihilation [dBT44b]. dans [Bau54, Bor17, Bri30a, Bri30b, Cos53a, Des55a, Géh33a, Géh33b, Geh37, GH05, Loc95, Loc04b, Mey33, Mum65, Pau53, Vrk51, dB21a, DdB21, dB27a, dB29a, dB30c, dB32k, dB32i, dB32b, dB32c, dB33d, dB34i, dB37b, dB37f, dB37d, dB40d, dB40c, dB41f, dB41b, dB42e, dB46b, dB48c, dB49f, dB49k, dB50g, dBR50, dB52i, dB52h, dB52g, dB52f, dB54a, dB54b, dB55b, dB56c, dBA57, dB63g, dB63f, dB65b, dB67b, dB67c, dB70e, dB79]. d’après [Geo32, dB32m]. Darkly [SDBP94]. Darmois [dB62a]. Darstellung [For80]. d’atome [dBdB21]. d’aujourd’hui [Boi57, Dan37, Bru48]. David [Dic96, Gol94a]. Davidson [Fes56, Bis57]. Davisson [Hil28]. dawn [Wei14]. days [Blo76, RC06]. DC [Kar05]. de-Broglie [SS01]. death [Ano87c, Ano87b, Ano87a]. Debate [BSM^{+}99, BPW^{+}71, DH77]. Debever [For80]. DeBroglie [Bis57]. débuts [dB41e, dB42d]. Debye [dBdB24]. Decay [RP53]. December [Som11]. décembre [Som11]. déclare [Ano62]. Decoherence [Zur91]. Découverte [dB09a, dB09b]. découvertes [Hum51, dB51g, S.52]. defended [Gri11]. defense [EGSR72]. Defiant [Kar05]. definition [Cos50, KdB51, Wig07, dB67h, Cos50, KdB51, dB67h]. Deflection [DAJZ96]. degenerate [Ste03]. degradation [dB21a, dB21a]. d’Einstein [Ano48, dB49g]. dela [dB47a]. delayed [Bel80]. delayed-choice [Bel80]. d’électricité [dB36f]. d’électrons [dBdB24]. Delevsky [Ano52b]. delivered [dB45c, dB75]. della [Car27]. Delorme [Ano52b]. demain [Ano51a]. demi [dB76b]. demi-siècle [dB76b]. d’Emile [dB57d]. démission [Ano75]. Demonstration [LSG^{+}08, SOG06, dB54f, dB54f]. denen [dB35b]. d’énergie [dB47d]. d’ensemble [dB53i, dB70g]. densité [dB33d, dB34i]. densités [dB32k]. densities [dB32k]. density [Bel80, Mar05, dB33d, dB34i]. déplacement [dB66c]. depuis [Bri54]. Derivation [Eva95, Rei44a, Rei98a, Sum88, Vig95]. dernières [dB46d]. Descartes [Vig93]. described [Vig91]. description [Bol50, Fie06, Uts02, dB77b]. design [Lub84]. designs [dB30h]. Destouches [MS54, Ros48, Hut54b].

Destouches-Février [MS54]. desynchronization [Bay07]. detect [CGS88]. detecting [dB48f]. detection [CGLM90, Lep92]. determining [vF82]. determinism [Ano55b, F^{+}76, dB29a]. Déterminisme [dB29a, Ano55b, Mey33]. déterministe [Bol50]. deterministic [Bol50]. Detouches [Pat45]. d’études [de 49]. Deux [dB27d, dB59a, Mum65, TP43, dB36f, dB98a]. Devant [dB70a]. Developed [RF69]. Development [JM67, ME76, BH82a, Smy15, Suv67, dB32a, dB48c]. développement [dB32a, dB48c]. Devices [CD90]. DeWitt [Fed08]. DeWitt-Morette [Fed08]. d’exposés [dB30e]. d’Heisenberg [dB82]. Dialog [LBB68]. Diamond [GAL66]. Diamond-like [GAL66]. Dicke [Gha06, Gha14]. did [Dar11, Dar13, Koj01]. dielectric [dB49k]. diélectrique [dB49k]. Different [IB93, Lev97b, dB51k, dB51k]. difficultés [dBT44b]. difficulties [dBT44b]. diffracted [vF82]. Diffraction [Bla02, CS67, Day03, dB67j, DEW93, Geh78, Ger64, Wit98, dB23c, dB26e]. diffractive [ERS^{+}17]. Diffusion [Cor01, Sma10, VV14, Win34, dBdB24]. difrakcija [dB67j]. digitalized [Lin65a]. Dilute [BEC99, Sch84]. Dimensional [EU26, SCC^{+}65, KMSU02, Kle27, Vig91, dB27i]. Dimensions [Fra05, Lub96a, Kle27, dB27i]. d’impulsion [Cos59, Geh37]. d’impulsion-énergie [Cos59]. d’incertitude [DF48, dB32m, dB49m, dB65d, dB69b]. d’individu [dB48e]. Diner [Gud85]. d’interaction [Vis51, dB72b]. d’interférence [dB26e]. d’interprétation [Abe57, Ros58, dB56e]. Dirac [Ble86, Deh36, Haa34, Sch85a, Tor87, dB51b, AW17, Ano35a, BVV84, Bou01, Col12, Dav98, Fab22, Haa34, Hol92, dB32l, dB32k, dB33b, dB34b, dB51b, dB52i, dB52c, dBDW^{+}84]. directed [dB56a, dB57a, dB58a]. Direction [SSL98, Car35]. dirigé [dB56a, dB57a, dB58a]. disappears [Ano62]. disc [dB59a]. discontinu [dBUC^{+}29, dBUC^{+}30, dB41a, Ano30b]. discontinuous [dBUC^{+}29, dBUC^{+}30, dB41a]. Discours [dB45c, dB75]. discover [Dar11, Dar13]. discovered [And70, Dar11, Dar13, SG78]. discoveries [Bra09, dB51g]. discovery [Kae48, Kub70, PCR73, dB09a]. discusses [Lub84]. discussions [LdB12, Som11, dB67h]. dislocated [Hol87]. Disordered [AL88, WW88]. disparait [Ano62]. dispersif [dB67b]. dispersion [Ben10, Met38]. dispersive [dB67b]. displacement [dB66c]. displacements [dBV72a]. disposal [McD79]. disque [dB59a]. distant [dB66c]. Distribution [Cos01, SSP^{+}91, DdB21, dB72b, Cos01]. divergence [Loc87a]. Divers [dB97b, dB97a, dB76d]. Diverses [dB95, LKF96]. Division [Mar64]. Dix [Ano52b]. Dix-huitième [Ano52b]. do [Rua69]. Documentary [For80]. documents [Hut54a, dB53b]. DOD [Fed09]. DOE [Lub84]. d’œil [dB34a, dB51a]. Does [Mol95, Whi62, Vrk51, dB50c]. Dokumente [For80]. Domain [CKS^{+}97]. dominated [SR13]. d’onde [dBdB34, dB52i]. d’onde-pilote [dB52i]. Donder [Géh33a]. d’ondes [dB34c, dB53f, dB53g, dB53d, dB56b, dB68f]. donne [Ano75]. Doppler [dB24a]. Dot [HK93, Lev96]. double [Bel80, CDW17, Des55b, Des55a, Dur17, Far17, Ros58, dB52i, dB52h, dB52a, dB54a, dB54b, dB54f, dB56b, dB56e, dB58d, dB65b, dB71b, dB98a, Mat17, dB56e]. double-slit [Bel80]. double-solution [Des55b, Des55a, dB52i, dB52h, dB54a, dB54b, dB54f, dB56b, dB58d, dB65b, dB71b, dB98a, dB56e]. d’Ouche [dL81]. Douglas [Kle14]. d’ouvrages [Aug49, Bou49, Bru48, Cos53b, Del52, Hum51, Spe68]. Down [ESI02, IB14]. Down-Conversion [ESI02]. Drach [Ano39a, Ano39b]. Draft [Som11]. drift [DH77]. Drop [Kae31]. dual [Mol95]. Dualism [BB68, Gud85, WN84, LBB68, Ley05, dB55b]. dualisme [dB55b]. Dualité [Bor17, dB47c, dB49g, dB62d]. Duality [NM12, Bod03, Bor17, GV82, dB47c, dB49g, dB62d]. Duc [Wil61, Ano66b]. Dugas [Cos53b, Gru54, Coh51, Gag53]. Duhem [Mil66]. Dukas [For80]. Duke [Ano32, Ano66b]. d’un [Cos01, Loc95, dB36b, dB49o, dB49p, dB72b, dB76b]. d’une [Dav98, Hut54a, TdB71, dBT44b, dB47h, dB49k, dBT50, dB51k, dB52e, dB53b, dB67c, dB70e]. Durand [Ano39b]. Durand-Viel [Ano39b]. durch [Hil28]. during [Koj96a, Koj96b, dBdB24, dB46d, dB51a]. Durtain [Del52]. Dutch [Ano51b]. dwarfs [Van79]. Dynamic [Kir98]. Dynamical [BFM80, GM96, Ham07, Mat17]. Dynamics [DS64, RFS^{+}91, ECT17, Mer52, ZB80, dB24b, dB25a, dB26f, dB28b, dB67e, dB67b]. Dynamics/Audretsch [RFS^{+}91]. dynamique [Mer52, dB24b, dB25a, dB26f, dB28b, dB67e, dB67b].

E. [DdB22]. E.-G [Ano39b]. E.d.B.B. [Vig95]. Early [Kha05, Mer02, Per18, Blo76, Cos82, RC06, SR13]. Earth [Cla67, Sch02]. Ease [Tay91]. èclairage [dB60a]. École [Bou01]. écrit [Ano55a]. Ed [Ano30b]. Edge [CL66]. Edinburgh [ABdB^{+}53]. edited [For80]. edition [Str65, Str66]. Éditions [Gru54, Lup49]. editor [Gar66]. Editorial [vdM82]. editors [Tor87]. Edmund [Kar05]. Eds [Mor54, Ble86, Sch85a]. Effect [RP53, And70, Bay07, Cos01, HTH^{+}14, SI32, TTS05, Vig71, Wer65, dBdB22, dB46g, dBV72b]. Effective [SV66]. effects [BT09, Sma10]. effet [Cos01, Vig71, dB46g, dBV72b, Cos01, dBV72b]. Efficiencies [SNG65]. Efficiency [SS92]. Effort [Ste97]. Efimov [Mil15]. eigenstates [Stu03]. Einführ [Ano35a]. Einführ-ung [Ano35a]. Einführung [Deh36, Haa34, RFS^{+}91, WBBE30, dB29b, dB65a]. Einführung/Gaisser [RFS^{+}91]. Einstein [For80, Kar05, L.58, Lev01b, Ano48, Ano55b, BH79, BB71a, BB71b, BEC99, Cas75, Col95, Elb13, FS98, Fes56, For80, Hof79, Hou16, Kar05, Ket99, Klo85, Lev98, Lev01a, Sch49a, Sch49b, Smo19a, Smo19b, Sto05, Vig93, Vig94, Vig95, Wei06a, Wil11, dB49b, dB49g, dB49c, dB55b, dB55c, dBAS^{+}66, dB72a, dBAS^{+}79, Bal72, For80, Kar05, Kle14, Sch07b]. Einstein. [For80]. Eject [Lub99]. elastic [Bor17]. élastique [Bor17]. elected [Ano58b]. electric [HTH^{+}14, dB49r, dB49i, Loc04b]. electric-field [HTH^{+}14]. Electrical [LB66, Ang52]. électricité [dB37a]. electricity [dB36f, dB37a]. electrified [dB32d, dB33a]. electrinos [dB47h, dB47h]. électrique [dB49i]. électriques [Loc04b, dB49r]. électrisés [dB32d, dB33a]. electro [Mer52, Mer52]. electro-dynamics [Mer52]. électro-dynamique [Mer52].

Electrodynamics [Var02, dB49a]. électrodynamique [dB49a]. Electromagnetic [Cos92, RFS^{+}91, BB92, Cos59, Géh33b, Vig91, dB27k, dB32l, dB32j, dB32e, dB35d, dB40a, dB41d, dB43g, dB49o, dB49p, dB49s, dB49r, dB49j, dB49i, de 49, dB49k, dB51d, dB65b, dB64e, dB68d, dB70b]. electromagnétique [dB49k, dB27k, dB32e, dB49p, Cos59, Géh33b, dB32l, dB32j, dB35d, dB40a, dB49o, dB49s, dB49r, dB49i]. Electromagnetiques [Lin69, dB70b, dB43g, de 49, dB51d, dB64e, dB65b, dB68d, dB41d]. electromagnetism [Cos53a, Cos59, Ton56]. électromagnétisme [Cos53a, Cos59, Ton56]. Electron [Bla02, CD90, CCG^{+}08, Day09, Geh78, Khu88, Mar64, Osc01, Osc11, RH67, Ton90, TWH03, Gau16, Géh38, Ger64, Hof17, Kie06, Mot79, PCR73, Pro38, Som34, SI32, Tho67, ZD15, dB25d, dB30f, dB30g, dB32l, dB32f, dB33b, dB34b, dB36c, dB47c, dB49o, dB49p, dBR50, dB50a, dB52i, Géh38, Pro38, dB25d, dB30f, dB30g, dB32l, dB32f, dB33b, dB36c, dB49o, dB49p, dBR50, dB52i]. electron-linked [dBR50]. Electron-Wave [RH67]. Electronic [AL88, HKK94, SS66, dBD21, dB50d]. électronique [dBD21, dB47c, dB50d]. Electrons [IS31a, RP97, Ano66d, SI31, DdB21, dBdB24, dB52c, Bri30a, Bri30b, Bri30a, Bri30b, DdB21, dB52c, dB51b, dB51b]. electrotechnical [Ang49]. électrotechnique [Ang49, Ang52]. Elects [Ano45]. Elektronenmikroskop [dB50a]. elektronowej [SI32]. élémentaires [dB42b, dBV63a, dBV63b]. Elementarteilchen [May46, dB43c, dB65a, HB50]. Elementary [FH64, Lin64, SL65, WSAS62, BH82b, HB50, dBV63a, dB63b, dBBH^{+}63, dBV63b, dB65a, dB42b, dB43c]. Eléments [Gui51, dB53a, DdB22, dB20b, Gui51, dB53a, dB59b, DdB22, dB20b, dB59b, Bac54]. Elettrica [MSB^{+}37]. Elie [For80]. elle [Ano54, Bla56, Hut54a, dB52b, dB53b, dB54e, dB55f]. éloge [Ano58a]. Elsasser [RO93]. élus [Ano58b]. Elwood [Tar00]. Emerging [DFW17]. emigrate [Muc79]. Emile [dB57d, dB43d, dB43d]. émises [dB66c]. emission [GM96]. emitted [dB66c]. Empirical [WN84]. emploi [dBT44b]. empty [ZGWM92]. Energetic [Sch88b]. Energia [MSB^{+}37]. energie [dB49o, dB48a, Ano51a, Cos59, Cos01, dB33d, dB42e, dB47d, dB49p, dB52f, dB63g, dB49a]. Energy [Gut67, SS92, Ano51a, Cos59, Cos01, Ger64, Lee79, YY99, dB33d, dB42e, dB47d, dB48a, dB49o, dB49p, dB49a, dB52f, dB63g]. engineering [ODHP00]. Engineers [SL69, Ang49, Ang52]. Englert [Sch13]. English [Hei74]. Entangled [ESI02, KRK10, SOG06]. Entanglement [Nor22]. Entdeckung [Kub70]. entier [Spe68]. entre [Cos50, DC48, dBD22b, dB26f, dB32l, dB35d, dB38a, dB38b, dB42a, dB42f, dB49s, dB49a, dB49r, dB49i, dB52g, dB58d, dB72b]. Envisioned [Bla02]. episodes [Bra09]. Equation [HDN02, Tap38, AW17, Ber04, Dar11, Dar13, Dav98, Eva95, Gom35a, Gom35b, Hol92, KKTY01, Kor30, Pro38, Rei44a, Rei98a, dB34c, dB58d, dB67g, Dav98, Gom35a, Gom35b, Pro38, dB58d, dB67g]. Equations [FS98, Kul69, Kul70, Stu04, BJK03, Bod97, Gar53, Hou16, Pro38, SV09, Stu03, Ton56, Vig91, dB30h, dB41f, Pro38, Ton56, dB30h, dB41f]. équilibre [dB22b]. Equilibrium [BK76, Ben10, Col12, dB22b]. Equipped [Ros15]. Equivalence [dH04]. equivalent [AW17]. ERA [WG79]. Ergebnisse [May46, dB39c, dB43c]. erinnert [For80]. Ernests [Oli66]. Ernst [BB73]. Errors [IB93]. Eruption [Lad96]. erweiterte [Kor30]. Erwin [Cro01b, Gri12, Gri13, Han77, RY13]. espace [dB52h, dB54b]. Essays [Ble86, CL66, Sch85a, Tor87, BVV84, dBDW^{+}84]. Essence [LM95]. est [Ano52a, dB54e, dB55f]. est-elle [dB54e, dB55f]. estestvennonauchnykh [Poi74]. était [Ano55b]. État [dB37a]. états [dB56c]. etc [Met38]. étendu [dBV63a]. Ether [DH77]. étude [Bor17, Win31, dB30a, Wer65, dB63a, dB98a]. Études [Win34]. Eugene [Ble86, Sch85a, Tor87, BVV84, dBDW^{+}84]. eulogy [Ano58a]. Europe [Rod19]. evaluation [Čap71b]. evanescent [DEW93, Wit98]. Evans [Eva95]. Even [Sch86b]. événement [TdB71]. evening [Ano51a]. event [TdB71]. Events [DFW17, Bar90b]. éventuelle [dB47h]. Everett [BW05]. évidence [dB48f, dB48g]. evolution [Bod97, Loc82, Pel47a, Poi25, dB27h, dB36c, dB45g, dB68e, vdM82, Pel47b, Poi25, dB27h, dB45g, dB68e]. Exact [Wit98]. Examines [Gut67]. example [dB35c, dB37h, dB56b]. excitation [Siv11]. exemple [dB35c, dB37h, dB56b]. Exhibit [For80, Sch02]. Exhibition [For80]. exhibitions [For80]. exist [Mol95, Vrk51]. Existe [Vrk51]. Existe- [Vrk51]. existence [Cro87, dB47h]. Exotic [Sno13]. expanded [Cos92]. expands [Sch80]. expected [dBdB24]. expériences [Vig71]. Experiment [Bau02, Bel80, CGLM90, CFGL97, VV03, dB68a]. Experimental [CGLM92, HV91, Ket99, Lyn02, Tap38, WZM91, ABdBC33, dBdB24, dB49d, dBV72a, Hil28]. expérimentale [dBdB24]. expérimentales [ABdBC33]. experimentaux [dB49d]. experimentelle [Hil28]. Experimenters [Lub96a]. Experiments [Wil11, Vig71, EFS31]. Explain [Kae31]. Explained [Dre97]. Explicit [Vig91]. Exploded [Kae31]. Exploring [GM07]. Exposé [Hut54a, dB53b]. Exposés [Car35, dB34j, dX33]. Exposure [dB53b]. expression [dB34i]. Extended [Vin22, Kor30, dBV63a, dBBH^{+}63, dBHH^{+}63]. Extension [Ber04, BB92, DH02]. Extreme [KMC05]. Eyes [K.52].

F [Bal65, Gud85]. F. [Sli76]. Fabry [dB46c]. Facility [Lub96b]. factories [Ros66, ZS63]. Faculté [dB58a, dB56a, dB57a]. Faculty [dB56a, dB57a, dB58a]. fali [SI32]. Falling [SNG65]. Family [Ano29b, Ano45]. famous [Wei06b]. Fargue [Gud85]. Fascicules [dX33]. Fashionable [BSM^{+}99]. fasi [Car27]. father [Ano87b]. fatty [dB25e]. February [Ano51b]. Feldes [Cap52]. Felix [Hof84]. Feller [RFS^{+}91]. Ferma [dB67i]. Fermat [dB23a, dB67i]. Fermi [Geo32, Ste03, Tho10]. Fermionic [Lev04]. fermions [Col12]. Ferrier [Ano52b]. Ferromagnets [PSZM02]. Février [MS54]. Few [DAJZ96, Dar11, Dar13]. Few-Photon [DAJZ96]. Feynman [Cos50, Fie06, Vis51]. Fichot [Ano39b]. Field [BKG94, FS98, Kul71, Sch02, SV66, dB51h, ACIB02, BB92, Cap52, Cos50, Eva95, Géh33b, HTH^{+}14, Kul70, Vis51, dB32j, dB35d, dB38c, dB43g, dB49s, dB49r, dB49h, dB49j, dB49i, dB49n, dBR50, dB50h, dB51m, dB51i, dB52g, dB52f, dB57c, dB98b]. Fields [Fab22, Bod03, HBR13, dB37d, dB40a, dB49q, dBR50, Wer65]. Fifty [Kae48, LvTW09, Med74, Tso06, Geh78, PCR73]. figure [Ano62]. Figures [Ste97, Del52]. Filippi [Lup49]. Filippo [Vig70]. Filosofi [Vig70]. filosofia [Cas75]. filter [ERS^{+}17]. Finally [Col95]. Finite [BBF95, RP53]. First [K.52, SLF69, Ano52a, Mar64, Mil15, PCR73, dB32g, dB32h, dB75, dB49l]. Fisica [Cas75]. Five [EU26, Kle27, dB27i]. Five-Dimensional [EU26, Kle27, dB27i]. fizika [Poi74]. fiziki [Suv67]. flatland [GM07]. flavour [Fie06]. flocculation [Met38]. floculation [Met38]. Flow [TWH03]. Fluctuations [WW88]. fluid [Des55a]. fluide [Des55a]. fluoresce [Ano66f]. Flux [Day03]. Focal [TP43]. focalisation [TP43, TP43]. foci [dB63f]. Focus [Sch86a]. Focusing [DGR^{+}99, ERS^{+}17, TP43]. followed [dB53b]. Folse [Shi85]. fonction [Vis51, dB49a]. fonctions [dB56b]. fondamental [Ano55b]. fondamentale [Gom35a, Gom35b]. fondamentales [Géh33a]. Fondamentaux [Ros48, Pat45]. Fondation [Ano73, dB75]. footsteps [FL09]. Force [II88]. Forced [Kae31]. Forces [Lev01b, Vig95]. Foreigners [Tay91]. Foreword [Hei39]. form [Muc74, SS01, dB32m, dB49r, dB50h, dB56b]. formation [dB47f]. forme [dB32m, dB49r, dB50h, dB56b]. Formed [Lev98, Bor17]. formée [Bor17]. Formel [And30]. Forms [Sch86a]. formula [And30, dB46e, dB54f, dB67e, dB67f]. formulas [dB70f]. Formulation [Fab22, Stu04]. formule [dB46e, dB54f, dB67e, dB67h, dB67f]. formules [dB70f]. Foundation [Ano73, dB75]. Foundations [HC98, Rei44b, Var02, Ano35a, Gid13, Rei98b]. Four [KJB^{+}10, SLH^{+}06, Fow29]. four-photon [SLH^{+}06]. Four-Wave [KJB^{+}10]. Fourth [Ano45]. foyers [dB63f]. Fr [Ano52b]. Framework [Shi85, HV13]. francaise [Koj04, dB45c, dB45c, Del52]. France [Com67, Koj96a, Koj96b, Nye97, VV14, dB43e, dB51a]. Francis [EP00]. franges [Bol50]. Frank [For80]. Fraunhofer [Bol50]. Frédéric [Ano58a, Ano59, dB60c]. free [Mot79, dB63g, Pel47a, dB49a, dBHH^{+}63]. free-electron [Mot79]. Freedom [SH69, dB79]. French [For80, Kae53, Wer65, Ang49, Ang52, Ano40, Ano45, Ano48, Ano51b, Ano51a, Ano52a, Ano53, Ano55b, Ano55a, Ano58b, Ano58a, Ano59, Ano62, Ano66b, Ano73, Ano75, Ano76a, Ano87c, Ano87b, Ano87a, Ano87d, Ano92a, ABdBC33, Bau54, Bol50, Bor17, Bou01, Bri30a, Bri30b, Bru48, Buh37, Cos50, Cos53a, Cos59, Cos85, Cos01, DdB22, Dav98, DF48, Des34, Des39, DC48, Des55b, Des55a, Dug50, EN52, Elb13, Enn87, FL09, Gam26, Géh33a, Géh33b, Geh37, Géh38, Geo32, Geo53, Ger87, Gom35c, Gom35a, Gom35b, GH05, Gui51, Hei53, Hei57, Hum36, Kle27, Koj96a, Koj96b, Koj01, Koj04, KdB51, LdB12, Loc95, Loc04a, Loc04b, Mer52, Met38, Mum65, Pau53, Pel47b, Pel47a, Poi25, Poi54, Pro38, Sch55b, Som11]. French [TP43, TdB71, Ton56, Vig71, VV14, Vis51, Vrk51, Win31, Win34, dB20a, dB20b, dB21a, DdB21, dBdB21, dBD21, dB21b, dBD22b, dB22c, dB22a, dBdB22, dB22b, dBD22a, dB23b, dB23c, dB23a, dB24b, dBD24, dB24c, dBdB24, dB25d, dB25e, dB25b, dB25c, dB25a, dB26c, dB26f, dB26b, dB26e, dB26d, dB27a, dB27c, dB27b, dB27d, dB27i, dB27h, dB27e, dB27k, dB27f, dB27j, dB27l, dB27g, dBdB28, dB28a, dB28b, dB28c, dBUC^{+}29, dB29c, dB29a, dBUC^{+}30, dB30h, dB30a, dB30f, dB30g, dB30d, dB30c, dB30e, dB32l, dB32j, dB32a, dB32k, dB32m, dB32e, dB32d, dB32g, dB32h, dB32f, dB32i, dB32b, dB32c, dB33d, dB33e, dB33a, dB33b, dB33c, dB34a, dB34i, dB34d, dB34b]. French [dB34c, dB34j, dB34k, dB34h, dBdB34, dB34f, dB34g, dBW34, dB34l, dB35c, dB35a, dB35d, dBD35, dB36c, dB36d, dB36e, dB36f, dB36a, dB36b, dB37a, dB37h, dB37b, dB37f, dB37g, dB37c, dB37d, dB37e, dB38a, dB38b, dB38c, dB39h, dB39a, dB39e, dB39g, dB39f, dB39b, dB40a, dB40b, dB40d, dB40c, dB41a, dB41c, dB41f, dB41b, dB41d, dB41e, dB42c, dB42a, dB42f, dB42e, dB42b, dB42d, dB43a, dB43e, dB43b, dB43g, dB43f, dB43h, dB43d, dBT44a, dB44b, dBT44b, dB44a, dB45a, dB45c, dB45d, dB45e, dB45g, dB45b, dB45f, dB46f, dB46g, dB46d, dB46b, dB46a, dB46e, dB46c, dB47a, dB47h, dB47f, dB47g, dB47c, dB47d, dB47e, dB47b, dB48e, dB48f]. French [dB48g, dB48c, dB48h, dB48b, dB48i, dB49o, dB49q, dB49p, dB49s, dB49a, dB49r, dB49e, dB49f, dB49g, dB49d, dB49h, dB49i, de 49, dB49c, dB49k, dB49l, dB49m, dB49n, dB50g, dBR50, dB50h, dB50c, dB50b, dB50d, dBT50, dB50e, dB50f, dB51a, dB51l, dB51m, dB51d, dB51k, dB51e, dB51f, dB51g, dB51i, dB51h, dB51j, dB51b, dB51c, dB52i, dB52h, dB52a, dB52e, dB52d, dB52b, dB52g, dB52f, dB52c, dB53a, dB53e, dB53f, dB53g, dB53d, dB53b, dB53h, dBJ53, dB53i, dB53c, dB54a, dB54e, dB54b, dB54f, dB54d, dB55b, dB55f, dB55c, dB55d, dB55e, dB56a, dB56b, dB56d, dB56c, dB56e, dB57a, dBA57, dB57c, dB57d, dB57b, dB58a, dB58c, dB58d]. French [dB59a, dB59b, dB59c, dBHV59, dB60a, dB60c, dB60d, dB61a, dB61c, dB61b, dB62b, dB62a, dB62c, dB62d, dBV63a, dB63a, dB63g, dB63e, dB63c, dB63h, dB63f, dBV63b, dB63d, dB64d, dB64e, dB64b, dB64c, dB65b, dB65c, dB65d, dB66a, dB66c, dBA66, dB66b, dB67e, dB67b, dB67h, dB67g, dB67f, dB67c, dB67d, dB68f, dB68d, dB68e, dB68b, dB68g, dB68c, dBA69, dB69b, dB69a, dB70a, dB70f, dB70b, dB70e, dB70d, dB70g, dB71a, dB71b, dB71c, dBA71, dB71e, dBV72b, dB72b, dB73, dB74b, dB75, dB76a, dB76b, dB76c, dB76d, dB77a, dB77b, dB78, dB79, dB82, dB95, dB97b, dB97a, dB98a, dB98b, dBxx, dB09a]. fréquence [dB21a, dB25d, dB47g]. fréquences [dB20b]. frequencies [dB20b]. frequency [Lab35, SKFS15, dB21a, dB25d, dB47g]. frequenza [Lab35]. Fresnel [Cos71, DGR^{+}99, dB27h, dB27f]. Friedrich [For80]. fringe [SKFS15]. fringes [Bol50]. Frisch [BW80, Dit80]. front [HTH^{+}14, ODHP00]. Frontiers [FCP85]. FRW [MM01]. Fueki [RFS^{+}91]. function [KKTY98, KKTY00b, KKTY00a, Siv11, Vis51, Wan09, dB49a]. functions [dB56b]. Fundamental [Sel93, Ano55b, Géh33a, Gom35a, Gom35b]. fünfdimensionalen [EU26]. Further [MSB^{+}37, BH82a]. Fusion [Mum65, Stu04, TJLB89, Bau54, EGSR72, dBV63a, Mum65]. Future [GMG^{+}79, dB41b, ZD15, dB50c].

G [Ano30b, Ano39b, Gud85, Hei74, Mor54]. Gaisser [RFS^{+}91]. Galaxies [HG65]. Galilean [All22]. Galileo [Cro01a]. Gamma [Sch05, dBdB28]. Gamma-Ray [Sch05]. Gammas [Sch88b]. gas [Ste03, Tho10, dB23a]. Gase [Kor30]. Gaseous [Col95]. Gases [BEC99, Lev04, Kor30, Sch84, dB23a, dB67i]. Gatekeepers [YF97]. Gauge [AHH84, TQ76, Cos92, KdB51]. Gauthier [Lin69, Tis65]. Gauthier-Villars [Lin69, Tis65]. gazov [dB67i]. geb [Deh36]. Geburtstag [For80]. Gedächtnisausstellung [For80]. Gedanken [dB35b]. General [For80, NM12, dB49b, dB42c, dB43h, dB54d, dBA71]. générale [dB42c, dB43h, dB54d]. générales [dB30h]. Généralisation [TP43]. generalization [Sut97, TP43]. Generalized [Stu04, Stu03]. Generated [ESI02]. généraux [dBA71, Ano71b]. geniu [Fed09]. Genius [Kar05]. Geometric [SC97, FS97, PW95, dB26f]. Geometrical [Syn54, Syn10, HV13, Mor54]. géométrique [dB26f]. géometriques [Hut54b]. Geometry [Hol87, Hut54b]. George [Lup49]. Georges [Car35, Pes88, Dan37, Hum51, dB62a]. Gerald [For80]. German [Str65, And30, Ano35a, Cap52, EU26, EFS31, EFS32, Gar53, Haa34, Hei39, Hil28, HW04, HB50, Kor30, Kub70, Nic54, Sch55a, Sie97, Som34, Ste35, WBBE30, dB27m, dB29b, dB35b, dB39c, dB43c, dB48d, dB48a, dB50a, dB53j, dB53k, dB65a]. Gets [Ano32]. Geĭzenberga [dB86]. ghost [Bru07]. Giant [BB86]. gibt [dB35b]. Ginzburg [Ber04]. given [Pel47a]. gives [dB35b]. Gleichungen [Gar53]. Glimpses [For80]. go [Rua69]. God [II88]. goes [Lub84]. Goos [dBV72b]. Gordon [BJK03, Enz85, HDN02, Pro38, Vig91]. gradients [HTH^{+}14]. Gramont [Ano39a]. grande [Ano62]. grandes [Del52]. Grandeur [dB45d, Vrk51]. grandeurs [Cos50, dB43g]. Graphene [GM07]. Graphical [EU26]. Graphische [EU26]. gras [dB25e]. gravifique [Géh33a, Géh33b]. gravitation [Dur17]. Gravitational [Sch02, Feo09, Géh33b]. gravity [Géh33a, Gha14, KMSU02, dH04]. great [Ano62, Kae48, Cro01a]. Greatness [dB45d]. Griffon [Gru54]. Ground [MLM^{+}00]. Group [Far01, dBBH^{+}62]. groupant [Pro38]. grouping [Pro38]. Groups [LT66]. Growth [TM68]. Grummt [RFS^{+}91]. Grund [Ano35a, Deh36, Haa34]. Grundlagen [Ber69]. guidage [Mum65, dB54f, dB67b]. guidance [Mum65, SV09, dB67b]. Guide [Kle97, Rod19]. Guided [Sch07a, CGLM92, HV91, WZM91, dB41d, dB51d]. guided-wave [CGLM92, HV91, WZM91]. Guidées [dB41d, dB51d]. guides [dB68f]. guiding [dB54f]. Günther [Sum88]. Gutton [dB65c].

H [Ano39a, Ano39b, Ano39c, Coh40, Din41, Mal39, Mor54]. H. [Hei74, Str65]. Haas [Deh36]. habit [Mer09]. hadron [Lub84]. Hafstad [AM94]. Hahn [For80]. half [dB76b]. half-century [dB76b]. Hamilton [AW17, Eva95]. Hamiltonian [dBA66]. Hamiltonians [SV09]. hamiltonien [dBA66]. Hanchen [dBV72b]. Handbook [MMJ89]. hardback [Dic96]. harmonic [Van94]. Harmony [dH04]. Hartmann [Str65]. harvest [Bra09]. Hasard [dB45e]. Hauriou [Ano30b]. haute [dB21a]. Hawking [Cro01a]. head [Sma10]. heat [dB66b, dB67e]. heavy [Pro38, dB20b, DdB21, dBD21]. Heidmann [Sho00]. Heisenberg [Ano35a, Cas75, Deh36, Haa34, Mil58b, RFS^{+}91, Ano35a, Blo76, DC48, Haa34, Loc07, Mil85, PLS02, PSZM02, dB58d, dB82, dB90]. held [LdB12]. Helen [For80]. Helium [DGR^{+}99, GS87, Lub96a, PR71, SP12]. Helium-4 [GS87]. Helmut [Lyn02]. Help [Bel72]. Hendrik [Ano51b, dB51c]. Henri [Ano52b, dB46d, dB97a, Poi74]. Henry [Kar05, ABP93, Shi85]. Herder [RFS^{+}91]. here [Ano55a, Rua69]. Herneck [For80]. Herta [For80]. Hertz [KdB51, Vrk51]. heterostructures [Nar84]. Hewlett [Lev93]. Hidden [Fra05, Hof17, Pau53, Van94, dB63g, dB64b, dB64c, dB66b, dB68c, dB76a, dBLBVP76]. hidden-variable [Van94]. Hideki [Bro86]. Higgs [Sch13]. High [Gar66, Gut67, Tha93, ABN01, dB21a]. High-Energy [Gut67]. high-frequency [dB21a]. high-temperature [ABN01]. highlighting [dB48g]. Hilbert [Sum88]. Hiley [Dic96]. him [Ano62]. Hiring [Tay91]. Histoire [Dug50, Bou49, Pel47a, dB34a, dB51a, dB70g, Ano39c, Coh51, Cos53b, Gag53, Gru54]. historically [Hun66]. History [DK93, Kha05, LH89, Mer02, Per18, Ber17, Dug88, ET07, KHFA67, Pel47a, dB34a, dB51a, dB70g, doB75, Dug50]. Hjalmar [DdB22]. Hoax [WW97]. Hoffman [For80]. Holds [Lev97b]. hole [KKTY00b, KKTY00a, Wan09]. holes [Gid13, KKTY98]. Holland [Dic96, Gar66, Gol94b]. hollandais [Ano51b]. Holography [DRC68, Ton90]. hommage [Ano59]. homme [dB51l]. Homodyne [Lep92]. honor [F^{+}76, Sel93]. Honored [Ano45, Ano93]. Honour [Ble86, Sch85a, Tor87, BVV84, dBDW^{+}84]. Hot [BB86]. Houches [Fed08]. Hüfner [RFS^{+}91]. Hughes [Ano84c]. huitième [Ano52b]. Human [For80, Pau79, SW79]. humanised [Ano55a]. Humanities [For80]. hurricane [Bel72]. Hurwitz [ABP93]. Hydrodynamics [SCC^{+}65]. Hydrogen [Kle99, Lev98]. Hypothese [And30, Ste35]. hypotheses [Des55b, Des55b]. hypothesis [And30, Ste35].

ICFA [Lub84]. ici [Ano55a]. Idea [Kae31, Bro78]. idéalisation [dB34f]. Idealistic [Kae31]. idealization [dB34f]. Ideas [RF69, BB92, Loc82, Loc87a, dB34d, dB35b, dB41c, dB48e, dB52i, dBA57, dB73, dB83, vdM82]. Idées [dBA57, dB34d, dB41c, dB48e, dB52i, dB73]. identical [BSB99]. identity [Fil02]. Ignored [Mil66]. ihre [Nic54]. II [For80, May46, Pat45, Ano39a, Bod03, Gom35b, Koj96b, Ton56, dB42a, dB42f, dBHH^{+}63]. III [Car35, Gui51, Pat45, dB32a, Poi74, dB32a, Poi74]. Illus [Mor54, Tis65]. Illustration [EU26, dB56b, dB56b]. image [dB47f]. imaged [Mil15]. images [dB47f, For80]. Imagination [TM68]. Imaging [TWH03]. Imbert [dBV72b]. immediate [ABdBC33]. immédiates [ABdBC33]. Impact [Jac87, PLS02]. Implications [NM12]. implies [SR13, dB77b]. implique [dB77b]. improving [LO10]. impulse [Geh37]. impulsion [TdB71, dB52f]. Inc. [Mal39]. incertitudes [dB66a, dB82]. incident [dBV72b]. incidente [dBV72b]. Independence [Bla86]. Indeterminism [Hut54a]. indéterministe [Ano54, Bla56, Hut54a, dB52b, dB53b]. indeterministic [dB52b, dB53b, dB53j, dB53k]. indeterministisch [dB53j, dB53k]. Index [Ano14]. India [Bla86]. Individual [YF97, dB48e]. Individualität [May46, dB43c]. Individualité [dB37b, dB27a]. Individuality [dB27a, dB37b, dB43c]. induce [ZGWM92]. induced [AKC^{+}89, HTH^{+}14]. industrielles [Car35]. Inelastic [Pal63]. inertia [dB47d, dB48a]. Inertial [Acu16, Vig95]. inertie [dB47d]. Infeld [Cas75]. influe [dB50c]. influe-t-il [dB50c]. Influence [IS31a, IS31b, Bil01, GM96, SI31, BB73]. information [Gid13]. ingénieurs [Ang49, Ang52]. Insight [Sto05]. Insights [GS12, BH82a]. Institut [dB97a]. Institute [Sch80, dB46d]. integral [Muc74]. intégrales [Ton56, dB32g, dB32h, dB50g]. integrals [Fed08, Ton56, dB32g, dB32h, dB50g]. integration [Cos92]. intellect [Mil66]. Intellectual [SH69]. Intensifies [Sch85c]. interacting [Ste03]. Interaction [RH67, Cap52, Vis51, dB35d, dB37b, dB43c, dB49s, dB49a, dB49r, dB49j, dB49i, dBHH^{+}63, dB72b, dB42f]. Interactions [TQ76, dB38a, dB38b, dB42a, dB42a]. Interference [WW88, dB26a, dB67j, dB23c, dB24b, dB26e, dB48b, dB48b]. interferences [dB22c, dB68a, dB22c, dB23c, dB24b]. interferencija [dB67j]. interferometer [HTH^{+}14, SKFS15]. Interferometers [Lev91, Lev95, SP12]. Interferometry [BK97, CKS^{+}97, GHZ93, JNA^{+}12, Lev04, Lyn02, Cro87, Wer80]. Internal [CCG^{+}08, Osc01, Osc11, Wer65]. International [Mei77, RM78, F^{+}76]. internes [Wer65]. Interpretation [BHG94, Dic96, Gol94b, Lin65c, Per18, RFS^{+}91, dB57e, dB68a, dB90, BH93, FF03, FS97, Gha06, GV82, Hes90, Hol93, Hol95, KKTY98, KKTY00b, KKTY00a, KKTY01, KMSU02, Loc82, NMEV12, SR13, Vig94, Vig95, Wan09, dB25e, dB41f, dB52h, dB52e, dB53e, dB53f, dB53g, dB53d, dB54e, dB55f, dB56c, dB56e, dBA57, dB57b, dB59c, dB60b, dB62d, dB64a, dBA66, dB69b, dB70g, dB71b, dB82, dB98a, doB75, vdM82, Mil58a, dB25e, dB41f, dB52h, dB52e, dB53e, dB53f, dB53g, dB54e, dB55f, dB56c, dBA57, dB57b, dB62d, dB63a, dBA66, dB70g, dB82, dB98a]. interpretations [dB63a]. interpretatsiya [dB86]. intervention [dB77b]. Interview [dB67a]. intrinsèque [Bou01]. intrinsic [Bou01]. intrinsic-particle [Bou01]. introduced [dB34d]. Introduction [BKG94, DFW17, LH89, LKF96, MLG67, Win31, dBdB28, dB30a, dB61a, dB63b, Ano35a, Cos53a, Haa34, dB30b, dB52i, dB63g, Lin64, WBBE30, Win31, dBdB28, dB29b, dB30a, dBT44a, dB61a, dB65a]. introduites [dB34d]. Invariance [All22, SL65, dB76a]. invariant [DH02]. Invention [JSSS71, dB37f]. inventory [KHFA67]. Investigation [Tap38]. Investigations [dB27m]. Invisible [FCP85]. Ion [Lub96b]. Ionics [RFS^{+}91]. Ionics/Kovetz [RFS^{+}91]. Ionization [HL94]. Isaacson [Sch07b]. ISBN [Ble86, Dic96, Eck09, Kar05, Kle14, Lyn02, Sch85a, Sch07b]. isobaric [dBBH^{+}62]. Isolated [dB64f, Hof17, dB61c, dB61b, dB62b, dB62c, dB63e, dB63f, dB64c]. isolé [dB61c]. isolée [Lin65b, Tis65, dB61b, dB62b, dB62c, dB63e, dB63f, dB64c]. issue [DFW17]. issues [Mat17]. Italian [Car27, Cas75, Lab35]. itinéraire [Pes88]. Itterbeek [Gar66]. IV [Far01, dB49l]. Izbrannye [Poi74].

J [Ano39a, Ano39b, Ano52b, Dic96, Hei74, Mor54, Sch85a, Ser58, Whi62]. J. [dB51e]. J.-P [Sch85a]. J.-P. [dB51e]. Jacobi [AW17, Eva95]. Jacques [Ano30b]. Jahre [For80]. Jahren [For80]. Jalons [dB78]. Japan [SS92]. jauge [KdB51]. Jean [Ano52b, Ano54, Ble86, Bou49, Hut54a, Hut54b, Len48, Ros48, Tor87, dB53b, dB61a, dB65a, Sho00, dB45f, dB64d, dB65a, dB71a]. Jean-Louis [Hut54b, Ros48]. Jean-Pierre [Ano54, Ble86, Hut54a, Tor87, dB53b, dB61a, dB65a, dB65a]. Joan [For80]. Johann [Ber69]. John [Gar66]. Johnston [Mal39, Coh40, Din41]. Joliot [Ano58a, Ano59, dB60c]. Joliot-Curie [Ano58a, Ano59]. Jones [Eck09]. Jordan [Gha14, Var02]. Joseph [Gol94a, Kar05]. Josephson [And70]. Jouguet [Ano39a, Ano39b]. journals [Pik77]. Journey [FR13b, FL09, Lev17]. Jr. [ABP93]. jubilee [F^{+}76]. July [Ano51b]. Jump [SSP^{+}91]. Justification [dB54b, dB54a, dB54b]. justify [Dre17].

Kaku [Kar05]. Kalinga [K.52]. Kanai [TBMC07]. Katinga [Ano52a]. Katze [RFS^{+}91]. Kemmer [Pro38]. kernel [Cos50]. Ketterle [Lev01a]. Key [Dre97]. kinds [dB36f]. kinetic [dB23a, dB63f, dB67i]. kineticheskaja [dB67i]. Kinetics [Fri62]. Klaus [Str66]. Klein [BJK03, EU26, HDN02, Pro38, Vig91, dB27j]. Knowledge [Kir98, Kir99, dB37a, dB49l]. Koenig [dBD35]. Kovetz [RFS^{+}91]. Kramers [Dre88]. Kronig [Bal65]. Kucharski [Ano52b]. Kudo [RFS^{+}91]. Kunz [Vig71]. kvantov [dB77c]. Kvanty [dB67i, dB67j, dB67k].

L [And30, Ano30b, Ano39a, Ano39b, Ano39c, Cos53a, Fow29, Gam26, Geh37, Kle27, Mor54, Pat45, Ton56]. L. [And30, CDW17, Cos50, Dur17, EN52, Géh33a, Géh33b, Pro38, Rob17]. Lab [Lev93]. L’activité [dB46d]. Lagrange [dB49a]. Lagrangian [dB51i, dB51h]. lagrangien [dB51i, dB51h]. Lalande [Ano52b]. Landau [dB32m, Ber04]. Landé [You76]. Langevin [Sch80, Koj96a, Koj96b, Koj01, Som11, dB47b]. Large [Day03]. Laser [BL89, Kle97, Lev96, Mar79, Sch85b, Sch86a, DEW93, Sch07a, Wit98, dBV72a]. Lasers [KMC05, Lub99, Mot79, dB63h]. Last [Bla02, Lev98, dB46d]. late [MSB^{+}37]. later [CDW17]. Lau [TTS05]. Laubeuf [Ano39b]. Laue [For80, Sch80]. launches [Fed09]. lauréat [Ano52a]. laureate [Ano52a, PCR73]. L’avenir [dB41b, dB50c]. Lawrence [AM94, Bro96]. laws [Géh33a, Kie06, dB44b]. lead [Vig71]. leading [Cro01a]. Leben [For80, RFS^{+}91]. Lecons [Bou49, Mer52, Pel47a]. Lecornu [Ano39a, Ano39b]. Lecture [EUG98, FL09]. Lectures [Fow29, Koj96a, Koj96b]. L’Effet [Wer65]. Legal [YF97]. Leipzig [Deh36]. L’électron [Som34, dB34b, Ano34]. Lenoir [Ano52b]. L’enseignement [dB49e]. Leon [B.36, dB45g]. Leprince [DFF13, Jac01]. Leprince-Ringuet [DFF13, Jac01]. Leptonen [Cap52]. leptons [Cap52]. L’équation [dB34c].

L’espace [dB49f]. Lessons [Lyn02, Mer52, Pel47a]. L’essor [dB43e]. L’état [dB32e]. letiju [Suv67]. Letter [Som11]. Letters [Bro96, For80, BB71a, BB71b, Hei74, Pik77, dB51l]. Lettre [Som11]. lettres [dB51l]. leur [dB27d]. L’évolution [Pel47b, Pel47a, dB36c, B.36, Bou49, Len48, Ser58]. Lewis [For80]. Lexikon [RFS^{+}91]. LHC [Wei11]. liberté [dB79]. libre [Bou49, Pel47a, dB49a, dB63g]. Licht [dB39c]. Lichtquanten [And30]. Lie [LT66]. liées [DF48, dBT44b]. Lies [Smo19b, dBR50]. Life [Bal72, BB73, Sch07b, Ano53, Cro01a, GF67, Hei74, dB44a, dB46c, dB47b, dB51c, dB53c, dB57d, dB60c, dB62a, dB63d, dB64d, dB65c, dB67d, dB69a, dB71c, dB43d]. Light [App99, Bla02, dB06, Coh40, Din41, Hei53, Joh91, K.52, Lad96, Mal39, Sch85b, SL69, Var02, dB26a, dB34e, dB37g, dB67j, And30, Cro04, Dav98, Gom35a, Gom35b, HBR13, Loc95, TTS05, dB22c, dB22a, dB24b, dB24d, dB26e, dB27d, dB32j, dB33d, dB34j, dB34k, dB34g, dB34l, dB35c, dB36d, dB36e, dB37h, dB37e, dB39d, dB40b, dB40d, dB42a, dB42f, dB42e, dBV72b, dB77c, dB23c, dB39c, dB40d, dB46b, dB60a, Swa39]. lighting [dB60a]. like [BB92, GAL66, Wai95]. Limit [DAJZ96, Lub96a, LO10]. limitant [dB46g]. limites [dB20b, dB47a]. limiting [dB20b, dB46g]. Limits [Day03, dB47a]. line [FF03, Lin65a, dB27k]. linéaire [Abe57, Ros58, dB53e, dB56e]. linéairement [dB49k]. Linear [RH67, Vig91, dB60b]. Linearized [FS98]. linearly [dB49k]. lines [dB20b, dB66c]. link [Fer96]. linked [dBR50]. linking [dB26e]. l’Institut [dB46d]. L’intéraction [dB42f]. L’interprétation [dB53d, dB59c, dB71b]. L’introduction [dBT44a]. L’invariance [dB76a]. L’invention [dB37f]. Lionnais [Ano52b]. Liquids [HMH88, Pal63]. Lise [For80]. Literaturberichte [Som34]. Literature [Som34]. Liverpool [Bro96]. Localization [Joh91, Cas12, GM96, LvTW09]. Localized [Bar90b, Bod03, Siv11]. Lochak [Gud85, RFS^{+}91, Pes88]. L’œuvre [dB27h, dB42c, dB49g]. lointain [dB66c]. lois [Géh33a, dB44b]. L’onda [Lab35]. L’onde [Des55a]. London [Dic96]. Long [Lev98, Mil13]. long-sought [Mil13]. longitudinal [dBV72a]. Look [SA65, dB34a, dB51a]. Loomis [Sli76]. loop [Hof17]. Lord [MSB^{+}37]. Lorentz [Ano51b, Ano51b, dB51c]. lors [Koj96a, Koj96b, dBdB24]. loses [BW05]. Louis [Abe57, Ano29b, Ano30b, Ano34, Ano35b, Ano37a, Ano37b, Ano48, Ano51b, Ano51a, Ano52a, Ano52b, Ano53, Ano54, Ano55a, Ano58b, Ano58a, Ano59, Ano62, Ano66b, Ano73, Ano75, Aug49, B.36, Bac54, Bau54, Bis57, Bla56, Bla63, Ble86, Boi57, Bou01, Bou49, Bru48, Car35, Coh40, Coh51, Cos53b, Dan37, Ano30a, Del52, Des34, Des55b, Des55a, Din41, Fes56, Fil47, For80, Gam54, Gat68, Gil66, Gom35a, Gom35b, Gru54, Gud85, Hum51, Hut54a, Hut54b, Kae53, L.58, Len48, Lin69, Lup49, Mal39, May46, MS54, Mil67, Mum65, Pat45, Pes88, PCR73, Ros58, Ros48, S.52, Sch85a, Ser58, Swa39, Tis65, Tor87, Vig70, Whi62, WBBE30, dX33, vdM82, Abr88, Ano29a, Ano45, Ano55b, Ano58b, Ano71b, Ano73, Ano76a, Ano82, Ano87c, Ano87b]. Louis [Ano87a, Ano87d, Ano92a, Ano92b, BVV84, Bol50, Čap71a, Čap91a, Cos59, Cos82, Cos85, Cos01, Cro01b, Dar11, Dar13, Dav98, DFF13, DF48, DC48, Elb13, EK51, Enn87, Far17, FR13a, F^{+}76, FL09, Geo53, Ger87, Gom35a, Gom35b, Han77, Hof17, Jac01, K.52, Kir99, Koj96a, Koj96b, Koj01, Koj04, Kub70, Lin64, Lin65c, Lin65b, Lin67, Loc82, Loc87b, Loc93, LKF96, Loc04a, Loc07, Pel47b, Pel47a, Sel93, Sie97, Smy15, Suv67, Val90, VV03, Vig82, VV14, Vis51, Wei06b, Win31, dB56a, dB57a, dB58a, dB75, dBDW^{+}84, And82, L.58, Pes88, Spe68]. Louis-de-Broglie [Ano73, Ano73]. Louis-Victor [Ano30a]. lourd [Pro38]. lourdes [DdB21, dBD21]. lourds [dB20b]. Low [Ger64, MS67]. Low-energy [Ger64]. l’U.R.S.S. [Ano58b]. Luc [Del52]. Ludwig [Sum88]. Lui [Suv67, Ano62]. lumiére [dB33d, Dav98, Gom35a, Gom35b, Hei53, Loc95, dB22c, dB22a, dB23c, dB24b, dB26e, dB27d, dB34j, dB34k, dB34g, dB34l, dB35c, dB36d, dB36e, dB37h, dB37g, dB37e, dB40b, dB40d, dB42a, dB42f, dB46b, dB60a, dBV72b, Ano35b, Dan37]. lumineuse [dB32j, dB42e]. L’univers [dB27i, Kle27]. l’Université [Bou49]. Lunn [WW97]. Lyot [dB69a].

M [Ano39b, Hut54a, RO93]. M. [Ano48, Ano51b, Ano52a, Ano53, Ano55a, Ano59, Ano62, Ano75, Bau54, Cos53a, Des34, Des55b, Des55a, Gam26, Geh37, Geo32, Gom35c, Kle27, Mum65, Som11, Ton56, Vrk51, Wil48, dB27j, dB30i, dB43d, dB47h, dB53b, dB53h, dB61a, LdB12]. même [dB52a]. Mach [BB73]. machines [Kor63]. Macroscopic [Bar90a]. Macrosystems [Sum88]. Made [Lev01b]. Magazine [Wei06b]. Magnetic [Lev01b, Loc04b, Som34, dB32f, dB33b, dB34b, dB48f, dB48g]. magnétique [dB32f, dB33b, dB34b, dB48f, dB48g, Som34, Ano34]. magnétiques [Loc04b]. Magnetism [SV66]. magnets [dB49a]. magnitude [Vrk51]. magnitudes [Cos50]. mai [dB45c]. Mainzer [RFS^{+}91]. Majorana [Var02]. Making [Arn14]. Man [Whe85, Wal06, dB51l]. manipulated [TTS05]. Manuscripts [For80]. Many [Cor01, Nor22]. Many-Body [Nor22]. March [Abr88]. Marie [Spe68]. Marie-Antoinette [Spe68]. Martin [Bis57, Fes56]. mass [And30, Bor17, PW95, Wig85, Wig07, dB52g, dB67e, dBV72a, Ste35, Vig71, dBV72b]. Masse [Vig71, dBV72b, dB52g, dB67e]. masselotte [Bor17]. Massenvariabilität [Ste35]. massive [AW17]. Masson [Ano52b]. Masson-Oursel [Ano52b]. matematicheskikh [Poi74]. Matematika [Poi74]. Materials [CS67, Bil01]. Materie [dB39c]. matériel [dB26f]. Materiewellen [Deh36, Ano35a, Haa34, Kub70]. mathémaliques [Car35]. Mathematical [Kae53, LS69b, Muc74, Poi74, Vig91, dB39e, dB54c]. Mathematician [dB51l]. mathématicien [dB51l]. Mathematics [Mor54, Ang49, Ang52, dB48c, Poi74]. mathématique [dB39e]. mathématiques [Ang49, Ang52, dB48c]. Mathematische [Ber69]. Matière [dB37g, Ano87c, dB20a, dB21b, dB27e, dB27g, dB35d, dB38a, dB38b, dB42a, dB42f, dB42b, Dan37]. matrice [DC48]. matrices [Bau54]. matrix [Bau54, Bel80, DC48, Mar05]. Matter [FS98, Kir99, dB39d, Ano87c, AKC^{+}89, Arn14, ERS^{+}17, Eva95, Fie06, KMM74, Kub70, Med74, SG78, dB20a, dB21b, dB26f, dB27e, dB27g, dB35d, dB38a, dB38b, dB39c, dB42a, dB42f, dB42b, Ano35a, Haa34, dB37g, Coh40, Din41, Mal39, Swa39].

Maurice [Ano30b, Bri54, Wil61]. Max [ABdB^{+}53, For80, BB71a, BB71b, dB48d]. maximally [KRK10, SOG06]. maximally-path-entangled [KRK10]. Maxwell [Hou16]. Maxwellian [KdB51]. maxwellien [KdB51]. May [dB45c, Ano66f]. mean [dB32k]. Meaning [MS67, Sch55a, Sch55b, dB56c]. Means [CCG^{+}08, Cap52]. Measurement [ESI02, FMP99, JKA91, SSL98, dB57b, dB57e, BW05, LSG^{+}08, LO10, SLH^{+}06, SOG06, And82]. Measurements [Lev95, Arn14, HTH^{+}14, Mil13, NMEV12, SC97]. measuring [TdB71]. Mécanique [Abe57, Ber69, Géh38, Geo32, dB39e, dB49h, dB57c, dB57b, Ano71b, Ano73, Ano87b, Coh51, Dug50, Kle27, LKF96, Mil58a, Pau53, Ros58, Sch55b, Spe68, VV14, Win31, dB26b, dB26d, dB27i, dB27e, dB27k, dB27l, dB27g, dB28a, dB30h, dB30a, dB32a, dB32g, dB32h, dB32c, dB34d, dBD35, dB36a, dB39h, dB39a, dB40b, dB40d, dB41e, dB42a, dB42f, dB42d, dB43a, dB43b, dB43g, dBT44a, dB45a, dB45b, dB46f, dB46a, dB47g, dB48h, dB48b, dB50b, dB52h, dB52a, dB52e, dB53a, dB53e, dB53f, dB53g, dB53d, dB54a, dB54e, dB54b, dB55f, dB55d, dB56c, dB56e, dBA57, dB59b, dB59c, dB63a, dBA66, dB67f, dB68f, dB68b, dB68g, dBA69, dB70f, dB70g, dB71b, dBA71, dB73]. mécanique [dB82, dB95, dB98a, Ano39a, Gru54, Cos53b, Gag53]. meccanica [Lab35]. Mechanical [Kae31]. Mechanics [BK76, BM88, BW00, Bri17, Cro01b, CFG96, Dre88, Esp99, Gat68, HT93, JM67, Joh02, Kir98, LS69b, Lin67, LS69c, Mil67, Mor54, RFS^{+}91, Rei44b, Sch55a, Sch55b, Sum88, Vin22, WW97, dB57b, dB57e, dB66d, dB90, Ano35a, Ano73, Ano87b, BPW^{+}71, Ben10, Blo76, CDE12, DH02, Dug50, Dug88, ECT17, ET07, FF03, Fie06, F^{+}76, Géh38, Geo32, Gri11, GV82, Haa34, Hes90, Hol93, Hol95, Kie10, Lab35, Loc82, Loc87a, NMEV12, Pau53, PCR73, RC06, Rei98b, Sie97, Syn54, Syn10, Vig94, Vig95, VV14, Win31, dB26b, dB26d, dB27i, dB27e, dB27k, dB27l, dB27g, dB28a, dBB28, dB29b, dB30h, dB30a, dB30b, dB32a, dB32g, dB32h, dB32c, dB34d, dB35b, dBD35, dB36a, dB39h, dB39a, dB39e, dB40b]. mechanics [dB40d, dB41e, dB42a, dB42f, dB42d, dB43a, dB43b, dB43g, dBT44a, dB45a, dB45b, dB46f, dB46a, dB47g, dB48h, dB48b, dB49h, dB50b, dB52h, dB52a, dB52e, dB53a, dB53e, dB53f, dB53g, dB53d, dB54a, dB54e, dB54b, dB55f, dB55d, dB56c, dB56e, dBA57, dB57c, dB59b, dB59c, dB60b, dB63a, dB64a, dBA66, dB67f, dB68f, dB68b, dB68g, dBA69, dB70f, dB70c, dB70g, dB71b, dBA71, dB73, dB82, dB83, dB95, dB98a, doB75, vdM82, Dic96, Fow29, Fow29, Gol94b, Lin65c, Lyn02, WBBE30, Kle27]. Mechanics/Spatschek [RFS^{+}91]. méchanique [Bac54]. Mechanisms [LB66]. mechanistic [Kor30]. mechanistischer [Kor30]. Medal [Ano84c]. media [dB42e]. Medium [SNG65, Bor17, dB49k, dB67b, dB70e]. Meet [Lev93]. meeting [LdB12]. Meetings [de 49]. Meets [Ano45]. Meillet [Ano30b]. meinen [For80]. Meitner [For80]. mekhaniki [dB86]. mémoire [Ano51b]. Memoirs [Kle96]. memories [dB41e, dB42d]. memory [Ano51b]. Men [HG65]. Merwe [Ble86, Sch85a, Tor87]. mes [dB53i]. Meson [Bro86, Ros66, ZS63, dB49d]. mesonic [dBR50]. mésoniques [dBR50]. mesoton [Pro38, Pro38]. message [Ano40]. Mesure [dB57b, Mil58a]. mesurer [TdB71]. Metals [WW88, Bri30a, Bri30b]. Metaphysics [dB58b]. métaux [Bri30a, Bri30b]. Meteorites [HÖ66]. Meteors [HÖ66]. meter [Arn14]. method [Bau54, dB55a, vF82]. méthode [Bau54]. Methodological [Hut54b]. Méthodologie [Hut54b]. Methods [CS67, LH89, RG66]. mettre [dB27k, dB48f, dB48g]. Metz [Gil66]. Michel [Lup49]. Michio [Kar05]. Microdosimetry [RZB97]. microfisica [Vig70]. microparticles [Fil02]. microphysics [BFM80, Bru48, dB47e, dB52d, dB56d, dB78, Bis57, Fes56]. microphysique [Bru48, dB47e, dB52d, dB56d, dB78, Fil47, Boi57]. Microscope [DGR^{+}99, Day09, dB46g, dB47c, dB50a]. Microscopes [Lev88]. Microscopic [Ton90, Wig07]. Microscopies [Sch87b]. microscopy [ZD15]. Middle [MLM^{+}00]. Mie [dH04]. Mike [EP00]. Mikhailov [Cos01]. Milestones [dB78]. milieu [Bor17, dB49k, dB67b, dB70e]. milieux [dB42e]. Mills [Hof17]. Mind [JKA91, HJS70]. Minkowski [dBBH^{+}63, dBHH^{+}63]. mirror [Uts02, ZW92]. Mirrors [LPM98]. mises [de 49]. Misleading [Kie10]. missing [Fer96]. mistakes [Wei06a]. Mitarbeitern [dB65a]. Mixing [KJB^{+}10]. Mlle [Ano52b]. MM [Ano52b]. MM. [Ano39a, Ano39b, Ano39c, Ano58b, dB32m]. Mode [DAJZ96]. Model [Lad96, SS92, Cro04, SC97, Squ92, Sut97, dBdB21, dBV63a, dBBH^{+}63, dBHH^{+}63, dBV63b]. modèle [dBdB21, dBV63a, dBV63b]. models [Kie06]. Modern [BKG94, Čap71b, HC98, Nye97, Stu04, WSAS62, dB27a, Boh57a, Boh57b, Boh61, Boh71, Boh84a, Boh84b, Boh99, Bra09, Gui51, Kar71, Kle79, dB30c, dB41a, dB49l, Des39, Poi25, dB27f]. moderne [Des39, Gui51, Poi25, dB27a, dB27f, dB30c, dB41a, dB49l]. Modified [BJK03, Joh02, JM13]. Modulation [SSDD95]. moléculaire [Met38]. Molecular [EUG98, Gat68, Lin67, Mil67, PDEK06, dB66d, EFS31, EFS32, Met38]. molecules [dB45f, dB45f]. Molekularstrahlen [EFS31, EFS32]. moment [dB32f, dB36b, dB48f, dB48g, dBA66, dB71e]. moments [Geh37, dB70d]. momentum [Bil01, TdB71, dB36b, dB49o, dB49p, dB52f, dBA66, dB70d, dB71e]. momentum-wavelength [Bil01]. Mon [dB70a, dBxx]. Monde [Hum51, Loc04a, Lup49, Spe68, dB37b, dB46b]. Monochromasierung [EFS32]. monochromatic [EFS31, dB47g]. monochromatiques [dB47g]. monochromatischen [EFS31]. Monochromatization [EFS32]. monographies [dBUC^{+}29, dBUC^{+}30]. Monographs [Mor54, Sch85a, dBUC^{+}29, dBUC^{+}30]. Montyon [Ano39a]. Moon [MH79]. moral [dB45d]. morale [dB45d]. Morette [Fed08]. mort [Ano87c, Ano87b, Ano87a]. mosaic [TP43]. mosaique [TP43]. Moseley [Hei74]. Mössbauer [Wer65]. Most [Kha20]. Motion [Cla67, Dic96, Gol94b, LPM98, AKC^{+}89, FF03, Hol92, Hol93, Hol95, Kie06, Met38, dB67c, dB70e]. motion-induced [AKC^{+}89]. motions [Vig94, dB26c]. mouvantes [dB47a]. mouvement [dB41c, dB49o, dB49p, dB67c, dB70e, dB70d, dB71e, Met38]. mouvements [dB26c]. movement [dB41c]. movements [Gar53]. moving [Enz85, Hof17]. moyennes [dB32k]. Mr. [Ano51b, Ano52a, Ano53, Ano55a, Ano59, Ano75, LdB12]. Multilayers [RP97]. Multiparticle [GHZ93]. Multiphoton [FMP99, Ste02]. Multiple [Ber88, Lee96]. Murder [KT66]. muséologie [dB09a]. museology [dB09a, dB09b]. my [Cos82, dB53i, dB70a, dBxx]. Mysteries [Fra05, Gol94b]. Mystery [FR13b]. Myths [GA98].

N [For80]. Nachweis [Hil28]. Nakano [dBHV59]. Nanofabrication [SC90]. nanomanufacturing [IB14]. Narratives [Per18]. Narrow [Khu88]. Nath [Sha74]. National [HW04]. Natural [ABdB^{+}53, Poi74, dB25d]. Nature [Bau02, HM66, Van05, PCR73, PR71, Wig85, dB27d, dB30f, dB30g, dB34h, dB52a]. Navigation [Ano39b]. near [AKC^{+}89, FF03, dBV72a]. nearly [Tho10]. Nécessité [dB79]. nécrologique [dB55c]. Need [Lev01b, Tha93, dB79]. Néel [Ano58b]. Negotiating [AGG^{+}93]. nella [Lab35]. neopredelennosteĭ [dB86]. Neuchatel [Gru54]. Neue [dB35b]. Neuen [May46, dB39c, dB43c]. Neumann [Ber69]. Neutral [BL89, Sch86a]. neutrinienne [Hei53]. Neutrino [Hei53, Moh79, Var02, Bro78, Fie06]. Neutron [Ano66c, Cro87, Day03, KT66, LP85, Sch85c, Wer80, AKC^{+}89, CBPW79, FS77, Sch82, Uts02, dB33e, dB47h, dB49q, Lyn02]. Neutrons [Sch02, Pal63, Sno13]. newcomer [dB51m]. News [Sch87a, Sch88a]. Newton [Joh77, Vig93, dBT44a]. Niels [Shi85, Ano62, Wei14, Whe85, dB63d]. Niemeyer [Gam54, Kae53]. ninety [HJS70]. Nobel [Ano29b, PCR73, Sch87b, Enn87, Lev01a, Sch13, Hum51]. nodal [FF03]. noir [dB22a]. nomenclature [dB58c]. Non [BK76, Kae53, RH67, dB60b, Abe57, Fil02, Ros58, dB53e, dB54c, dB56e]. Non-equilibrium [BK76]. non-identity [Fil02]. Non-Linear [RH67, dB60b]. Non-Mathematical [Kae53, dB54c]. nondispersive [Mac78]. Nonlinear [BB90, KMC05, Bod97, Vig91, YY99, dB53e, dB56e]. Nonlocal [FPN^{+}01]. Nonlocality [Tap22]. nonrelativistic [BB90]. Nonsense [BSM^{+}99]. Noonday [Kae53]. Nordström [KKTY00b, KKTY00a]. North [Gar66]. North-Holland [Gar66]. Norton [Kar05, Mal39]. nos [dB37a]. Notable [Kha20]. Note [EK51, Wil48, dB27j, dB30i, dB51e]. Noted [Ano29b]. Notes [Cas12, EUG98, FL09, Gag53, K.52, Koj96a, Koj96b, For80, Koj96a, Koj96b, dB26d]. Notice [dB55c, dB57d, dB60c, dB62a, dB64d, dB65c, dB67d, dB69a, dB71c, dB71d, dB57d, dB60c, dB62a, dB64d, dB65c, dB67d, dB69a, dB71c]. notion [Bou49, Len48, Pel47b, Pel47a, Ser58, dB47d]. Notions [Hut54b, dBdB34, dB44b]. nous [Ano62]. nouveau [dB33e, dB51m, dBV63a, dBV63b]. Nouvelle [Gom35a, Gom35b, dB62b, dB63e, Ano35b, Ano37a, Ano37b, dB26b, dB26d, dB27k, dB28b, dB32i, dB32b, dB32c, dB34i, dB34j, dB34k, dB37c, dB40b, dB40d, dB42a, dB42f, dB49s, dB49r, dBR50, dB50h, dB54e, dB54f, dB55f, dB61a, dB63c, dB70f, dB78]. Nouvelles [dB36d, dB36e, dB49i, dB56d, dB34d, dB34l, dB42b, dBA57, Boi57]. Novelties [Čap91b, Kim15]. November [LdB12]. novembre [LdB12]. noyau [dB42b, dB43b, dB45b, dB46a]. noyaux [Cos50]. nucléaire [dB51m]. Nuclear [BVH69, GMG^{+}79, McD79, Mil85, Sch84, AKC^{+}89, Kae48, dB43b, dB45b, dB46a, dB51m]. Nuclei [BB86]. Nucleus [FR13b, dB42b]. Numbers [DK93, SNG65, Yan12]. Numerical [MLG67]. Nuove [Vig70]. Nur [RFS^{+}91].

O [And30, Ber69, Ble86, EU26, Sch85a, SI32, Tor87, dB27j]. O. [And30, EU26]. Oak [Lub96b]. obituary [Dit80, Enn87, dB55c]. object [dB66c]. objective [dB52e]. objet [dB66c]. Observation [AKC^{+}89, SLH^{+}06]. Observatory [SM69]. Observed [Bla02, Col95]. Observers [Gol98]. Observing [KRK10]. occasion [Cos85, F^{+}76]. October [Ano53, Far01, LdB12]. octobre [LdB12]. oddities [Kim15]. Odyssey [NPG01]. œuvre [Ano48, dB27f, dB43d, dB44a, dB45f, dB46c, dB47b, dB49c, dB51c, dB53c, dB55b, dB57d, dB60c, dB62a, dB63d, dB64d, dB65c, dB67d, dB69a, dB71c, dB71d]. Œuvres [Poi54]. off [Rub50]. Offer [Lev88, Lev04]. offered [dB97a]. Old [Sch87b, dB34l]. Omission [Per18]. on-line [Lin65a]. Once [Lev97b]. onde [Bor17, Géh33a, Pau53, dB32l, dB32j, dB49k, dB51f, dB58d, dB67c]. Ondes [dB23b, dB26c, dB30d, dB30c, dB41d, dB55d, dB64e, dB68d, dB70b, Bri30a, Bri30b, Gam26, Géh33b, Win34, dB27c, dB27b, dB27l, dB30e, dB43a, dB45a, dB47g, dB47c, dB49g, de 49, dB51d, dB55b, dB62d, dB65b, dB97b, dB97a, Lin69]. Ondulatoire [dB57b, Ano71b, Ano73, Ano87b, Géh38, Mil58a, Ros58, Sch55b, Spe68, Win31, dB26b, dB26d, dB27i, dB27e, dB27k, dB27l, dB27g, dB28a, dB29c, dB30h, dB30a, dB30f, dB30g, dB32a, dB32g, dB32h, dBD35, dB36a, dB39h, dB39a, dB39e, dB40b, dB40d, dB41e, dB42a, dB42f, dB42d, dB43a, dB43b, dB43g, dBT44a, dB45a, dB45b, dB46f, dB46a, dB47g, dB48h, dB48b, dB49h, dB50b, dB52h, dB52a, dB52e, dB53a, dB53e, dB53f, dB53g, dB53d, dB54a, dB54e, dB54b, dB55f, dB55d, dB56c, dB56e, dBA57, dB57c, dB59b, dB59c, dB63a, dB67f, dB68f, dB68b, dB68g, dBA69, dB70f, dB70g, dB71b, dBA71, dB73, dB82, dB98a, Abe57, Bac54, Kle27]. ondulatoires [Buh37]. One [Kae31, Ano92a]. Ontological [BHG94, Dic96, BH93]. opérateur [Gom35c, dBA66]. operator [Gom35c, dBA66]. Opinions [CSR^{+}90]. Oppenheimer [Var02]. opposing [dB27d]. opposition [Loc07]. Optical [CEC72, HKK94, Lev01b, RFS^{+}91, Sch86a, CGLM90, Lee96, dBD22b]. Optics [BL89, Ebe02, KMC05, KJB^{+}10, PSZM02, RMW74, Fie06, dB25a, dB26f, dB29c, dB34a, dB47f, dB50d]. Optique [dB50d, dB25a, dB26f, dB29c, dB34a, dB47f]. optiques [dBD22b]. Orbital [DS64]. Order [CDE12, ERS^{+}17]. Ordered [Lip00]. orientation [Bri54]. origin [doB75]. orthodoxy [Squ92]. oscillante [Bor17]. oscillating [Bor17]. oscillation [Kar29]. Oscillations [BBF95, Fie06]. oscillator [Enz85, Van94]. Oswald [EP00]. Other [Bar90a, Fed08, dB40c]. Otto [And30, For80, BW80].

Our [Kar05, Ber17]. Oursel [Ano52b]. Ouvrages [Ano71b]. Overcomes [Day03]. Overview [HC98, dB53i, dB70g]. Oxford [Eck09, Lyn02].

P [Ano52b, For80, Sch85a, Vis51]. P. [dB51e]. P/b [Sch85a]. Packages [Kae53]. Packard [Lev93]. Packet [App99, FMP99, Mac78, Ste02, Uts02]. Pact [Ste97]. page [Gou74]. pair [ABN01]. pair-breaking [ABN01]. Pairs [ESI02]. Palace [dB09a]. Palais [dB09a, dB09b]. Palo [Lev93]. Paper [Tis65, Wei06b]. paperback [Dic96, Dic96]. Papers [Fow29, ABdB^{+}53, Sel93, dBB28, Bal65]. Paradigm [Kir98]. paradox [Moh79]. Parallelism [For80, dB26f]. parallélisme [dB26f]. paramètres [Pau53]. Parametric [ESI02]. Paris [Lin69, Tis65, dB58a, dB56a, dB57a, dB58a]. parle [Ano48]. parlera [Ano51a]. Part [Gol98, Mor54, dBA71]. Particle [AHH84, BB92, Bod03, CCG^{+}08, DF48, Gud85, RFS^{+}91, Wei11, WN84, dB64f, AD95, AW17, Bou01, BH82b, FF03, FPN^{+}01, Gha06, GH05, GV82, Ley05, NM12, TdB71, Vig94, Vrk51, Wai95, YY99, dBdB21, dBdB22, dB33e, dBdB34, dB39a, dB47f, dB50b, dB51b, dB52h, dB52a, dBA57, dB58c, dB61a, dB61c, dB61b, dB62b, dB62c, dB63e, dBHH^{+}63, dB63f, dB64c, dB66b, dB67c, dB70e, dB27b, Bor17]. Particle-like [BB92]. Particle-wave [Bod03]. particle/waves [Wai95]. Particles [FH64, SL65, BSB99, F^{+}76, HB50, KdB51, Mum65, dB30d, dB33a, dB39h, dB39e, dB39g, dB39f, dB40c, dB41f, dB42b, dB43c, dB43h, dB47c, dB48f, dB48g, dB49s, dBT50, dB50e, dB50f, dB51k, dB52d, dB52c, dB54d, dB62d, dBV63a, dB63g, dB63b, dBBH^{+}63, dBHH^{+}63, dBV63b, dB64b, dB64c, dB65a, dB68c, dBA69, dB71f, dB72b, dB76a, dB97b, dB97a, dB43a, dB45a, Lin64]. particular [Rob17]. Particule [Lin65b, Bou01, GH05, TdB71, Tis65, Vrk51, dB61b, dB62b, dB62c, dB63e, dB63f, dB64c, dB66b, dB67c, dB70e]. Particules [Lin65b, Mum65, Tis65, dB39h, dB39g, dB39f, dB40c, dB41f, dB42b, dB43h, dB48f, dB48g, dB49s, dBT50, dB50e, dB50f, dB51k, dB51b, dB52a, dB52d, dB52c, dB54d, dB58c, dB61a, dBV63a, dB63g, dBV63b, dB64b, dB64c, dB68c, dBA69, dB72b, dB76a]. partie [Ano71b, dBA71]. partisan [Ano55b]. Passage [Pei85, dB33a, dB32d, dB33a]. Passages [Fra05]. passé [TdB71]. passing [AKC^{+}89]. Passion [Eck09, Jon08, Ger87]. past [TdB71]. Path [Fed08, Lev04, Fie06, KRK10]. Pathological [CSR^{+}90]. Pathologies [HDN02]. Paths [Hun66, dB60d]. Patterns [PDEK06]. Paul [Ble86, Sch85a, Tor87, BVV84, Koj96a, Koj96b, dB47b, dBDW^{+}84]. Paulette [MS54]. Pauli [Bal65, vMS01]. Pauling [Pau79]. Pay [Gou74]. Pays [dL81, Ano59]. Pedestrians [LT66]. Peierls [dB32m]. Pelseneer [Bou49, Len48, Ser58]. pendant [dB46d, dB51a]. Penetrating [Hol29]. Pénétration [dB49k, dB49k]. pensait [Koj01]. pensée [Ano87d, Bou49, Pel47a]. penseur [Ano87c, Geo53, L.58]. Pérard [dB67d]. père [Ano87b]. Peres [TBMW05]. perevodov [Suv67]. Pérez [Ano39c]. perfect [Tho10]. Perihelion [SS00]. period [dB51a]. période [dB51a]. periodic [Fer96, dB46e]. périodiques [dB46e]. permanent [dB49a, Ano75]. permanents [dB49a]. Perpétuel [Ano75]. Perrier [Ano39b]. Perrin [dB45f, dB71a]. personal [dB68e, dB41e, dB42d]. personnelles [dB68e]. personnels [dB41e, dB42d]. perspective [Gha14, PNF13]. Perspectives [HC98, Whi62, Boi57, ET07, dB34l, dB56d, dB34l]. perturbations [PNSS12, PNSS14]. Peter [Dic96, Gol94b]. petite [dB47h]. Phase [Ebe02, EU26, NPG01, SSDD95, Sut97, AKC^{+}89, Gam26, LO10, Mil13, SC97, Tak27, Vig94, dB47g, Gar66]. phase-measurement [LO10]. Phasenwellen [EU26]. Phases [dH04, Gam26, vF82]. Phenomena [Bar90a, RH67, Bil01, dB26e, dB32i]. phénomène [Bou49, Len48, Pel47b, Pel47a, Ser58]. phenoménes [dB32i, dB26e]. phenomenon [Fer96, Pel47b, Pel47a]. Philipp [For80]. Philosopher [Sch49a, Sch49b]. Philosopher-Scientist [Sch49a, Sch49b]. Philosophic [Rei44b, Rei98b]. Philosophical [SSL98, dBJ53, dB51j, Wei06b]. Philosophie [dB49l, Ano39c, Car35, Des39, dBJ53, B.36]. philosophique [dB51j]. Philosophy [ABdB^{+}53, CL66, Des39, Fre07, Shi85, dB49l, Cas75, ET07]. Phimsy [EM68, ELF68]. Phonons [RP97]. Photo [Ano82, Ano92b]. photoelectric [Vig71, dBdB22]. photoélectrique [Vig71, dBdB22]. Photoelectron [HS85]. Photoemission [Mar88]. Photon [DAJZ96, ESI02, Esp99, Lub96a, dBV72a, Cos53a, EN52, Gau16, Geh37, Géh38, LSG^{+}08, LO10, Loc95, Loc04b, Pro38, Stu03, SLH^{+}06, SOG06, Vig71, Wil48, dB34i, dB34c, dB34h, dBW34, dB36a, dB37d, dB40b, dB40d, dB42a, dB42f, dB43g, dBT44b, dB49h, dB57c, dBV72b, dB36a]. Photonic [ESI02, JBCY95, KRK10, Sch03]. Photons [Loc04b, Sch88b, CGLM92, Des34, HV91, WZM91, Wil11, dB40c, dB42a, dB42f, dB64e, dB65b, dB68a, dB68d, dB70b, Lin69]. Physical [Bau02, BO67, CMW65, HW04, HC98, LH89, LB66, SSL98, Whi62, dB56a, Fie06, GV82, Loc04a, Pel47b, Pel47a, dB25e, dB37b, dB46b, dB55e, dB57a, dB58a, Hei57]. physicien [Ano51b, Ano87c, Geo53, L.58, dB71a]. Physicist [Enn87, Pei85, Vig82, Ano51b, Ano87c, Geo53, Hei74, dB71a]. Physicists [CvF71, Lee79, Cro01a, GP72]. Physics [AS70, AHH84, Ano29b, Ano32, BVH69, BSS98, BBSR69, Čap71b, Com67, Des39, EUG98, Fre07, GD88, HÖ66, Hut54b, Jac87, Kae53, Kim15, Mal39, McW77, ME76, MS67, Mil85, RFS^{+}91, Sch87a, Sch88a, SSB96, Sch99, Sch85a, SS95, Val92, WSAS62, Whi62, dB27a, dBdB28, dB28c, dB41b, dB46d, dB49e, Ano62, Boh57a, Boh57b, Boh61, Boh71, Boh84a, Boh84b, Boh99, Bra09, BH82b, DFF13, ET07, FS77, Gam85, Ger87, Gid13, Gru74, Gui51, Kae48, Kar71, Kle79, Koj04, KHFA67, Loc93, Lub80, Poi25, Poi74, RM78, Sel93, Sla68, Smy15, Sno13, Suv67, Van79, Van64, Wei11, dB27h, dB27f, dB29a, dB30c, dB32i, dB32b, dB34j, dB35c, dB35a, dB37h, dB37c, dB38c, dB39c, dB39e, dB41a, dB41c, dB42c, dB43c, dB43e, dB45e]. physics [dB48c, dB49f, dB49c, dB49l, dB51m, dB52b, dB53b, dB53j, dB53k, dB54c, dB63c, dB68e, dB76c, dB76d, Cas75, Gar66, Mar64, ME76, Sch87b, Sch13, dB47e, dB58b, Coh40, Din41, Gam54, Mal39, Bis57, Bru48, Fes56].

Physics/Hüfner [RFS^{+}91]. Physik [RFS^{+}91, dB43c, dB39c, May46]. Physik/Kudo [RFS^{+}91]. Physikalische [Str66]. Physique [Ano37a, Ano37b, Ano54, B.36, Des39, DL66b, Hut54a, Pat45, Ros48, dL66a, dB28c, dB38c, dB47e, dB52b, dX33, Ano62, Bla56, Bou49, Bri54, Ger87, Gui51, Koj04, Len48, Loc04a, Mey33, Pel47b, Pel47a, Poi25, Ser58, dB25e, dB27a, dB27h, dB27f, dBdB28, dB29a, dB30c, dB32i, dB32b, dB34j, dB35c, dB35a, dB37h, dB37b, dB37c, dB39e, dB41a, dB41c, dB41b, dB42c, dB43e, dB45e, dB46d, dB46b, dB48c, dB49e, dB49f, dB49c, dB49l, dB51m, dB53b, dB56a, dB63c, dB68e, dB76c, dB76d, Lup49, Bru48, Fil47]. physiques [Hei57, Mil58b, dB55e, dB57a, dB58a, MS54]. Picard [Ano39a, Ano39b, Ano39c, dB43d]. picometer [Day09]. picture [Kor30]. Pierre [Abr88, Ano54, Ble86, Hut54a, Tor87, dB53b, dB61a, dB65a, Mil66, dB65a]. Pilot [All22, HDN02, Nor22, Per18, Tap22, Val92, BH82a, Bus15, Car16, Col12, Dre17, Gha14, Pau53, SS01, ZGWM92, dB51f, dB52i]. Pilot-Wave [All22, HDN02, Nor22, Tap22, Val92, Bus15, Col12, Pau53, dB51f, dB52i]. pilote [Pau53, dB51f, dB52i]. piloted [Vig91]. Pioneer [Bla86]. Planck [dB48d, Cas75, dB48d]. plane [dB47g]. planes [dB47g]. plasma [Fri62]. Plasmaphysik [RFS^{+}91]. plastic [Bil01]. Plates [DGR^{+}99]. plomb [Vig71]. Plumey [Ano39b]. plus [dB32m]. Poincaré [dB46d, Poi74, dB46d, dB97a]. Poincelot [dB53h]. point [dB26f, de 49, dB54b, dB63h, dB38c]. polarisation [dB50g]. polarisée [dBV72b]. Polariton [SL10]. polarization [dB50g]. Polarized [Sch05, Sch84, dBV72b]. policy [MB79]. Polish [SI32]. Politics [GMG^{+}79]. Pomerans [Whi62]. Poncelet [Ano39a]. ponctuelle [dB38c]. Popytka [dB77c]. pore [Sma10]. portée [dB51j]. Portraits [For80, Far01]. Position [Dre97, NMEV12, RM78]. Positron [Lev88]. possibilité [TdB71, dB26e, dB27k, dB48f, dB48g, dBT50, dB51k, dB52e]. possibilités [dB32b, dB46g]. possibilities [dB32b, dB46g]. possibility [TdB71, dB26e, dB27k, dB48f, dB48g, dBT50, dB51k, dB52e]. Possible [Vig94, Bil01, CGS88, Fil02, dB27d, dB47h, dB54e, dB55f]. Postmodern [HT93]. postroenija [dB77c]. postscript [vdM82]. postulate [Kar29]. potential [Cos01, dB32d, dB33a, dB47d, Hum36]. potentials [dB72b]. potentiel [dB32d, dB33a]. potentielle [Cos01, dB47d]. Potentiels [Hum36, dB72b]. pour [GH05, TP43, TdB71, dBT50, dB56c, dB78]. Power [Per18, Kob07]. pp [Dic96, Eck09, Gar66, Gru54, Kae53, Kar05, Kle14, Lin69, Lyn02, Mal39, Mor54, Sch07b, Str65, Str66, Tis65]. Practical [dB51j]. pratique [dB51j]. Pre [Bla86]. Pre-Independence [Bla86]. précédente [dB51e]. Precession [SS00]. Précis [Ber69]. precision [HTH^{+}14]. predicted [dBLBVP76]. prediction [Cos01, Cos01]. predictions [CGLM90]. Preface [Gru54, Coh51]. Preis [Ble86, Deh36]. Premier [dB49l, Ano52a]. première [dB75]. premières [dB32g, dB32h]. Premiers [Bri54]. préparant [dB54a]. preparing [dB54a]. prepotential [Hum36]. prepotentiels [Hum36]. presence [Loc95, Loc95]. Present [dBLBVP76, dB50c]. presentation [dB62b, dB63e, dB70f, dB62b, dB63e, dB70f]. presentations [dB30e, dB34j]. presented [Ano74, ABdB^{+}53]. Présentés [Ano71b]. presents [Ano84c]. President [Ano40]. Press [Ble86, Dic96, Eck09, Kae53, Kar05, Kle14, Lyn02, Mor54, Sch85a]. pression [dB67h]. pressure [dB67h]. Pressures [Gar66]. previous [dB51e]. Prévisionnelle [DL66b]. prévues [dBdB24]. Primitifs [Bou49, Len48, Pel47b, Pel47a, Ser58]. primitives [Pel47b, Pel47a]. primordial [PNSS12, PNSS14]. Prince [Ano51a, Ano58a, Ano29b, Ano45, Ano30a]. Princeton [Kle14]. princip [dB67i]. principe [TP43, dB21b, dB23a, dB38a, dB38b, dB47d]. Principes [Ano71b, dBA71, Hei57, Mer52, Mil58b, dB26b, Pat45, Ros48]. Principle [GHZ93, dB24a, Lib84, Rei44a, Rei98a, TP43, dB23a, dB38a, dB38b, dB47d, dB67i, dB21b, dB90, dH04]. Principles [LS69b, RFS^{+}91, SL65, Hei57, Mer52, dBA71, dB26b]. prises [Koj96a, Koj96b]. Prism [DAJZ96]. privat [For80]. Prix [Ano39a, Ano39b, Ano39c, Ano52a]. Prize [Ano32, Sch87b, Sch13, Ano74, Ano29b, Ano52a, K.52, Lev01a]. Prize-Winner [Ano29b]. probabiliste [dB82]. Probabilistic [RFS^{+}91, dB90, dB82]. probabilités [dB46f, dB48b]. probabilities [dB46f, dB48b, dBLBVP76]. Probing [FCP85]. Problem [Acu16, Nor22, Sto05, BW05, Lee79, PW95, Pau53, dB50g, dB53b, dB59a, dB70a, dB70e, dBxx]. Problème [dB70a, dB50g, Hut54a, Pau53, dB53b, dB59a, dB70e, dBxx]. Problèmes [dB41d, dB49l, dB51d, dB97b, dB97a, dB49l]. Problems [Gut67, RMW74, SCC^{+}65, SS01, dB97b, dB97a, dB41d, dB49l, dB51d]. Processes [LD93, dB74a]. prof [Gom35a, Gom35b]. professé [dB97a]. professées [Bou49, Pel47a]. Professeur [Som11]. Professor [Gom35a, Gom35b, PCR73, Som11]. program [CDW17, MW79]. programmes [Elb13]. Programming [MLG67]. programs [Elb13]. progrès [dB35a, dB40c]. Progress [SH69, dB35a]. Projection [SOG06, LSG^{+}08, SLH^{+}06]. projections [dBdB24]. Projet [Som11]. Prolongation [GF67]. promising [Far17]. prononce [Ano58a, dB45c, dB75]. pronounces [Ano58a]. proof [Hil28]. Propaganda [MLM^{+}00]. Propagation [Rub50, dB41d, dB42e, dB51d]. Propagations [dB41d, dB51d]. propagator [TBMC07]. Properties [BO67, HKK94, Stu03, dBHV59]. property [Gom35c]. Proposed [Ste97]. propre [dB25d, dB48f, dB48g, dB67e]. propriété [Gom35c]. Propriétés [dBHV59]. Prospects [JNA^{+}12, dB56d]. prospettive [Vig70]. Protein [Day03]. proton [dB49q]. Prove [Lev95, Cro87]. Pruniev [Ano52b]. Puankare [Poi74]. publication [Suv67]. publications [For80]. publiés [Car35]. publikacii [Suv67]. pulse [Cos59]. Pulsed [CBPW79]. pure [dB48h, dB48b]. purely [dB49o, dB49p]. purement [dB49o, dB49p]. purs [dB48h, dB48b]. Pursuit [Nye97]. Put [Kae31]. Puzzle [HS05].

QELS [Ano01]. Quandary [Kae31]. Quanta [dB06, JL73, dB23c, dB23d, dB70a, dB04, dB67i, dB67k, And30, ABdBC33, Bac54, Hei57, Koj96a, Koj96b, LdB12, dB22c, dB22a, dB23b, dB23a, dB24d, dB25b, dB25c, dB25a, dB26e, dB28b, dB33c, dB37c, dB53a, dB54c, dB59b, dB60a, dB63c, dBxx, dB77c, dB23a, dB67j, Kae53, Ano37a, Ano37b, Mil58b]. Quantenmechanik [Ano35a, Deh36, Haa34, dB35b, Ber69, Sie97]. Quantenphysik [dB53j, dB53k]. Quantentheorie [dB27m]. quantification [dB32c, dB37d, dB49m]. quantile [dB66b]. Quantique [Ber69, Ano54, Bla56, Cos53a, Geo32, Hut54a, Mey33, Pau53, VV14, dB34d, dB39b, dB40a, dB43g, dB45e, dB49f, dB49h, dB52b, dB53b, dB57c, dBA66]. quantiques [Bor17, Gui51]. quantité [dB49o, dB49p, dB70d, dB71e]. quantities [dB43g]. quantity [dB66b]. quantization [dB49m]. Quantized [GS87, Khu88]. Quantizing [Sto05]. Quantum [Acu16, AGG^{+}93, BPW^{+}71, Bar90a, BVV84, Bar90b, BM88, BBF95, BHG94, BW00, Bri17, BKG94, CD90, CGLM90, CvF71, CFG96, DK93, Dic96, DAJZ96, DFW17, Ebe02, Eck09, ET07, Fie06, F^{+}76, Gol94b, Gol98, Gui51, Han77, HKK94, HK93, HT93, HC98, Jäh10, JM67, Joh02, Jon08, Kar05, Kas13, Kie06, Kle14, Kob07, LS69b, LPM98, Lev96, Lyn02, Mer02, Per18, PNSS12, PNF13, PNSS14, PR71, Rei44b, SK11, Sch85a, SSL98, Sch02, SS00, Sla68, SJ74, Sma10, Smo19b, Sum88, Val92, Var02, Vin22, WW88, WW97, Yan12, ZW92, Zur91, dB24a, dBDW^{+}84, dH04, Ano35a, ABdBC33, Ben10, Blo76, BH93, Bor17, BB68, Col12, CDE12, Cos53a, Cro04, DH02, ECT17, FF03, Fil02, Gam85, Gha14, Gid13, Gri12, Gri13, Gri11, GV82]. quantum [Haa34, Hei57, Hes90, Hol93, Hol95, Hun66, HV13, JM13, KKTY98, KMSU02, Kie10, Koj96a, Koj96b, KHFA67, LO10, MM01, Mat17, NMEV12, Pau53, RC06, Rei98b, Sel93, Sie97, SR13, SC97, Squ92, Vig94, Vig95, VV14, Wei14, dB21a, dB22c, dB24b, dB25b, dB25c, dB26e, dB27m, dB34d, dB35b, dB40a, dB43g, dB45e, dB48d, dB49f, dB49h, dB52b, dB53a, dB53b, dB53j, dB53k, dB57c, dB59b, dBA66, And82, Ble86, Geo32, Tor87, dB25a, dB39b]. Quantum-Dot [Lev96]. Quantum-mechanics [BPW^{+}71]. Quantum-optical [CGLM90]. Quantum-to-classical [PNSS12, PNSS14]. quark [Fie06]. Que [Koj01]. Quebec [Ano74]. quelconque [Géh33b, dB39g, dB39f]. Quelques [Gom35a, Gom35b, KdB51, dB33b, dBdB34, dB62c, dB68e, dBT44b]. Quest [AH65, Ble86, Kle14, MS67, Sch85a, Tor87, BVV84, dBDW^{+}84]. Question [JKA91]. questioned [Gri11]. questions [LKF96, Sel93, dB95, dB98a]. Quiet [AB76].

R [Ano39b, Bal65, Bal72, Dic96, Gol94b]. R. [Vis51]. rôle [dB27l, dB48c]. Rabi [Bru07]. rabot [Poi74]. raccord [dB58d]. Radiation [LD93, Hol29, SR13, dB21a, dB22a, dB27e, dB27g, dB38a, dB38b, dB39b, dB40a, dB65b, LdB12]. radiations [dB21a]. Radioactive [Lub96b]. radioastronomy [BB79]. radioélectricité [dB70b]. radioelectricity [dB70b]. Radium [Kae48]. raies [dB66c]. raised [SS01]. Ralph [Gam54, Kae53]. Randolph [AM94]. Randomness [Kir98]. rang [Bau54]. rank [Bau54]. Raoul [Ano52b]. Rapport [dB28c, Som11]. rapports [LdB12]. Rational [Bri17]. Rauch [Lyn02]. Ray [Sch05, DFF13, dB21b, dBD22b, dB25e, dBdB28]. Raymond [Abr88, Ano52b]. Rayonnement [dB22a, LdB12, dB27e, dB27g, dB38a, dB38b, dB39b, dB40a]. rayonnements [dB65b]. Rayons [dB22b, dB20a, dB21b, dBD22a, dBD24, dBdB24, dBdB28]. Rays [RFS^{+}91, KMC05, Kor63, dB20a, dB22b, dBD22a, dBD24, dBdB24, vF82]. razvitie [Suv67]. Re [Čap71b]. Re-evaluation [Čap71b]. Reaction [Han77, ZB80]. Reading [FL09]. Real [JL73, dB40a]. realism [Smo19a]. Realité [dB34f, dB45f, dB77b, Gil66]. reality [Car16, CFGL97, Cro04, Vig94, dB45f, dB77b, dB34f]. Realization [DGR^{+}99]. Really [Sch88b, Koj01]. realm [Kae48]. reasons [Dar11, Dar13]. Received [Ano64a, Ano64b, Ano65a, Ano65b, Ano66a]. récente [dB29c]. récentes [dB37e]. Récents [dB40c]. réception [dB45c, dB45c]. Recherche [DL66b, dB43f, dB79]. Recherches [dB25b, dB25c, dB76b, dB36d, dB36e]. Recoil [Lub96a]. Recollections [Pei85]. Reconsidering [Jäh10]. Reconstructing [LPM98]. Reconstruction [NM12]. Recover [Ros15]. Recueil [dB30e, dBUC^{+}29, dBUC^{+}30]. recus [Ano71b]. reducibility [dB39h]. réductibilité [dB39h]. reduction [SKFS15, Ste02]. Redundance [Kir98]. réécriture [Dav98]. Réel [Mey33]. réels [dB40a]. Reflection [DEW93, Wit98, dBV72a]. Reflections [dB36f, dB76c, dB77a]. Réflexions [dB36f, dB77a, dB76c]. refractive [dB67b, dB70e]. réfringent [dB67b, dB70e]. refutation [dB74b, dB74b]. refute [Cro87]. regime [Mat17]. region [dB53f, dB53g, dB55d, dB53f, dB53g, dB55d]. regions [dB53d, dB53d]. Regular [dB55d]. régulières [dB55d]. Reinterpretation [Čap71b, dB68b, dB70c, dBA71, dB68b, Ano71b, dBA71]. Reissner [KKTY00b, KKTY00a]. rejected [Bon94]. Related [DF48, dBT44b, dB59a]. relation [BB90, Bil01, dB49m, dB59a, dB65d]. Relations [DF48, Bod97, Bod03, dB32m, dB52g, dB69b]. relationship [Kul70]. Relative [Ben10, dB46e, dB67h]. relatives [dB67h]. relativiste [dB36a, dB36b, dB48i, dB66b, dB67e, dB67h, dB67g, dB68g, dB68c, Ber69]. relativistes [dB95, LKF96]. relativistic [AW17, Ano66d, BB90, Gar53, SS01, Vig91, dB36a, dB36b, dB48i, dBBH^{+}63, dBHH^{+}63, dB66b, dB67e, dB67h, dB67g, dB95, dB68g, dB68c]. Relativistically [DH02]. relativistischen [Gar53]. Relativité [dB33c, dB32a, dB59a]. relativitédans [dB32a]. Relativity [HM84, Jac87, Kar29, Kas13, LS69a, Sel93, Tak27, dB32a, dB59a, For80, dB32a, dB33c]. Relaxation [Col12, ECT17]. relier [dB26e]. remain [dB52b, dB53b, dB53j, dB53k]. remark [Bol50, dB35d, dB51e]. Remarks [BSB99, dB33b, dB49i, dB50e, dB50f, dB59a, dB76d, DdB22, Hei53, Pau53, dBdB22, dB32g, dB32h, dB32f, dB34g, dBT44b, dB46e, dB51f, dB62d]. Remarque [dB51e, Bol50, dB35d]. Remarques [DdB22, Hei53, Pau53, dBdB22, dB26d, dB32g, dB32h, dB32f, dB34g, dBT44b, dB46e, dB50e, dB50f, dB51f, dB62d, dB33b, dB49i, dB59a, dB76d]. Remembering [Bru07]. Reminiscences [Cos82]. Renaissance [Kir99]. rend [Ano59]. René [Cos53b, Gag53, Gru54, Coh51]. répartition [dB72b]. reply [Ano66b]. Réponse [dB27j, Ano66b]. report [KHFA67, Som11, Som34, dB28c]. reports [LdB12]. representation [dB32i, dB32b, dB43g, dBHV59, dB32i, dB32b, dB43g, dBHV59]. Reprint [Ste97]. reproduction [Hut54a, dB53b]. Research [For80, dB36d, dB36e, dB43f, dB79, dB25b, dB25c, dB76b]. Researchers [Lev96]. resigns [Ano75]. résolu [Ano55b]. resolution [Day09]. Resolve [Day03]. Resolved [RFS^{+}91]. Resonance [HL94, AKC^{+}89, Osc01, Osc11]. Resonances [BB86]. Response [dB27j]. rest [And30]. restera [Ano54, Bla56, Hut54a, dB52b, dB53b]. restera-t-elle [Ano54, Bla56, Hut54a, dB52b, dB53b]. restrictive [dB32m]. Results [dB43c, dBV72a, dH04, dB39c]. retirement [ABdB^{+}53]. retrace [Ano53]. returns [PPF^{+}81]. réunion [LdB12]. Réunions [de 49]. Reusing [Ste97]. reveal [Mil13, Wil11]. Review [Abe57, Ano30b, Ano34, Ano35b, Ano37a, Ano37b, Ano52b, Ano54, B.36, Bac54, Bal65, Bal72, Ber69, Bis57, BB73, Bla56, Bla63, Ble86, Boi57, Bri54, Bru48, Car35, Coh40, Coh51, Dan37, Din41, Eck09, Fes56, Fil47, Fow29, Gam54, Gar66, Gat68, Gil66, Gol94b, Gru54, Gud85, Hut54a, Hut54b, Kae53, Kle14, L.58, Len48, Lin64, Lin65c, Lin65b, Lin67, Lin69, Lup49, Lyn02, Mal39, May46, MS54, Mil67, Mil58b, Mil58a, Mor54, Pat45, Pes88, Ros58, Ros48, S.52, Sch85a, Sch07b, Ser58, Shi85, Str65, Str66, Sum88, Swa39, Tis65, Tor87, Vig70, Wer65, Whi62, dX33, HJS70]. Reviewer [WG67]. Reviews [Dic96, For80, Kar05, WBBE30]. Revisited [Kha05, Cos92, Lib84, Sur82]. Revisiting [Wei06b]. Revolution [Gam54, Kae48, Kae53, Smo19b, Gri12, Gri13, dB54c, Kar05]. rewriting [Dav98]. Rey [B.36]. Reynolds [SNG65]. Rhetorical [Per18]. Ridge [Lub96b]. Riemann [Kul70]. Right [Muc79]. rights [Pau79, SW79]. rigorous [dB44b]. rigoureuses [dB44b]. Ring [Sch85c]. Ringuet [DFF13, Jac01]. rise [dB35b, dB43e]. road [Kie10]. Robert [BW80, For80]. Rochot [Ano52b]. Rœntgen [dBD22a]. Roger [Tar00]. Rogue [DFW17]. role [Smy15, dB27l, dB48c]. Roles [BW00]. Röntgen [TP43, dBD22b, dBD24]. Roots [HM84, WN84]. rotateur [dBHV59]. rotating [dB36b, dB59a]. rotation [dB32f, dB36b]. rotatoire [TP43]. Rotator [dBBH^{+}63, dBHV59]. rotatory [TP43]. Routledge [Dic96]. Roy [Ano30b]. Royal [Ano84c]. Ruhemasse [And30]. rule [Dre17]. Run [Lub99]. Russian [Poi74, Suv67, dB67i, dB67j, dB67k, dB77c]. Rutherford [MSB^{+}37, Wei11].

S [Deh36, Gud85, Sch85a]. sa [Ano75, Loc04a]. Sagdeev [Ano93]. same [dB52a]. Samuel [Lyn02]. Satyendra [Sha74]. Savants [Hum51, Spe68, dB51g, S.52]. Saving [Ste97]. says [Ano62]. scalar [Gha06]. Scattering [LP85, Pal63, Sch82]. scenario [PNSS14]. Schäfer [Str66].

Schatzmann [Ano52b]. Schema [dB51h, dB51i]. scheme [dB51i]. Schilpp [For80]. Schladming [Gut67]. scholarships [Ano73]. School [Bou01, Tha93, Smy15, Gut67]. Schrödinger [Ano35a, Deh36, Fow29, Gar53, Haa34, Hou16, Ano35a, BJK03, Bar90b, Cro01b, Dar11, Dar13, Fow29, Gar53, Gri12, Gri13, Haa34, Kle96, Loc87a, RF69, Rei44a, Rei98a, RY13, Vig91]. Schrödingers [RFS^{+}91, Han77]. Schuette [EP00]. Schuster [Sch07b]. Schwinger [Cos53a]. Science [BSM^{+}99, CMW65, CS67, CL66, CSR^{+}90, Del52, Jon08, K.52, Kae31, Kha20, MB79, Nye97, SLF69, TM68, YF97, dB49l, dB56a, dB57a, Ano55a, Ano74, FL09, Lev17, dB41b, dB45d, dB45g, dB47a, dB51a, dB60d, dB66a, Ano55a, Ano75, Kae31, Bla63, Eck09, Gil66]. Sciences [Aug49, Boi57, Bru48, Car35, Dan37, dB49l, dB56a, dB57a, dB58a, Ano39c, Ano58b, Poi74, dB37f, dB45g, Ano75]. Scientific [Bal65, Rod19, Smy15, dB49b, dB09b, Pel47a, dB53i, dB79, ABdB^{+}53, dB43f, dB09a]. scientifique [Bou49, Pel47a, dB43f, dB79, dB09a, Pes88]. scientifiques [Car35, dB53i]. Scientist [Sch49a, Sch49b, Bal72]. Scientists [Bla86, Kae31, Ano40, dB51g]. Scope [dB51j]. se [Bau54]. séance [dB75]. Search [CCG^{+}08, Smo19b]. Second [CSR^{+}90, SLF69, SA65, Loc95, Str66, dB49m]. seconde [dB49m]. Secrétaire [Ano75]. Secretary [Ano75]. seeing [Tso06]. seen [Sch84]. Sein [For80]. seine [For80]. Selected [For80, dBB28, Fow29, Poi74]. self [Dur17]. self-gravitation [Dur17]. Selfintersecting [Hof17]. Selleri [Gud85]. selon [Cos01]. Selvaggi [Vig70]. Semaine [Ano52b]. semiclassical [Mat17]. Semiconductor [RP97, Nar84]. Semiconductors [GAL66, HKK94]. Séminaire [dB55e, dB58a, dB56a, dB57a]. Seminar [dB55e, dB56a, dB57a, dB58a]. Sens [dB51j]. Sense [dB51j, Enz85]. sensitivity [LO10]. Sentiers [Bla63, dB60d]. septuagenarian [Tho67]. Sergescu [Ano52b]. série [DdB22]. Series [SLF69, DdB22, dBD22b, dBD22b]. Serrus [B.36]. ses [Ano51a, Wer65, dB61a]. session [dB75]. Seventh [Far01]. Shadows [LPM98]. shape [Ano66c]. Share [Lev01a]. Shark [WN84]. Sheds [Lad96]. Sheilla [Eck09]. shift [AKC^{+}89]. shifting [dB47a]. shock [dBA69]. shook [Gam85]. Shot [SSP^{+}91]. Should [YF97]. s’humanise [Ano55a]. sich [For80]. Side [For80]. siècle [FL09, dB76b]. signification [Sch55b, dB56c]. signposts [Kie10]. similarities [Bor17]. similitudes [Bor17]. Simon [Sch07b]. Simple [HMH88]. simultanée [dB32b]. simultanément [TdB71]. simultaneous [dB32b]. simultaneously [TdB71]. since [ME76]. sine [Enz85]. sine-Gordon [Enz85]. Single [DAJZ96, Lub96a, Bar90b, FF03]. Single-Photon [Lub96a]. singular [Bal72, Mum65, dB53f, dB53g, dB53d, dB55d, dB56b]. singular-region [dB53f, dB55d]. singular-wave [dB56b]. singulière [dB53f, dB53g, dB55d]. singulières [Mum65, dB53d, dB56b]. Sitter [KKTY00b, KKTY00a]. Six [LO10]. Six-photon [LO10]. Sixty [FR13b]. slit [Bel80]. Slow [Ano66d, Sno13]. Small [HS05, Nor22, dB47h]. Smooth [Wai95]. Socialism [HW04]. sociétés [dBJ53]. societies [dBJ53]. Society [SLF69, Ano84c, HW04]. soir [Ano51a]. Sokal [WW97]. Solar [Ano66e, AB76]. Solid [Lub80, RFS^{+}91, RT75]. Solids [HM66, Pal63]. solitaire [Ano92a]. solitary [Ano92a]. soliton [Siv11]. solitons [BB90, Vig91]. Solution [Ber88, CDW17, Des55b, Des55a, Dur17, Far17, Hou16, PW95, Rob17, Ros58, Uts02, Wit98, dBT44b, dB52i, dB52h, dB52a, dB54a, dB54b, dB54f, dB56b, dB56e, dB58d, dB65b, dB71b, dB98a, Mat17, dB56e]. solutions [KKTY01, KMSU02, Mum65, Vig91]. Solvay [Bri54, Far01, Jäh10, dB28c, LdB12, ME76]. Solving [BW05]. some [SS01, dB41f, dBT44b, Gom35a, Gom35b, KdB51, dB33b, dBdB34, dB62c, dB68e]. son [dB67c, Poi25]. Sootnosheniya [dB86]. Sorbonne [Ano51a]. sortes [dB36f]. sought [Mil13]. Sound [SL69]. Source [Sch85c]. Sources [JSSS71, KHFA67, LP85, CBPW79, Sch82]. sous [Car35, LdB12, dB45c]. soustractif [EK51, Vis51, dB49n, dBR50, dB50h, dB51i, dB51h, dB52g, dB52f, dB98b]. Souvenirs [dB41e, dB42d]. Space [Bau02, Ble86, DS64, Ebe02, HDN02, IS31b, Kar05, Kul71, Sch85a, Sum88, Tor87, dBHH^{+}63, ABN01, ACIB02, BVV84, Fie06, KKTY01, Sut97, Vig94, dB52h, dB54b, dBBH^{+}63, dBDW^{+}84, dB49f]. Space-time [dBHH^{+}63, ACIB02, Fie06, KKTY01]. Spallation [LP85, Sch82, CBPW79]. spark [Lin65a]. spark-chamber [Lin65a]. späten [For80]. spatially [Bod03]. spatially-localized [Bod03]. Spatschek [RFS^{+}91]. speak [Ano51a]. Speakable [BM88]. speaks [Ano48]. Special [Gut67, Jac87, Sch88a, DFW17, Tak27]. spécialistes [Ano73]. specialists [Ano73]. Specially [Ros15]. Speckle [PDEK06]. spectra [dBdB21, dBD22b, dBdB22, dB25e]. spectral [dBD22a, dBD24]. spectres [dBdB21, dBdB22, dB25e]. spectrographes [TP43]. spectrographie [TP43]. spectrographs [TP43]. Spectroscopy [HK93, HS85, HL94, TP43, RFS^{+}91]. speech [dB45c, dB75]. Spezialinventar [For80]. Spheres [SNG65]. Spherical [Kul69, WCÖ67]. Spherically [Gha14, KKTY01]. Spin [Lev97b, Vin22, dB71e, KdB51, Mum65, Sch84, Vrk51, dBW34, dB39h, dB39g, dB39f, dB41f, dB43h, dB48f, dB48g, dBT50, dB50e, dB50f, dB51k, dB51b, dB52c, dB54d, dBBH^{+}62, dB71e]. spin- [KdB51, Mum65, dB39h, dBT50, dB50e, dB50f, dB51b, dB52c]. spinning [AW17]. Spins [dB70d, dB70d]. Spontaneous [ESI02, KJB^{+}10, GM96]. spread [VV14]. Springer [Str65, Str66]. squared [dBA66]. SSC [Lub84]. Stabilität [Gar53]. stability [Gar53]. stages [SR13]. Stand [GD88, Whi62]. standard [LO10]. standing [Lee96]. Starflight [MMJ89]. start [dB41e]. State [RFS^{+}91, ZB80, Kor30, KRK10, Lub80, SKFS15, SLH^{+}06, SOG06, dB32e, dB37a]. state-projection [SLH^{+}06]. State-to-state [ZB80]. statej [Suv67]. States [LPM98, Lev97b, Sch02, JM13, dB56c, dB74a]. static [Cap52]. stationary [dB56c]. stationnaires [dB56c]. statische [Cap52]. Statistical [BK76, Dre88, dB44b]. statistics [Gui51, Wal06, dB48h, dB48b]. statistique [dB48h, dB48b]. Statistiques [Gui51, dB44b]. Stern [And30, And30]. stick [Arn14]. Stifle [YF97]. Still [Vys71, Kae48]. stochastic [GV82, Vig94, Vig95]. Stone [Kle14]. Stopped [Sch85b]. Storage [Sch85c]. Störfall [RFS^{+}91]. Stories [RC06]. Story [Eck09, Jon08, Bru07, Gam85]. strain [Bil01]. Stresa [Cla67]. stress [dB55a]. string [MM01]. Strings [Bos88]. Strong [dB74a]. strongly [Ste03]. structural [dBD22b]. Structure [Bau02, Day03, IS31a, IS31b, MS54, RT75, SS66, SI31, SI32, TP43, dBD21, dBD22b, dBD24, dB27e, dB27g, dB37a, dBT50, dB50e, dB50f, dB51k, Sum88]. structures [dBBH^{+}63]. strukture [SI32]. Studies [Ble86, Ket99, Sch85a, Tor87, BVV84, Win34, de 49, dBDW^{+}84]. Study [Cla67, YF97, Bor17, Wer65, Win31, dB30a, dB30b, dB63a, dB98a]. studying [Ber17]. subcritical [Uts02]. Subjective [LS69a]. subjects [dB76d]. Submissions [Ber88]. Subquantum [Val92]. subtle [BT09]. Subtractive [dB51h, Vis51, dB49n, dBR50, dB50h, dB51i, dB52g, dB52f, dB98b]. successive [dB21a, dB35c, dB37h]. successives [dB21a, dB35c, dB37h]. suite [dB97b]. suivi [Hut54a, dB53b]. sujets [dB76d]. Sul [Car27]. Sun [AB76]. Super [YY99]. superconducting [Mil13, Vig71]. Superconductivity [Bar90a]. superconductors [ABN01]. Superfluid [GS87, SP12, PR71]. Superlattices [LS74]. superluminal [Vig94]. Supernovas [Ano66f]. Superposition [GHZ93]. superpositions [ZW92]. superquantification [Des34]. supersonic [Ben12]. Superstrings [Bos88]. Supplement [Sch88a]. support [MW79]. supraconducteur [Vig71]. surfaces [RT75]. Survey [Kae53, dB49b, dB54c]. sveta [dB67j]. svetovyh [dB77c]. Swabian [Kle14]. symbolic [Pro38]. symbolique [Pro38]. symmetric [Gha14, Gri11, KKTY01]. Symmetrie [Nic54]. Symmetry [BO67, Kul69, Stu03, Nic54]. symposium [dB49l, Cla67]. Synge [Mor54]. synthèse [dB25a, dB27d, Ano52b]. syntheses [dB35c, dB37h, dB35c, dB37h]. synthesis [dB25a, dB27d]. System [DF48, YF97, FPN^{+}01, dBD22a, dBD24, dB48e, dB72b]. système [DF48, dBD22a, dBD24, dB48e, dB54a, dB72b]. systemes [dB46e, dB36a, dB39a, dB39e, dB50b, dB52h, dB52a, dB54b, dBA57]. Systems [AL88, BBF95, RFS^{+}91, BFM80, Mat17, dB36a, dB39a, dB46e, dB50b, dB52h, dB52a, dB54a, dB54b, dBA57].

T [Pat45, Ano54, Bla56, Hut54a, dB50c, dB52b, dB53b]. Table [dBV63b, dBV63b]. tachyon [Cos71]. tachyons [Vys71]. Tait [ABdB^{+}53]. taken [Koj96a, Koj96b]. Takes [GD88]. Talbot [TTS05]. Taton [Ano52b]. Tauber [For80]. Teachers [Tha93]. Teaching [MLM^{+}00, dB49e]. technical [dB43f]. technique [Lin65a, dB43f, dB60a]. technology [McW77]. Teil [For80, May46]. telecommunications [Ang49, Ang52, Ang49, Ang52]. telescope [Ano66e]. Temperature [MS67, ABN01, dB48i, dB66b, dB48i, dB66b]. temps [dB41c, dB49f]. Ten [Eck09, Jon08]. tenseur [Cos59, dB52f]. tenseurs [KdB51]. tensor [Cos59, Gha06, Kul70, dB52f]. tensors [KdB51]. Tentative [dB06, Ros58, dB58d, Abe57, dB24d, dB56e]. tenue [LdB12]. teorema [Car27]. Teoreticheskaya [Poi74]. teorii [dB77c]. teorija [dB67i]. Terms [Dre97, Kar29]. Test [Dre97, CGLM92, CFGL97, HV91, Vig94, WZM91]. testable [GV82]. Textbooks [IB93, NM12]. Th. [Géh33a]. Their [Lev95, LB66, Kul70, dB27d]. theme [BT09]. theorem [Car27, Mum65, dB24c, dBD35, dB46f, dB53h, dB74b]. théorème [Mum65, dB24c, dBD35, dB46f, dB53h, dB74b]. Theoretical [Jac87, MIS69, dB46d, Bor17, Gui51, Smy15, dB34j, dB37f, dB37e, dB48c, dB68e, dB76d, Poi74, Win34, dB49d]. Theoretische [RFS^{+}91]. Theorie [HB50, dB65a, Bac54, DL66b, dB43h, dB54d, dB56a, dB57b, dB98b, dB43b, dB45b, dB46a, ABdBC33, Bou01, Cos53a, Dav98, Des34, Des55b, Des55a, EN52, Gam26, Geh37, Hei53, Hei57, Koj96a, Koj96b, KdB51, LdB12, Loc95, Loc04b, Met38, Mil58b, Mil58a, Pau53, Ros58, Vis51, Vrk51, dB21b, dB22c, dB23a, dB25b, dB25c, dB25a, dB26e, dB27k, dB32k, dB32e, dB32c, dB33d, dB33b, dB34i, dB34b, dB34g, dB34l, dB36e, dB36a, dB37d, dB39f, dB39b, dB40a, dB40b, dB40d, dB40c, dB41f, dB42a, dB42f, dB43g, dBT44b, dB49h, dB49n, dBR50, dB50h, dB51f, dB51i, dB51h, dB51b, dB52i, dB52h, dB52a, dB52g, dB52f, dB52c, dB53a, dB54a, dB54b, dB54f, dB56b, dB56e, dB57c]. théorie [dB58d, dB59a, dB59b, dB61a, dBV63a, dB63h, dB63f, dB65b, dB67b, dB68f, dB71b, dB98a]. Theorien [Ano35a, Deh36, Haa34]. Theories [AHH84, Kae31, SV66, TQ76, MS54, ACIB02, Ano35a, Ano76a, Haa34, dB35c, dB37h, dB42b, dB55e, dB57a, dB58a, Ano76a, Car35, dB35c, dB37h, dB42b, dB55e, dB57a, dB58a]. théorique [Bor17, Gui51, dB34j, dB46d, dB48c, dB68e, dB76d, Ros48, dX33, Pat45]. théoriques [Win34, dB37f, dB37e, dB49d]. Theorists [Lip00]. Theory [Acu16, All22, Bau02, BHG94, Bri17, dB06, Bro86, BKG94, BEC99, CEC72, CFG96, DK93, DRC68, Dic96, Enn87, For80, Gol94b, Gol98, Ham07, Han77, HMH88, HKK94, Hei53, HDN02, Jäh10, Kas13, Lin64, Per18, RFS^{+}91, Ros15, RP53, SS00, Stu04, Tap22, Val92, Var02, WW97, Whi62, dB24a, dB51h, dB56a, dB57b, dB57e, dB04, AW17, ABdBC33, Bar90b, Ber04, BH82a, BH93, Bon94, BB68, Bou01, BSB99, Bus15, Col12, Cos53a, CGLM92, Dav98, Des34, Des55b, Des55a, DH02, Dre17, EN52, Far17, Fil02, Fri62, Gam26, Gam85, Geh37, Gha06, Hei57, HV91, Hol92, Hol93, Hol95, HB50, Hun66, HV13, Kob07, Koj96a, Koj96b, KdB51, LdB12, Lev17, Loc95, Loc04b, Mat17, Met38, Pau53, PNSS12, PNSS14, SS01, TBMC07, Van94, Vis51]. theory [Vrk51, WZM91, Wei14, Wil48, YY99, dB21b, dB22c, dB23a, dB24d, dB25b, dB25c, dB25a, dB26e, dB27k, dB27m, dB32k, dB32e, dB32c, dB33d, dB33b, dB34i, dB34b, dB34g, dB34l, dB36e, dB36a, dB37d, dB39f, dB39b, dB40a, dB40b, dB40d, dB40c, dB41f, dB42a, dB42f, dB43b, dB43g, dB43h, dBT44b, dB45b, dB46a, dB49h, dB49n, dBR50, dB50h, dB51f, dB51i, dB52i, dB52h, dB52a, dB52g, dB52f, dB52c, dB53a, dB54a, dB54b, dB54f, dB54d, dB56b, dB56e, dB57c, dB58d, dB59a, dB59b, dB61a, dBV63a, dB63b, dB63h, dBHH^{+}63, dB63f, dB65b, dB65a, dB67b, dB68f, dB71b, dB98a, dB67i, dB77c, dH04, And82, Dic96, dB51b, dB98b]. Theory/Brückner [RFS^{+}91]. there [Vrk51]. thermal [Ben10, dB55a]. thermodynamical [dB22b]. Thermodynamics [dB64f, Hof17, dB61c, dB62b, dB62c, dB63g, dB63e, dB63f, dB64b, dB64c, dB66b, dB67h, dB67g, dB68g, dB68c, dB76a, dB95, dB61b, dB64c]. Thermodynamique [Lin65b, Tis65, dB68g, LKF96, dB22b, dB61c, dB61b, dB62b, dB62c, dB63g, dB63e, dB63f, dB64b, dB64c, dB66b, dB67h, dB67g, dB68c, dB76a, dB95]. Thermophysics [KR62]. thèse [Koj01]. Thesis [Han77, F^{+}76, Koj01]. They’re [Sch86b]. Thibaud [dB47h]. Things [Lev88]. think [Koj01]. thinker [Ano87c, Geo53, Vig82]. Thirteen [dB76d]. Thirty [Gam85]. Those [YF97, Pro38]. thought [Ano87d, Pel47a]. Thoughts [TJLB89]. Three [Lub96a, LSG^{+}08, Poi74, Vig91]. three-dimensional [Vig91]. three-photon [LSG^{+}08]. Tiger [WN84]. Tilho [Ano39b]. Time [Ble86, CKS^{+}97, Čap91b, Gri11, HDN02, RFS^{+}91, Sch85a, SSL98, SDBP94, Ste97, Tak27, Tor87, ABN01, ACIB02, BVV84, Fie06, KKTY01, Kul71, Lev17, Mil15, dB41c, dB49f, dBHH^{+}63, dBDW^{+}84, Kar05]. Time-Domain [CKS^{+}97]. Time-Resolved [RFS^{+}91]. Time-symmetric [Gri11]. times [Cro01a, Bal72]. tired [Cro04]. Tl [dB43c]. Today [SSL98, Whi62, Bru48]. Tom [Poi74]. tomakh [Poi74]. Tome [dB40d, dB42f]. tomorrow [Ano51a]. Tonnelat [Spe68]. Top [IB14, Kha20]. Top-down [IB14]. torch [EGSR72]. toroidal [Wan09]. total [YY99, dBV72a]. tournant [dB59a]. Toute [dB77b]. traces [Ano53, FL09]. Tracking [DFF13]. Tragedy [Eck09, Jon08, Lad96]. Trägheit [dB48a]. train [Rub50]. Trajectoires [GH05]. Trajectories [Acu16, Bar90b, GH05, JM13]. transformation [Kar29, dB66b, dB67e]. transformations [KdB51, dB21a]. Transformed [Kar05]. transient [dB74a]. Transition [Zur91, Mil13, PNSS12, PNSS14]. Translated [Kae53, Mal39, For80]. translations [Suv67]. translator [Gam54]. transmission [Uts02]. Transport [Khu88]. Trap [Lev97a, Lev97b, Lev01b, Sch86a]. Trapped [Sch85b]. Trapping [Sch86b]. travail [DdB22]. Travaux [dBJ53, dB53i]. traveling [Kar29]. Travelling [Rod19, Wit98]. travers [dB32d, dB33a]. Treaty [Dre97]. Treize [dB76d]. trekh [Poi74]. très [dB47h]. Tribute [Gud85, Ano59, PCR73, MSB^{+}37]. Tributes [MSB^{+}37]. Tricky [AGG^{+}93]. trimers [Mil15]. trouvant [Bau54]. trudy [Poi74]. true [Ger87, dB83]. Tunneling [Mer02]. turn [Sma10]. Two [Gol98, Lev97b, SCC^{+}65, SNG65, SA65, FPN^{+}01, Mum65, Oli66, TP43, dB36f, dB98a, dB27d, dB59a]. two-crystal [TP43]. two-particle [FPN^{+}01]. Type [Lad96, Bod97, Bod03, dB33e].

U [Eck09, Kle14, Lyn02, Haa34, RFS^{+}91]. Ulmo [Ano52b]. Ultimate [II88, Kir98]. Ultra [Sch88b]. Ultra-Energetic [Sch88b]. Ultracold [Ben12, Lev04, Sch02]. ultrahigh [SKFS15]. ultrashort [SKFS15]. Ultrasound [Mil13]. Ulysse [Lup49]. Unbottled [TJLB89]. unbounded [Kae48]. Uncertainties [RFS^{+}91, dB66a, dB82]. Uncertainty [Kae31, dB90, dB32m, dB49m, dB65d, dB69b, DF48]. underlying [dB83].

Understanding [Kar05]. Undivided [BHG94, Dic96, BH93]. Unexpected [Acu16]. Unfinished [Smo19b, Far17]. ung [Ano35a]. Unification [Kir98]. unifié [dB98b]. unified [dB98b]. uniformly [Hof17]. Unifying [II88]. uniqueness [AD95]. Univ [Mor54]. Universal [Kir98]. Universe [BHG94, Dic96, FCP85, Fra05, GA98, Kae31, SDBP94, BH93, Kle27, SR13, dB27i, Sch07b]. universes [MM01]. Université [Pel47a]. University [ABdB^{+}53, Ble86, Dic96, HM79, Pel47a, Sch85a]. Unknown [Sto05, dB30j, dB62e, dB62f]. Unraveling [Fra05]. Unravelling [FR13b]. Unspeakable [BM88]. Untersuchungen [dB27m]. Update [BSS98, SSB96, Sch99, SS95]. upon [LO10]. Urbain [Ano30b]. usage [Ang49, Ang52]. use [Ang49, Ang52, Cos59, dBT44b]. Uses [Cor01]. Using [DGR^{+}99, Dre17, dB53f, dB53g, dB53d]. USSR [Ano58b]. Usual [dB57e, dB57b]. usuelle [Mil58a, dB57b]. utilisation [Cos59].

V [Ano30b, Bal65, Poi74, Suv67, Ano39b]. vacuum [dB40d, dB50g]. valeur [dB45d]. valeurs [dB32k]. Valiant [Kle14]. value [Wei06a, dB45d]. values [dB32k]. variability [Ste35]. variable [PW95, Van94, dB67e]. variables [Pau53]. variance [dB36b, dB48i]. variation [dB48i]. Variations [BT09, BT10]. varie [dB49k]. varies [dB49k]. various [dB76d, dB95, dB97b, dB97a]. Vatican [Klo85]. vecteur [Vrk51]. vector [Vrk51]. Vehicles [DS64]. velocity [Kar29, Tak27, dB47g]. venu [dB51m]. Veranschaulichung [EU26]. verb [Haa34]. verification [dBdB24, dBdB24]. véritables [dB73]. Verlag [Str65, Str66]. Verlagsgesellschaft [Deh36]. verm [Haa34]. vermittels [Cap52]. veroyatnostnaya [dB86]. Versuche [EFS31]. Versus [Kar05, DT74]. very [dB47h]. vi [Mor54, Tis65]. Vialleton [Ano30b]. victim [Bel72]. Victor [Abr88, Ano30a]. vide [dB40d, dB50g]. Vie [Lev96, Ano53, dB43d, dB44a, dB46c, dB47b, dB51c, dB53c, dB57d, dB60c, dB62a, dB63d, dB64d, dB65c, dB67d, dB69a, dB71c, dB71d]. Viel [Ano39b]. View [Lev88, Ton90, dB54b]. viewed [Hun66]. views [dB34l, dB68e]. Vigier [Ano54, Ble86, Lin64, Sch85a, Tor87, dB65a, Eva95, dB51e, dB53b, dB61a, dB63b, dB65a, Hut54a]. Vii [Sch85a]. Villars [Lin69, Tis65]. Villat [Ano39a, Ano39b, Ano39c]. Viscous [SNG65]. Vision [Kar05]. vitesse [dB47g]. Vklad [Suv67]. Vois [dB34l]. Vol [BVH69]. volnovoĭ [dB86]. Volny [dB67k]. Volume [KS62, Sum88, dB43c, Poi74, Sum88, dB40d, dB42f, For80]. volumes [Poi74]. Vorlesungskurs [Str66]. Vorstellung [Kor30]. vortex [Hof17, Wai95]. vortex-like [Wai95]. Vortices [GS87]. Vorwort [Hei39]. Voyage [FL09]. vraie^{≫} [Ger87]. vraiment [Koj01]. vs [Anoxx]. Vue [dB53i, dB70g, dB54b]. vues [dB68e].

W [Bal72, Coh40, Din41, For80, Gam54, Kae53, Kar05, Mal39, Mil58b]. Walter [Sch07b, RO93]. Ward [Ano84c]. Warnow [For80]. Warped [Fra05]. Wars [BSM^{+}99, Sla68]. Was [RF69, Ano55b]. Washington [Kar05, SLF69, SLF69]. Wave [AD95, All22, CKS^{+}97, Cro01b, DAJZ96, Enn87, Esp99, FMP99, Fow29, Gat68, Gud85, HDN02, Kir98, Kir99, KJB^{+}10, Kul71, Lev91, Lin65c, LS69c, NM12, Nor22, Per18, PCR73, RFS^{+}91, RH67, Sch55a, Sch55b, Tap22, Val92, Van05, WN84, WBBE30, dB57b, dB57e, dB66d, dB90, Ano73, Ano87b, Bod03, BH82a, Buh37, Bus15, Car16, Col12, CGLM92, Des55a, DEW93, Dre17, Eva95, Gam26, Géh33a, Gha14, GV82, HV91, Hou16, HTH^{+}14, Kar29, KKTY98, KKTY00b, KKTY00a, Kle27, Kul70, Lab35, Ley05, Loc82, Loc87a, Mac78, Mol95, ODHP00, Pau53, SS01, Ste02, Sur82, SI32, Uts02, Vig91, WZM91, Wan09, Win31, Wit98, YY99, ZGWM92, dB26b, dB26d, dB27i, dB27k, dB27l, dB27g, dBB28, dB29c, dB29b, dB30h, dB30a, dB30b, dB30f]. wave [dB30g, dB32l, dB32j, dB32a, dB32g, dB32h, dB34c, dBdB34, dBD35, dB36a, dB39h, dB39a, dB40b, dB40d, dB41e, dB42f, dB42d, dB43a, dB43b, dB43g, dBT44a, dB45a, dB45b, dB46f, dB46a, dB47g, dB48h, dB48b, dB49k, dB51d, dB51f, dB52i, dB52h, dB52e, dB53a, dB53e, dB53f, dB53g, dB53d, dB54a, dB54e, dB54b, dB55f, dB55d, dB56b, dB56c, dB56e, dBA57, dB58d, dB59b, dB59c, dB60b, dB63a, dB64a, dB67f, dB67c, dB68f, dB68b, dB68g, dBA69, dB70f, dB70c, dB70g, dB71b, dBA71, dB73, dB82, dB83, dB98a, doB75, vdM82, Bor17, Géh38, Lin67, dB27e, dB28a, dB39e, dB42a, dB49h, dB50b, dB52a, dB57c, Mil67]. wave-front [HTH^{+}14, ODHP00]. Wave-Particle [Gud85, AD95, NM12, Ley05]. Wavefield [Kul69]. wavefunction [Ley05]. waveguides [dB68f]. Wavelength [App99, ESI02, FMP99, RP53, Bil01, FPN^{+}01, Gau16, LSG^{+}08, LO10, SKFS15, Sch03, Ste02, SLH^{+}06, SOG06]. wavelets [Bar90b]. Waves [Bau02, Cor01, DFW17, EU26, FS98, IS31a, IS31b, JBCY95, Kae53, Kas13, KJB^{+}10, PDEK06, PSZM02, SSDD95, dB23d, dB71f, Ano35a, AKC^{+}89, Arn14, BB90, Bar90b, Bay07, Bri30a, Bri30b, Cro87, CGS88, CGLM90, CFGL97, Cro04, Dar11, Dar13, DEW93, ERS^{+}17, EFS31, EFS32, FS97, Feo09, Fie06, Fil02, GM96, Haa34, Hil28, Hol87, Hol29, KMM74, Kor30, Kub70, KRK10, Lee96, Lep92, Med74, Rob17, Rub50, SG78, Sch84, Syn54, Syn10, SI31, TTS05, Vig91, Vig94, Wai95, Wig85, Win34, Wit98, ZW92, dB27c, dB27b, dB27l, dB41d, dB43a, dB45a, dB47g, dB47c, dB49g, de 49, dB53f, dB53g, dB53d, dB55b, dB55d, dB62d, dB64e, dB65b, dB68d, dB70b, dB97b, dB97a, dB23b, dB26c, dB30d, dB30c, dB67k, Mor54]. Way [Tay91, CGS88]. Weak [TQ76]. Weapons [GMG^{+}79, McD79]. Wechselwirkung [Cap52, May46, dB43c]. Weisbuch [RFS^{+}91]. Weisskopf [Bal65, Ano93]. Welfare [Mal77]. Wellen [EFS31, EFS32, Hil28, Kor30]. Wellenmechanik [Sch55a, dB29b, WBBE30]. Welt [EU26]. Werner [Lyn02, Loc07, Mil85, PLS02]. wes [Haa34]. Wheeler [Sli76, KKTY01]. Where [Rua69, Whi62, dB49k]. which [dB35b]. white [Van79]. Who [Lip00, RF69, SG78]. width [SKFS15]. Wieman [Lev01a]. Wieviele [RFS^{+}91]. Wigner [Sch85a, BVV84, Ble86, SSP^{+}91, Stu03, Stu04, dBDW^{+}84, Tor87]. Wiley [Gar66]. will [Ano51a, Ano73, dB52b, dB53b, dB53j, dB53k]. Winner [Ano29b]. Wird [dB53j, dB53k]. Wirkungsquantum [dB48d]. Without [Tap22, Gol98, KRK10]. Witness [GD88]. Witt [KKTY01]. Wolfgang [Bal65, vMS01]. Won [Enn87]. Work [BB73, SSL98, dB49b, DdB22, dB27h, dB27f, dB42c, dB43d, dB45f, dB46c, dB49g, dB49c, dB51c, dB53i, dB53c, dB55b, dB57d, dB60c, dB62a, dB64d, dB65c, dB67d, dB69a, dB71c]. workers [dB65a]. works [Ano48, PLS02, Poi74, dB44a, dB47b, dB63d, Poi54, dBJ53]. World [EU26, FH64, GMG^{+}79, Loc04a, dB37b, dB46b]. Worth [Lev95]. wpływie [SI32]. Wrinkles [SDBP94]. writes [Ano55a].

x [vF82, KMC05, Kar05, dB20a, dB21b, dBD22b, dB22b, dBD22a, dBD24, dBdB24, dB25e, dBdB28]. X-ray [dB21b, dBD22b, dB25e, dBdB28]. X-rays [KMC05, dB20a, dB22b, dBD22a, dBD24, dBdB24]. xii [Dic96]. XIII [dB34j]. XVII [Ano39c]. xx [Dic96]. XXe [FL09]. XXXV [dB06].

Yang [Kle99, Hof17]. Year [FR13b, dB55e, dB56a, dB58a, Ano40]. Years [Bla02, Kir99, Mar88, PLS02, RY13, CDW17, Gam85, Geh78, HJS70, Kae48, LvTW09, Mar64, Med74, PCR73, Tso06, dB46d]. Yin [Kle99]. York [Eck09, Gar66, Kae53, Kar05, Lyn02, Mal39, Mor54, Sch07b, Str66]. Yukawa [Bro86].

Zajac [dB30i]. Zeit [For80]. Zero [ERS^{+}17, MS67]. Zero-order [ERS^{+}17]. Zitterbewegung [Hes90]. Zone [DGR^{+}99]. zum [For80]. zur [dB27m, HB50]. Zustandsgleichung [Kor30].

References

Athay:1976:SCC

[AB76] R. G. Athay and Jacques M. Beckers. The solar chromosphere and corona: Quiet sun. Physics Today, 29(11):74–76, November 1, 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

            (electronic). URL http://physicstoday.scitation.org/doi/
            abs/10.1063/1.3024520.

Appleton:1953:SPP

[ABdB^{+}53] Sir Edward Appleton, David Bohm, Louis de Broglie, Richard Courant, Albert Einstein, Pascual Jordan, Theodore V. Kármán, S. S. Penner, Alfred Landé, Erwin Schrödinger, and Hermann Weyl, editors. Scientific papers presented to Max Born on his retirement from the Tait Chair of Natural Philosophy in the University of Edinburgh. Oliver and Boyd, Edinburgh, UK; London, UK, 1953. vi + 94 pp. LCCN QC71 .S35.

Auger:1933:BEI

[ABdBC33] P. Auger, E. Bauer, Louis de Broglie, and M. Courtines. Les bases expérimentales immédiates de la théorie des quanta. (French) [The immediate experimental bases of quantum theory]. Hermann, Paris, France, 1933. 29 pp.

Abele:1957:RTI

[Abe57] Jean Abelé. Review: Une tentative d’interprétation causale et non linéaire de la Mécanique ondulatoire by Louis de Broglie. Archives de Philosophie, 20(2):306–309, 1957. ISSN 0003-9632 (print), 1769-681X (electronic). URL http://www.jstor.org/

stable/43035890.

Agop:2001:CST

[ABN01] M. Agop, C. Gh. Buzea, and P. Nica. Cantorian E^{∞} space–time, de Broglie and the pair-breaking time of high-temperature superconductors. Chaos, Solitons & Fractals, 12(3):571–577, 2001. CO-DEN CSFOEH. ISSN 0960-0779 (print), 1873-2887 (electronic).

Alpher:1993:HHJ

[ABP93] Ralph Alpher, Joseph E. Burke, and Herbert C. Pollock. Henry Hurwitz, Jr. Physics Today, February 1, 1993. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2808823.

Abragam:1988:LVP

[Abr88] A. Abragam. Louis Victor Pierre Raymond de Broglie. 15 August 1892–19 March 1987. Biographical Memoirs of Fellows of the Royal Society, 34:22–41, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://

www.jstor.org/stable/770045. See correction [].

Agop:2002:CST

[ACIB02] M. Agop, G. Ciobanu, P. D. Ioannou, and C. Gh. Buzea. Cantorian E^{∞} space-time in Cartan, de Broglie and field theories. Chaos, Solitons & Fractals, 14(6):863–890, 2002. CODEN CS-FOEH. ISSN 0960-0779 (print), 1873-2887 (electronic).

Acuna:2016:ITB

[Acu16] Pablo Acuña. Inertial trajectories in de Broglie–Bohm quantum theory: An unexpected problem. International Studies in the Philosophy of Science, 30(3):201–230, 2016. CODEN ???? ISSN 0269-8595 (print), 1469-9281 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/02698595.2017.
1316107.

Agnese:1995:WPA

[AD95] A. G. Agnese and E. Di Salvo. Wave-particle association: uniqueness of the de Broglie assumption. Annales de la Fondation Louis de Broglie, 20(3):281–285, 1995. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

Anderson:1993:NTB

[AGG^{+}93] James L. Anderson, GianCarlo Ghirardi, Renata Grassi, Philip Pearle, Nicolas Gisin, David Z. Albert, Gerald Feinberg, Peter R. Holland, Vinay Ambegaokar, Kenneth J. Epstein, and Wojciech H. Zurek. Negotiating the tricky border between quantum and classical. Physics Today, April 1, 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2808860.

Allen:1965:Q

[AH65] Tom Allen and J. Allen Hynek. The quest. Physics Today, 18(11): 74–76, November 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3047003.

Aitchison:1984:GTP

[AHH84] I. J. R. Aitchison, A. G. J. Hey, and Peter Higgs. Gauge theories in particle physics. Physics Today, 37(3):93–95, March 1, 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2916169.

Arif:1989:OMI

[AKC^{+}89] M. Arif, H. Kaiser, R. Clothier, S. A. Werner, W. A. Hamilton, A. Cimmino, and A. G. Klein. Observation of a motion-induced phase shift of neutron de Broglie waves passing through matter near a nuclear resonance. Physical Review A (Atomic, Molecular, and Optical Physics), 39(??):931, ???? 1989. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Altshuler:1988:DES

[AL88] Boris L. Al’tshuler and Patrick A. Lee. Disordered electronic systems. Physics Today, December 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.881139.

Allori:2022:GIP

[All22] Valia Allori. On the Galilean invariance of the pilot-wave theory. Foundations of Physics, 52(5):??, October 2022. CO-DEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

URL https://link.springer.com/article/10.1007/s10701-
022-00631-4.

Abelson:1994:LRH

[AM94] Philip H. Abelson and Nils L. Muench. Lawrence Randolph Hafstad. Physics Today, 47(8):73–74, August 1, 1994. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2808619.

Anderson:1930:FSH

[And30] W. Anderson. Die Formel von O. Stern und die Hypothese von L. de Broglie über die “Ruhemasse” der Lichtquanten. (German) [The formula of O. [Otto] Stern and the hypothesis of L. de Broglie on the “rest mass” of light quanta]. Zeitschrift für Physik, 66:709– 711, 1930. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic).

Anderson:1970:HJD

[And70] Philip W. Anderson. How Josephson discovered his effect. Physics Today, 23(11):23–29, November 1, 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3021826.

eSilva:1982:CLB

[And82] João Andrade e Silva. On the contribution of Louis de Broglie to the quantum theory of measurement. Foundations of Physics, 12 (10):977–987, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889271.

Angot:1949:CMU

[Ang49] Andre Angot. Complements de mathématiques à l’usage des ingénieurs de l’électrotechnique et des télécommunications. (French) [Complements of mathematics for the use of electrotechnical and telecommunications engineers]. Revue d’Optique théorique et Instrumentale, ??(??):??, 1949. Supplement to the January 1949 number. Preface de Louis de Broglie.

Angot:1952:CMU

[Ang52] André Angot. Compléments de mathématiques à l’usage des ingénieurs de l’électrotechnique et des télécommunications. (French) [Complements of mathematics for the use of electrical and telecommunications engineers]. Collection Technique du C.N.E.T. ????, Paris, France, second edition, 1952. vi + 688 pp. Preface by Louis de Broglie.

Anonymous:1929:LBB

[Ano29a] Anonymous. Louis de Broglie — biographical. Nobel Prize Foun-

dation biography., 1929. URL http://www.nobelprize.org/
nobel_prizes/physics/laureates/1929/broglie-bio.html.

Anonymous:1929:PLB

[Ano29b] Anonymous. Prince Louis de Broglie Nobel Prize-winner for physics a noted family. The Observer (1901–2003) [London, UK], ??(??):??, November 17, 1929.

Debroglie:1930:BF

[Ano30a] Anonymous. Prince Louis-Victor de Broglie. Scientific American, 142(3):183, March 1930. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v142/n3/pdf/
scientificamerican0330-183.pdf.

Anonymous:1930:RCD

[Ano30b] Anonymous. Review: Continu et discontinu by Jacques Chevalier; V. Carlhian; Louis de Broglie; G. Urbain; L. Vialleton; Maurice Hauriou; A. Meillet; Ed. Le Roy. Revue de métaphysique et de morale, 37(4):4–5, 1930. ISSN 0035-1571 (print), 2102-5177 (elec-

tronic). URL http://www.jstor.org/stable/40896836.

Anonymous:1932:DBG

[Ano32] Anonymous. Duke de Broglie gets physics prize. New York Times, ??(??):5, December 7, 1932. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.proquest.

com/hnpnewyorktimes/docview/99783079.

Anonymous:1934:RLM

[Ano34] Anonymous. Review: L’électron magnétique by Louis de Broglie. Revue de métaphysique et de morale, 41(3):3–4, 1934. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.jstor.

org/stable/40897292.

Anonymous:1935:MQE

[Ano35a] Anonymous. Materiewellen und Quantenmechanik: eine Einführung auf Grund der Theorien von de Broglie, Schrödinger, Heisenberg und Dirac. (German) [Matter waves and quantum mechanics: an introduction the foundations of the theories of de Broglie, Schrödinger, Heisenberg and Dirac]. Nature, 135(3403):88, January 19, 1935. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v135/n3403/pdf/135088d0.pdf.

Anonymous:1935:RNC

[Ano35b] Anonymous. Review: Une nouvelle conception de la lumière by Louis de Broglie. Revue de métaphysique et de morale, 42(1):1, 1935. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http:

//www.jstor.org/stable/40897356.

Anonymous:1937:RPNa

[Ano37a] Anonymous. Review: La Physique nouvelle et les quanta by Louis de Broglie. Revue de métaphysique et de morale, 44(3):1, 1937. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://

www.jstor.org/stable/40898656.

Anonymous:1937:RPNb

[Ano37b] Anonymous. Review: La Physique nouvelle et les quanta by Louis de Broglie. Revue Philosophique de la France et de l’Étranger, 124 (9–10):118, 1937. ISSN 0035-3833 (print), 2104-385X (electronic).

URL http://www.jstor.org/stable/41084351.

Anonymous:1939:IMP

[Ano39a] Anonymous. II. Mécanique: Prix Montyon, Poncelet, Boileua — MM. É. Picard, L. Lecornu, É. Borel, J. Drach, É. Jouguet, A. de Gramont, H. Villat, L. de Broglie, A. Caquot. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 208(10):695, March 6, 1939. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic). URL http://gallica.bnf.fr/

ark:/12148/bpt6k6238828t/f11/.

Anonymous:1939:VNP

[Ano39b] Anonymous. V. Navigation: Prix Plumey — MM. É. Picard, L. Lecornu, R. Bourgeois, M. Laubeuf, É. Borel, E. Fichot, G. Perrier, J. Drach, É. Jouguet, H. Villat, J. Tilho, L. de Broglie, A. Caquot, G. Durand-Viel, E.-G. Barrillon. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 208(10): 695–696, March 6, 1939. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic). URL http://gallica.bnf.fr/

ark:/12148/bpt6k6238828t/f11/.

Anonymous:1939:XHP

[Ano39c] Anonymous. XVII. Histoire et philosophie des sciences: Prix Binoux — MM. É. Picard, L. Bouvier, É. Borel, E. Caullery, H. Villat, L. de Broglie, C. Pérez. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 208(10):697, March 6, 1939. CODEN COREAF. ISSN 0001-4036 (print), 2419-

6304 (electronic). URL http://gallica.bnf.fr/ark:/12148/
bpt6k6238828t/f13/.

Anonymous:1940:PNY

[Ano40] Anonymous. The President’s New Year message to French scientists. Notes and Records of the Royal Society of London, 3(1): 158, April 1, 1940. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic).

Anonymous:1945:FAM

[Ano45] Anonymous. French Academy meets: Elects Prince Louis, fourth of [de] Broglie family to be honored. New York Times, ??(??): 3, June 2, 1945. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/107296174.

Anonymous:1948:MLB

[Ano48] Anonymous. M. Louis de Broglie parle de l’œuvre d’Einstein. (French) [M. Louis de Broglie speaks of the works of Einstein]. Le Monde [Paris], ??(??):??, December 20, 1948. ISSN 0395-

2037. URL http://www.lemonde.fr/archives/article/1949/
12/20/m-louis-de-broglie-parle-de-l-ouvre-d-einstein_
1931475_1819218.html.

Anonymous:1951:PLB

[Ano51a] Anonymous. Le Prince Louis de Broglie parlera demain soir à la Sorbonne de l’‘énergie atomique et ses applications’. (French) [The Prince Louis de Broglie will speak tomorrow evening at the Sorbonne on ‘atomic energy and its applications’]. Le Monde [Paris], ??(??):??, March 9, 1951. ISSN 0395-2037. URL http:/

/www.lemonde.fr/archives/article/1951/03/09/le-prince-
louis-de-broglie-parlera-demain-soir-a-la-sorbonne-de-
l-energie-atomique-et-ses-applications_2063764_1819218.
html.

Anonymous:1951:MLB

[Ano51b] Anonymous. M. Louis de Broglie célèbre la mémoire du physicien hollandais Lorentz. (French) [Mr. Louis de Broglie celebrates the memory of the Dutch physicist Lorentz [Hendrik Antoon Lorentz (18 July 1853–4 February 1928)]]. Le Monde [Paris], ??(??):??, December 11, 1951. ISSN 0395-2037. URL http:

//www.lemonde.fr/archives/article/1951/12/11/m-louis-
de-broglie-celebre-la-memoire-du-physicien-hollandais-
lorentz_2071476_1819218.html.

Anonymous:1952:MLB

[Ano52a] Anonymous. M. Louis de Broglie est le premier lauréat du Prix Katinga. (French) [Mr. Louis de Broglie is the first laureate of the Katinga Prize]. Le Monde [Paris], ??(??):??, May 30, 1952. ISSN 0395-2037. URL http://www.lemonde.fr/

archives/article/1952/05/30/m-louis-de-broglie-est-le-
premier-laureat-du-prix-katinga_1991186_1819218.html.

Anonymous:1952:RDH

[Ano52b] Anonymous. Review: Dix-huitième Semaine de Synthèse by MM. Kucharski; P. Masson-Oursel; B. Rochot; J. Delevsky; P. Sergescu; Fr. Le Lionnais; E. Schatzmann; Louis de Broglie; Henri Berr; André Lalande; Mlle Delorme; MM. Raoul Ferrier; Raymond Lenoir; Pruniev; Taton; Jean Ulmo. Revue de métaphysique et de morale, 57(3):373, July/September 1952. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.jstor.org/

stable/40899670.

Anonymous:1953:MLB

[Ano53] Anonymous. M. Louis de Broglie retrace la vie d’Aimé Cotton. (French) [Mr. Louis de Broglie traces the life of Aimé Cotton [Aimé Auguste Cotton (9 October 1869–16 April 1951)]].

Le Monde [Paris], ??(??):??, December 15, 1953. ISSN 0395-

2037. URL http://www.lemonde.fr/archives/article/1953/
12/15/m-louis-de-broglie-retrace-la-vie-d-aime-cotton_
1979596_1819218.html.

Anonymous:1954:RPQ

[Ano54] Anonymous. Review: La Physique quantique restera-t-elle indéterministe ? by Louis de Broglie; Jean-Pierre Vigier . Revue de métaphysique et de morale, 59(1):93–94, 1954. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.jstor.

org/stable/40900195.

Anonymous:1955:SIS

[Ano55a] Anonymous. “La science, ici s’humanise’ écrit M. Louis de Broglie. (French) [Science is humanised here, writes Mr. Louis de Broglie]. Le Monde [Paris], ??(??):??, February 24, 1955. ISSN

0395-2037. URL http://www.lemonde.fr/archives/article/
1955/02/24/la-science-ici-s-humanise-ecrit-m-louis-de-
broglie_1945185_1819218.html.

Anonymous:1955:LBE

[Ano55b] Anonymous. Louis de Broglie: Einstein était le partisan résolu du déterminisme fondamental. (French) [Louis de Broglie: Einstein was the partisan of fundamental determinism]. Le Monde [Paris], ??(??):??, April 20, 1955.

ISSN 0395-2037. URL http://www.lemonde.fr/archives/
article/1955/04/20/louis-de-broglie-einstein-etait-le-
partisan-resolu-du-determinisme-fondamental_1959819_1819218.
html.

Anonymous:1958:PLB

[Ano58a] Anonymous. Le Prince Louis de Broglie prononce l’éloge de Frédéric Joliot-Curie. (French) [The Prince Louis de Broglie pronounces the eulogy for Frédéric Joliot-Curie]. Le Monde [Paris], ??(??):??, August 27, 1958. ISSN 0395-2037. URL http:/

/www.lemonde.fr/archives/article/1958/08/27/le-prince-
louis-de-broglie-prononce-l-eloge-de-frederic-joliot-
curie_2299834_1819218.html.

Anonymous:1958:MLB

[Ano58b] Anonymous. MM. Louis de Broglie et Louis Néel élus à l’Académie des sciences de l’U.R.S.S.. (French) [MM. Louis de Broglie and Louis Néel elected at the Academy of Sciences of the USSR]. Le Monde [Paris], ??(??):??, June 24, 1958. ISSN 0395-2037. URL

http://www.lemonde.fr/archives/article/1958/06/24/mm-
louis-de-broglie-et-louis-neel-elus-a-l-academie-des-
sciences-de-l-u-r-s-s_2306557_1819218.html.

Anonymous:1959:MLB

[Ano59] Anonymous. M. Louis de Broglie rend hommage à Frédéric Joliot-Curie. (French) [Mr. Louis de Broglie pays tribute to Frédéric Joliot-Curie]. Le Monde [Paris], ??(??):??, December 15, 1959.

ISSN 0395-2037. URL http://www.lemonde.fr/archives/
article/1959/12/15/m-louis-de-broglie-rend-hommage-a-
frederic-joliot-curie_2164879_1819218.html.

Anonymous:1962:ALD

[Ano62] Anonymous. Avec lui disparait une grande figure de la physique contemporaine nous déclare M. Louis de Broglie. (French) [With him [Niels Bohr] disappears a great figure of contemporary physics, says M. Louis de Broglie]. Le Monde [Paris], ??(??):??, November 20, 1962. ISSN

0395-2037. URL http://www.lemonde.fr/archives/article/
1962/11/20/avec-lui-disparait-une-grande-figure-de-la-
physique-contemporaine-nous-declare-m-louis-de-broglie_
2360411_1819218.html.

Anonymous:1964:BRa

[Ano64a] Anonymous. Books received. Physics Today, 17(2):70–71, February 1, 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.3051423.

Anonymous:1964:BRb

[Ano64b] Anonymous. Books received. Physics Today, 17(4):82–86, April 1, 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3051557.

Anonymous:1965:BRa

[Ano65a] Anonymous. Books received. Physics Today, 18(1):118–122, January 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.3047097.

Anonymous:1965:BRb

[Ano65b] Anonymous. Books received. Physics Today, 18(9):86–90, September 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.3047736.

Anonymous:1966:BR

[Ano66a] Anonymous. Books received. Physics Today, October 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3047792.

Anonymous:1966:RDD

[Ano66b] Anonymous. La réponse du duc Louis de Broglie. (French) [The reply of the Duke Louis de Broglie]. Le Monde [Paris], ??(??):??, October 21, 1966. ISSN 0395-2037. URL http://

www.lemonde.fr/archives/article/1966/10/21/la-reponse-
du-duc-louis-de-broglie_2688062_1819218.html.

Anonymous:1966:NS

[Ano66c] Anonymous. Neutron shape. Physics Today, 19(5):82–85, May 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3048270.

Anonymous:1966:SRE

[Ano66d] Anonymous. Slow relativistic electrons. Physics Today, 19(5): 82, May 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3048269.

Anonymous:1966:ST

[Ano66e] Anonymous. Solar telescope. Physics Today, 19(5):82, May 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3048268.

Anonymous:1966:SMF

[Ano66f] Anonymous. Supernovas may fluoresce. Physics Today, 19(5): 81–82, May 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3048267.

Anonymous:1971:NB

[Ano71a] Anonymous. New books. Physics Today, 24(11):57–59, November 1, 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3022439.

Anonymous:1971:OPO

[Ano71b] Anonymous. Ouvrages présentés ou reçus: Louis de Broglie, La réinterprétation de la mécanique ondulatoire. 1re partie: Principes généraux. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 275(17):3, October 27, 1971. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic). URL

http://gallica.bnf.fr/ark:/12148/bpt6k5618990q/f3/.

Anonymous:1973:FLB

[Ano73] Anonymous. La Fondation Louis-de-Broglie accordera des bourses à des spécialistes de la mécanique ondulatoire. (French) [The Louis-de-Broglie Foundation will award scholarships to specialists in wave mechanics]. Le Monde [Paris], ??(??):??, November 23, 1973. ISSN 0395-2037. URL http://www.lemonde.

fr/archives/article/1973/11/23/la-fondation-louis-de-
broglie-accordera-des-bourses-a-des-specialistes-de-la-
mecanique-ondulatoire_2545509_1819218.html.

Anonymous:1974:QSP

[Ano74] Anonymous. Quebec science prize presented to Bachynski. Physics Today, 27(4):129, April 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3128567.

Anonymous:1975:MLB

[Ano75] Anonymous. M. Louis de Broglie donne sa démission de Secrétaire Perpétuel de l’Académie des Sciences. (French) [Mr. Louis de Broglie resigns as Permanent Secretary of the Academy of Science]. Le Monde [Paris], ??(??):??,

September 24, 1975.  ISSN 0395-2037.  URL http://

www.lemonde.fr/archives/article/1975/09/24/m-louis-de-
broglie-donne-sa-demission-de-secretaire-perpetuel-de-
l-academie-des-sciences_2591883_1819218.html.

Anonymous:1976:TLB

[Ano76a] Anonymous. Les théories de Louis de Broglie. (French) [The theories of Louis de Broglie]. Le Monde [Paris], ??(??):??, July 14, 1976. ISSN 0395-2037. URL http://www.lemonde.fr/

archives/article/1976/07/14/les-theories-de-louis-de-
broglie_2952450_1819218.html.

Anonymous:1976:NB

[Ano76b] Anonymous. New books. Physics Today, 29(6):58–59, June 1, 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3023528.

Anonymous:1980:C

[Ano80] Anonymous. Calendar. Physics Today, 33(5):97–106, May 1, 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2914106.

Anonymous:1982:PLB

[Ano82] Anonymous. Photo of Louis de Broglie. Foundations of Physics, 12(10):ii, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889265.

Anonymous:1984:C

[Ano84a] Anonymous. calendar. Physics Today, 37(1):101–106, January 1, 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2916029.

Anonymous:1984:NB

[Ano84b] Anonymous. New books. Physics Today, 37(5):89–90, May 1, 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2916252.

Anonymous:1984:RSP

[Ano84c] Anonymous. Royal Society presents Hughes Medal to Ward. Physics Today, 37(3):113–115, March 1, 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2916127.

Anonymous:1987:MLBb

[Ano87a] Anonymous. La mort de Louis de Broglie. (French) [The death of Louis de Broglie]. Le Monde [Paris], ??(??):

??, March 21, 1987. ISSN 0395-2037. URL http://

www.lemonde.fr/archives/article/1987/03/21/la-mort-de-
louis-de-broglie_4060719_1819218.html.

Anonymous:1987:MLBa

[Ano87b] Anonymous. La mort de Louis de Broglie, le père de la mécanique ondulatoire. (French) [The death of Louis de Broglie, the father of wave mechanics]. Le Monde [Paris], ??(??):??, March 20, 1987. ISSN 0395-2037. URL http://

www.lemonde.fr/archives/article/1987/03/20/la-mort-de-
louis-de-broglie-le-pere-de-la-mecanique-ondulatoire_
4060668_1819218.html.

Anonymous:1987:MDP

[Ano87c] Anonymous. La mort du physicien Louis de Broglie: Un penseur de la matière. (French) [The death of the physicist Louis de Broglie: A thinker on matter]. Le Monde [Paris], ??(??):??, March 20, 1987. ISSN 0395-2037. URL http://

www.lemonde.fr/archives/article/1987/03/20/la-mort-du-
physicien-louis-de-broglie-un-penseur-de-la-matiere_4051388_
1819218.html.

Anonymous:1987:PLB

[Ano87d] Anonymous. La pensée de Louis de Broglie. (French) [The thought of Louis de Broglie]. Le Monde [Paris], ??(??):??, October 31, 1987. ISSN 0395-2037. URL http://www.lemonde.

fr/archives/article/1987/10/31/la-pensee-de-louis-de-
broglie_4079441_1819218.html.

Anonymous:1992:LBS

[Ano92a] Anonymous. Louis de Broglie, le solitaire. (French) [Louis de Broglie, the solitary one]. Le Monde [Paris], ??(??): ??, December 4, 1992. ISSN 0395-2037. URL http:/

/www.lemonde.fr/archives/article/1992/12/04/louis-de-
broglie-le-solitaire_3924344_1819218.html.

Anonymous:1992:PLB

[Ano92b] Anonymous. Photo of Louis de Broglie. Foundations of Physics, 22(10):i, October 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889709.

Anonymous:1993:WSC

[Ano93] Anonymous. Weisskopf, Sagdeev and Cassidy honored with AIP Awards. Physics Today, 46(2):78, February 1, 1993. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2808812.

Anonymous:1994:NBa

[Ano94a] Anonymous. New books. Physics Today, 47(3):61, March 1, 1994. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2808447.

Anonymous:1994:NBb

[Ano94b] Anonymous. New books. Physics Today, 47(7):69, July 1, 1994. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2808580.

Anonymous:1995:NB

[Ano95] Anonymous. New books. Physics Today, 48(10):71, October 1, 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2808218.

Anonymous:2001:CQC

[Ano01] Anonymous. CLEO/QELS convenes in Baltimore. Physics Today, May 1, 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/full/10.1063/1.1381103.

Anonymous:2013:NB

[Ano13] Anonymous. New books. Physics Today, 66(8):52–53, August 1, 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/PT.3.2088.

Anonymous:2014:AI

[Ano14] Anonymous. Annual index. Physics Today, 67(12):84–94, December 1, 2014. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/PT.3.2632.

Anonymous:20xx:BVB

[Anoxx] Anonymous. de Broglie vs Bohm. Undated Web site, 20xx. URL

http://www.yoga6d.org/debroglie_vs_bohm/.

Appell:1999:BWP

[App99] David Appell. The de Broglie wavelength of a packet of light. Physical Review Focus, 3(??):21, ???? 1999. CODEN PRFHAQ.

ISSN 1539-0748. URL http://link.aps.org/doi/10.1103/
PhysRevFocus.3.21.

Arndt:2014:BMS

[Arn14] Markus Arndt. de Broglie’s meter stick: Making measurements with matter waves. Physics Today, 67(5):30–36, May 1, 2014. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/PT.3.2381.

Adair:1970:CP

[AS70] R. K. Adair and Joel A. Snow. Concepts in physics. Physics Today, 23(7):69–70, July 1, 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3022252.

Auger:1949:AOA

[Aug49] Léon Auger. Analyses d’ouvrages: Almanach des Sciences 1948 par Louis de Broglie; Almanach des Sciences 1949 par Louis de Broglie. Revue d’Histoire des Sciences et de Leurs Applications, 2(4):382–383, September 1949. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23905448.

Alves:2017:RHJ

[AW17] Waldyr Alves, jun.Rodrigues and Samuel A. Wainer. The relativistic Hamilton–Jacobi equation for a massive, charged and spinning particle, its equivalent Dirac equation and the de Broglie– Bohm theory. Adv. Appl. Clifford Algebr., 27(2):1779–1799, 2017. ISSN 0188-7009 (print), 1661-4909 (electronic).

B:1936:RLP

[B.36] G. B. Review: L’Évolution de la Physique et la Philosophie by E. Bauer; Louis de Broglie; Charles Serrus; Leon Brunschvicg; Abel Rey. The Journal of Philosophy, 33(11):305, 1936. ISSN 0022-362X (print), 1939-8549 (electronic). URL http://www.jstor.

org/stable/2016457.

Bacca:1954:RET

[Bac54] Juan David García Bacca. Review: Eléments de Théorie des quanta et de méchanique ondulatoire by Louis de Broglie. Theoria: An International Journal for Theory, History and Foundations of Science, 2(7/8):197–198, 1954. ISSN 0495-4548 (print), 2171-679X

(electronic). URL http://www.jstor.org/stable/23912822.

Balazs:1965:BRW

[Bal65] Nandor L. Balazs. Book review: Wolfgang Pauli and R. Kronig and V. F. Weisskopf, Collected Scientific Papers. Physics Today, 18(6):51–52, June 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3047484.

Balazs:1972:BRS

[Bal72] Nandor Balazs. Book review: A singular scientist: Einstein: The Life and Times, by R. W. Clark. Physics Today, 25(11):51–53, November 1, 1972. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3071093.

Bardeen:1990:SOM

[Bar90a] John Bardeen. Superconductivity and other macroscopic quantum phenomena. Physics Today, 43(12):25–31, December 1, 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881218.

Barut:1990:QTS

[Bar90b] A. O. Barut. Quantum theory of single events: Localized de Broglie wavelets, Schrödinger waves, and classical trajectories. Foundations of Physics, 20(10):1233–1240, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01889467.

Bauer:1954:RAM

[Bau54] Friedrich L. Bauer. Le rang de l’algébre de matrices se trouvant dans la méthode de fusion de M. Louis de Broglie. (French)

[The rank of the matrix algebra in the fusion method of M. Louis de Broglie]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 238(??):1477–1478, ???? 1954. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Baurov:2002:SPS

[Bau02] Yu. A. Baurov. Structure of physical space and nature of de Broglie waves (theory and experiment). Annales de la Fondation Louis de Broglie, 27(2–3):443–462, 2002. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://aflb.

ensmp.fr/AFLB-272/aflb272p443.htm.

Baylis:2007:BWE

[Bay07] William E. Baylis. de Broglie waves as an effect of clock desynchronization. Canadian Journal of Physics = Journal canadien de physique, 85(12):1317–1323, 2007. CODEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (electronic). URL http://www.

nrcresearchpress.com/doi/abs/10.1139/p07-121.

Born:1968:DQT

[BB68] Max Born and Walter Biem. Dualism in quantum theory. Physics Today, 21(8):51–55, August 1, 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3035102.

Born:1971:BELa

[BB71a] Max Born and Hedwig Born, editors. The Born–Einstein letters: correspondence between Albert Einstein and Max and Hedwig Born from 1916–1955. Macmillan Publishing Company, New York, NY, USA, 1971. ISBN 0-333-11267-9. xi + 240 + 5 pp. LCCN QC16.E5 A4513 1971b. Preface by Bertrand Russell and foreword by Werner Heisenberg. Translated by Irene Born.

Born:1971:BELb

[BB71b] Max Born and Hedwig Born, editors. The Born–Einstein letters; correspondence between Albert Einstein and Max and Hedwig Born from 1916 to 1955. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 1971. ISBN 0-8027-0326-7. x + 240 pp. LCCN QC16.E5 A4513.

Blackmore:1973:BRE

[BB73] John T. Blackmore and Jed Z. Buchwald. Book review: Ernst Mach: His Life, Work and Influence. Physics Today, 26(8):56– 57, August 1, 1973. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3128185.

Brecher:1979:BR

[BB79] Kenneth Brecher and Bernard F. Burke. Benefit of radioastronomy. Physics Today, 32(10):15–96, October 1, 1979. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2995235.

Bertsch:1986:GRH

[BB86] George F. Bertsch and Ricardo A. Broglia. Giant resonances in hot nuclei. Physics Today, 39(8):44–52, August 1, 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881038.

Barut:1990:NBW

[BB90] A. O. Barut and B. V. Baby. Nonlinear de Broglie waves and the relation between relativistic and nonrelativistic solitons. Proc. Nonlinear Sci., ??(??):579–588, 1990. Nonlinear evolution equations: integrability and spectral methods, Proc. Workshop, Como/Italy 1988.

Barut:1992:PLC

[BB92] A. O. Barut and A. J. Bracken. Particle-like configurations of the electromagnetic field: An extension of de Broglie’s ideas. Foundations of Physics, 22(10):1267–1285, October 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889713.

Bertsch:1995:OFQ

[BBF95] G. F. Bertsch, R. A. Broglia, and Herman Feshbach. Oscillations in finite quantum systems. Physics Today, 48(5):68–69, May 1, 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2808026.

Born:1969:AP

[BBSR69] Max Born, R. J. (Roger John) Blin-Stoyle, and J. M. Radcliffe. Atomic Physics. Dover books on physics and chemistry. Dover Publications, Inc., New York, NY, USA, eighth edition, 1969. ISBN 0-486-65984-4. xiv + 495 + 11 pp. LCCN QC776

.B5713 1989. US$11.95. URL http://www.loc.gov/catdir/
description/dover031/89012033.html.

Burnett:1999:TBE

[BEC99] Keith Burnett, Mark Edwards, and Charles W. Clark. The theory of Bose–Einstein condensation of dilute gases. Physics Today, December 1, 1999. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.882899.

Bellas:1972:HHV

[Bel72] Fred Bellas. Help for hurricane victim. Physics Today, 25(10):13– 15, October 1, 1972. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3071034.

Bell:1980:BBD

[Bel80] J. S. Bell. De Broglie–Bohm, delayed-choice, double-slit experiment, and density matrix. International Journal of Quantum Chemistry, 18(S14):155–159, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Bennett:2010:RDQ

[Ben10] Andrew F. Bennett. Relative dispersion and quantum thermal equilibrium in de Broglie–Bohm mechanics. Journal of Physics A: Mathematical and Theoretical, 43(19):19, 2010. CODEN JPAMB5. ISSN 1751-8113 (print), 1751-8121 (electronic).

Benka:2012:UCS

[Ben12] Stephen G. Benka. Ultracold chemistry in supersonic beams. Physics Today, 65(12):21, December 1, 2012. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/PT.3.1811.

Bergmann:1969:BRJ

[Ber69] Peter G. Bergmann. Book review: Johann von Neumann, Mathematische Grundlagen der Quantenmechanik and O. Costa de Beauregard, Précis de Mécanique Quantique Relativiste. Physics Today, 22(5):85, 87, May 1, 1969. CODEN PHTOAD.

ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3035594.

Berezin:1988:MSA

[Ber88] Alexander A. Berezin. Multiple submissions: An abstract solution. Physics Today, 41(1):128, January 1, 1988. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2811305.

Berger:2004:EBB

[Ber04] Jorge Berger. Extension of the de Broglie–Bohm theory to the Ginzburg–Landau equation. Foundations of Physics Letters, 17 (3):287–294, 2004. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Berry:2017:ASO

[Ber17] Michael Berry. Approaches to studying our history. Physics Today, 70(3):11–12, March 1, 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/PT.3.3478.

Blaquiere:1980:DSM

[BFM80] A. Blaquiere, F. Fer, and A. Marzollo, editors. Dynamical systems and microphysics, volume 261 of CISM Courses and Lectures (International Centre for Mechanical Sciences). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ix + 412 pp. Foreword by Louis de Broglie.

Bacry:1979:EB

[BH79] Henri Bacry and Banesh Hoffmann. Einstein and de Broglie. Physics Today, 32(8):11–13, August 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2995657.

Bohm:1982:BPW

[BH82a] D. J. Bohm and B. J. Hiley. The de Broglie pilot wave theory and the further development of new insights arising out of it. Foundations of Physics, 12(10):1001–1016, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889273;
http://scalettar.physics.ucdavis.edu/p298/pilotwavetheory.
pdf.

Brown:1982:BEP

[BH82b] Laurie M. Brown and Lillian Hoddeson. The birth of elementary particle physics. Physics Today, April 1, 1982. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2915009.

Bohm:1993:UUO

[BH93] David Bohm and B. J. (Basil J.) Hiley. The undivided universe: an ontological interpretation of quantum theory. Routledge & Kegan Paul, London, UK and New York, NY, USA, 1993. ISBN 0-415-12185-X (paperback), 0-415-06588-7 (hardcover). xii + 397 pp. LCCN QC174.12 .B6326 1993. URL http://www.loc.gov/

catdir/enhancements/fy0916/91021387-d.html.

Bohm:1994:UUO

[BHG94] David Bohm, Basil J. Hiley, and Sheldon Goldstein. The undivided universe: An ontological interpretation of quantum theory. Physics Today, 47(9):90, September 1, 1994. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2808635.

Billingsley:2001:PIB

[Bil01] J. P. Billingsley. The possible influence of the de Broglie momentum-wavelength relation on plastic strain “autowave” phenomena in “active materials”. International Journal of Solids and Structures, 38(24–25):4221–4234, 2001. CODEN IJSOAD. ISSN 0020-7683 (print), 1879-2146 (electronic).

Biser:1957:BRP

[Bis57] Erwin Biser. Book review: Physics and Microphysics by Louis De-Broglie; Martin Davidson. Philosophy of Science, 24(3):281–282, July 1957. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X

(electronic). URL http://www.jstor.org/stable/185149.

Baker-Jarvis:2003:MBA

[BJK03] James Baker-Jarvis and Pavel Kabos. Modified de Broglie approach applied to the Schrödinger and Klein–Gordon equations. Physical Review A (Atomic, Molecular, and Optical Physics), 68 (??):042110, ???? 2003. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Balescu:1976:ENE

[BK76] R. Balescu and Allan N. Kaufman. Equilibrium and nonequilibrium statistical mechanics. Physics Today, 29(11):76, November 1, 1976. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3024522.

Berman:1997:AI

[BK97] Paul R. Berman and Vasili Kharchenko. Atom interferometry. Physics Today, 50(12):68–70, December 1, 1997. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.882030.

Brown:1994:QFT

[BKG94] Lowell S. Brown, Michio Kaku, and O. W. Greenberg. Quantum field theory and quantum field theory: A modern introduction. Physics Today, 47(2):104–106, February 1, 1994. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2808409.

Balykin:1989:LON

[BL89] Victor I. Balykin and Vladilen S. Letokhov. Laser optics of neutral atomic beams. Physics Today, 42(4):23–28, April 1, 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881212.

Blanche:1956:RPQ

[Bla56] R. Blanché. Review: La physique quantique restera-t-elle indéterministe? by Louis de Broglie. Revue Philosophique de la France et de l’Étranger, 146:421–422, 1956. ISSN 0035-3833 (print), 2104-385X (electronic). URL http://www.jstor.org/

stable/41088435.

Blanche:1963:RSS

[Bla63] R. Blanché. Review: Sur Les Sentiers de la Science by Louis de Broglie. Revue Philosophique de la France et de l’Étranger, 153: 378, 1963. ISSN 0035-3833 (print), 2104-385X (electronic). URL

http://www.jstor.org/stable/41089327.

Blanpied:1986:PSP

[Bla86] William A. Blanpied. Pioneer scientists in pre-independence India. Physics Today, 39(5):36–44, May 1, 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.881025.

Blau:2002:EDL

[Bla02] Steven K. Blau. Electron diffraction by light, envisioned 70 years ago, is observed at last. Physics Today, January 1, 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/1.1457252.

Bleyer:1986:BRA

[Ble86] Ulrich Bleyer. Book review: Asim O. Barut, Alwyn Van der Merwe and Jean-Pierre Vigier (eds.): Quantum, Space and Time — The Quest Continues. Studies and Essays in Honour of Louis de Broglie, Paul Dirac and Eugene Wigner. Cambridge University Press, Cambridge 1984, ISBN 0-521-31911-0, Preis: £25. Astronomische Nachrichten, 307(3), 1986. CODEN ASNAAN. ISSN 0004-6337 (print), 1521-3994 (electronic).

Bloch:1976:HED

[Blo76] Felix Bloch. Heisenberg and the early days of quantum mechanics. Physics Today, 29(12):23–27, December 1, 1976. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3024633.

Bell:1988:SUQ

[BM88] J. S. Bell and N. David Mermin. Speakable and unspeakable in quantum mechanics. Physics Today, 41(10):89, October 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2811599.

Bhagavantam:1967:CSP

[BO67] S. Bhagavantam and Henry M. Otte. Crystal symmetry and physical properties. Physics Today, 20(6):75–77, June 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3034370.

Bodurov:1997:BTR

[Bod97] T. G. Bodurov. de Broglie-type relations from nonlinear evolution equations. International Journal of Theoretical Physics, 36(8): 1771–1785, 1997. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Bodurov:2003:PWD

[Bod03] Theodore Bodurov. Particle-wave duality for spatially-localized complex fields. II. de Broglie-type relations. Annales de la Fondation Louis de Broglie, 28(2):163–172, 2003. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://aflb.

ensmp.fr/AFLB-282/aflb282p163.htm.

Bohm:1957:CCMa

[Boh57a] David Bohm. Causality and chance in modern physics. Routledge and Paul, London, UK, 1957. 170 pp. LCCN QC6 .B597 1957.

Bohm:1957:CCMb

[Boh57b] David Bohm. Causality and chance in modern physics. Van Nostrand, Princeton, NJ, USA, 1957. 170 pp. LCCN QC6 .B597 1957a.

Bohm:1961:CCM

[Boh61] David Bohm. Causality and chance in modern physics. Harper torchbooks. The science library. Harper, New York, NY, USA, 1961. 170 pp. LCCN QC6.4.C3 B63x 1961.

Bohm:1971:CCM

[Boh71] David Bohm. Causality and chance in modern physics. University of Pennsylvania Press, Philadelphia, 1971. 170 pp. LCCN QC6 .B6755 1971.

Bohm:1984:CCMa

[Boh84a] David Bohm. Causality and chance in modern physics. Routledge and Kegan Paul, London, UK, 1984. ISBN 0-7102-0031-5 (paperback). xv + 170 pp. LCCN QC6.4.C3 B63x 1984.

Bohm:1984:CCMb

[Boh84b] David Bohm. Causality and chance in modern physics. University of Pennsylvania Press, Philadelphia, third edition, 1984. ISBN 0-8122-1002-6. xi + 170 pp. LCCN QC6 .B6755 1984.

Bohm:1999:CCM

[Boh99] David Bohm. Causality and chance in modern physics. University of Pennsylvania Press, Philadelphia, 1999. ISBN 0-585-12650-X (e-book), 0-8122-1002-6. xi + 170 pp. LCCN QC6.4.C3 B64 1999eb.

Boirel:1957:RNP

[Boi57] René Boirel. Review: Nouvelles perspectives en microphysique, collection Sciences d’aujourd’hui by Louis de Broglie. Les Études philosophiques, 12(1):76–77, 1957. ISSN 0012-2017 (print),

1746-8361 (electronic). URL http://www.jstor.org/stable/
20842216.

Boll:1950:DCD

[Bol50] Marcel Boll. Une description corpusculaire et déterministe des franges de Fraunhofer. Avec une remarque de Louis de Broglie. (French) [A corpuscular and deterministic description of Fraunhofer fringes. With a remark by Louis de Broglie]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 230 (??):1034–1036, ???? 1950. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Bonk:1994:WBT

[Bon94] Thomas Bonk. Why has de Broglie’s theory been rejected? Studies in History and Philosophy of Science Part A, 25(3):375–396, June 1994. CODEN SHPSB5. ISSN 0039-3681 (print), 1879-2510

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0039368194900582.

Borghesi:2017:DOC

[Bor17] C. Borghesi. Dualité onde–corpuscule formée par une masselotte oscillante dans un milieu élastique: étude théorique et similitudes quantiques. (French) [Wave–particle duality formed by an oscillating mass in an elastic medium: theoretical study and quantum similarities]. Annales de la Fondation Louis de Broglie, 42(1):161–196, 2017. CODEN ???? ISSN 0182-4295 (print),

2108-6397 (electronic). URL http://aflb.ensmp.fr/AFLB-421/
aflb421m7.htm.

Bostick:1988:CSS

[Bos88] Winston H. Bostick. Cosmic strings and superstrings. Physics Today, 41(1):128–129, January 1, 1988. CODEN PHTOAD.

ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2811306.

Bouvier:1949:AOL

[Bou49] Robert Bouvier. Analyses d’ouvrages: L’Évolution de la notion de phénomène physique, des Primitifs à Bohr et Louis de Broglie. Leçons sur l’histoire de la pensée scientifique, professées à l’Université libre de Bruxelles par Jean Pelseneer. Revue d’Histoire des Sciences et de Leurs Applications, 2(3):283–286, mai–août 1949. CODEN RHSAAM. ISSN 0048-7996 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23904777.

Boudet:2001:TIP

[Bou01] Roger Boudet. La théorie intrinsèque de la particule de Dirac et “l’École Louis de Broglie”. (French) [The intrinsic-particle theory of Dirac and ‘the School of Louis de Broglie’]. Annales de la Fondation Louis de Broglie, 26(1–3):95–114, 2001. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://

aflb.ensmp.fr/AFLB-26j/aflb26jp095.htm.

Ballentine:1971:QMD

[BPW^{+}71] Leslie E. Ballentine, Philip Pearle, Evan Harris Walker, Mendel Sachs, Toyoki Koga, Joseph Gerver, and Bryce DeWitt. Quantum-mechanics debate. Physics Today, 24(4):36–44, April 1, 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3022676.

Brandt:2009:HCD

[Bra09] Siegmund Brandt. The harvest of a century: discoveries of modern physics in 100 episodes. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-954469-7 (hardcover). xiv + 500 pp. LCCN QC7 .B64 2009.

Brillouin:1930:EMC

[Bri30a] L. Brillouin. Les électrons dans les métaux et le classement des ondes de de Broglie correspondantes. (French) [Electrons in metals and classification of the corresponding de Broglie waves]. C. E., 191(??):292–294, 1930.

Brillouin:1930:EMCb

[Bri30b] L. Brillouin. Les électrons dans les métaux et le classement des ondes de de Broglie correspondantes. (French) [Electrons in metals and the classification of the corresponding de Broglie waves]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 191(??):292–294, July 28, 1930. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic). URL http://

gallica.bnf.fr/ark:/12148/bpt6k31445/f292/.

Bridgman:1954:BRP

[Bri54] P. W. Bridgman. Book review: Les Premiers congrès de physique Solvay et l’orientation de la physique depuis 1911 by Maurice de Broglie. Isis, 45(3):314–315, September 1954. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/226726.

Bricmont:2017:BBT

[Bri17] Jean Bricmont. The de Broglie–Bohm theory as a rational completion of quantum mechanics. Canadian Journal of Physics = Journal canadien de physique, ??(??):??, 2017. CO-DEN CJPHAD. ISSN 0008-4204 (print), 1208-6045 (elec-

tronic). URL http://www.nrcresearchpress.com/doi/abs/10.
1139/cjp-2017-0192.

Brown:1978:IN

[Bro78] Laurie M. Brown. The idea of the neutrino. Physics Today, 31(9): 23–28, September 1, 1978. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2995181.

Brown:1986:HYM

[Bro86] Laurie M. Brown. Hideki Yukawa and the meson theory. Physics Today, 39(12):55–62, December 1, 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.881048.

Brown:1996:LBC

[Bro96] Andrew P. Brown. Liverpool and Berkeley: The Chadwick– Lawrence letters. Physics Today, 49(5):34–40, May 1, 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881497.

Brunet:1948:AOP

[Bru48] Pierre Brunet. Analyses d’ouvrages: Physique et microphysique, Collection Sciences d’aujourd’hui par Louis de Broglie. (French) [Book review: Physics and microphysics, Collection Sciences of today by Louis de Broglie]. Revue d’Histoire des Sciences et de Leurs Applications, 1(3):286–287, January 1948. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL

http://www.jstor.org/stable/23903897.

Brush:2007:RRC

[Bru07] Stephen G. Brush. Remembering Rabi: A challenge and a ghost story. Physics Today, June 1, 2007. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.2754581.

Brown:1999:RIP

[BSB99] Harvey R. Brown, Erik Sjöqvist, and Guido Bacciagaluppi. Remarks on identical particles in de Broglie–Bohm theory. Physics Letters A, 251(4):229–235, 1999. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Bricmont:1999:SWC

[BSM^{+}99] Jean Bricmont, Alan Sokal, Simon C. Moss, Hector J. Sussmann, Dan T. Abell, Jonathan Katz, and David Mermin. The science wars continue, with debate on “fashionable nonsense”. Physics Today, October 1, 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.882844.

Benka:1998:PU

[BSS98] Stephen G. Benka, Phillip F. Schewe, and Benjamin P. Stein. Physics update. Physics Today, 51(8):9, August 1, 1998. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2805883.

Batelaan:2009:ABE

[BT09] Herman Batelaan and Akira Tonomura. The Aharonov–Bohm effects: Variations on a subtle theme. Physics Today, 62(9):38–43, September 1, 2009. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/1.3226854.

Batelaan:2010:MVA

[BT10] Herman Batelaan and Akira Tonomura. More variations on Aharonov–Bohm. Physics Today, April 1, 2010. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://physicstoday.scitation.org/doi/full/10.1063/1.
4797322.

Buhl:1937:ACO

[Buh37] A. Buhl. Analogies corpusculaires et ondulatoires. (French) [Corpuscular and wave analogies], volume 34 of Mém. Sci. physiques. ????, ????, 1937. vi + 62 pp. Préface de L. de Broglie.

Bush:2015:NWP

[Bus15] John W. M. Bush. The new wave of pilot-wave theory. Physics Today, August 1, 2015. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/PT.3.2882.

Baranger:1969:ANP

[BVH69] Michel Baranger, Erich Vogt, and Evans Hayward. Advances in nuclear physics, vol. 1. Physics Today, April 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3035500.

Barut:1984:QST

[BVV84] A. O. (Asim Orhan) Barut, Alwyn Van der Merwe, and Jean-Pierre Vigier, editors. Quantum, space, and time–the quest continues: studies and essays in honour of Louis de Broglie, Paul Dirac, and Eugene Wigner. Cambridge monographs on physics. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-31911-0 (paperback). vii + 659 pp. LCCN QC174.125 .Q376 1984.

Bethe:1980:ORF

[BW80] H. A. Bethe and George Winter. Otto Robert Frisch. Physics Today, 33(1):99–100, January 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/physicstoday.scitation.org/doi/abs/10.1063/1.2913924.

See correction [Dit80].

Born:2000:CRQ

[BW00] Gustav Born and Alvin M. Weinberg. Copenhagen — roles in quantum mechanics. Physics Today, 53(7):74–75, July 1, 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.1292494.

Brown:2005:SMP

[BW05] Harvey R. Brown and David Wallace. Solving the measurement problem: de Broglie–Bohm loses out to Everett. Foundations of Physics, 35(4):517–540, 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Cap:1952:SWL

[Cap52] Ferdinand Cap. Über die statische Wechselwirkung von Leptonen vermittels eines de Broglie-Feldes. (German) [On the static interaction of leptons by means of a de Broglie field]. Acta Physica Austriaca, 6:35–44, 1952. CODEN APASAP. ISSN 0001-6713.

Capek:1971:BLB

[Čap71a] Milič Čapek. Bergson and Louis de Broglie. In Bergson and Modern Physics: A Reinterpretation and Re-evaluation [Čap71b], chapter 38, pages 292–301. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346. LCCN ???? URL https://link.

springer.com/chapter/10.1007/978-94-010-3096-0_38.

Capek:1971:BMP

[Čap71b] Milič Čapek. Bergson and Modern Physics: A Reinterpretation and Re-evaluation, volume 7 of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1971. ISBN 94-010-3098-7, 94-010-3096-0 (e-book). ISSN 0068-0346. xvi + 418 pp. LCCN

???? URL https://link.springer.com/book/10.1007/978-
94-010-3096-0.

Capek:1991:BLB

[Čap91a] Milič Čapek. Bergson and Louis de Broglie. In The New Aspects of Time: Its Continuity and Novelties [Čap91b], chapter 15, pages 286–295. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-

0346. LCCN ???? URL https://link.springer.com/chapter/
10.1007/978-94-009-2123-8_15.

Capek:1991:NAT

[Čap91b] Milič Čapek. The New Aspects of Time: Its Continuity and Novelties, volume 125 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1991. ISBN 94-010-7455-0, 94-009-2123-3 (e-book). ISSN 0068-0346. 380 (est.) pp. LCCN

???? URL https://link.springer.com/book/10.1007/978-
94-009-2123-8.

Carrelli:1927:STD

[Car27] A. Carrelli. Sul teorema della concordanza delle fasi di de Broglie. (Italian) [On a concordance theorem of de Broglie]. Nuovo Cimento, n. Ser., 4:137–141, 1927. ISSN 0029-6341 (print), 1827-6121 (electronic).

Carre:1935:BRA

[Car35] J.-R. Carré. Book review: Actualités scientifiques et industrielles. 184. Exposés de philosophie des Sciences, publiés sous la direction de Louis de Broglie.... III. La causalité des théories mathémaliques by Georges Bouligand. Revue philosophique de la France et de l’Étranger, 119(3–4):273–274, March/April 1935. ISSN 0035-3833 (print), 2104-385X (electronic). URL http://www.jstor.org/

stable/41084151.

Cardoso:2016:RBP

[Car16] António Cardoso. The reality of de Broglie’s pilot wave. Annales de la Fondation Louis de Broglie, 41(1):1–7, 2016. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://

aflb.ensmp.fr/AFLB-411/aflb411m830.htm.

Casertano:1975:FFE

[Cas75] Giovanni Casertano. Fisica e filosofia: Einstein, Infeld, de Broglie, Heisenberg, Planck. (Italian) [Physics and philosophy: Einstein, Infeld, de Broglie, Heisenberg, Planck]. Edizione Il Tripode, Napoli, Italia, 1975. 151 pp. LCCN QC6 .F55.

Castner:2012:NAL

[Cas12] Theodore Castner. Notes on Anderson localization. Physics Today, 65(5):10–11, May 1, 2012. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/PT.3.1538.

Carpenter:1979:PSN

[CBPW79] John M. Carpenter, Thomas H. Blewitt, David L. Price, and Samuel A. Werner. Pulsed spallation neutron sources. Physics Today, 32(12):42–49, December 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2995314.

Catillon:2008:SBP

[CCG^{+}08] P. Catillon, N. Cue, M. J. Gaillard, R. Genre, M. Gouanère, R. G. Kirsch, J.-C. Poizat, J. Remillieux, L. Roussel, and M. Spighel. A search for the de Broglie particle internal clock by means of electron channeling. Foundations of Physics, 38(7):659–664, July 2008. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10701-008-9225-1.

Capasso:1990:QED

[CD90] Federico Capasso and Supriyo Datta. Quantum electron devices. Physics Today, February 1, 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.881226.

Contopoulos:2012:OBB

[CDE12] G. Contopoulos, N. Delis, and C. Efthymiopoulos. Order in de Broglie–Bohm quantum mechanics. Journal of Physics A: Mathematical and Theoretical, 45(16):21, 2012. CODEN JPAMB5. ISSN 1751-8113 (print), 1751-8121 (electronic).

Colin:2017:BDS

[CDW17] S. Colin, T. Durt, and R. Willox. L. de Broglie’s double solution program: 90 years later. Annales de la Fondation Louis de Broglie, 42(1):19–72, 2017. CODEN ???? ISSN 0182-4295 (print),

2108-6397 (electronic). URL http://aflb.ensmp.fr/AFLB-421/
aflb421m3.htm.

Caldwell:1972:TOA

[CEC72] D. J. Caldwell, H. Eyring, and Tai Yup Chang. The theory of optical activity. Physics Today, 25(11):53–54, November 1, 1972. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3071094.

Cushing:1996:BMQ

[CFG96] James T. Cushing, Arthur Fine, and Sheldon Goldstein. Bohmian Mechanics and Quantum Theory: an Appraisal, volume 184 of Boston Studies in the Philosophy of Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 90-481-4698-4, 94-015-8715-9 (ebook). ISSN 0068-0346. viii + 408 pp. LCCN QC173.96-

174.52. URL https://link.springer.com/book/10.1007/978-
94-015-8715-0.

Croca:1997:ETR

[CFGL97] J. R. Croca, M. Ferrero, A. Garuccio, and V. L. Lepore. An experiment to test the reality of de Broglie waves. Foundations of Physics Letters, 10(5):441–447, October 1997. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Croca:1990:QOP

[CGLM90] J. R. Croca, A. Garuccio, V. L. Lepore, and R. N. Moreira. Quantum-optical predictions for an experiment on the de Broglie waves detection. Foundations of Physics Letters, 3(6):557–564, December 1990. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Croca:1992:CET

[CGLM92] J. R. Croca, A. Garuccio, V. L. Lepore, and R. N. Moreira. Comment on “Experimental test of the de Broglie guided-wave theory for photons”. Physical Review Letters, 68(??):3813, ???? 1992. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. See [WZM91].

Croca:1988:PWD

[CGS88] J. R. Croca, A. Garuccio, and F. Selleri. On a possible way to detect de Broglie’s waves. Foundations of Physics Letters, 1(2): 101–112, June 1988. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Cahn:1997:TDB

[CKS^{+}97] S. B. Cahn, A. Kumarakrishnan, U. Shim, T. Sleator, P. R. Berman, and B. Dubetsky. Time-domain de Broglie wave interferometry. Physical Review Letters, 79(??):784, ???? 1997. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

Colodny:1966:BEC

[CL66] Robert G. Colodny and R. B. Lindsay. Beyond the edge of certainty: Essays in contemporary science and philosophy. Physics Today, 19(9):102–103, September 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3048437.

Clancy:1967:ASE

[Cla67] Edward P. Clancy. Astronomers study Earth motion at Stresa Symposium. Physics Today, 20(7):107–109, July 1, 1967. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3034385.

Cleveland:1965:PS

[CMW65] John M. Cleveland, Charles E. Merrill, and Robert L. Weber. Physical science. Physics Today, 18(2):62, February 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3047194.

Cohen:1940:BRM

[Coh40] I. Bernard Cohen. Book review: Matter and Light. The New Physics by Louis de Broglie; W. H. Johnston. Isis, 32(2):378–380, ???? 1940. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/226269.

Cohen:1951:BRH

[Coh51] I. Bernard Cohen. Book review: Histoire de la mécanique. Préface de Louis de Broglie by Réne Dugas. Isis, 42(3):271–272, October 1951. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/226588.

Collins:1995:GBE

[Col95] Graham P. Collins. Gaseous Bose–Einstein condensate finally observed. Physics Today, 48(8):17–20, August 1, 1995. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2808119.

Colin:2012:RQE

[Col12] Samuel Colin. Relaxation to quantum equilibrium for Dirac fermions in the de Broglie–Bohm pilot-wave theory. Proceedings of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 468(2140):1116–1135, 2012. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Combrisson:1967:PF

[Com67] Jean Combrisson. Physics in France. Physics Today, 20(11):55–61, November 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3034023.

Corngold:2001:MUD

[Cor01] Noel Corngold. Many uses for diffusion waves. Physics Today, March 1, 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/full/10.1063/1.4796264.

CostadeBeauregard:1950:CED

[Cos50] O. Costa de Beauregard. Connexion entre la définition des noyaux de Feynman et celle des grandeurs de champ de L. de Broglie. (French) [Connection between the kernel definition of Feynman and that of the field magnitudes of L. of Broglie]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 230 (??):2073–2075, ???? 1950. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

CostadeBeauregard:1953:ITD

[Cos53a] Olivier Costa de Beauregard. Sur l’introduction de la théorie du photon de M. L. de Broglie dans l’électromagnétisme quantique de Schwinger. (French) [On the introduction of the photon theory of M. L. de Broglie in the quantum electromagnetism of Schwinger]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 236(??):2215–2217, ???? 1953. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Costabel:1953:AOH

[Cos53b] Pierre Costabel. Analyses d’ouvrages: Histoire de la mécanique par René Dugas; Louis de Broglie. Revue d’Histoire des Sciences et de Leurs Applications, 6(1):72–74, January 1953. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23903511.

CostadeBeauregard:1959:UDT

[Cos59] Olivier Costa de Beauregard. Sur l’utilisation du tenseur d’impulsion-énergie électromagnétique de Louis de Broglie en électromagnétisme classique. (French) [On the use of the pulse– energy electromagnetic tensor of Louis de Broglie in classical electromagnetism]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 249(??):2285–2286, ???? 1959. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

CostadeBeauregard:1971:FT

[Cos71] O. Costa de Beauregard. Fresnel and the tachyon. Physics Today, 24(3):14, March 1, 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3022614.

CostadeBeauregard:1982:RME

[Cos82] O. Costa de Beauregard. Reminiscences on my early association with Louis de Broglie. Foundations of Physics, 12(10):963–969, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF01889269.

CostadeBeauregard:1985:ODA

[Cos85] O. Costa de Beauregard. A l’occasion du 90^{e} anniversaire de Louis de Broglie. (French) [On the occasion of the 90th birthday of Louis de Broglie]. Revue d’Histoire des Sciences, 38(3–4): 365–367, July/December 1985. CODEN RHSAAM. ISSN 0151-4105 (print), 1969-6582 (electronic). URL http://www.jstor.

org/stable/23632504.

CostadeBeauregard:1992:EGI

[Cos92] O. Costa de Beauregard. Electromagnetic gauge as an integration condition: de Broglie’s argument revisited and expanded. Foundations of Physics, 22(12):1485–1494, December 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01883735.

CostadeBeauregard:2001:DNP

[Cos01] O. Costa de Beauregard. Distribution de l’énergie potentielle selon Louis de Broglie. Effet Mikhailov, prédiction d’un effet angulaire analogue. (French) [Distribution of potential energy according to Louis de Broglie. Mikhailov effect, prediction of an analogous angular effect]. Annales de la Fondation Louis de Broglie, 26(4):765– 773, 2001. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (elec-

tronic). URL http://aflb.ensmp.fr/AFLB-264/aflb264p765.
htm.

Courtoy:1967:C

[Cou67] C. Courtoy. A correction. Physics Today, 20(1):12, January 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3034119.

Croca:1987:NIC

[Cro87] J. R. Croca. Neutron interferometry can prove (or refute) the existence of de Broglie’s waves. Foundations of Physics, 17(10): 971–980, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF00938007.

Cropper:2001:GPL

[Cro01a] William H. Cropper. Great physicists: the life and times of leading physicists from Galileo to Hawking. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-513748-5. xii + 500 pp. LCCN QC15 .C76 2001. URL http:

//www.loc.gov/catdir/enhancements/fy0637/2001021611-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2001021611-b.html.

Cropper:2001:WME

[Cro01b] William H. Cropper. Wave mechanics: Erwin Schrödinger and Louis de Broglie. In Great physicists: the life and times of leading physicists from Galileo to Hawking [Cro01a], pages 275– 292. ISBN 0-19-513748-5. LCCN QC15 .C76 2001. URL http:

//www.loc.gov/catdir/enhancements/fy0637/2001021611-d.
html;   http://www.loc.gov/catdir/enhancements/fy0724/
2001021611-b.html.

Croca:2004:BTL

[Cro04] J. R. Croca. de Broglie tired light model and the reality of the quantum waves. Foundations of Physics, 34(12):1929–1954, 2004. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Cohen:1967:DMM

[CS67] J. B. Cohen and Martin E. Straumanis. Diffraction methods in materials science. Physics Today, 20(6):79–81, June 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3034373.

Cooper:1990:SOP

[CSR^{+}90] Christopher Cooper, Peter A. Sturrock, K. Ramakrishna Rao, Peter R. Phillips, Dean Radin, Robert G. Fleagle, Albert A. Bartlett, and Michael Wixom. Second opinions on ‘pathological science’. Physics Today, 43(3):13–112, March 1, 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2810480.

Cropper:1971:QP

[CvF71] W. Cropper and Thomas von Foerster. The quantum physicists. Physics Today, 24(2):43–44, February 1, 1971. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3022566.

Domokos:1996:ABW

[DAJZ96] P. Domokos, P. Adam, J. Janszky, and A. Zeilinger. Atom de Broglie wave deflection by a single cavity mode in the few-photon limit: Quantum prism. Physical Review Letters, 77(??):1663, ???? 1996. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Danic:1937:RML

[Dan37] D. Danic. Review: Matière et lumière. (Collection Sciences d’aujourd’hui) by Louis de Broglie; A. Georges. Thalès, 4:282,

1937. ISSN 0398-7817. URL http://www.jstor.org/stable/
43861511.

Darrigol:2011:FRW

[Dar11] Olivier Darrigol. A few reasons why Louis de Broglie discovered [de] Broglie’s waves and yet did not discover Schrödinger’s equation. Mitschnitt einer gemeinsamen Veranstaltung des Erwin Schrödinger-Instituts für Mathematische Physik (ESI) und der Fakultät für Physik der Universität Wien am Donnerstag, dem 13. Jänner 2011 im Kleinen Festsaal der Universität Wien [Recording of a joint event at the Erwin Schrödinger Institute for Mathematical Physics (ESI) and the Department of Physics, University of Vienna on Thursday, the 13th January 2011 in the Small Hall of the University of Vienna]., 2011.

Darrigol:2013:FRW

[Dar13] Olivier Darrigol. A few reasons why Louis de Broglie discovered Broglie’s waves and yet did not discover Schrödinger’s equation. In Reiter and Yngvason [RY13], pages 165–174. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). LCCN QC16.S265 I58 2011.

Daviau:1998:ATL

[Dav98] Claude Daviau. Application à la théorie de la lumière de Louis de Broglie d’une réécriture de l’équation de Dirac. (French) [Application to Louis de Broglie’s theory of light in a rewriting of the Dirac equation]. Annales de la Fondation Louis de Broglie, 23(3–4):121–127, 1998. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

Day:2003:NDO

[Day03] Charles Day. Neutron diffraction overcomes flux limits to resolve a large protein structure. Physics Today, 56(11):17–19, November 1, 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/1.1634521.

Day:2009:EMA

[Day09] Charles Day. Electron microscope attains 50-picometer resolution. Physics Today, 62(4):17, April 1, 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.4797094.

deBroglie:1920:ARX

[dB20a] Louis de Broglie. Sur l’absorption des rayons X par la matière. (French) [On the absorption of X-rays by matter]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 171 (??):1137–??, ???? 1920. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1920:CFL

[dB20b] Louis de Broglie. Sur le calcul des fréquences limites d’absorption K et L des éléments lourds. (French) [On the calculation of the limiting frequencies of the K and L lines of heavy elements]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 170(??):585–??, ???? 1920. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1921:DDQ

[dB21a] Louis de Broglie. Sur la dégradation du quantum dans les transformations successives des radiations de haute fréquence. (French) [On the degradation of the quantum in the successive transformations of high-frequency radiation]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 173(??):1160– ??, ???? 1921. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1921:TAR

[dB21b] Louis de Broglie. Sur la théorie de l’absorption des rayons X par la matière et le principe de correspondance. (French) [On the theory of X-ray absorption by matter and the Correspondence Principle]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 173(??):1456–??, ???? 1921. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1922:RNQ

[dB22a] Louis de Broglie. Rayonnement noir et quanta de lumière. (French) [Blackbody radiation and light quanta]. Journal de Physique et le Radium, 3(11):422–428, ???? 1922. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1922:RXE

[dB22b] Louis de Broglie. Rayons X et équilibre thermodynamique. (French) [X-rays and thermodynamical equilibrium]. Journal de Physique et le Radium, 3(2):33–45, 1922. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1922:ITQ

[dB22c] Louis de Broglie. Sur les interférences et la théorie des quanta de lumière. (French) [On the interferences and quantum theory of light]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 175(??):811–813, November 6, 1922. CO-DEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

URL http://gallica.bnf.fr/ark:/12148/bpt6k3128v/f811.
item.r=de%20Broglie.

deBroglie:1923:QTC

[dB23a] Louis de Broglie. Les quanta, la théorie cinétique des gas et le principe de Fermat. (French) [Quanta, the kinetic theory of gases, and the Fermat principle]. Comp. Rend., 177(??):630–??, ???? 1923.

deBroglie:1923:OQ

[dB23b] Louis de Broglie. Ondes et quanta. (French) [Waves and quanta]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 177(??):517–??, ???? 1923. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1923:QLD

[dB23c] Louis de Broglie. Quanta de lumière, diffraction et interférences. (French) [Light quanta, diffraction, and interference]. Comp. Rend., 177(??):548–??, ???? 1923.

deBroglie:1923:WQ

[dB23d] Louis de Broglie. Waves and quanta. Nature, 112(2815):540, October 13, 1923. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

deBroglie:1924:BTO

[dB24a] Louis de Broglie. de Broglie’s theory of the quantum and the Doppler principle. Nature, 114(2854):51–52, July 1924. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

deBroglie:1924:DDQ

[dB24b] Louis de Broglie. Sur la dynamique du quantum de lumière et les interférences. (French) [On quantum light dynamics and interference]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 179(??):1039–??, ???? 1924. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1924:TB

[dB24c] Louis de Broglie. Sur un théorème de Bohr. (French) [On a theorem of Bohr]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 179(??):676–??, ???? 1924. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1924:TTL

[dB24d] Louis de Broglie. A tentative theory of light quanta. Philosophical Magazine (6), 47(??):446–458, ???? 1924. CODEN PHMAA4. ISSN 1941-5982 (print), 1941-5990 (electronic).

deBroglie:1925:TQS

[dB25a] Louis de Broglie. La théorie des quanta, synthèse de la dynamique et de l’optique. (French) [Quantum theory, the synthesis of dynamics and optics]. Revue gènèrale des Sciences, 35(22):629–??, ???? 1925.

deBroglie:1925:RTQa

[dB25b] Louis de Broglie. Recherches sur la théorie des quanta. (French) [Research on quantum theory]. Doctoral thesis, La Sorbonne, Paris, France, 1925. ???? pp. URL https://www.annphys.org/

articles/anphys/abs/1925/03/anphys19251003p22/anphys19251003p22.

html. Presented at the Sorbonne on 25 November 1924. German translation in [dB27m]. English translation in [dB04].

deBroglie:1925:RTQb

[dB25c] Louis de Broglie. Recherches sur la théorie des quanta. (French) [Research on quantum theory]. Ann. Physique, X. Sér, 3(??):22– 128, January/February 1925. ISSN 0003-4169 (print), 1286-4838 (electronic). This is the author’s doctoral dissertation.

deBroglie:1925:FPE

[dB25d] Louis de Broglie. Sur la fréquence propre de l’électron. (French) [On the natural frequency of the electron]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 180(??):498, ???? 1925. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1925:IPS

[dB25e] Louis de Broglie. Sur l’interprétation physique des spectres X d’acides gras. (French) [On the physical interpretation of the Xray spectra of fatty acids]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 180(??):1485–??, ???? 1925. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1926:ICL

[dB26a] Louis de Broglie. Interference and corpuscular light. Nature, 118 (2969):441–442, September 25, 1926. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

deBroglie:1926:PNM

[dB26b] Louis de Broglie. Les principes de la nouvelle mécanique ondulatoire. (French) [Principles of the new wave mechanics]. Journal de Physique et le Radium, 7(11):321–337, 1926. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1926:OMF

[dB26c] Louis de Broglie. Ondes et mouvements. (French) [Waves and motions]. Number 1 in Collection de Physique mathématique. Gauthier-Villars, Paris, France, 1926. vi +133 pp.

deBroglie:1926:RNM

[dB26d] Louis de Broglie. Remarques sur la nouvelle mécanique ondulatoire. (French) [Notes on the new wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 183(??), ???? 1926. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1926:PRP

[dB26e] Louis de Broglie. Sur la possibilité de relier les phénomènes d’interférence et de diffraction à la théorie des quanta de lumière. (French) [On the possibility of linking the phenomena of interference and diffraction in the quantum theory of light]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 183(??):447–??, ???? 1926. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1926:PED

[dB26f] Louis de Broglie. Sur le parallélisme entre la dynamique du point matériel et l’optique géométrique. (French) [On the parallelism between the dynamics of the matter point and geometric optics]. Journal de Physique et le Radium, 7(1):1–6, 1926. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1927:CIP

[dB27a] Louis de Broglie. Continuité et individualité dans la physique moderne. (French) [Continuity and individuality in modern physics]. Cahier de la nouvelle journée, 15(??):60–??, ???? 1927.

deBroglie:1927:COF

[dB27b] Louis de Broglie. Corpuscule et ondes Ψ. (French) [Particle and waves]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 185(??):1180–??, ???? 1927. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1927:CO

[dB27c] Louis de Broglie. Corpuscules et ondes. (French) [Corpuscles and waves]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 185(??), ???? 1927. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1927:DCA

[dB27d] Louis de Broglie. Deux conceptions adverses sur la nature de la lumière et leur synthèse possible. (French) [Two opposing conceptions about the nature of light and their possible synthesis]. Scientia (Milan), 42:127–134, September 1927. CODEN SCIMAI. ISSN 0036-8687 (print), 1825-4373 (electronic).

deBroglie:1927:MOS

[dB27e] Louis de Broglie. La mécanique ondulatoire et la structure atomique de la matière et du rayonnement. (French) [Wave mechanics and the atomic structure of matter and radiation]. Journal de Physique et le Radium, 8(5):225–241, 1927. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1927:PMO

[dB27f] Louis de Broglie. La physique moderne et l’œuvre de Fresnel. (French) [Modern physics and the work of Fresnel]. Revue de métaphysique et de morale, 34(4):421–440, 1927. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.jstor.org/

stable/40896501.

deBroglie:1927:SAM

[dB27g] Louis de Broglie. La structure atomique de la matière et du rayonnement et la mécanique ondulatoire. (French) [The atomic structure of matter and radiation and wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 184 (??):273–??, ???? 1927. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1927:LUF

[dB27h] Louis de Broglie. L’œuvre de Fresnel et l’évolution actuelle de la physique. (French) [The work of Fresnel and the current evolution of physics]. Revue d’Optique théorique et Instrumentale, 6(??): 493–??, ???? 1927. Lecture on the occasion of the centenary of Fresnel, 29 October 1927.

deBroglie:1927:LCD

[dB27i] Louis de Broglie. L’univers à cinq dimensions et la mécanique ondulatoire. (French) [The five-dimensional universe and wave mechanics]. Journal de Physique et le Radium, 8(2):65–73, 1927. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1927:RNM

[dB27j] Louis de Broglie. Réponse à la note de M. O. Klein. (French) [Response to the note of M. O. Klein]. Journal de Physique et le Radium, 8(5):244, 1927. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1927:PMA

[dB27k] Louis de Broglie. Sur la possibilité de mettre en accord la théorie électromagnetique avec la nouvelle mécanique ondulatoire. (French) [On the possibility of bringing electromagnetic theory into line with the new wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 184(??):81–??, ???? 1927. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1927:ROC

[dB27l] Louis de Broglie. Sur le rôle des ondes continues en mécanique ondulatoire. (French) [On the role of continuous waves in wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 185(??):380–??, ???? 1927. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1927:UQG

[dB27m] Louis de Broglie. Untersuchungen zur Quantentheorie. (German) [Investigations on quantum theory]. Akademische Verlagsgesellschaft, Leipzig, Germany, 1927. ii + 88 pp. Translation by German W. Becker.

deBroglie:1928:MOF

[dB28a] Louis de Broglie. La mécanique ondulatoire. (French) [Wave mechanics]. Number 1 in Mémorial des Sciences physiques. Gauthier-Villars, Paris, France, 1928. iii + 54 pp.

deBroglie:1928:NDQ

[dB28b] Louis de Broglie. La nouvelle dynamique des quanta. (French) [The new dynamics of quanta]. In V. Conseil de Physique Solvay, pages 105–132. ????, ????, 1928.

deBroglie:1928:RAC

[dB28c] Louis de Broglie. Rapport au 5ème Conseil de Physique Solvay, Bruxelles, Belgium.(French) [Report to the 5th Solvay Physics Congress, Brussels, Belgium]. Gauthier-Villars, Paris, France, 1928. 105- pp.

deBroglie:1929:DCP

[dB29a] Louis de Broglie. Déterminisme et causalité dans la physique contemporaine. (French) [Determinism and causality in contemporary physics]. Revue de métaphysique et de morale, 36(4):433– 443, October/December 1929. ISSN 0035-1571 (print), 2102-5177

(electronic). URL http://www.jstor.org/stable/40896722.

deBroglie:1929:EW

[dB29b] Louis de Broglie. Einführung in die Wellenmechanik. (German) [Introduction to wave mechanics]. Akademische Verlagsgesellschaft, Leipzig, Germany, 1929. iv + 221 pp. Translation to German by Rudolf Peierls. See [Pei85, pages 40–41] for the story of the translation experience, and how de Broglie’s views later changed. The German translation appeared before the French original was in print, because French publishers viewed the material as too speculative. Their views changed after de Broglie received the Nobel Prize in Physics for 1929 “for his discovery of the wave nature of electrons”.

deBroglie:1929:CRO

[dB29c] Louis de Broglie. La crise récente de l’optique ondulatoire. (French) [The recent crisis of wave optics]. Revue Scientifique, 67(12):353–360, ???? 1929. ISSN 0370-4556. Lecture given at the Conservatoire des Arts et Métiers, 17 April 1929.

deBroglie:1930:IEM

[dB30a] Louis de Broglie. Introduction à l’étude de la mécanique ondulatoire. (French) [Introduction to the study of wave mechanics]. Hermann, Paris, France, 1930. xvi + 292 + 9 + 2 pp.

deBroglie:1930:ISW

[dB30b] Louis de Broglie. An introduction to the study of wave mechanics. Dutton, London, UK, 1930. vi + 249 + 14 pp. Translation to English by H. T. Flint.

deBroglie:1930:OCP

[dB30c] Louis de Broglie. Ondes et corpuscules dans la physique moderne. (French) [Waves and corpuscles in modern physics]. Revue générale des Sciences, 41(4):101–108, 1930. Lecture given at the Conservatoire des Arts et Métiers under the presidency of M. Paul Painlevé, 26 January 1930.

deBroglie:1930:OCF

[dB30d] Louis de Broglie. Ondes et corpuscules. (French) [Waves and particles]. Hermann, Paris, France, 1930. ???? pp.

deBroglie:1930:REO

[dB30e] Louis de Broglie. Recueil d’exposés sur les ondes et corpuscules. (French) [Collection of presentations on airwaves and corpuscles]. Hermann, Paris, France, 1930. 80 pp.

deBroglie:1930:NOE

[dB30f] Louis de Broglie. Sur la nature ondulatoire de l’électron. (French) [On the wave nature of the electron]. Révue scientifique, 68(??): 1–7, 1930.

deBroglie:1930:NOEb

[dB30g] Louis de Broglie. Sur la nature ondulatoire de l’électron. (French) [On the wave nature of the electron]. Revue Scientifique, 68(1): 1–??, ???? 1930. Nobel Lecture given in Stockholm, Sweden, on 11 December 1929.

deBroglie:1930:ECG

[dB30h] Louis de Broglie. Sur les équations et les conceptions générales de la mécanique ondulatoire. (French) [On the equations and designs of wave mechanics]. Bulletin de la Société Mathématique de France, 58:1–28, May 1930. CODEN BSMFAA. ISSN 0037-9484 (print), 2102-622x (electronic). URL http://www.numdam.org/

item?id=BSMF_1930__58__1_0.

deBroglie:1930:NMZ

[dB30i] Louis de Broglie. [Sur une note de M. Zajac]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 191(??): 1332, December 8, 1930. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic). URL http://gallica.bnf.fr/

ark:/12148/bpt6k31445/f1332.item.r=de%20Broglie.

deBroglie:1930:ZZZ

[dB30j] Louis de Broglie. [unknown]. Akademische Verlagsgesellschaft, Leipzig, Germany, 1930. ???? pp.

deBroglie:1932:CRD

[dB32a] Louis de Broglie. Conséquences de la relativitédans le développement de la mécanique ondulatoire (La relativité III). (French) [Consequences of relativity in the development of wave mechanics (Relativity III)], volume 42 of Actualités scientifiques et industrielles. ????, ????, 1932. 14 pp.

deBroglie:1932:RSP

[dB32b] Louis de Broglie. La représentation simultanée des possibilités dans la nouvelle physique. (French) [The simultaneous representation of possibilities in the new physics]. Revue de métaphysique et de morale, 39(2):141–151, 1932. ISSN 0035-1571 (print),

2102-5177 (electronic). URL http://www.jstor.org/stable/
40897010.

deBroglie:1932:TQN

[dB32c] Louis de Broglie. La théorie de la quantification dans la nouvelle mécanique. (French) [The theory of quantification in the new mechanics]. Hermann, Paris, France, 1932. xxviii + 250 pp.

deBroglie:1932:PCE

[dB32d] Louis de Broglie. Le passage des corpuscules électrisés à travers les barrières de potentiel. (French) [The passage of electrified corpuscles through the potential barriers]. Annales de l’Institut Henri Poincaré, 3(??):339–446, ???? 1932. CODEN AIHPA2. ISSN 0365-320X (print), 2400-4855 (electronic).

deBroglie:1932:LAT

[dB32e] Louis de Broglie. L’état actuel de la théorie électromagnetique. (French) [The current state of electromagnetic theory]. Report to the Congrés international d’électricité, Paris, France., 1932.

deBroglie:1932:RMM

[dB32f] Louis de Broglie. Remarques sur le moment magnétique et le moment de rotation de l’électron. (French) [Remarks on the electron’s magnetic moment and moment of rotation]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 195 (??):577–??, ???? 1932. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1932:RIP

[dB32g] Louis de Broglie. Remarques sur les intégrales premières de la mécanique ondulatoire. (French) [Remarks on the first integrals of wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 194(??):693–??, ???? 1932. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1932:RIPb

[dB32h] Louis de Broglie. Remarques sur les intégrales premières en mécanique ondulatoire. (French) [Remarks on the first integrals in wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 194(??):693–??, ???? 1932. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1932:RPN

[dB32i] Louis de Broglie. Sur la représentation des phenoménes dans la nouvelle physique. (French) [On the representation of phenomena in the new physics]. Revue de l’université de Bruxelles, 3:277–??, ???? 1932.

deBroglie:1932:CEO

[dB32j] Louis de Broglie. Sur le champ électromagnétique de l’onde lumineuse. (French) [On the electromagnetic field of the light wave]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 195(??):862–??, ???? 1932. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1932:DVM

[dB32k] Louis de Broglie. Sur les densités de valeurs moyennes dans la théorie de Dirac. (French) [On the densities of mean values in Dirac’s theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 194(??):1062–??, ???? 1932. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1932:AEE

[dB32l] Louis de Broglie. Sur une analogie entre l’électron de Dirac et l’onde électromagnétique. (French) [On an analogy between Dirac’s electron and the electromagnetic wave]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 195 (??):536–??, ???? 1932. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1932:FPR

[dB32m] Louis de Broglie. Sur une forme plus restrictive des relations d’incertitude d’après MM. Landau et Peierls. (French) [On a more restrictive form of the uncertainty relations (MM. Landau and Peierls)]. Number 31 in Actualités scientifiques et industrielles. ????, ????, 1932. 24 pp. Exposés de physiques théorique I.

deBroglie:1933:PCE

[dB33a] Louis de Broglie. Passage des corpuscules électrisés à travers les barriéres de potentiel. (French) [Passage of electrified particles through potential barriers]. Annales de l’Institut Henri Poincaré, 3, 1933. CODEN AIHPA2. ISSN 0365-320x (print), 2400-4855 (electronic).

deBroglie:1933:QRT

[dB33b] Louis de Broglie. Quelques remarques sur la théorie de l’électron magnétique de Dirac. (French) [Some remarks on the theory of Dirac’s magnetic electron]. Archives Sc. physiques (5), 15(??): 465–483, 1933.

deBroglie:1933:RQ

[dB33c] Louis de Broglie. Relativité et quanta. (French) [Relativity and quanta]. Revue de métaphysique et de morale, 40(3):269–279, 1933. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http:

//www.jstor.org/stable/40897161.

deBroglie:1933:DET

[dB33d] Louis de Broglie. Sur la densité de l’énergie dans la théorie de la lumiére. (French) [On energy density in the theory of light]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 197(??):1377–??, ???? 1933. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1933:NTC

[dB33e] Louis de Broglie. Un nouveau type de corpuscule, le neutron. (French) [A new type of particle, the neutron]. Revue Questions scient. (4), 23:23–32, 1933. ISSN 0035-2160.

deBroglie:1934:CHO

[dB34a] Louis de Broglie. Coup d’œil sur l’histoire de l’optique. (French) [A look at the history of the optics]. Thalès, 1:3–8, 1934. ISSN

0398-7817. URL http://www.jstor.org/stable/43861140.

deBroglie:1934:LMT

[dB34b] Louis de Broglie. L’électron magnétique (théorie de Dirac). (French) [The magnetic electron (Dirac theory)]. Hermann, Paris, France, 1934. viii + 315 pp. LCCN ???? URL http://hooke.

lib.cam.ac.uk/cgi-bin/bib_seek.cgi?cat=ul&bib=4782869.

deBroglie:1934:LOD

[dB34c] Louis de Broglie. L’équation d’ondes du photon. (French) [The photon wave equation]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 199(??):445–??, ???? 1934. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1934:INI

[dB34d] Louis de Broglie. Les idées nouvelles introduites par la mécanique quantique. (French) [New ideas introduced by quantum mechanics]. L’Enseignement mathématique, 32:137–150, 1934. CODEN ENMAAR. ISSN 0013-8584 (print), 2309-4672 (electronic).

deBroglie:1934:NCLc

[dB34e] Louis de Broglie. A New Conception of Light, volume 181 of Actualités Scientifiques et Industrielles. Hermann, Paris, France,

1934. 32 pp. LCCN ???? URL http://neo-classical-
physics.info/uploads/3/0/6/5/3065888/de_broglie_-_new_

conception_of_light.pdf. Translated by D. H. Delphenich.

deBroglie:1934:RIF

[dB34f] Louis de Broglie. Realité et idéalisation. (French) [Reality and idealization]. Revue de Synthèse, 8(??):125–??, ???? 1934. ISSN 0035-1776 (print), 1955-2343 (electronic).

deBroglie:1934:RTL

[dB34g] Louis de Broglie. Remarques sur la théorie de la lumière. (French) [Remarks on the theory of light]. Mémoires de l’Académie royale des Sciences de Liége, 3e série, 19(4):1–28, 1934.

deBroglie:1934:NDP

[dB34h] Louis de Broglie. Sur la nature du photon. (French) [On the nature of the photon]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 198(??):135–??, ???? 1934. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1934:EDN

[dB34i] Louis de Broglie. Sur l’expression de la densité dans la nouvelle théorie du photon. (French) [On the expression of density in the new photon theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 199(??):1165–??, ???? 1934. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1934:NCLa

[dB34j] Louis de Broglie. Une nouvelle conception de la lumière. (exposés de physique théorique, xiii). (French) [A new conception of light. (Presentations of theoretical physics, xiii)]. Actualités Scientifiques et Industrielles, 181:48–??, 1934.

deBroglie:1934:NCLb

[dB34k] Louis de Broglie. Une nouvelle conception de la lumière. (French) [A new conception of light], volume 13. Hermann, Paris, France, 1934. ?? pp. LCCN ????

deBroglie:1934:VAP

[dB34l] Louis de Broglie. Vois anciennes et perspectíves nouvelles en théorie de lumière. (French) [Old views and new perspectives on the theory of light]. Revue de métaphysique et de morale, 41(4): 445–457, 1934. ISSN 0035-1571 (print), 2102-5177 (electronic).

URL http://www.jstor.org/stable/40897320.

deBroglie:1935:PPC

[dB35a] Louis de Broglie. Les progrès de la physique contemporaine. (French) [Progress in contemporary physics]. Revue Française de Prague, 68(??):93–??, ???? 1935.

deBroglie:1935:NGD

[dB35b] Louis de Broglie. Neue Gedanken, zu denen die Quantenmechanik Anlaß gibt. (German) [New ideas to which quantum mechanics gives rise]. Časopis Praha, 64:D134–D143, 1935.

deBroglie:1935:ESS

[dB35c] Louis de Broglie. Un exemple des synthèses successives de la physique: les théories de la lumière. (French) [An example of the successive syntheses of physics: theories of light]. Thalès, 2(??): 9–22, ???? 1935. ISSN 0398-7817. URL http://www.jstor.org/ stable/43861525. Reprinted in [dB37h].

deBroglie:1935:RIE

[dB35d] Louis de Broglie. Une remarque sur l’interaction entre la matière et le champ électromagnétique. (French) [A remark on the interaction between matter and the electromagnetic field]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 200(??):361–??, ???? 1935. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1936:TDP

[dB36a] Louis de Broglie. La théorie du photon et la mécanique ondulatoire relativiste des systèmes. (French) [Photon theory and relativistic wave mechanics of systems]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 203(??):473–??, ???? 1936. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1936:VRD

[dB36b] Louis de Broglie. La variance relativiste du moment cinétique d’un corps en rotation. (French) [The relativistic variance of the angular momentum of a rotating body]. Journal de Mathématiques Pures et Appliquées, 15:89–95, 1936. CODEN JMPAAM. ISSN 0021-7824 (print), 1776-3371 (electronic).

deBroglie:1936:LE

[dB36c] Louis de Broglie. L’évolution de l’électron. (French) [The evolution of the electron]. In Livre du Centenaire d’Ampère, Lyon, page ?? ????, ????, March 1936.

deBroglie:1936:NRL

[dB36d] Louis de Broglie. Nouvelles recherches sur la lumière. (French) [New research on light], volume 20. Hermann, Paris, France, 1936. ???? pp.

deBroglie:1936:NRT

[dB36e] Louis de Broglie. Nouvelles recherches sur la théorie de la lumière. (French) [New research on the theory of light], volume 411 of Actual. sci. industr. ????, ????, 1936. 54 pp.

deBroglie:1936:RDS

[dB36f] Louis de Broglie. Réflexions sur les deux sortes d’électricité. (French) [Reflections on the two kinds of electricity]. Revue de métaphysique et de morale, 43(2):173–185, 1936. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.jstor.org/

stable/40897833.

deBroglie:1937:EAN

[dB37a] Louis de Broglie. État actuel de nos connaissances sur la structure de l’électricité. (French) [Current state of knowledge of the structure of electricity]. Nuovo Cimento, 14(9):??, November 1937.

deBroglie:1937:IIM

[dB37b] Louis de Broglie. Individualité et interaction dans le monde physique. (French) [Individuality and interaction in the physical world]. Revue de métaphysique et de morale, 44(2):353–368, 1937. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://

www.jstor.org/stable/40897962.

deBroglie:1937:PNQ

[dB37c] Louis de Broglie. La physique nouvelle et les quanta. (French) [The new physics and quanta]. Bibliothèque de philosophie scientifique, dirigée par Paul Gaultier. Flammarion fils, Paris, France, 1937. 312 pp.

deBroglie:1937:QCT

[dB37d] Louis de Broglie. La quantification des champs dans la théorie du photon. (French) [The quantification of fields in photon theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 205(??):345–??, ???? 1937. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1937:RCT

[dB37e] Louis de Broglie. Les récentes conceptions théoriques sur la lumière. (French) [Recent theoretical concepts on light]. Annales de la Societe scientifique de Bruxelles, 1^{re} série, 157(??):99–119, 1937. ISSN 0037-959X.

deBroglie:1937:LST

[dB37f] Louis de Broglie. L’invention dans les sciences théoriques. (French) [Invention in the theoretical sciences]. Sciences (Paris), 2(14):??, June 1937.

deBroglie:1937:MLF

[dB37g] Louis de Broglie. Matière et lumière. (French) [Matter and Light]. Albin Michel, Paris, France, 1937. ???? pp.

deBroglie:1937:ESS

[dB37h] Louis de Broglie. Un exemple des synthèses successives de la physique: les théories de la lumière. (French) [An example of the successive syntheses of physics: theories of light]. Revue des Questions Scientifiques, ??(??):361–381, ???? 20, 1937. CODEN RQS-

CAN. ISSN 0398-7817. URL http://www.jstor.org/stable/

43861525. Reprint of [dB35c].

deBroglie:1938:PCIa

[dB38a] L. de Broglie. Le principe de correspondance et les interactions entre la matière et le rayonnement. (French) [The principle of correspondence and interactions between matter and radiation], volume 704 of Actual. sci. industr. ????, ????, 1938. 170 pp.

deBroglie:1938:PCIb

[dB38b] Louis de Broglie. Le principe de correspondance et les interactions entre la matière et le rayonnement. (French) [The principle of correspondence and the interactions between matter and radiation]. Hermann, Paris, France, 1938. 704–?? pp.

deBroglie:1938:PPP

[dB38c] Louis de Broglie. Physique ponctuelle et physique du champ. (French) [Point physics and field physics]. Revue de métaphysique et de morale, 45(3):325–338, 1938. ISSN 0035-1571 (print),

2102-5177 (electronic). URL http://www.jstor.org/stable/
40898762.

deBroglie:1939:MOSa

[dB39a] Louis de Broglie. La mécanique ondulatoire des systèmes de corpuscules. (French) [The wave mechanics of particle systems], volume 5 of Collection de physique mathématique. Gauthier-Villars, Paris, France, 1939. vi + 223 pp.

deBroglie:1939:TQD

[dB39b] Louis de Broglie. La théorie quantique du rayonnement. (French) [Quantum radiation theory]. Revue de métaphysique et de morale, 46(2):199–210, 1939. ISSN 0035-1571 (print), 2102-5177 (elec-

tronic). URL http://www.jstor.org/stable/40898315.

deBroglie:1939:LME

[dB39c] Louis de Broglie. Licht und Materie. Ergebnisse der neuen Physik. (German) [Light and matter. Results of the new physics]. H. Goverts-Verlag, Hamburg, Germany, 1939. 331 + 3 pp.

deBroglie:1939:ML

[dB39d] Louis de Broglie. Matter and light. W. W. Norton & Co., New York, NY, USA, 1939. 300 pp. Translation to English by W. H. Johnston.

deBroglie:1939:MOSb

[dB39e] Louis de Broglie. Mécanique ondulatoire des systèmes de corpuscules de physique mathématique. (French) [Wave mechanics of particles in mathematical physics]. Gauthier-Villars, Paris, France, 1939. vi + 224 pp.

deBroglie:1939:TPS

[dB39f] Louis de Broglie. Sur la théorie des particules de spin quelconque. (French) [On the theory of arbitrary-spin particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 209 (??), ???? 1939. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1939:PSQ

[dB39g] Louis de Broglie. Sur les particules de spin quelconque. (French) [On particles of arbitrary spin]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 209(??):265–??, ???? 1939. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1939:CRM

[dB39h] Louis de Broglie. Sur un cas de réductibilité en mécanique ondulatoire des particules de spin 1. (French) [On a case of reducibility in wave mechanics of spin-1 particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 208(??):1697– ??, ???? 1939. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1940:CRC

[dB40a] Louis de Broglie. Champs réels et champs complexes en théorie électromagnétique quantique du rayonnement. (French) [Real fields and complex fields in electromagnetic quantum theory of radiation]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 211(??):41–??, ???? 1940. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1940:MOD

[dB40b] Louis de Broglie. La mécanique ondulatoire du photon. Une nouvelle théorie de la lumière. I. (French) [The wave mechanics of the photon. A new theory of light. I]. Gauthier-Villars, Paris, France, 1940. 281 pp.

deBroglie:1940:RPT

[dB40c] Louis de Broglie. Récents progrès dans la théorie des photons et autres particules. (French) [Recent advances in the theory of photons and other particles]. Revue de métaphysique et de morale, 47(1):1–16, 1940. ISSN 0035-1571 (print), 2102-5177 (electronic).

URL http://www.jstor.org/stable/40898475.

deBroglie:1940:NTL

[dB40d] Louis de Broglie. Une nouvelle théorie de la lumière, la mécanique ondulatoire du photon. Tome I. La lumière dans le vide. (French) [A new theory of light, the wave mechanics of the photon. Volume I. Light in the vacuum]. Hermann, Paris, France, 1940. ???? pp.

deBroglie:1941:CDP

[dB41a] Louis de Broglie. Continu et discontinu en physique moderne. (French) [Continuous and discontinuous in modern physics]. Sci. d’aujourd’hui. Albin Michel, Paris, France, 1941. 268 pp.

deBroglie:1941:LPA

[dB41b] Louis de Broglie. L’avenir de la physique dans l’avenir de la science. (French) [The Future of Physics in the Future of Science].

Collection Présences, Plon, Paris, France, 1941. ???? pp. LCCN ????

deBroglie:1941:CPC

[dB41c] Louis de Broglie. Les conceptions de la physique contemporaine et les idées de Bergson sur le temps et sur le mouvement. (French) [Conceptions of contemporary physics and Bergson’s ideas on time and movement]. Revue de métaphysique et de morale, 48 (4):241–257, October 1941. ISSN 0035-1571 (print), 2102-5177

(electronic). URL http://www.jstor.org/stable/40898604.

deBroglie:1941:PPG

[dB41d] Louis de Broglie. Problèmes de Propagations Guidées des Ondes Électromagnétiques. (French) [Problems of guided propagation of electromagnetic waves]. Gauthier-Villars, Paris, France, 1941. vi + 114 + 14 pp.

deBroglie:1941:SPD

[dB41e] Louis de Broglie. Souvenirs personnels sur les débuts de la mécanique ondulatoire. (French) [Personal memories of the start of wave mechanics]. Revue de métaphysique et de morale, 48(1): 1–23, 1941. ISSN 0035-1571 (print), 2102-5177 (electronic). URL

http://www.jstor.org/stable/40898568.

deBroglie:1941:ICE

[dB41f] Louis de Broglie. Sur l’interprétation de certaines équations dans la théorie des particules de spin 2. (French) [On the interpretation of some equations in the theory of particles of spin 2]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 212(??):657–??, ???? 1941. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1942:MOD

[dB42a] Louis de Broglie. La mécanique ondulatoire du photon. Une nouvelle théorie de la lumière. II. Les interactions entre les photons et la matière. (French) [Wave mechanics of the photon. A new theory of light. II. Interactions between photons and matter]. Hermann, Paris, France, 1942. 150 pp.

deBroglie:1942:PEM

[dB42b] Louis de Broglie. Les particules élémentaires de la matière et les nouvelles théories du noyau de l’atome. (French) [Elementary particles of matter and the new theories of the nucleus of the atom]. L’Astronomie, 56(??):1–6, 27–34, 51–57, ???? 1942.

deBroglie:1942:LUA

[dB42c] Louis de Broglie. L’œuvre d’André Blondel en physique générale. (French) [The work of Andrè Blondel in general physics]. In Commemoration de l’œuvre d’Andrè-Eugène Blondel, pages 7–19. Gauthier-Villars, Paris, France, 1942.

deBroglie:1942:SPD

[dB42d] Louis de Broglie. Souvenirs personnels sur les débuts de la mécanique ondulatoire. (French) [Personal memories on the beginnings of wave mechanics]. Revue de métaphysique et de morale, 53(??):1–23, ???? 1942. ISSN 0035-1571 (print), 2102-5177 (electronic).

deBroglie:1942:PEL

[dB42e] Louis de Broglie. Sur la propagation de l’énergie lumineuse dans les milieux anisotropes. (French) [On the propagation of light energy in anisotropic media]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 215(??):153–156, ???? 1942. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1942:NTL

[dB42f] Louis de Broglie. Une nouvelle théorie de la lumière, la mécanique ondulatoire du photon. Tome II. L’intéraction entre les photons et la matière. (French) [A new theory of light, the wave mechanics of the photon. Volume II. Interaction between photons and matter]. Hermann, Paris, France, 1942. ???? pp.

deBroglie:1943:COM

[dB43a] Louis de Broglie. Corpuscules, ondes et mécanique ondulatoire. (French) [Particles, waves and wave mechanics]. Centre de documentation universitaire, Paris, France, 1943. ???? pp. LCCN ???? Conférences à l’École supérieure d’Électricité.

deBroglie:1943:MOT

[dB43b] Louis de Broglie. De la mécanique ondulatoire à la theorie du noyau. (French) [From wave mechanics to nuclear theory], volume 1. Hermann, Paris, France, 1943. ???? pp.

deBroglie:1943:EIW

[dB43c] Louis de Broglie. Die Elementarteilchen. Individualität und Wechselwirkung. Ergebnisse der neuen Physik. 2. Tl. (German) [Elementary particles. Individuality and interaction. Results of the new physics. Volume 2]. H. Goverts-Verlag, Hamburg, Germany, 1943. 281 pp.

deBroglie:1943:VOU

[dB43d] Louis de Broglie. La vie et l’œuvre de M. Émile Picard. (French) [Life and work of Émile Picard]. Mem. Acad. Sci. Inst. France, II. Sér., 66(??):45–??, 1943. Lecture of 21 December 1942.

deBroglie:1943:LPF

[dB43e] Louis de Broglie. L’essor de la physique en France de 1815 à 1825. (French) [The rise of physics in France from 1815 to 1825]. Collection Hier et Demain, 4:80–108, 1943.

deBroglie:1943:RSR

[dB43f] Louis de Broglie. Recherche scientifique et recherche technique. (French) [Scientific research and technical research]. Revue générale de caoutchouc, 20(??):45–49, ???? 1943.

deBroglie:1943:RGE

[dB43g] Louis de Broglie. Sur la représentation des grandeurs électromagnétiques en théorie quantique des champs et en mécanique ondulatoire du photon. (French) [On the representation of electromagnetic quantities in quantum field theory and in wave mechanics of the photon]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 217(??):89–??, ???? 1943. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1943:TGP

[dB43h] Louis de Broglie. Théorie générale des particules à spin. (French) [General theory of spin particles]. Gauthier-Villars, Paris, France, 1943. ???? pp.

deBroglie:1944:VOU

[dB44a] Louis de Broglie. La vie et l’œuvre d’André Blondel. (French) [The life and works of] André Blondel. Public lecture of 18 December 1944, 1944.

deBroglie:1944:NLR

[dB44b] Louis de Broglie. Sur les notions de lois rigoureuses et de lois statistiques. (French) [On the notions of rigorous laws and statistical laws]. Revue de l’Économie contemporaine, 3(25):1–3, May 1944.

deBroglie:1945:COM

[dB45a] Louis de Broglie. Corpuscules, ondes et mécanique ondulatoire. (French) [Particles, waves and wave mechanics]. Albin Michel, Paris, France, 1945. ???? pp. LCCN ???? Conférences à l’École supérieure d’Électricité.

deBroglie:1945:MOT

[dB45b] Louis de Broglie. De la mécanique ondulatoire à la theorie du noyau. (French) [From wave mechanics to nuclear theory], volume 2. Hermann, Paris, France, 1945. ???? pp.

deBroglie:1945:DRA

[dB45c] Louis de Broglie. Discours de réception à l’Académie française prononcé sous la Coupole le 31 mai 1945. (French) [Reception speech at the Académie française delivered under the Cupola on 31 May 1945]. Albin Michel, Paris, France, de luxe edition, 1945.

deBroglie:1945:GVM

[dB45d] Louis de Broglie. Grandeur et valeur morale de la science. (French) [Greatness and moral value of science]. N. R. F., ??(??):41–??, March 1945.

deBroglie:1945:HCP

[dB45e] Louis de Broglie. Hasard et contingence en physique quantique. (French) [Chance and contingency in quantum physics]. Revue de métaphysique et de morale, 50(4):241–252, 1945. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.jstor.

org/stable/40899173.

deBroglie:1945:RMO

[dB45f] Louis de Broglie. La réalité des molécules et l’œuvre de Jean Perrin. (French) [The reality of molecules and the work of Jean Perrin]. Public lecture of 17 December 1945., 1945.

deBroglie:1945:LBE

[dB45g] Louis de Broglie. Léon Brunschvicg et l’évolution des sciences. (French) [Léon Brunschvicg and the evolution of science]. Revue de métaphysique et de morale, 50(1/2):73–76, 1945. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.jstor.

org/stable/40899136.

deBroglie:1946:MOT

[dB46a] Louis de Broglie. De la mécanique ondulatoire à la theorie du noyau. (French) [From wave mechanics to nuclear theory], volume 3. Hermann, Paris, France, 1946. ???? pp.

deBroglie:1946:LMP

[dB46b] Louis de Broglie. La lumière dans le monde physique. (French) [Light in the physical world]. Cahiers du monde nouveau, 2(3): 1–46, March 1946.

deBroglie:1946:VOU

[dB46c] Louis de Broglie. La vie et l’œuvre de Charles Fabry. (French) [The life and work of Charles Fabry]. Public lecture help on 16 December 1946, 1946.

deBroglie:1946:LDC

[dB46d] Louis de Broglie. L’activité du centre de physique théorique de l’Institut Henri Poincaré pendant les dernières années. (French) [The activity of the Center of Theoretical Physics of the Henri Poincaré Institute during the last years]. Experientia, 2(1):??, ???? 1946. CODEN EXPEAM. ISSN 0014-4754.

deBroglie:1946:RFB

[dB46e] Louis de Broglie. Remarques sur la formule de Boltzmann relative aux systemes périodiques. (French) [Remarks on the Boltzmann formula for periodic systems]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 223:298–??, ???? 1946. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1946:ADT

[dB46f] Louis de Broglie. Sur l’application du théorème des probabilités composées en mécanique ondulatoire. (French) [On the application of the theorem of composite probabilities in wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 223(??):874–877, ???? 1946. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1946:ELP

[dB46g] Louis de Broglie. Sur un effet limitant les possibilités du microscope corpusculaire. (French) [On an effect limiting the possibilities of the corpuscular microscope]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 222(??):1017– ??, ???? 1946. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1947:ADM

[dB47a] Louis de Broglie. Au dela des mouvantes limites de la science. (French) [Beyond the shifting limits of science]. Revue de métaphysique et de morale, 52(3/4):277–289, 1947. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.jstor.

org/stable/40899292.

deBroglie:1947:VOU

[dB47b] Louis de Broglie. La vie et l’œuvre de Paul Langevin. (French) [The life and works of Paul Langevin]. L’Académie des Sciences, Paris, France, 1947. ???? pp.

deBroglie:1947:MED

[dB47c] Louis de Broglie. Le microscope électronique et la dualité des ondes et des corpuscules. (French) [The electron microscope and the duality of waves and particles]. Revue de métaphysique et de morale, 52(1):3–14, 1947. ISSN 0035-1571 (print), 2102-5177

(electronic). URL http://www.jstor.org/stable/40899251.

deBroglie:1947:PIE

[dB47d] Louis de Broglie. Le principe de l’inertie de l’énergie et la notion d’énergie potentielle. (French) [The principle of energy inertia and the notion of potential energy]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 225(??):163–??, ???? 1947. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1947:PM

[dB47e] Louis de Broglie. Physique et microphysique. (French) [Physics and microphysics]. Collection “Sciences d’aujourd’hui”. Albin Michel, Paris, France, 1947. 370 + 8 pp.

deBroglie:1947:FIO

[dB47f] Louis de Broglie. Sur la formation des images en optique corpusculaire. (French) [On image formation in particle optics]. Rev. Optique Theor. Instrum., 26(??):397–410, 1947.

deBroglie:1947:FVP

[dB47g] Louis de Broglie. Sur la fréquence et la vitesse de phase des ondes planes monochromatiques en mécanique ondulatoire. (French) [On the frequency and phase velocity of monochromatic plane waves in wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 225(??):361–363, ???? 1947. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1947:EMT

[dB47h] Louis de Broglie. Sur les électrinos de M. Thibaud et l’existence éventuelle d’une très petite charge du neutron. (French) [On the electrinos of M. Thibaud and the possible existence of a very small neutron charge]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 224(??):615–??, ???? 1947. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1948:TEG

[dB48a] Louis de Broglie. Die Trägheit der Energie. (German) [The inertia of energy]. Euclides, Madrid, 8(??):1–4, 1948.

deBroglie:1948:SCPb

[dB48b] Louis de Broglie. La statistique des cas purs en mécanique ondulatoire et l’interférence des probabilités. (French) [The statistics of pure cases in wave mechanics and the interference of probabilities]. Revue Scientifique, 86(??):259–264, ???? 1948. ISSN 0370-4556.

deBroglie:1948:RMD

[dB48c] Louis de Broglie. Le rôle des mathématiques dans le développement de la physique théorique contemporaine. (French) [The role of mathematics in the development of contemporary theoretical physics]. Les grands courants de la pensée mathématique, Cahiers du Sud, ??(??):??, ???? 1948.

deBroglie:1948:MPW

[dB48d] Louis de Broglie. Max Planck und das Wirkungsquantum. (German) [Max Planck and quantum of action]. Physikalische Blätter, 4(4):138–142, April 1948. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19480040402/abstract. Special

issue on Max Planck.

deBroglie:1948:CII

[dB48e] Louis de Broglie. Sur la complémentarité des idées d’individu et de système. (French) [On the complementarity of ideas of the individual and the system]. Dialectica: International Review of Philosophy of Knowledge, 2(3/4):325–330, 1948. ISSN 0012-2017 (print), 1746-8361 (electronic). URL http://www.jstor.org/

stable/42963891.

deBroglie:1948:PMEa

[dB48f] Louis de Broglie. Sur la possibilité de mettre en évidence le moment magnétique propre des particules à spin. (French) [On the possibility of detecting the magnetic moment of spin particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 226(??):1765–??, ???? 1948. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1948:PMEb

[dB48g] Louis de Broglie. Sur la possibilité de mettre en évidence le moment magnétique propre des particules de spin 1/2. (French) [On the possibility of highlighting the characteristic magnetic moment of particles of spin 1/2]. Journal de Physique et le Radium, 9(11): 265–272, 1948. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1948:SCPa

[dB48h] Louis de Broglie. Sur la statistique des cas purs en mécanique ondulatoire. (French) [On the statistics of pure cases in wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 226(??):1056–1058, ???? 1948. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1948:VRT

[dB48i] Louis de Broglie. Sur la variance relativiste de la température. (French) [On the relativistic variation of temperature]. Cahiers de Physique, 1948(31–32):1–11, January 1948. CODEN CAPHAI. ISSN 0366-5291.

deBroglie:1949:ELF

[dB49a] Louis de Broglie. Énergie libre et fonction de Lagrange. Application à l’électrodynamique et à l’interaction entre courants et aimants permanents. (French) [Free energy and function of Lagrange. Application to electrodynamics and the interaction between currents and permanent magnets]. Portugaliae Phys., 3: 1–19, 1949.

deBroglie:1949:GSS

[dB49b] Louis de Broglie. A general survey of the scientific work of Albert Einstein. In Schilpp [Sch49a], chapter 2, pages 107–127. ISBN 0-87548-286-4. ISSN 0075-9139. LCCN QC16.E5 S3 1970. Translated from the French manuscript by Forrest W. Williams.

deBroglie:1949:PCO

[dB49c] Louis de Broglie. La physique contemporaine et l’œuvre d’Albert Einstein. (French) [Contemporary physics and the work of Albert Einstein]. Public lecture of 19 December 1949., 1949.

deBroglie:1949:MAT

[dB49d] Louis de Broglie. Le meson. Aspects théoriques et experimentaux. (French) [The meson. Theoretical and experimental aspects]. Éditions de la Revue d’Optique, Paris, Fran, 1949. 192 pp.

deBroglie:1949:LP

[dB49e] Louis de Broglie. L’enseignement de la physique. (French) [Teaching physics]. Revue de métrologie pratique et légale, 27(??):5–??, ???? 1949.

deBroglie:1949:LTP

[dB49f] Louis de Broglie. L’espace et le temps dans la physique quantique. (French) [Space and time in quantum physics]. Revue de métaphysique et de morale, 54(2):113–125, 1949. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.jstor.org/

stable/40899404.

deBroglie:1949:LUE

[dB49g] Louis de Broglie. L’œuvre d’Einstein et la dualité des ondes et des corpuscules. (French) [The work of Einstein and the duality of waves and corpuscles]. Reviews of Modern Physics, 21(3):345–347, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.21.345;   http://rmp.aps.org/abstract/

RMP/v21/i3/p345_1. Special issue in honor of Albert Einstein on his 70th birthday.

deBroglie:1949:MOD

[dB49h] Louis de Broglie. Mécanique ondulatoire du photon et théorie quantique des champs. (French) [Wave mechanics of the photon and quantum field theory]. Gauthier-Villars, Paris, France, 1949. vi + 208 pp.

deBroglie:1949:NRI

[dB49i] Louis de Broglie. Nouvelles remarques sur l’interaction entre une charge électrique et le champ électromagnétique. (French) [New remarks on the interaction between an electric charge and the electromagnetic field]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 229(??):269–271, ???? 1949. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1949:NCI

[dB49j] Louis de Broglie. On a new conception of the interaction between charges and electromagnetic field. Physical Review (2), 76:862– 863, 1949. CODEN PHRVAO. ISSN 0031-899x (print), 1536-6065 (electronic).

deBroglie:1949:POE

[dB49k] Louis de Broglie. Pénétration d’une onde electromagnétique dans un milieu ou la constante diélectrique varie linéairement. (French) [Penetration of an electromagnetic wave in a medium where the dielectric constant varies linearly]. Preliminary note 129, Laboratoire national de radioélectricité, Paris, France, 1949.

deBroglie:1949:PPS

[dB49l] Louis de Broglie. Problèmes de Philosophie des Sciences. (Premier symposium, Bruxelles 1947.) IV: Problèmes de connaissance en physique moderne. (French) [Problems of Philosophy of Science. (First Symposium, Brussels 1947.) IV: Problems of knowledge in modern physics], volume 1066 of Actualités scientifiques et industrielles. Hermann, Paris, France, 1949. 103 pp. Arch. Inst. Internat. Sci. Theor. Ser. A, Bull. Acad. Internat. Philos. Sci. 4.

deBroglie:1949:RIS

[dB49m] Louis de Broglie. Sur la relation d’incertitude de la seconde quantification. (French) [On the uncertainty relation of second quantization]. Revue Internationale de Philosophie, 3(8):166–175, 1949. ISSN 0048-8143 (print), 2033-0138 (electronic). URL http://

www.jstor.org/stable/23932673.

deBroglie:1949:TDC

[dB49n] Louis de Broglie. Sur la théorie du champ soustractif. (French) [On subtractive field theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 229(??):401–404, ???? 1949. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1949:CCE

[dB49o] Louis de Broglie. Sur le calcul classique de l’energie et de la quantité de mouvement d’un électron purement électromagnétique. (French) [On the classical calculation of the energy and the momentum of a purely electromagnetic electron]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 228(??): 1264–1268, ???? 1949. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1949:CEQ

[dB49p] Louis de Broglie. Sur le calcul de l’énergie et de la quantité de mouvement d’un électron purement électromagnetique. (French) [On the calculation of the energy and the momentum of a purely electromagnetic electron]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 228(??), ???? 1949. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1949:CCP

[dB49q] Louis de Broglie. Sur les champs crées par le proton et par le neutron. (French) [On the fields created by the proton and by the neutron]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 229(??):640–??, ???? 1949. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1949:FNI

[dB49r] Louis de Broglie. Sur une forme nouvelle de l’interaction entre les charges électriques et le champ électromagnétique. (French) [On a new form of the interaction between electric charges and the electromagnetic field]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 229(??):157–161, ???? 1949. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1949:CNI

[dB49s] Louis de Broglie. Une conception nouvelle de l’interaction entre les particules chargées et le champ électromagnétique. (French)

[A new conception of the interaction between charged particles and the electromagnetic field]. Portugaliae Mathematica, 8:37– 58, 1949. ISSN 0032-5155 (print), 1662-2758 (electronic).

deBroglie:1950:E

[dB50a] Louis de Broglie. Das Elektronenmikroskop. (German) [The electron microscope]. Physik Journal, 6(7):311–312, July 1950. CO-DEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

deBroglie:1950:MOS

[dB50b] Louis de Broglie. La mécanique ondulatoire des systèmes de corpuscules. (French) [Wave mechanics of particle systems]. Gauthier-Villars, Paris, France, 1950. vi + 223 pp. 2d ed.

deBroglie:1950:LIP

[dB50c] Louis de Broglie. L’avenir influe-t-il sur le present ?. (French) [Does the future affect the present?]. L’orientation medicale, 16 (2):11–??, ???? 1950.

deBroglie:1950:OEC

[dB50d] Louis de Broglie. Optique électronique et corpusculaire. (French) [Electronic and corpuscular optics]. Hermann, Paris, France, 1950. 176 + 60 pp.

deBroglie:1950:RCS

[dB50e] Louis de Broglie. Remarques complémentaires sur la structure complexe des particules de spin 1. (French) [Additional remarks on the complex structure of spin-1 particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 230(??), ???? 1950. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1950:RCSb

[dB50f] Louis de Broglie. Remarques complémentaires sur la structure complexe des particules de spin 1. (French) [Additional remarks on the complex structure of spin-1 particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 230(??): 1434–??, ???? 1950. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1950:CIP

[dB50g] Louis de Broglie. Sur la convergence des intégrales dans le probléme de la polarisation du vide. (French) [On the convergence of integrals in the problem of vacuum polarization]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 230 (??):2061–??, ???? 1950. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1950:FNT

[dB50h] Louis de Broglie. Sur une forme nouvelle de la théorie du “champ soustractif”. (French) [On a new form of the “subtractive field” theory]. Journal de Physique et le Radium, 11(8–9):481–489, 1950. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1951:COH

[dB51a] Louis de Broglie. Coup d’œil sur l’histoire de la science en France pendant la période 1799–1810. (French) [A look at the history of science in France during the period 1799–1810]. Revue d’Histoire des Sciences et de Leurs Applications, 4(2):105–108, April/June 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582

(electronic). URL http://www.jstor.org/stable/23904857.

deBroglie:1951:TPS

[dB51b] Louis de Broglie. La théorie des particules de spin 1/2 (Électrons de Dirac). (French) [Theory of spin-1/2 particle (Dirac electrons)]. Gauthier-Villars, Paris, France, 1951. ???? pp. LCCN ????

deBroglie:1951:VOU

[dB51c] Louis de Broglie. La vie et l’œuvre de Hendrik Antoon Lorentz. (French) [The life and work of Hendrik Antoon Lorentz]. Public lecure of 10 December 1951, 1951.

deBroglie:1951:PPG

[dB51d] Louis de Broglie. Problèmes de propagations guidées des ondes électromagnétiques. (French) [Problems in guided electromagnetic wave propagation]. Gauthier-Villars, Paris, France, second edition, 1951. vii + 118 pp.

deBroglie:1951:RNP

[dB51e] Louis de Broglie. Remarque sur la note précédente de J.-P. Vigier. (French) [Remark on the previous note of J.-P. Vigier]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 233(??):1012–1013, ???? 1951. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1951:RTO

[dB51f] Louis de Broglie. Remarques sur la théorie de l’onde pilote. (French) [Remarks on pilot-wave theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 233(??):20–??, ???? 1951. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1951:SDF

[dB51g] Louis de Broglie. Savants et découvertes. (French) [Scientists and discoveries]. Albin Michel, Paris, France, 1951. ???? pp.

deBroglie:1951:SLTb

[dB51h] Louis de Broglie. Schema lagrangien de la théorie du champ soustractif. (French) [Lagrangian schema of subtractive field theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 232(??):1269–??, ???? 1951. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1951:SLT

[dB51i] Louis de Broglie. Schéma lagrangien de la théorie du champ soustractif. (French) [Lagrangian scheme of subtractive field theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 232(??):1269–1272, ???? 1951. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1951:SPP

[dB51j] Louis de Broglie. Sens philosophique et portée pratique de la cybernétique. (French) [Philosophical sense and practical scope of cybernetics]. Lecture given at the Conservatoire des Arts et Métiers on 5 October 1951., 1951.

deBroglie:1951:PSC

[dB51k] Louis de Broglie. Sur la possibilité d’une structure complexe des particules de spin différent de \frac{1}{2}. (French) [On the possibility of a complex structure of particles of spin different from 1/2]. Journal de Physique et le Radium, 12:509–516, 1951. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1951:MHL

[dB51l] Louis de Broglie. Un mathématicien, homme de lettres: d’Alembert. (French) [Mathematician and man of letters: d’Alembert]. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):204–212, juillet–décembre 1951. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://

www.jstor.org/stable/23905018.

deBroglie:1951:NVP

[dB51m] Louis de Broglie. Un nouveau venu en physique: le champ nucléaire. (French) [A newcomer in physics: the nuclear field]. Revue de métaphysique et de morale, 56(2):117–127, 1951. ISSN 0035-1571 (print), 2102-5177 (electronic). URL http://www.

jstor.org/stable/40899559.

deBroglie:1952:MOS

[dB52a] Louis de Broglie. La mécanique ondulatoire des systèmes de particules de même nature et la théorie de la double solution. (French) [Wave mechanics of particle systems of the same nature and the theory of the double solution]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 235(??):1453–1455, ???? 1952. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1952:PQR

[dB52b] Louis de Broglie. La physique quantique restera-t-elle indéterministe?. (French) [Will quantum physics remain indeterministic?]. Revue d’Histoire des Sciences et de Leurs Applications, 5(4):289–311, October 1952. CODEN RHSAAM. ISSN 0048-7996 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23905080.

deBroglie:1952:TPS

[dB52c] Louis de Broglie. La théorie des particules de spin 1/2: électrons de Dirac. (French) [The theory of spin-1/2 particles: Dirac electrons]. Gauthier-Villars, Paris, France, 1952. 164 pp. LCCN ????

deBroglie:1952:PM

[dB52d] Louis de Broglie. Les particules de la microphysique. (French) [The particles of microphysics]. Lecture given at the Palais de la Découverte on 20 December 1952., 1952.

deBroglie:1952:PIC

[dB52e] Louis de Broglie. Sur la possibilité d’une interprétation causale et objective de la mécanique ondulatoire. (French) [On the possibility of a causal and objective interpretation of wave mechanics].

Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 234(??):265–268, ???? 1952. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1952:TEI

[dB52f] Louis de Broglie. Sur le tenseur énergie–impulsion dans la théorie du champ soustractif. (French) [On the energy–momentum tensor in subtractive field theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 234(??):20–22, ???? 1952. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1952:REC

[dB52g] Louis de Broglie. Sur les relations entre les coefficients de charge et de masse dans la théorie du champ soustractif. (French) [On the relations between charge and mass coefficients in subtractive field theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 234(??):1505–1507, ???? 1952. CO-DEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1952:IMO

[dB52h] Louis de Broglie. Sur l’interprétation de la mécanique ondulatoire des systèmes de corpuscules dans l’espace de configuration par la théorie de la double solution. (French) [On the interpretation of wave mechanics of particle systems in configuration space by double-solution theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 235(??):1345–1349, ???? 1952. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1952:IIO

[dB52i] Louis de Broglie. Sur l’introduction des idées d’onde-pilote et de double solution dans la théorie de l’électron de Dirac. (French) [On the introduction of pilot-wave and double-solution ideas in Dirac’s electron theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 235(??):557–??, ???? 1952. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1953:ETQ

[dB53a] Louis de Broglie. Eléments de théorie des quanta et de mécanique ondulatoire. (French) [Elements of quantum theory and wave mechanics], volume III of Traité de Physique Théorique et de Physique Mathématique. Gauthier-Villars, Paris, France, 1953. viii + 302 pp.

deBroglie:1953:PQR

[dB53b] Louis de Broglie. La physique quantique restera-t-elle indéterministe? Exposé du problème suivi de la reproduction de certains documents et d’une contribution de M. Jean-Pierre Vigier. (French) [Will quantum physics remain indeterministic? Exposure of the problem followed by the reproduction of certain documents and a contribution from M. Jean-Pierre Vigier], volume I of Les grands problèmes des Sciences. Gauthier-Villars, Paris, France, 1953. vii + 113 pp.

deBroglie:1953:VOU

[dB53c] Louis de Broglie. La vie et l’œuvre d’Aimé Cotton. (French) [The life and work of Aimé Cotton]. Public lecture of 14 December 1953, 1953.

deBroglie:1953:LMO

[dB53d] Louis de Broglie. L’interprétation de la mécanique ondulatoire à l’aide d’ondes à régions singulières. (French) [The interpretation of wave mechanics using waves with singular regions], 1953.

deBroglie:1953:ICN

[dB53e] Louis de Broglie. Sur l’interprétation causale et non linéaire de la mécanique ondulatoire. (French) [On the causal and nonlinear interpretation of wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 237(??):441–444, ???? 1953. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1953:IMO

[dB53f] Louis de Broglie. Sur l’interprétation de la mécanique ondulatoire à l’aide d’ondes à région singulière. (French) [On the interpretation of wave mechanics using singular-region waves]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 236(??):1453–1456, ???? 1953. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1953:IMOb

[dB53g] Louis de Broglie. Sur l’interprétation de la mécanique ondulatoire à l’aide d’ondes à région singulière. (French) [On the interpretation of wave mechanics using singular region waves]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 236(??):1459–??, ???? 1953. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1953:TMP

[dB53h] Louis de Broglie. Sur un théorème de M. Poincelot. (French) [On a theorem of M. Poincelot]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 237(??):586–588, ???? 1953. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1953:VEM

[dB53i] Louis de Broglie. Vue d’ensemble sur mes travaux scientifiques. (French) [Overview of my scientific work]. In Louis de Broglie, physicien et penseur, pages 457–493. Éditions Albin Michel, Paris, France, 1953.

deBroglie:1953:WQIa

[dB53j] Louis de Broglie. Wird die Quantenphysik indeterministisch bleiben?. (German) [Will quantum physics remain indeterministic?]. Physik Journal, 9(11):488–497, November 1953. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

deBroglie:1953:WQIb

[dB53k] Louis de Broglie. Wird die Quantenphysik indeterministisch bleiben?. (German) [Will quantum physics remain indeterministic?]. Physik Journal, 9(12):541–548, December 1953. CODEN PJHOB2. ISSN 1617-9439 (print), 1619-6597 (electronic).

deBroglie:1954:CMC

[dB54a] Louis de Broglie. Considérations de mécanique classique préparant la justification de la mécanique ondulatoire des système dans la théorie de la double solution. (French) [Considerations of classical mechanics preparing the justification of wave mechanics of systems in double-solution theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 239(??):521– 524, ???? 1954. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1954:JDP

[dB54b] Louis de Broglie. Justification, du point de vue de la théorie de la double solution, de la mécanique ondulatoire des systèmes dans l’espace de configuration. (French) [Justification, from the point of view of double-solution theory, of the wave mechanics of systems in configuration space]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 239(??):565–567, ???? 1954. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1954:RPN

[dB54c] Louis de Broglie. The revolution in physics. A non-mathematical survey of quanta. Routledge and Kegan Paul Ltd., London, UK, 1954. 310 pp.

deBroglie:1954:TGP

[dB54d] Louis de Broglie. Théorie générale des particules à spin. (French) [General theory of spin particles]. Gauthier-Villars, Paris, France, second edition, 1954. vi + 209 + 7 pp.

deBroglie:1954:INM

[dB54e] Louis de Broglie. Une interprétation nouvelle de la mécanique ondulatoire est-elle possible ?. (French) [Is a new interpretation of wave mechanics possible?]. Nuovo Cimento, 1(37):50–??, ???? 1954.

deBroglie:1954:NDF

[dB54f] Louis de Broglie. Une nouvelle démonstration de la formule du guidage en théorie de la double solution. (French) [A new demonstration of the guiding formula in the double-solution theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 239(??):737–739, ???? 1954. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1955:BMT

[dB55a] Louis de Broglie. Balance method in thermal stress analysis. In Proceedings of the Conference on High-Speed Aeronautics, January 20–22, 1955, pages 189–211. ????, ????, 1955.

deBroglie:1955:DOC

[dB55b] Louis de Broglie. Le dualisme des ondes et des corpuscules dans l’œuvre d’Albert Einstein. (French) [The dualism of waves and corpuscles in the work of Albert Einstein]. Public lecture of 5 December 1955, 1955.

deBroglie:1955:NNA

[dB55c] Louis de Broglie. Notice nécrologique sur Albert Einstein. (French) [Obituary of Albert Einstein]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 240(??):1741– 1745, ???? 1955. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1955:ORO

[dB55d] Louis de Broglie. Ondes régulières et ondes à région singulière en mécanique ondulatoire. (French) [Regular waves and singularregion waves in wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 241(??):345–348, ???? 1955. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1955:STP

[dB55e] Louis de Broglie. Séminaire de théories physiques. 24e année 1954/55. (French) [Seminar of physical theories. 24th year 1954/55]. Secrétariat mathématique (Faculté des Sciences de Paris), 1955. Paris, France pp.

deBroglie:1955:INM

[dB55f] Louis de Broglie. Une interprétation nouvelle de la mécanique ondulatoire est-elle possible?. (French) [Is a new interpretation of wave mechanics possible?]. Il Nuovo Cimento (10), 1:37–50, 1955. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

deBroglie:1956:FSP

[dB56a] Louis de Broglie. Faculté des Sciences de Paris. Séminaire de Théorie physique, dirigé par Louis de Broglie. 25e année 1955/56. (French) [Faculty of Science of Paris. Seminar of Physical Theory, directed by Louis de Broglie. 25th year 1955/56]. Secrétariat mathématique, Paris, France, 1956. ???? pp.

deBroglie:1956:IPE

[dB56b] Louis de Broglie. Illustration par un exemple de la forme des fonctions d’ondes singulières de la théorie de la double solution. (French) [Illustration by an example of the form of singular-wave functions in double-solution theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 243(??):617– 620, ???? 1956. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1956:SDP

[dB56c] Louis de Broglie. La signification du |Ψ|^{2} pour les états stationnaires dans l’interprétation causale de la mécanique ondulatoire. (French) [The meaning of |Ψ|^{2} for stationary states in the causal interpretation of wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 243(??):689– 692, ???? 1956. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1956:NPM

[dB56d] Louis de Broglie. Nouvelles perspectives en microphysique. (French) [New prospects in microphysics]. Collection Sciences d’aujourd’hui, dirigèe par Andrè George. Albin Michel, Paris, France, 1956. ???? pp.

deBroglie:1956:TIC

[dB56e] Louis de Broglie. Une tentative d’interprétation causale et non linéaire de la mécanique ondulatoire. (La théorie de la double solution). (French) [An attempt at causal and nonlinear interpretation of wave mechanics. (Double-solution theory)]. Gauthier-Villars, Paris, France, 1956. vii + 297 pp. URL http://www.yoga6d. org/debroglie_vs_bohm/. English translation in [dB60b].

deBroglie:1957:FSP

[dB57a] Louis de Broglie. Faculté des Sciences de Paris. Séminaire de théories physiques, dirigé par Louis de Broglie. (French) [Faculty of Science of Paris. Seminar of physical theories, directed by Louis de Broglie]. Secrétariat mathématique, Paris, France, 1957. ???? pp.

deBroglie:1957:TMM

[dB57b] Louis de Broglie. La Théorie de la Mesure en Mécanique Ondulatoire (interprétation usuelle et interprétation causale). (French) [The Theory of Measurement in Wave Mechanics (usual interpretation and causal interpretation)]. Gauthier-Villars, Paris, France, 1957. ???? pp. LCCN ????

deBroglie:1957:MOD

[dB57c] Louis de Broglie. Mécanique ondulatoire du photon et théorie quantique des champs. (French) [Wave mechanics of the photon and quantum field theory]. Gauthier-Villars, Paris, France, second edition, 1957. vi + 208 pp.

deBroglie:1957:NVO

[dB57d] Louis de Broglie. Notice sur la vie et l’œuvre d’Emile Borel. (French) [Notice on the life and work of Emile Borel]. Public lecture of 9 December 1957, 1957.

deBroglie:1957:TMW

[dB57e] Louis de Broglie. The Theory of Measurement in Wave Mechanics (Usual Interpretation and Causal Interpretation), volume 7 of The Great Problems of Science. Gauthier-Villars, Paris, France, 1957. iv + 97 pp. LCCN ???? URL http://

neo-classical-physics.info/uploads/3/0/6/5/3065888/de_
broglie_-_hidden_thermodynamics_book.pdf. Translated by

D. H. Delphenich.

deBroglie:1958:FSP

[dB58a] Louis de Broglie. Faculté des Sciences de Paris. Séminaire de théories physiques, dirigé par Louis de Broglie. (27e année 1957/58). (French) [Faculty of Sciences of Paris. Seminar of physical theories, directed by Louis de Broglie. (27th year 1957/58)], volume 15. Secrétariat mathématique, Paris, France, 1958.

deBroglie:1958:PM

[dB58b] Louis de Broglie. “Physics and metaphysics”. Nature, 181(4626): 1814, June 28, 1958. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

deBroglie:1958:NPF

[dB58c] Louis de Broglie. Sur la nomenclature des particules. (French) [On particle nomenclature]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 247(??):1069–??, ???? 1958. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1958:TRE

[dB58d] Louis de Broglie. Tentative de raccord entre l’équation de Heisenberg et l’équation de l’onde u en théorie de la double solution. (French) [Attempt at connecting the Heisenberg equation to the wave equation u in the double-solution theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 246(??): 2077–??, ???? 1958. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1959:DRR

[dB59a] Louis de Broglie. Deux remarques en relation avec le problème du disque tournant en théorie de la relativité. (French) [Two remarks related to the problem of the rotating disc in the theory of relativity]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 249(??):1426–??, ???? 1959. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1959:ETQ

[dB59b] Louis de Broglie. Éléments de théorie des quanta et de mécanique ondulatoire. (French) [Elements of quantum theory and wave mechanics]. Traité de Physique Théorique et de Physique Mathématique, III. Gauthier-Villars, Paris, France, 1959. xiii + 302 pp.

deBroglie:1959:LMO

[dB59c] Louis de Broglie. L’interprétation de la mécanique ondulatoire. (French) [The interpretation of wave mechanics]. Journal de Physique et le Radium, 20:963–979, 1959. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1960:LQT

[dB60a] Louis de Broglie. La lumière, les quanta et la technique de l’èclairage. (French) [Light, quanta and lighting technique]. Lecture given at the Palais de la Decouverte on 16 January 1960., 1960.

deBroglie:1960:NLW

[dB60b] Louis de Broglie. Non-linear wave mechanics. A causal interpretation. Elsevier, Amsterdam, The Netherlands, 1960. xii + 304 + 20 pp. English translation of [dB56e].

deBroglie:1960:NVO

[dB60c] Louis de Broglie. Notice sur la vie et l’œuvre de Frédéric Joliot. (French) [Notice on the life and work of Frédéric Joliot]. Public lecture of 14 December 1960., 1960.

deBroglie:1960:SSF

[dB60d] Louis de Broglie. Sur les sentiers de la science. (French) [On the Paths of Science]. Albin Michel, Paris, France, 1960. ???? pp. TO DO: Find Italian translation from Paolo Boringhieri, Turin,

deBroglie:1961:INT

[dB61a] Louis de Broglie. Introduction à la nouvelle théorie des particules de M. Jean-Pierre Vigier et de ses collaborateurs. (French) [Introduction to the new particle theory of M. Jean-Pierre Vigier and his collaborators]. Gauthier-Villars, Paris, France, 1961. xx + 108 pp. English translation in [dB63b].

deBroglie:1961:TPI

[dB61b] Louis de Broglie. La thermodynamique de la particule isolée. (French) [Thermodynamics of the isolated particle]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 253 (??):1078–??, ???? 1961. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1961:TDC

[dB61c] Louis de Broglie. Sur la thermodynamique du corpuscule isolé. (French) [On the thermodynamics of the isolated particle]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 253(??):1078–1081, ???? 1961. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1962:NVO

[dB62a] Louis de Broglie. Notice sur la vie et l’œuvre de Georges Darmois. (French) [Notice on the life and work of Georges Darmois]. Public lecture of 9 December 1962., 1962.

deBroglie:1962:NPT

[dB62b] Louis de Broglie. Nouvelle présentation de la thermodynamique de la particule isolée. (French) [New presentation of the thermodynamics of the isolated particle]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 255(??):87–??, ???? 1962. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1962:QCT

[dB62c] Louis de Broglie. Quelques conséquences de la thermodynamique de la particule isolée. (French) [Some consequences of the thermodynamics of the isolated particle]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 255(??):1052– ??, ???? 1962. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1962:RID

[dB62d] Louis de Broglie. Remarques sur l’interprétation de la dualité des ondes et des corpuscules. (French) [Remarks on the interpretation of the duality of waves and particles]. Cahiers de Phys., 16(147): 425–445, October 1962.

deBroglie:1962:Ua

[dB62e] Louis de Broglie. [unknown]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 255(??):807–??, ???? 1962. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1962:Ub

[dB62f] Louis de Broglie. [unknown]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 255(??):1052–??, ???? 1962. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1963:ECB

[dB63a] Louis de Broglie. Étude critique des bases de l’interprétation actuelle de la mécanique ondulatoire. (French) [Critical study of the bases of the current interpretations of wave mechanics]. Gauthier-Villars, Paris, France, 1963. ???? pp.

deBroglie:1963:IVT

[dB63b] Louis de Broglie. Introduction to the Vigier theory of elementary particles. With a chapter by Jean-Pierre Vigier. Translation to English by Arthur J. Knodel of [dB61a]. Elsevier, Amsterdam, The Netherlands, 1963. vi + 139 pp. English tr.

deBroglie:1963:PNQ

[dB63c] Louis de Broglie. La physique nouvelle et les quanta. (French) [The new physics and quanta]. Staatsverlag für Literatur des Staatlichen Ausschusses zur Ausnutzung der Atomenergie der UdSSR, Moscow, USSR, 1963. 231 pp. Translation from French by S. P. Bakanov and L. M. Kovrizhnykh.

deBroglie:1963:VOU

[dB63d] Louis de Broglie. La vie et l’œuvre de Niels Bohr. (French) [The life and works of Niels Bohr]. Annals of Physics, 13(8):117–120, 1963. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035X (electronic).

deBroglie:1963:NPT

[dB63e] Louis de Broglie. Nouvelle présentation de la thermodynamique de la particule isolée. (French) [New presentation of the thermodynamics of the isolated particle]. Bul. Inst. Politeh. Iaşi, N. Ser, 9(3–4):91–94, 1963. ISSN 0032-6100.

deBroglie:1963:TFC

[dB63f] Louis de Broglie. Sur la théorie des foyers cinétiques dans la thermodynamique de la particule isolée. (French) [On the theory of kinetic foci in the the thermodynamics of the isolated particle]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 251(??):1822–??, ???? 1963. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1963:IEL

[dB63g] Louis de Broglie. Sur l’introduction de l’énergie libre dans la thermodynamique cachée des particules. (French) [On the introduction of free energy in the hidden thermodynamics of particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 257(??):1430–1433, ???? 1963. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1963:PTL

[dB63h] Louis de Broglie. Sur un point de la théorie des lasers. (French) [On a point of the theory of lasers]. Communication to l’Academie des Sciences de Lisbonne., November 21, 1963.

deBroglie:1964:CIW

[dB64a] Louis de Broglie. The current interpretation of wave mechanics. Elsevier, Amsterdam, The Netherlands, 1964. 94 pp.

deBroglie:1964:TCP

[dB64b] Louis de Broglie. La thermodynamique “cachée” des particules. (French) [The “hidden” thermodynamics of particles]. Ann. Inst. H. Poincaré Sect. A (N.S), 1:1–19, 1964.

deBroglie:1964:TPI

[dB64c] Louis de Broglie. La thermodynamique de la particule isolée (thermodynamique cachée des particules). (French) [Thermodynamics of the isolated particle (hidden thermodynamics of particles)]. Gauthier-Villars, Paris, France, 1964. ???? pp. LCCN ????

deBroglie:1964:NVO

[dB64d] Louis de Broglie. Notice sur la vie et l’œuvre de Jean Becquerel. (French) [Notice on the life and work of Jean Becquerel]. Public lecture of 9 December 1964., 1964.

deBroglie:1964:OEP

[dB64e] Louis de Broglie. Ondes électromagnétiques et photons. (French) [Electromagnetic waves and photons]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 258(??):6345– 6347, ???? 1964. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1964:TIP

[dB64f] Louis de Broglie. The Thermodynamics of the Isolated Particle, volume 17 of The Great Problems of Science. Gauthier-Villars, Paris, France, 1964. iv + 98 pp. LCCN ???? URL http://

neo-classical-physics.info/uploads/3/0/6/5/3065888/de_
broglie_-_hidden_thermodynamics_book.pdf. Translated by

D. H. Delphenich.

deBroglie:1965:ETE

[dB65a] Louis de Broglie. Einführung in die Theorie der Elementarteilchen von Jean-Pierre Vigier und Mitarbeitern. (German) [Introduction to the theory of elementary particles by Jean-Pierre Vigier and co-workers]. Mit einem Kapitel von Jean-Pierre Vigier. Übersetzt von Bodo W. Zimme rmann. Verlag G. Braun, Karlsruhe, West Germany, 1965. viii + 124 pp.

deBroglie:1965:CPO

[dB65b] Louis de Broglie. La coexistence des photons et des ondes dans les rayonnements électromagnétiques et la théorie de la double solution. (French) [The coexistence of photons and waves in electromagnetic radiation and double-solution theory]. Énergie nuclèaire, 7(3):??, May 1965.

deBroglie:1965:NVO

[dB65c] Louis de Broglie. Notice sur la vie et l’œuvre de Camille Gutton. (French) [Notice on the life and work of Camille Gutton]. Public lecture of 11 December 1965., 1965.

deBroglie:1965:RI

[dB65d] Louis de Broglie. Sur la relation d’incertitude δn · δφ ≥ 2π. (French) [On the uncertainty relation δn · δφ ≥ 2π]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 260(??):6041–6043, ???? 1965. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1966:CIS

[dB66a] Louis de Broglie. Certitudes et incertitudes de la science. (French) [Certainties and uncertainties of science]. Albin Michel, Paris, France, 1966. ???? pp.

deBroglie:1966:TRQ

[dB66b] Louis de Broglie. Sur la transformation relativiste de la quantile de chaleur et de la température et la thermodynamique cachée de la particule. (French) [On the relativistic transformation of the heat quantity and temperature and the hidden thermodynamics of the particle]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 262(??):1235–??, ???? 1966. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1966:DRE

[dB66c] Louis de Broglie. Sur le déplacement des raies émises par un objet astronomique lointain. (French) [On the displacement of the lines emitted by a distant astronomical object]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 263(??):589–??, ???? 1966. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1966:WMM

[dB66d] Louis de Broglie. Wave Mechanics and Molecular Biology. Adiwes international series in biology. Addison-Wesley, Reading, MA, USA, 1966. v + 186 pp. LCCN QH506 .B763.

deBroglie:1967:I

[dB67a] Louis de Broglie. Interview. Video (13m50s), January 1, 1967.

URL http://www.ina.fr/video/AFE04002106. In French.

deBroglie:1967:DDG

[dB67b] Louis de Broglie. La dynamique du guidage dans un milieu réfringent et dispersif et la théorie des antiparticules. (French) [The dynamics of guidance in a refractive and dispersive medium and the theory of antiparticles]. J. Phys. France, 28(5–6):481–486, May/June 1967.

deBroglie:1967:MBP

[dB67c] Louis de Broglie. Le mouvement brownien d’une particule dans son onde. (French) [Brownian motion of a particle in its wave]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 264(??):1041–??, ???? 1967. CODEN CHD-BAN. ISSN 0366-6077.

deBroglie:1967:NVO

[dB67d] Louis de Broglie. Notice sur la vie et l’œuvre d’Albert Pérard. (French) [Notice on the life and work of Albert Pérard]. Public lecture of 11 December 1967., 1967.

deBroglie:1967:DCM

[dB67e] Louis de Broglie. Sur la dynamique des corps à masse propre variable et la formule de transformation relativiste de la chaleur. (French) [On the dynamics of bodies with variable mass and the relativistic transformation formula of heat]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 264(??):1173–??, ???? 1967. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1967:FBM

[dB67f] Louis de Broglie. Sur la formule Q = Q_{0}\sqrt{1 − β^{2}} et les bases de la mécanique ondulatoire. (French) [On the formula Q = Q_{0}\sqrt{1 − β^{2}} and the bases of wave mechanics]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 264(??): 1041–??, ???? 1967. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1967:ETR

[dB67g] Louis de Broglie. Sur l’équation ΔW = ΔQ + ΔL en thermodynamique relativiste. (French) [On the equation ΔW = ΔQ + ΔL in relativistic thermodynamics]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 265(??): 437–??, ???? 1967. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1967:DRF

[dB67h] Louis de Broglie. Sur les discussions relatives à la formule Q = Q_{0}\sqrt{1 − β^{2}} et la définition de la pression en thermodynamique relativiste. (French) [On discussions relative to the formule Q = Q_{0}\sqrt{1 − β^{2}} and the definition of pressure in relativistic thermodynamics]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 265(??):889–?, ???? 1967. CODEN CHDBAN. ISSN 0366-6077.

deBrojl:1967:KKT

[dB67i] L. de Brojl’. Kvanty, kineticheskaja teorija gazov i princip Ferma. (Russian) [quanta, kinetic theory of gases, and the Fermat principle]. Physics-Uspekhi, 93(??):182–183, September 1967. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL

http://ufn.ru/ru/articles/1967/9/k/. Russian translation of

Les quanta, la théorie cinétique des gas et le principe de Fermat, Comp. Rend. 177, 630 (1923) [dB23a].

deBrojl:1967:KSD

[dB67j] L. de Brojl’. Kvanty sveta, difrakcija i interferencija. (Russian) [Quanta of light, diffraction and interference]. Physics-Uspekhi, 93(??):180–181, September 1967. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL http://ufn.ru/ru/ articles/1967/9/j/. Translation to Russian of [dB23c].

deBrojl:1967:VKR

[dB67k] L. de Brojl’. Volny i kvanty. (Russian) [Waves and quanta]. Physics-Uspekhi, 93(??):178–180, September 1967. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL

http://ufn.ru/ru/articles/1967/9/i/. Translation to Rus-

sian of [dB23b].

deBroglie:1968:IRE

[dB68a] Louis de Broglie. Interpretation of a recent experiment on interferences of photons beams. Physical Review (2), 172(5):1284–??, August 25, 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

deBroglie:1968:RMO

[dB68b] Louis de Broglie. La réinterpretation de la mécanique ondulatoire. (French) [The reinterpretation of wave mechanics]. Physics Bulletin, 19(??):133–??, May 1968. CODEN PHSBB4. ISSN 0031-9112.

deBroglie:1968:TRT

[dB68c] Louis de Broglie. La thermodynamique relativiste et la thermodynamique cachée des particules. (French) [Relativistic thermodynamics and hidden thermodynamics of particles]. International Journal of Theoretical Physics, 1(1):1–23, May 1968. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

deBroglie:1968:OEP

[dB68d] Louis de Broglie. Ondes électromagnétiques et photons. (French) [Electromagnetic waves and photons]. Gauthier-Villars, Paris, France, 1968. ???? pp. LCCN ????

deBroglie:1968:QVP

[dB68e] Louis de Broglie. Quelques vues personnelles sur l’évolution de la physique théorique. (French) [Some personal views on the evolution of theoretical physics]. In La philosophic contemporaine, chronique, pages 217–222. La Nuova Italia, Florence, Italy, 1968.

deBroglie:1968:AMO

[dB68f] Louis de Broglie. Sur l’application de la mécanique ondulatoire à la théorie des guides d’ondes. (French) [On the application of wave mechanics to the theory of waveguides]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 266(??):1233, ???? 1968. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1968:TRM

[dB68g] Louis de Broglie. Thermodynamique relativiste et mécanique ondulatoire. (French) [Relativistic thermodynamics and wave mechanics]. Annales de l’Institut Henri Poincaré, 9(2):89–108, ???? 1968. CODEN AIHPA2. ISSN 0365-320X (print), 2400-4855 (electronic).

deBroglie:1969:NVO

[dB69a] Louis de Broglie. Notice sur la vie et l’œuvre Bernard Lyot. (French) [Notice on the life and work of Bernard Lyot]. Public lecture of 8 December 1969., 1969.

deBroglie:1969:LRI

[dB69b] Louis de Broglie. Sur l’interpretation des relations d’incertitude. (French) [On the interpretation of uncertainty relations]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 268(??):277–??, ???? 1969. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1970:MAD

[dB70a] Louis de Broglie. Mon anxiété devant le problème des quanta. (French) [My anxiety about the problem of quanta]. In Homberger et al. [HJS70], pages 181–188. ISBN 0-224-61914-4. LCCN AC5 .H64.

deBroglie:1970:OEP

[dB70b] Louis de Broglie. Ondes électromagnetiques et photons en radioélectricité. (French) [Electromagnetic waves and photons in radioelectricity]. Onde électrique, ??(??):657–??, September 1970.

deBroglie:1970:RWM

[dB70c] Louis de Broglie. The reinterpretation of wave mechanics. Foundations of Physics, 1(1):5–15, ???? 1970. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF00708650.

deBroglie:1970:SMQ

[dB70d] Louis de Broglie. Spins et moments de quantité de mouvement. (French) [Spins and moments of momentum]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 272(??):349, ???? 1970. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1970:PDM

[dB70e] Louis de Broglie. Sur un problème du mouvement d’une particule dans un milieu réfringent. (French) [On a problem of the motion of a particle in a refractive medium]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 272(??): 1333–1335, June 14, 1970. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1970:NPF

[dB70f] Louis de Broglie. Sur une nouvelle présentation des formules de la mécanique ondulatoire. (French) [On a new presentation of the formulas of the wave mechanics]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 271(??): 559–??, ???? 1970. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1970:VEH

[dB70g] Louis de Broglie. Vue d’ensemble sur l’histoire et l’interprétation de la mécanique ondulatoire. (French) [Overview of the history and interpretation of wave mechanics]. Revue française de l’Électricité, 43(228):13–??, ???? 1970.

deBroglie:1971:JPP

[dB71a] Louis de Broglie. Jean Perrin physicien. (French) [Jean Perrin physicist]. Revue d’Histoire des Sciences et de Leurs Applications, 24(2):99–105, 1971. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://www.jstor.org/

stable/24468260.

deBroglie:1971:LMO

[dB71b] Louis de Broglie. L’interprétation de la mécanique ondulatoire par la théorie de la double solution. (French) [The interpretation of wave mechanics by double-solution theory]. In Réunion de Varenna Foundations of Quantum Mechanics, page ?? Il Corso Academic Press, New York, NY, USA, 1971.

deBroglie:1971:NVOa

[dB71c] Louis de Broglie. Notice sur la vie et l’œuvre d’André Danjon. (French) [Notice on the life and work of André Danjon]. Public lecture of 13 December 1971., 1971.

deBroglie:1971:NVOb

[dB71d] Louis de Broglie. Notice sur la vie et l’œuvre de André Danjon. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 274(24):147, December 13, 1971. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic). URL http://

gallica.bnf.fr/ark:/12148/bpt6k5622401d/f23/.

deBroglie:1971:SMQ

[dB71e] Louis de Broglie. Spin et moment de quantité de mouvement. (French) [Spin and momentum moment]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 272(??):349–352, February 8, 1971. ISSN 0366-6077. URL http:

//gallica.bnf.fr/ark:/12148/bpt6k480293k/f354/.

deBroglie:1971:WP

[dB71f] Louis de Broglie. Waves and particles. Physics Bulletin, 22(??): ??, ???? 1971. CODEN PHSBB4. ISSN 0031-9112.

deBroglie:1972:E

[dB72a] Louis de Broglie. Einstein. Réalités Hachette, Paris, France, 1972. ISBN ???? ???? pp. LCCN ????

deBroglie:1972:RPI

[dB72b] Louis de Broglie. Sur la répartition des potentiels d’interaction entre les particules d’un système. (French) [On the distribution of the interaction potentials between the particles of a system]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 275(??):899–??, ???? 1972. CODEN CHD-BAN. ISSN 0366-6077.

deBroglie:1973:VIB

[dB73] Louis de Broglie. Sur les véritables idées de base de la mécanique ondulatoire. (French) [On the basic ideas of wave mechanics]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 277(??):71–??, ???? 1973. CODEN CHD-BAN. ISSN 0366-6077.

deBroglie:1974:SPT

[dB74a] Louis de Broglie. Strong processes and transient states. Foundations of Physics, 4(3):321–333, September 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708539.

deBroglie:1974:RDT

[dB74b] Louis de Broglie. Sur la réfutation du théorème de Bell. (French) [On the refutation of Bell’s Theorem]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 278(??):721–??, ???? 1974. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1975:DPP

[dB75] Louis de Broglie. Discours prononcé à la première séance de la Fondation Louis de Broglie. (French) [Speech delivered at the first session of the Louis de Broglie Foundation]. Annales de la Fondation Louis de Broglie, 1(??):??, ???? 1975. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

deBroglie:1976:LAT

[dB76a] Louis de Broglie. L’invariance adiabatique et la thermodynamique cachèe des particules. (French) [Adiabatic invariance and hidden thermodynamics of particles]. Annales de la Fondation Louis de Broglie, 1(1):1–??, ???? 1976. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

deBroglie:1976:RDS

[dB76b] Louis de Broglie. Recherches d’un demi-siècle. (French) [Research of a half-century]. Albin Michel, Paris, France, 1976. ???? pp.

deBroglie:1976:RPC

[dB76c] Louis de Broglie. Rèflexions sur la physique contemporaine. (French) [Reflections on contemporary physics]. Annales de la Fondation Louis de Broglie, 1(2):49–??, ???? 1976. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

deBroglie:1976:TRD

[dB76d] Louis de Broglie. Treize remarques sur divers sujets de physique théorique. (French) [Thirteen remarks on various subjects of theoretical physics]. Annales de la Fondation Louis de Broglie, 1 (3):116–??, ???? 1976. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

deBroglie:1977:RC

[dB77a] Louis de Broglie. Réflexions sur la causalité. (French) [Reflections on causality]. Annales de la Fondation Louis de Broglie, 2(2):69– ??, ???? 1977. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

deBroglie:1977:TDC

[dB77b] Louis de Broglie. Toute description complète de la réalité implique l’intervention de la causalité. (French) [Any complete description of reality implies the intervention of causality]. Annales de la Fondation Louis de Broglie, 2(2):133–??, ???? 1977. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

deBrojl:1977:PPT

[dB77c] L. de Brojl’. Popytka postroenija teorii svetovyh kvantov. (Russian) [Attempt to construct a theory of light quanta]. Physics-Uspekhi, 122(??):572–571, August 1977. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). URL http://ufn.ru/

ru/articles/1977/8/a/.

deBroglie:1978:JPN

[dB78] Louis de Broglie. Jalons pour une nouvelle microphysique. (French) [Milestones for a new microphysics]. Bordas, Paris, France, 1978. ISBN 2-04-010147-0. v + 182 pp. Exposé d’ensemble sur l’interprétation de la mécanique ondulatoire, Discours de la Méthode.

deBroglie:1979:NLR

[dB79] Louis de Broglie. Nécessité de la liberté dans la recherche scientifique. (French) [The need for freedom in scientific research]. Annales de la Fondation Louis de Broglie, 4(1):62–??, ???? 1979. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

deBroglie:1982:IHI

[dB82] Louis de Broglie. Les incertitudes d’Heisenberg et l’interprétation probabiliste de la mécanique ondulatoire. (French) [Heisenberg uncertainties and the probabilistic interpretation of wave mechanics]. Gauthier-Villars, Paris, France, 1982. ISBN 2-04-015411-6. xlii + 304 pp. With a preface and with notes by Georges Lochak.

deBroglie:1983:TIU

[dB83] Louis de Broglie. On the true ideas underlying wave mechanics. In Old and new questions in physics, cosmology, philosophy, and theoretical biology, pages 83–86. Plenum Press, New York, NY, USA; London, UK, 1983.

deBroil:1986:CSN

[dB86] L. de Broĭl’. Sootnosheniya neopredelennosteĭ Geĭzenberga i veroyatnostnaya interpretatsiya volnovoĭ mekhaniki. “Mir”, Moscow, USSR, 1986. 341 pp. With a preface and notes by Zh. Loshak. Translated from the French by N. V. Samsonenko. Translation edited and with a preface by G. Z. Zaĭtsev.

deBroglie:1990:HUP

[dB90] Louis de Broglie. Heisenberg’s Uncertainty Principle and the Probabilistic Interpretation of Wave Mechanics, volume 40 of Fundamental Theories of Physics. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-0929-4. xliv + 302 pp. LCCN QC174.2 .B653513 1990. With critical notes of the author. With a preface and supplementary notes by Georges Lochak. Translated from the French by Alwyn van der Merwe. With a preface by Asim Barut.

deBroglie:1995:DQM

[dB95] Louis de Broglie. Diverses questions de mécanique et de thermodynamique classiques et relativistes. (French) [Various questions of mechanics and classical relativistic thermodynamics], volume 32 of Lecture Notes in Physics. New Series m: Monographs. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-59446-9. xii + 198 pp. Based on an unpublished manuscript by de Broglie, Edited and with a preface and an introduction by Georges Lochak, Michel Karatchentzeff and Daniel Fargue.

deBroglie:1997:DPCb

[dB97a] Louis de Broglie. Divers problèmes concernant les ondes et les corpuscules. Cours professé à l’Institut Henri Poincaré en 1956– 1957. (French) [Various problems concerning waves and particles. Course offered at the Institut Henri Poincaré in 1956–1957]. Annales de la Fondation Louis de Broglie, 22(4):335–359, 1997. ISSN 0182-4295 (print), 2108-6397 (electronic).

deBroglie:1997:DPCa

[dB97b] Louis de Broglie. Divers problèmes concernant les ondes et les corpuscules (suite). (French) [Various problems concerning waves and particles (continued)]. Annales de la Fondation Louis de Broglie, 22(3):221–236, 1997. ISSN 0182-4295 (print), 2108-6397 (electronic). Cours professé à l’institute Henri Poincaré en 1956–1957 [Course given at the Institut Henri Poincaré in 1956–1957].

deBroglie:1998:EDQ

[dB98a]  Louis de Broglie.  Étude de deux questions concernant

l’interprétation de la mécanique ondulatoire par la théorie de la double solution. (French) [A study of two questions concerning the interpretation of wave mechanics by double-solution theory]. Annales de la Fondation Louis de Broglie, 23(1):1–10, 1998. ISSN 0182-4295 (print), 2108-6397 (electronic).

deBroglie:1998:TDC

[dB98b] Louis de Broglie. Théorie du “champ soustractif” ou “champ unifié”. (French) [Theory of the “subtractive field” or “unified field”]. Annales de la Fondation Louis de Broglie, 23(2):57–69, 1998. ISSN 0182-4295 (print), 2108-6397 (electronic).

deBroglie:19xx:MAP

[dBxx] Louis de Broglie. Mon anxiété sur le problème des quanta. (French) [My anxiety about the problem of quanta]. Audio record-

ing (10m9s)., 19xx. URL http://gallica.bnf.fr/ark:/12148/
bpt6k88298546?rk=236052;4.

deBroglie:2004:TQ

[dB04] Louis de Broglie. On the Theory of Quanta. Annales de la Fondation Louis de Broglie, Paris, France, 2004. vi + 73

pp. URL http://aflb.ensmp.fr/LDB-oeuvres/De_Broglie_

Kracklauer.pdf. English translation of doctoral thesis [dB25b].

Broglie:2006:XTT

[dB06] Louis de Broglie. XXXV. a tentative theory of light quanta. Philosophical Magazine Letters, 86(7):411–423, July 2006. CODEN PMLEEG. ISSN 0950-0839 (print), 1362-3036 (electronic). Republication of [dB24d].

deBroglie:2009:MSP

[dB09a] Louis de Broglie. La muséologie scientifique et le Palais de la Découverte. (French) [Scientific museology and the Palace of Discovery]. Museum International (Edition Française), 2(3):141–149, April 2009.

deBroglie:2009:SMP

[dB09b] Louis de Broglie. Scientific museology and the Palais de la Découverte. Museum International, 2(3):141–149, April 2009.

deBroglie:1957:INC

[dBA57] Louis de Broglie and João Andrade e Silva. Idées nouvelles concernant les systèmes de corpuscules dans l’interprétation causale de la mécanique ondulatoire. (French) [New ideas concerning particle systems in the causal interpretation of wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 244(??):529–533, ???? 1957. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1966:IOH

[dBA66] Louis de Broglie and João Andrade e Silva. Sur l’interprétation de l’opérateur hamiltonien H_{0p} et de l’opérateur ^{≪}carré du moment angulaire ^{≪}M_{op} de la mécanique quantique. (French) [On the interpretation of the Hamiltonian operator H_{0p} and the “angular-momentum squared” operator M_{op} of quantum mechanics]. Comptes Rendus des Séances de l’Académie des Sciences, Série A: Sciences Mathématiques, 263(??):645–??, ???? 1966. ISSN 0366-6034.

deBroglie:1969:CPM

[dBA69] Louis de Broglie and João Andrade e Silva. Sur le choc des particules en mécanique ondulatoire. (French) [On the shock of particles in wave mechanics]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 268(??):1449–1451, June 9, 1969. CODEN CHDBAN. ISSN 0366-6077. URL http:

//gallica.bnf.fr/ark:/12148/bpt6k480289h/f1468/.

deBroglie:1971:RMO

[dBA71] Louis de Broglie and Joãao Luis Andrade e Silva. La réinterprétation de la mécanique ondulatoire. 1re partie: Principes généraux. (French) [The reinterpretation of wave mechanics. Part 1: General principles]. Gauthier-Villars, Paris, France, 1971. ISBN 2-04-006118-5 (paperback). 194 pp. LCCN QC174.2 B654.

deBroglie:1966:E

[dBAS^{+}66] Louis de Broglie, Louis Armand, Pierre Henri Simon, et al., editors. Einstein. Hachette, Paris, France, 1966. ???? pp.

deBroglie:1979:E

[dBAS^{+}79] Louis de Broglie, Louis Armand, Pierre Henri Simon, et al., editors. Einstein. Peebles Press, New York, NY, USA, 1979. ISBN 0-85690-070-2. 219 pp. LCCN QC16.E5 E34E 1979; QC16.E5 E3613; QC16.E5 D4. Translation to English of [dBAS^{+}66].

deBroglie:1928:SPW

[dBB28] Louis de Broglie and Léon Brillouin. Selected papers on wave mechanics. Blackie & Son, London, UK, 1928. vi + 151 pp. Authorised translation to English by W. M. Deans.

deBroglie:1962:NIS

[dBBH^{+}62] L. de Broglie, D. Bohm, F. Halbwachs, T. Takabayasi, P. Phillion, and J. P. Vigier. New isobaric spin group. In Proceedings, Conférence Internationale d’Aix-en-Provence sur les Particules Elémentaires: Aix-en-Provence, France, September 14–20, 1961, volume 1, pages 503–506. ????, ????, 1962. URL http://

inspirehep.net/record/1382687/files/Pages_from_C61-09-
14_503.pdf.

deBroglie:1963:RME

[dBBH^{+}63] Louis de Broglie, David Bohm, Pierre Hillion, Francis Halbwachs, Takehiko Takabayasi, and Jean-Pierre Vigier. Rotator model of elementary particles considered as relativistic extended structures in Minkowski space. Physical Review (2), 129:438–450, 1963. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

deBroglie:1921:SEA

[dBD21] Louis de Broglie and A. Dauvillier. Sur la structure électronique des atomes lourdes. (French) [On the electronic structure of heavy atoms]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 172(??):1650–1653, June 27, 1921. CO-DEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

URL http://gallica.bnf.fr/ark:/12148/bpt6k6237828h/f22/

.

deBroglie:1922:SSR

[dBD22a] Louis de Broglie and A. Dauvillier. Sur le système spectral des rayons Rœntgen. (French) [On the spectral system of X-rays]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 175(??):685–688, October 23, 1922. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic). URL http://

gallica.bnf.fr/ark:/12148/bpt6k3128v/f685/.

deBroglie:1922:ASE

[dBD22b] Louis de Broglie and A. Dauvillier. Sur les analogies de structure entre les séries optiques et les séries Röntgen. (French) [On structural analogies between optical spectra and X-ray series]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 175(??):755–758, July/December 1922. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic). URL http://gallica.

bnf.fr/ark:/12148/bpt6k3128v/f755.item.r=de%20Broglie.

deBroglie:1924:SSR

[dBD24] Louis de Broglie and A. Dauvillier. Le système spectral des rayons Röntgen et structure de l’atome. (French) [The spectral system of X-rays and structure of the atom]. Journal de Physique et le Radium, 5(1):1–19, January 1924. CODEN JPRAAJ. ISSN 0368-3842.

deBroglie:1935:TKM

[dBD35] Louis de Broglie and Jean-Louis Destouches. Sur le théorème de Koenig en mécanique ondulatoire. (French) [On Koenig’s theorem in wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 201(??):369–??, ???? 1935. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1921:MAB

[dBdB21] Maurice de Broglie and Louis de Broglie. Sur le modèle d’atome de Bohr et les spectres corpusculaires. (French) [On the Bohr atomic model and particle spectra]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 172(??):746–748, March 21, 1921. CODEN COREAF. ISSN 0001-4036 (print), 2419-

6304 (electronic). URL http://gallica.bnf.fr/ark:/12148/
bpt6k6237814g/f18/.

deBroglie:1922:RSS

[dBdB22] Maurice de Broglie and Louis de Broglie. Remarques sur les spectres corpusculaires et l’effect photoélectrique. (French) [Remarks on particle spectra and the photoelectric effect]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 175(??): 1139–1141, December 4, 1922. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic). URL http://gallica.bnf.

fr/ark:/12148/bpt6k3128v/f1139.

deBroglie:1924:VEP

[dBdB24] Maurice de Broglie and Louis de Broglie. Sur la vérification expérimentale des projections d’électrons prévues lors de la diffusion des rayons X par les considérations de Compton et Debye. (French) [On the experimental verification of projections of electrons expected during diffusion of X-rays by the considerations of Compton and Debye]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 178(??):383–384, January 21, 1924. CODEN COREAF. ISSN 0001-4036 (print), 2419-

6304 (electronic). URL http://gallica.bnf.fr/ark:/12148/
bpt6k3131z/f383/.

deBroglie:1928:IPR

[dBdB28] Louis de Broglie and Maurice de Broglie. Introduction à la physique des rayons X et des rayons γ. (French) [Introduction to X-ray and Gamma Ray Physics]. Gauthier-Villars, Paris, France, 1928. 201 pp. German translation in [dB30j].

deBroglie:1934:QCN

[dBdB34] Maurice de Broglie and Louis de Broglie. Quelques considérations sur les notions d’onde et de corpuscule. (French) [Some considerations on the notions of wave and particle]. Scientia (Milan), 55: 177–185, 1934. CODEN SCIMAI. ISSN 0036-8687 (print), 1825-4373 (electronic).

deBroglie:1984:QST

[dBDW^{+}84] Louis de Broglie, P. A. M. (Paul Adrien Maurice) Dirac, Eugene Paul Wigner, A. O. (Asim Orhan) Barut, Alwyn Van der Merwe, and Jean-Pierre Vigier, editors. Quantum, space, and time — the quest continues: studies and essays in honour of Louis de Broglie, Paul Dirac, and Eugene Wigner. Cambridge monographs on physics. Cambridge University Press, Cambridge, UK, 1984. ISBN 0-521-31911-0 (paperback). vii + 659 pp. LCCN QC174.125 .Q376 1984.

deBroglie:1963:STM

[dBHH^{+}63] Louis de Broglie, Francis Halbwachs, Pierre Hillion, Takehiko Takabayasi, and Jean-Pierre Vigier. Space-time model of relativistic extended particles in Minkowski space. II. Free particle and interaction theory. Physical Review (2), 129:451–466, 1963. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

deBroglie:1959:PCR

[dBHV59] Louis de Broglie, Pierre Hillion, and Jean-Pierre Vigier. Propriétés classiques et représentation bilocale du rotateur de Nakano. (French) [Classical properties and bilocal representation of the Nakano rotator]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 249(??):2255–2258, ???? 1959. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1953:TSP

[dBJ53] Louis de Broglie and Georges Jamati. Travaux des sociétés de philosophie. (French) [Works of philosophical societies]. Les Études philosophiques, 8(2):195–197, 1953. ISSN 0014-2166 (print), 2101-0056 (electronic). URL http://www.jstor.org/

stable/20841572.

deBroglie:1976:PPH

[dBLBVP76] Louis de Broglie, Georges Lochak, Juan Alberto Beswick, and José Vassalo-Pereira. Present, predicted, and hidden probabilities. Foundations of Physics, 6(1):3–14, February 1976. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00708659.

deBroglie:1950:CML

[dBR50] Louis de Broglie and René Reulon. Sur les champs mésoniques liés à l’électron dans la nouvelle théorie du champ soustractif. (French) [On electron-linked mesonic fields in the new subtractive field theory]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 230(??), ???? 1950. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1944:LCC

[dBT44a] Louis de Broglie and Marie-Antoinette Tonnelat. L’introduction des constantes de Coulomb et de Newton en mécanique ondulatoire. (French) [Introduction of the Coulomb and Newton constants into wave mechanics]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 218(??):373–376, ???? 1944. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1944:RQD

[dBT44b] Louis de Broglie and Marie-Antoinette Tonnelat. Remarques sur quelques difficultés de la théorie du photon, liées à l’emploi d’une solution d’annihilation. (French) [Remarks on some difficulties of photon theory related to the use of an annihilation solution]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 218(??):889–892, ???? 1944. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1950:PSC

[dBT50] Louis de Broglie and Marie-Antoinette Tonnelat. Sur la possibilité d’une structure complexe pour les particules de spin 1. (French) [On the possibility of a complex structure of spin-1 particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 230(??):1329–??, ???? 1950. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1929:CDR

[dBUC^{+}29] L. de Broglie, G. Urbain, V. Carlhian, et al. Continu et discontinu. Recueil de monographies. (French) [Continuous and discontinuous. Collection of monographs]. Bloud et Gay, Paris, France, 1929. 197 pp.

deBroglie:1930:CDR

[dBUC^{+}30] L. de Broglie, G. Urbain, V. Carlhian, et al. Continu et discontinu. Recueil de monographies. (French) [Continuous and discontinuous. Collection of monographs]. Bloud et Gay, Paris, France, 1930. 199 pp.

deBroglie:1963:ATF

[dBV63a] Louis de Broglie and Jean-Pierre Vigier. Application de la théorie de la fusion au nouveau modèle étendu des particules élémentaires. (French) [Application of fusion theory to the new extended model of elementary particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 256(??):3390–??, ???? 1963. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1963:TPE

[dBV63b] Louis de Broglie and Jean-Pierre Vigier. Table des particules élémentaires associées au nouveau modèle des particules élémentaires. (French) [Table of elementary particles associated with the new model of elementary particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 256(??): 3551–??, ???? 1963. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1972:PMN

[dBV72a] L. de Broglie and J. P. Vigier. Photon mass and new experimental results on longitudinal displacements of laser beams near total reflection. Physical Review Letters, 28:1001–1004, 1972. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

deBroglie:1972:MDP

[dBV72b] Louis de Broglie and Jean-Pierre Vigier. Masse du photon. Effet Imbert et effet Goos–Hanchen en lumière incidente polarisée. (French) [Mass of the photon. Imbert effect and Goos–Hanchen effect in polarized incident light]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 273(??): 1069–??, ???? 1972. CODEN CHDBAN. ISSN 0366-6077.

deBroglie:1934:SDP

[dBW34] Louis de Broglie and Jacques Winter. Sur le spin du photon. (French) [On the photon spin]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 199(??):813–??, ???? 1934. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Destouches:1948:CEC

[DC48] Jean-Louis Destouches and Michel Cazin. Corrélation entre la complémentarité de Louis de Broglie et la matrice caractéristique de Heisenberg. (French) [Correlation between the complementarity of Louis de Broglie and the characteristic matrix of Heisenberg]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 226(??):566–567, ???? 1948. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

deBroglie:1921:DEA

[DdB21] A. Dauvillier and Louis de Broglie. Sur la distribution des électrons dans les atomes lourdes. (French) [On the distribution of electrons in heavy atoms]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 173(??):137–139, July 18, 1921. CODEN COREAF. ISSN 0001-4036 (print), 2419-

6304 (electronic). URL http://gallica.bnf.fr/ark:/12148/
bpt6k31267/f137/.

Dauvillier:1922:RTH

[DdB22] A. Dauvillier and Louis de Broglie. Remarques sur le travail de E. Hjalmar concernant la série M des éléments. (French) [Remarks on the work of E. Hjalmar on the M series of the elements]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 175(??):1198–1201, December 4, 1922. CODEN COREAF. ISSN 0001-4036 (print), 2419-

6304 (electronic). URL http://gallica.bnf.fr/ark:/12148/
bpt6k3128v/f1198.item.r=de%20Broglie.

deBroglie:1949:OEC

[de 49] Louis de Broglie, editor. Les ondes électromagnétiques centimétriques. Réunions d’études et de mises du point. (French) [Centimetric electromagnetic waves. Meetings and studies]. Éditions de la Revue d’optique théorique et Instrumentale, Paris, France, 1949. 274 pp.

Dehlinger:1936:BMQ

[Deh36] U. Dehlinger. Buchbesprechung: Materiewellen und Quantenmechanik. Eine Einführung auf Grund der Theorien von de Broglie, Schrödinger, Heisenberg und Dirac. 4. und 5. Aufl. Von Arthur Haas. Leipzig 1934, Akademische Verlagsgesellschaft. 299 S., 7 Abb. Preis geb. RM 7.80. Zeitschrift für Elektrochemie und Angewandte Physikalische Chemie, 42(1):43, 1936. CODEN ZEA-PAA. ISSN 0005-9021.

Delorme:1952:AOG

[Del52] Suzanne Delorme. Analyses d’ouvrages: Les grandes figures de la Science française par Luc Durtain; Louis de Broglie. Revue d’Histoire des Sciences et de Leurs Applications, 5(3):274, July/ September 1952. CODEN RHSAAM. ISSN 0048-7996 (print),

1969-6582 (electronic). URL http://www.jstor.org/stable/
23904393.

Destouches:1934:ATPa

[Des34] Jean-Louis Destouches. Axiomatique de la théorie des photons de M. Louis de Broglie et superquantification. (French) [Axiomatic theory of photons of M. Louis de Broglie and superquantification]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 199(??):779–781, ???? 1934. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Destouches:1939:PMP

[Des39] Jean-Louis Destouches. Physique moderne et philosophie. (French) [Modern Physics and Philosophy]. Hermann, Paris, France, 1939. 76 pp. Preface de Louis de Broglie.

Destouches:1955:LFA

[Des55a] Jean-Louis Destouches. L’onde u et le fluide associé dans la théorie de la double solution de M. Louis de Broglie. (French) [The wave u and the associated fluid in the double-solution theory of M. Louis de Broglie]. Journal de Physique et le Radium, 16:81–85, 1955. CODEN JPRAAJ. ISSN 0368-3842.

Destouches:1955:CCH

[Des55b] Jean-Louis Destouches. Sur la compatibilité de certaines hypothèses de la théorie de la double solution de M. Louis de Broglie. (French) [On the compatibility of certain hypotheses of the double-solution theory of M. Louis de Broglie]. Journal de Physique et le Radium, 16:86–91, 1955. CODEN JPRAAJ. ISSN 0368-3842.

Deutschmann:1993:RDA

[DEW93] R. Deutschmann, W. Ertmer, and H. Wallis. Reflection and diffraction of atomic de Broglie waves by an evanescent laser wave. Physical Review A (Atomic, Molecular, and Optical Physics), 47 (??):2169, ???? 1993. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Destouches-Fevrier:1948:RIL

[DF48] Paulette Destouches-Fevrier. Relations d’incertitude liées à la complémentarité corpuscules — système de Louis de Broglie. (French) [Uncertainty relations related to particle complementarity — system of Louis de Broglie]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 226(??):468– 470, ???? 1948. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Degrange:2013:TLL

[DFF13] Bernard Degrange, Gérard Fontaine, and Patrick Fleury. Tracking Louis Leprince-Ringuet’s contributions to cosmic-ray physics.

Physics Today, 66(6):8, June 1, 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/PT.3.1989.

Durt:2017:ISI

[DFW17] T. Durt, D. Fargue, and R. Willox. Introduction [to the special issue on “quantum rogue waves as emerging quantum events”]. Annales de la Fondation Louis de Broglie, 42(1):1–8, 2017. CO-DEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL

http://aflb.ensmp.fr/AFLB-421/aflb421m1.htm.

Doak:1999:TRA

[DGR^{+}99] R. B. Doak, R. E. Grisenti, S. Rehbein, G. Schmahl, J. P. Toennies, and Ch. Wöll. Towards realization of an atomic de Broglie microscope: Helium atom focusing using Fresnel zone plates. Physical Review Letters, 83(??):4229, ???? 1999. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Dudley:1977:EDD

[DH77] H. C. Dudley and J. W. Haslett. Ether drift debate. Physics Today, 30(5):101–102, May 1, 1977. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3037565.

Dewdney:1996:BBB

[DH96] Chris Dewdney and George Horton. De Broglie, Bohm and the boson. In Bohmian Mechanics and Quantum Theory: an Appraisal [CFG96], chapter 12, pages 169–189. ISBN 90-481-4698-4, 94-015-8715-9 (e-book). ISSN 0068-0346. LCCN QC173.96-

174.52. URL https://link.springer.com/chapter/10.1007/
978-94-015-8715-0_12.

Dewdney:2002:RIE

[DH02] Chris Dewdney and George Horton. Relativistically invariant extension of the de Broglie–Bohm theory of quantum mechanics. Journal of Physics A (Mathematical and General), 35(47):10117– 10127, 2002. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

deHaas:2004:CBH

[dH04] P. de Haas. The combination of de Broglie’s harmony of the phases and Mie’s theory of gravity results in a Principle of Equivalence for quantum gravity. Annales de la Fondation Louis de Broglie, 29(4):707–726, 2004. CODEN ???? ISSN 0182-4295 (print),

2108-6397 (electronic). URL http://aflb.ensmp.fr/AFLB-294/
aflb294m171.htm.

Dickson:1996:BRA

[Dic96] Michael Dickson. Book reviews: Antidote or theory?: David Bohm and Basil J. Hiley, The Undivided Universe: An Ontological Interpretation of Quantum Theory (London: Routledge, 1993), xii + 397 pp. ISBN 0-415-06588-7. Peter R. Holland, The Quantum Theory of Motion: An Account of the de Broglie–Bohm Causal Interpretation of Quantum Mechanics (Cambridge: Cambridge University Press, 1993 hardback, 1995 paperback), xx + 598 pp. ISBN 0-521-35404-8 Hardback; 0-521-48543-6 Paperback. Studies in History and Philosophy of Modern Physics, 27(2):229– 238, June 1996. CODEN ???? ISSN 1355-2198 (print), 1879-2502

(electronic). URL http://www.sciencedirect.com/science/
article/pii/1355219896000135. See [BH93, Hol93].

Dingle:1941:RML

[Din41] Herbert Dingle. Review: Matter and Light. The New Physics by Louis de Broglie; W. H. Johnston. Philosophy, 16(62):210– 211, 1941. ISSN 0031-8191 (print), 1469-817X (electronic). URL

http://www.jstor.org/stable/3748336.

Ditchburn:1980:FO

[Dit80] R. W. Ditchburn. Frisch obituary. Physics Today, 33(5):86, May 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). Correction to [BW80], noting that Frisch did not succeed to Rutherford’s chair, but rather, was Jacksonian Professor of Natural Philosophy.

Darrigol:1993:NNC

[DK93] Olivier Darrigol and Arthur B. Komar. From c-numbers to qnumbers: The classical analogy in the history of quantum theory. Physics Today, 46(9):67–68, September 1, 1993. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2809045.

dEspagnat:1966:CPC

[dL66a] Bernard d’Espagnat and R. B. Lindsay. Conceptions de la physique contemporaine. Physics Today, 19(7):117–120, July 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3048373.

Destouches:1966:TPR

[DL66b] J.-L. L. Destouches and R. Bruce Lindsay. Théorie physique et recherche prévisionnelle. Physics Today, 19(1):111, January 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3047913.

LaVarende:1981:VPO

[dL81] Jean de La Varende. Pays d’Ouche: 1940–1933. Association des amis de La Varende, Paris, France, 1981. ???? pp. LCCN PQ2623.A895 Z88 1981. 10.00F. URL http://gallica.bnf.

fr/ark:/12148/bpt6k48094212/f11.image. Preface by Louis de

Broglie (pages 7–8).

oBroil:1975:HOI

[doB75] Lui d′ o Broĭl. On the history of the origin and interpretation of wave mechanics. Fiz.-Mat. Spis. B’lgar. Akad. Nauk., 18(51) (4):263–267, 1975. ISSN 0015-3265. Translated from the Russian (Voprosy Istor. Estestvoznan. i Tehn. 1974, no. 2–3(47–48), 3–8).

DeVelis:1968:TAH

[DRC68] John B. DeVelis, George O. Reynolds, and Robert J. Collier. Theory and applications of holography. Physics Today, 21(6):87–89, June 1, 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.3035028.

Dresden:1988:KCS

[Dre88] Max Dresden. Kramers’ contributions to statistical mechanics. Physics Today, September 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.881132.

Drell:1997:UPT

[Dre97] Sidney Drell. Us position on test ban treaty explained, key terms clarified. Physics Today, June 1, 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.881791.

Drezet:2017:HJB

[Dre17] A. Drezet. How to justify Born’s rule using the pilot wave theory of de Broglie? Annales de la Fondation Louis de Broglie, 42(1):103–132, 2017. CODEN ???? ISSN 0182-4295 (print),

2108-6397 (electronic). URL http://aflb.ensmp.fr/AFLB-421/
aflb421m5.htm.

Deutsch:1964:ODS

[DS64] Ralph Deutsch and R. L. Street. Orbital dynamics of space vehicles. Physics Today, 17(4):69–70, April 1, 1964. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3051545.

Diemer:1974:CVA

[DT74] G. Diemer and T. Tsang. Creativity versus age. Physics Today, 27(8):9, August 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3128793.

Dugas:1950:HMF

[Dug50] René Dugas. Histoire de la mécanique. (French) [History of mechanics]. Éditions du Griffon and Éditions Dunod, Neuchatel, Switzerland and Paris, France, 1950. 650 pp. Préface de Louis de Broglie.

Dugas:1988:HM

[Dug88] René Dugas. A history of mechanics. Dover Publications, Inc., New York, NY, USA, reprint of the 1957 translation edition, 1988. ISBN 0-486-65632-2. 662 pp. With a foreword by Louis de Broglie. Translated from the French by J. R. Maddox. Reprint of the 1957 translation.

Durt:2017:BDS

[Dur17] T. Durt. L. de Broglie’s double solution and self-gravitation. Annales de la Fondation Louis de Broglie, 42(1):73–102, 2017. CO-DEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL

http://aflb.ensmp.fr/AFLB-421/aflb421m4.htm.

deXivrey:1933:REP

[dX33] G. de Xivrey. Review: Exposés de Physique Théorique. Fascicules 1 et 2 by Louis de Broglie. Les études philosophiques, 7(1/2):

67, 1933. ISSN 0014-2166 (print), 2101-0056 (electronic). URL

http://www.jstor.org/stable/20840528.

Eberly:2002:QOP

[Ebe02] Joseph H. Eberly. Quantum optics in phase space. Physics Today, August 1, 2002. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/1.1510285.

Eckert:2009:BRQ

[Eck09] Michael Eckert. Book review: The Quantum Ten: A Story of Passion, Tragedy, Ambition and Science, Sheilla Jones, Oxford U. Press, New York, 2008. $24.95 (323 pp.). ISBN 978-0-19-536909-0. Physics Today, 62(5):55, May 1, 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.3141943.

Efthymiopoulos:2017:CBB

[ECT17] C. Efthymiopoulos, G. Contopoulos, and A. C. Tzemos. Chaos in de Broglie–Bohm quantum mechanics and the dynamics of quantum relaxation. Annales de la Fondation Louis de Broglie, 42(1):133–160, 2017. CODEN ???? ISSN 0182-4295 (print),

2108-6397 (electronic). URL http://aflb.ensmp.fr/AFLB-421/
aflb421m6.htm.

Estermann:1931:VMB

[EFS31] I. Estermann, R. Frisch, and O. Stern. Versuche mit monochromatischen de Broglie-Wellen der Molekularstrahlen. (German) [Experiments with monochromatic de Broglie waves of molecular beams]. Zeitschrift für Physik, 72(??):370–674, 1931. CODEN ZEPYAA. ISSN 0044-3328.

Estermann:1932:MBW

[EFS32] I. Estermann, R. Frisch, and O. Stern. Monochromasierung der de Broglie-Wellen von Molekularstrahlen. (German) [Monochromatization of de Broglie waves of molecular beams]. Zeitschrift für Physik, 73(5–6):348–365, May 1932. CODEN ZEPYAA. ISSN 0044-3328.

Eastlund:1972:DFT

[EGSR72] Bernard J. Eastlund, William C. Gough, Zeinab A. Sabri, and David J. Rose. In defense of the fusion torch. Physics Today, 25 (10):9–13, October 1, 1972. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3071032.

Endo:1951:NCS

[EK51] S. Endo and H. Kanazawa. Note on the “champ soustractif” of Louis de Broglie. Progress of Theoretical Physics, 6:436–437, 1951. CODEN PTPKAV. ISSN 0033-068X (print), 1347-4081 (electronic).

Elbaz:2013:PAE

[Elb13] Claude Elbaz. Sur les programmes d’Albert Einstein et de Louis de Broglie. Une contribution. (French) [On the programs of Albert Einstein and Louis de Broglie. A contribution]. Annales de la Fondation Louis de Broglie, 38(1):195–217, 2013. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://

aflb.ensmp.fr/AFLB-381/aflb381m758.htm.

Ellis:1968:Pb

[ELF68] R. Hobart Ellis, Jr., Felix Lighthead, and Theodore H. Freilich. Phimsy. Physics Today, 21(12):17–19, December 1, 1968. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3034643.

Ellis:1968:Pa

[EM68] R. Hobart Ellis, Jr. and Alan Mackay. Phimsy. Physics Today, 21(9):21–23, September 1, 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3035149.

El-Nadi:1952:TDP

[EN52] M. El-Nadi. Sur la théorie du photon de L. de Broglie. (French) [On the photon theory of L. of Broglie]. Journal de Physique et le Radium, 13:540–542, 1952. CODEN JPRAAJ. ISSN 0368-3842.

Ennis:1987:OLB

[Enn87] Thomas W. Ennis. [Obituary]: Louis de Broglie, French physicist; won ’29 Nobel for wave theory. New York Times, ??(??):B6, March 20, 1987. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/110809191.

Enz:1985:SGB

[Enz85] U. Enz. The sine-Gordon breather as a moving oscillator in the sense of de Broglie. Physica. D, Nonlinear phenomena, 17:116– 119, 1985. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic).

Edge:2000:OFM

[EP00] Ronald D. Edge and Charles P. Poole. Oswald Francis “Mike” Schuette. Physics Today, 53(12):85–86, December 1, 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.1342836.

Eder:2017:ZOF

[ERS^{+}17] S. D. Eder, A. K. Ravn, B. Samelin, G. Bracco, A. Salvador Palau, T. Reisinger, E. B. Knudsen, K. Lefmann, and B. Holst. Zeroorder filter for diffractive focusing of de Broglie matter waves. Physical Review A (Atomic, Molecular, and Optical Physics), 95 (??):023618, ???? 2017. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Edamatsu:2002:MPB

[ESI02] Keiichi Edamatsu, Ryosuke Shimizu, and Tadashi Itoh. Measurement of the photonic de Broglie wavelength of entangled photon pairs generated by spontaneous parametric down-conversion. Physical Review Letters, 89(??):213601, ???? 2002. CODEN PRL-TAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Esposito:1999:PWM

[Esp99] S. Esposito. Photon wave mechanics: a de Broglie–Bohm approach. Foundations of Physics Letters, 12(6):533–545, December 1999. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Evans:2007:QMC

[ET07] James Evans and A. S. (Alan S.) Thorndike, editors. Quantum mechanics at the crossroads: new perspectives from history, philosophy and physics. The frontiers collection. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 3-540-32663-4 (hardcover), 3-540-32665-0 (e-book). ISSN 1612-3018. x + 249 pp.

LCCN QC174.12 .Q346 2007. URL http://www.loc.gov/

catdir/enhancements/fy0824/2006934045-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0824/2006934045-d.
html;   http://www.loc.gov/catdir/toc/fy0708/2006934045.
html.

Ehrenfest:1926:GVB

[EU26] P. Ehrenfest and G. E. Uhlenbeck. Graphische Veranschaulichung der De Broglieschen Phasenwellen in der fünfdimensionalen Welt von O. Klein. (German) [Graphical illustration of de Broglie’s phase waves in the five-dimensional world by O. Klein]. Zeitschrift für Physik, 39(7–8):495–498, July 1926. CODEN ZEPYAA. ISSN 0044-3328.

Erkof:1998:LNA

[EUG98] Sakir Erkof, Turgay Uzer, and J. D. Garcia. Lecture notes on atomic and molecular physics. Physics Today, April 1, 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.882217.

Evans:1995:EVF

[Eva95] M. W. Evans. The Evans–Vigier field, B^{(3)}: Derivation of the de Broglie matter–wave equation from the Hamilton–Jacobi equation. Foundations of Physics Letters, 8(5):481–489, October 1995. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Flato:1976:QMD

[F^{+}76] M. (Moshé) Flato et al., editors. Quantum mechanics, determinism, causality, and particles: an international collection of contributions in honor of Louis de Broglie on the occasion of the jubilee of his celebrated thesis, volume 1 of Mathematical physics and applied mathematics. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1976. ISBN 90-277-0623-9. x + 252 pp. LCCN QC174.125 .Q36.

Fabbri:2022:BBF

[Fab22] Luca Fabbri. de Broglie–Bohm formulation of Dirac fields. Foundations of Physics, 52(6):??, December 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://

link.springer.com/article/10.1007/s10701-022-00641-2.

Fara:2001:GPI

[Far01] Patricia Fara. Group portraits IV — the Seventh Solvay Conference [Bruxelles, 22–29 October 1933]. Endeavour, 25(4):137– 138, December 2001. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Fargue:2017:LBD

[Far17] D. Fargue. Louis de Broglie’s “double solution”, a promising but unfinished theory. Annales de la Fondation Louis de Broglie, 42(1):9–18, 2017. CODEN ???? ISSN 0182-4295 (print),

2108-6397 (electronic). URL http://aflb.ensmp.fr/AFLB-421/
aflb421m2.htm.

Field:1985:IUP

[FCP85] George B. Field, Eric J. Chaisson, and Jay M. Pasachoff. The invisible universe: Probing the frontiers of astrophysics. Physics Today, October 1, 1985. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2814738.

Feder:2008:PIH

[Fed08] Toni Feder. Path integrals, Les Houches, and other adventures of Cécile DeWitt-Morette. Physics Today, 61(8):28–30, August 1, 2008. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/1.2970207.

Feder:2009:DLG

[Fed09] Toni Feder. DOD launches ‘genius’ awards. Physics Today, 62(2): 29–30, February 1, 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/full/10.1063/1.3086095.

Feoli:2009:ABG

[Feo09] Antonio Feoli. The amplitude of the de Broglie gravitational waves. Modern Physics Letters A (MPLA), 24(31):2497–2505, 2009. CODEN MPLAEQ. ISSN 0217-7323 (print), 1793-6632 (electronic).

Ferber:1996:MLW

[Fer96] R. Ferber. A missing link: What is behind de Broglie’s “periodic phenomenon”? Foundations of Physics Letters, 9(6):575–586, December 1996. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Feshbach:1956:RPM

[Fes56] Herman Feshbach. Review: Physics and Microphysics by Louis de Broglie; Martin Davidson; A. Einstein. Science, 123(3196): 550, 1956. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1750554.

Falsaperla:2003:MSP

[FF03] P. Falsaperla and G. Fonte. On the motion of a single particle near a nodal line in the de Broglie–Bohm interpretation of quantum mechanics. Physics Letters A, 316(6):382–390, 2003. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Ford:1964:WEP

[FH64] Kenneth W. Ford and W. G. Holladay. The world of elementary particles. Physics Today, 17(4):68–69, April 1, 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3051544.

Field:2006:QMS

[Fie06] J. H. Field. Quantum mechanics in space-time: the Feynman path amplitude description of physical optics, de Broglie matter waves and quark and neutrino flavour oscillations. Annals of Physics, 321(3):627–707, 2006. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Filippi:1947:RPM

[Fil47] U. Filippi. Review: Physique et Microphysique by Louis de Broglie. Les Études philosophiques, 2(3–4):314, July/December 1947. ISSN 0014-2166 (print), 2101-0056 (electronic). URL http:

//www.jstor.org/stable/20841087.

Filippov:2002:PNI

[Fil02] R. V. Filippov. On possible non-identity of microparticles and de Broglie waves in quantum theory. Mat. Strukt. Model., 10: 130–142, 2002. ISSN 2222-8772 (print), 2222-8799 (electronic).

Francis:2009:NLV

[FL09] G. Francis and Ch. Lorin. Notes de lecture: Voyage au centre de la science au xxe siècle — sur les traces de Louis de Broglie.

(French) [Reading notes: Journey to the center of science in the 20th century — in the footsteps of Louis de Broglie]. Annales de la Fondation Louis de Broglie, 34(1):111–??, 2009. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http:

//aflb.ensmp.fr/AFLB-341/aflb341m901.htm.

Fonseca:1999:MBW

[FMP99] E. J. S. Fonseca, C. H. Monken, and S. Pádua. Measurement of the de Broglie wavelength of a multiphoton wave packet. Physical Review Letters, 82(??):2868, ???? 1999. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Forman:1980:BRE

[For80] Paul Forman. Book reviews: Einstein: A Centenary Volume, edited by A. P. French; Albert Einstein’s Theory of General Relativity, edited by Gerald E. Tauber; Einstein, by Louis de Broglie and others; Einstein. Sein Leben und seine Zeit, by Philipp Frank; Documentary publications: Albert Einstein Autobiographical Notes, translated and edited by Paul Arthur Schilpp; Elie Cartan, Albert Einstein, Letters on Absolute Parallelism 1929– 1932, edited by Robert Debever; Aus meinen späten Jahren, by Albert Einstein; Albert Einstein in Berlin 1918–1933, Teil I: Darstellung und Dokumente; Teil II: Spezialinventar; Images of Einstein: a Catalog, compiled by Joan N. Warnow; Albert Einstein: the Human Side. New Glimpses from his Archives, edited by Helen Dukas and Banesh Hoffman; Einstein privat. Herta W. erinnert sich an die Jahre 1927 bis 1933, by Friedrich Herneck; Centenary exhibitions: Albert Einstein, 1879–1955. A Centenary Exhibit of Manuscripts, Books, and Portraits Selected from the Humanities Research Center Collections, by Albert C. Lewis; Gedächtnisausstellung zum 100. Geburtstag von Albert Einstein, Otto Hahn, Max von Laue, Lise Meitner; Einstein = mc^{2}; Einstein 1879–1979. Exhibition. Bulletin of the Atomic Scientists, 36(3):49–55, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fowler:1929:RCP

[Fow29] R. H. Fowler. Review: Collected Papers on Wave Mechanics by E. Schrödinger; Selected Papers on Wave Mechanics by L. de Broglie and L. Brillouin; Four Lectures on Wave Mechanics by E. Schrödinger. Mathematical Gazette, 14(199):369–370, March 1929. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL http://www.jstor.org/stable/3606729.

Fonseca:2001:NBW

[FPN^{+}01] E. J. S. Fonseca, Zoltan Paulini, P. Nussenzveig, C. H. Monken, and S. Pádua. Nonlocal de Broglie wavelength of a two-particle system. Physical Review A (Atomic, Molecular, and Optical Physics), 63(??):043819, ???? 2001. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Fernandez:2013:LB

[FR13a] Bernard Fernandez and Georges Ripka. Louis de Broglie. In Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956 [FR13b], pages 131–132. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). LCCN QC773 .F47 2013.

Fernandez:2013:UMA

[FR13b] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (e-book). xviii + 522 pp. LCCN QC773 .F47 2013.

Frampton:2005:WPU

[Fra05] Paul H. Frampton. Warped passages: Unraveling the mysteries of the universe’s hidden dimensions. Physics Today, December 1, 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/1.2169446.

French:2007:PP

[Fre07] Steven French. On physics and philosophy. Physics Today, June 1, 2007. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/1.2754610.

Frieman:1962:BPT

[Fri62] E. A. Frieman. Basic plasma theory — kinetics. Physics Today, 15(12):28–30, December 1, 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3057903.

Feshbach:1977:RAN

[FS77] Herman Feshbach and Eric Sheldon. Recent advances in neutron physics. Physics Today, 30(2):40–51, February 1, 1977.

CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3037411.

Feoli:1997:GIB

[FS97] Antonio Feoli and Gaetano Scarpetta. A geometric interpretation of de Broglie waves. Rend. Accad. Sci. Fis. Mat., Napoli (4), 64: 397–404, 1997. ISSN 0370-3568.

Feoli:1998:BMW

[FS98] A. Feoli and G. Scarpetta. De Broglie matter waves from the linearized Einstein field equations. Foundations of Physics Letters, 11(4):395–403, August 1998. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Gleiser:1998:DUC

[GA98] Marcelo Gleiser and Don L. Anderson. The dancing universe: From creation myths to the big bang. Physics Today, 51(4):72– 73, April 1, 1998. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.882216.

Gagnebin:1953:NBH

[Gag53] S. Gagnebin. Notes bibliographiques: Histoire de la mécanique par René Dugas. Dialectica: International Review of Philosophy of Knowledge, 7(1):52–54, March 1953. CODEN ???? ISSN 0012-2017 (print), 1746-8361 (electronic). URL http://www.jstor.

org/stable/42968355.

Goryunova:1966:CDL

[GAL66] N. A. Goryunova, J. C. Anderson, and Lars C. Luther. The chemistry of diamond-like semiconductors. Physics Today, July 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3048374.

Gamow:1926:TOP

[Gam26] G. Gamow. Sur la théorie des ondes de phases de M. L. de Broglie. (French) [On the phase wave theory of M. L. de Broglie]. C. B., 183:875–876, 1926.

Gamow:1954:RRP

[Gam54] George Gamow. Review: The Revolution in Physics by Louis de Broglie; Ralph W. Niemeyer (translator). The Scientific Monthly, 78(1):48, January 1954. CODEN SCMOAA. ISSN 0096-3771 (print), 2327-7513 (electronic). URL http://www.jstor.org/

stable/21151.

Gamow:1985:TYS

[Gam85] George Gamow. Thirty years that shook physics: the story of quantum theory. Dover Publications, Inc., New York, NY, USA, 1985. ISBN 0-486-24895-X (paperback). xiv + 224 + 9 pp. LCCN QC174.12 .G35 1985. US$4.95. URL http://www.loc.

gov/catdir/description/dover032/85006797.html.

Garcia:1953:RGB

[Gar53] Godofredo García. Die relativistischen Gleichungen von de Broglie und Schrödinger. Die Stabilität der Bewegungen. (German) [The relativistic equations of de Broglie and Schrödinger. The stability of the movements]. Actas Acad. Nac. Ci. Exact., Fis. Natur. Lima, 16(3–4):36–55, 1953.

Garland:1966:BRV

[Gar66] Carl W. Garland. Book review: A. Van Itterbeek, editor, Physics of High Pressures and the Condensed Phase, 598 pp. North-Holland, Amsterdam and John Wiley, New York, 1965. Physics Today, 19(1):126–128, January 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3047926.

Gates:1968:RCB

[Gat68] David M. Gates. Review: Chemical Biology: Wave Mechanics and Molecular Biology by Louis de Broglie. BioScience, 18(1):55– 56, 1968. CODEN BISNAS. ISSN 0006-3568 (print), 1525-3244

(electronic). URL http://www.jstor.org/stable/1293965.

Gauthier:2016:ECP

[Gau16] Richard Gauthier. The electron is a charged photon with the de Broglie wavelength. In Unified field mechanics: natural science beyond the veil of spacetime. Proceedings of the IX symposium honoring noted French mathematical physicist Jean-Pierre Vigier, Morgan State University, Baltimore, MD, USA, November 16–19, 2014, pages 122–130. World Scientific Publishing Co. Pte. Ltd., P.

O. Box 128, Farrer Road, Singapore 9128, 2016. ISBN 981-4719-05-6 (hardcover); 981-4719-07-2 (e-book).

Goldsmith:1988:PTW

[GD88] Werner Goldsmith and Arthur C. Damask. Physics takes the witness stand. Physics Today, 41(1):127–128, January 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2811304.

Geheniau:1933:LFOa

[Géh33a] Jules Géhéniau. Les lois fondamentales de l’onde de L. de Broglie dans la gravifique de Th. de Donder. (French) [The fundamental laws of the wave of L. De Broglie in the gravity of Th. Of Donder]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 197(??):1393–1395, ???? 1933. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Geheniau:1933:OBC

[Géh33b] Jules Géhéniau. Sur les ondes de L. de Broglie dans un champ gravifique et électromagnétique quelconque. (French) [On the air of L. De Broglie in any gravitational and electromagnetic field]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 197(??):1305–1307, ???? 1933. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Geheniau:1937:MIT

[Geh37] Jules Geheniau. Les moments d’impulsion dans la théorie du photon de M. L. de Broglie. (French) [The moments of impulse in the photon theory of M. L. de Broglie]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 204(??):665– 668, ???? 1937. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Geheniau:1938:MOE

[Géh38] J. Géhéniau. Mécanique ondulatoire de l’électron et du photon. (French) [Wave mechanics of electron and photon]. Gauthier-Villars, Paris, France, 1938. viii + 142 pp. Préface de M. L. de Broglie.

Gehrenbeck:1978:EDF

[Geh78] Richard K. Gehrenbeck. Electron diffraction: fifty years ago. Physics Today, 31(1):34–41, January 1, 1978. CODEN PHTOAD.

ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3001830.

George:1932:MQCa

[Geo32] André George. Mécanique quantique et causalité d’après M. Fermi. (French) [Quantum mechanics and causality according to M. Fermi], volume 38 of Actualités scientifiques et industrielles. Hermann, Paris, France, 1932. 18 pp. Exposés de physique théoriques, fasc. 5. With remarks by L. de Broglie.

George:1953:LBP

[Geo53] André George. Louis de Broglie, physicien et penseur (French) [Louis de Broglie, physicist and thinker]. Éditions Albin Michel, Paris, France, 1953. xi + 497 pp. LCCN QC71 .L88.

Germer:1964:LEE

[Ger64] Lester H. Germer. Low-energy electron diffraction. Physics Today, 17(7):19–23, July 1, 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3051686.

Germain:1987:LBO

[Ger87] Paul Germain. Louis de Broglie ou la passion de la^{≪}vraie^{≫}physique. (French) [Louis de Broglie or the passion of “true” physics]. Comptes Rendus de l’Académie des Sciences, Série Générale: la Vie des Sciences, 4(6):569–593, 1987. CODEN CRAGET. ISSN 0762-0969 (print), 2419-8730 (electronic). URL http://

www.ordiecole.com/broglie_eloge.pdf.

Grob:1967:PL

[GF67] Victor E. Grob and Ephraim Fischbach. Prolongation of life. Physics Today, 20(1):10–12, January 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3034102.

Gondran:2005:TBP

[GH05] M. Gondran and R. Hoblos. Trajectoires de de Broglie pour la particule dans la boîte. (French) [de Broglie’s trajectories for the particle in the box]. Annales de la Fondation Louis de Broglie, 30(3–4):273–282, 2005. CODEN ???? ISSN 0182-4295 (print),

2108-6397 (electronic). URL http://aflb.ensmp.fr/AFLB-303/
aflb303m306.htm.

Ghafarnejad:2006:BDS

[Gha06] Hosein Ghafarnejad. Brans–Dicke scalar tensor theory and de Broglie–Bohm particle interpretation. Astrophysics and Space Science, 301(1–4):145–148, 2006. CODEN APSSBE. ISSN 0004-640X (print), 1572-946X (electronic).

Ghaffarnejad:2014:SSJ

[Gha14] Hossein Ghaffarnejad. Spherically symmetric Jordan–Brans– Dicke quantum gravity with de Broglie–Bohm pilot wave perspective. International Journal of Theoretical Physics, 53(8):2616– 2622, 2014. CODEN IJTPBM. ISSN 0020-7748 (print), 1572-9575 (electronic).

Greenberger:1993:MIS

[GHZ93] Daniel M. Greenberger, Michael A. Horne, and Anton Zeilinger. Multiparticle interferometry and the superposition principle. Physics Today, 46(8):22–29, August 1, 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.881360.

Giddings:2013:BHQ

[Gid13] Steven B. Giddings. Black holes, quantum information, and the foundations of physics. Physics Today, April 1, 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/PT.3.1946.

Gillet:1966:RSR

[Gil66] Marcel Gillet. Review: Science et réalité by André Metz; Louis de Broglie. Archives de Philosophie, 29(4):634–635, 1966. ISSN 0003-9632 (print), 1769-681X (electronic). URL http://www.jstor.

org/stable/43032864.

Graham:1996:DLA

[GM96] R. Graham and S. Miyazaki. Dynamical localization of atomic de Broglie waves: The influence of spontaneous emission. Physical Review A (Atomic, Molecular, and Optical Physics), 53(??):2683, ???? 1996. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Geim:2007:GEC

[GM07] Andrey K. Geim and Allan H. MacDonald. Graphene: Exploring carbon flatland. Physics Today, August 1, 2007. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/1.2774096.

Gompert:1979:NWW

[GMG^{+}79] D. C. Gompert, M. Mandelbaum, R. L. Garwin, J. H. Barton, and Sidney D. Drell. Nuclear weapons and world politics: Alternatives for the future. Physics Today, 32(4):55–57, April 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2995494.

Goldstein:1994:DJB

[Gol94a] Sheldon Goldstein. David Joseph Bohm. Physics Today, 47(8):72– 73, August 1, 1994. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2808618.

Goldstein:1994:RQM

[Gol94b] Sheldon Goldstein. Review: Quantum mysteries: The Quantum Theory of Motion. An Account of the de Broglie-Bohm Causal Interpretation of Quantum Mechanics, by Peter R. Holland. Science, 263(5144):254–255, January 14, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/2882392.

Goldstein:1998:QTO

[Gol98] Sheldon Goldstein. Quantum theory without observers — part two. Physics Today, 51(4):38–42, April 1, 1998. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.882241.

Gomes:1935:QCEa

[Gom35a] R. L. Gomes. Quelques considérations sur l’équation fondamentale de la “nouvelle conception de la lumière” du prof. Louis de Broglie. I. (French) [Some considerations on the fundamental equation of the “new conception of light” of Professor Louis de Broglie. I]. Atti Accad. Naz. Lincei, Rend., VI. Ser., 21:358–364, 1935. ISSN 0001-4435.

Gomes:1935:QCEb

[Gom35b] R. L. Gomes. Quelques considérations sur l’équation fondamentale de la “nouvelle conception de la lumière” du prof. Louis de Broglie.

II. (French) [Some considerations on the fundamental equation of the “new conception of light” of Professor Louis de Broglie. II]. Atti Accad. Naz. Lincei, Rend., VI. Ser., 21:443–447, 1935. ISSN 0001-4435.

Gomes:1935:POM

[Gom35c] R. L. Gomes. Sur une propriété de l’opérateur H de M. de Broglie. (French) [On a property of the H operator of M. de Broglie]. Atti Accad. Naz. Lincei, Rend., VI. Ser., 21:499–501, 1935. ISSN 0001-4435.

Goudsmit:1974:PYP

[Gou74] S. A. Goudsmit. Pay your page charges. Physics Today, 27(8):9– 11, August 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3128794.

Golay:1972:CCA

[GP72] Marcel J. E. Golay and Jean Piaget. Can children aid physicists? Physics Today, 25(10):13, October 1, 1972. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3071033.

Griffiths:2011:TSQ

[Gri11] Robert B. Griffiths. Time-symmetric quantum mechanics questioned and defended. Physics Today, 64(5):62, May 1, 2011. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/1.3595149.

Gribbin:2012:ESQ

[Gri12] John R. Gribbin. Erwin Schrödinger and the quantum revolution. Bantam Press, London, UK, 2012. ISBN 0-593-06776-2 (hardcover), 0-593-06865-3 (paperback), 1-4464-6571-3 (e-book). xii + 322 + 16 pp. LCCN QC16.S265 G75 2012.

Gribbin:2013:ESQ

[Gri13] John R. Gribbin. Erwin Schrödinger and the quantum revolution. Wiley, New York, NY, USA, 2013. ISBN 1-118-29926-4 (hardback), 1-118-33411-6 (pdf), 1-118-33188-5 (epub), 1-118-33519-8 (mobi). xii + 321 pp. LCCN QC16.S265 G75 2012.

URL http://catalogimages.wiley.com/images/db/jimages/
9781118299265.jpg.

Grunbaum:1954:RHM

[Gru54] Adolf Grunbaum. Review: Histoire de la Mecanique. by René Dugas; Preface by Louis de Broglie. Neuchatel, Éditions du Griffon, 1950, pp. 649. Philosophy and Phenomenological Research, 15 (1):119–121, September 1954. ISSN 0031-8205 (print), 1933-1592

(electronic). URL http://www.jstor.org/stable/2103748.

Gruber:1974:CP

[Gru74] Gary R. Gruber. Competition in physics. Physics Today, 27 (8):74–75, August 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3128845.

Glaberson:1987:QVS

[GS87] William I. Glaberson and Klaus W. Schwarz. Quantized vortices in superfluid helium-4. Physics Today, 40(2):54–60, February 1, 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.881099.

Gruji:2012:ICA

[GS12] Petar Gruji and Nenad Simonovič. Insights from the classical atom. Physics Today, May 1, 2012. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/PT.3.1559.

Gudder:1985:BRW

[Gud85] Stanley P. Gudder. Book review: The Wave-Particle Dualism: A Tribute to Louis de Broglie on His 90th Birthday by S. Diner; D. Fargue; G. Lochak; F. Selleri. Philosophy of Science, 52(1):169– 170, March 1985. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/
187613.

Guinier:1951:EPM

[Gui51] Georges Guinier. Eléments de physique moderne théorique. III: Statistiques quantiques. (French) [Elements of modern theoretical physics. III: Quantum statistics]. Bibliothèque de la Science moderne. Bordas, Paris, France, 1951. 174 pp. Préface de L. de Broglie.

Guth:1967:SSE

[Gut67] Eugene Guth. Schladming School examines special high-energy problems. Physics Today, 20(7):109–111, July 1, 1967. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3034386.

Gueret:1982:BWP

[GV82] Ph. Gueret and J.-P. Vigier. De Broglie’s wave particle duality in the stochastic interpretation of quantum mechanics: A testable physical assumption. Foundations of Physics, 12(11):1057–1083, November 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF01300546.

Haas:1934:MQE

[Haa34] Arthur Haas. Materiewellen und Quantenmechanik. Eine Einführung auf Grund der Theorien von de Broglie, Schrödinger, Heisenberg und Dirac. 4. u. 5. verb. u. abermals wes. verm. Aufl. (German) [Matter waves and quantum mechanics. An introduction based on the theories of de Broglie, Schrödinger, Heisenberg and Dirac]. Akad. Verlagsges., Leipzig, Germany, 1934. viii + 299 + 7 pp.

Hamdan:2007:DBT

[Ham07] N. Hamdan. The dynamical de Broglie theory. Annales de la Fondation Louis de Broglie, 32(1):11–24, 2007. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://aflb.

ensmp.fr/AFLB-321/aflb321m402.htm.

Hanle:1977:ESR

[Han77] Paul A. Hanle. Erwin Schrödinger’s reaction to Louis de Broglie’s thesis on the quantum theory. Isis, 68(4):606–609, December 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230015.

Honl:1950:BTE

[HB50] Helmut Hönl and Hermann Boerner. Zur de Broglie’schen Theorie der Elementarteilchen. (German) [On the de Broglie theory of elementary particles]. Z. Naturforsch., A, 5:353–366, 1950. ISSN 0340-4811.

Haroche:2013:ACC

[HBR13] Serge Haroche, Michel Brune, and Jean-Michel Raimond. Atomic clocks for controlling light fields. Physics Today, 66(1):27–32, January 1, 2013. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/PT.3.1856.

Home:1998:CFQ

[HC98] Dipankar Home and James T. Cushing. Conceptual foundations of quantum physical an overview from modern perspectives. Physics Today, 51(10):79, October 1, 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.882411.

Horton:2002:BPW

[HDN02] George Horton, Chris Dewdney, and Ulrike Ne’eman. de Broglie’s pilot-wave theory for the Klein–Gordon equation and its space– time pathologies. Foundations of Physics, 32(3):463–476, March 2002. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1023/A%
3A1014865430144.

Heisenberg:1939:VGF

[Hei39] Werner Heisenberg. Vorwort. (German) [Foreword]. In Louis de Broglie, editor, Licht und Materie. (German) [Light and matter], pages 7–8. H. Groverts-Verlag, Hamburg, Germany, 1939. LCCN ????

Heisenberg:1953:RTN

[Hei53] Werner Heisenberg. Remarques sur la théorie neutrinienne de la lumière. (French) [Remarks on the neutrino theory of light]. In ????, editor, Louis de Broglie, Physicien et Penseur. (French) [Louis de Broglie, Physicist and Thinker], pages 283–286. Albin Michel, Paris, France, 1953. LCCN ????

Heisenberg:1957:PPT

[Hei57] W. Heisenberg. Les principes physiques de la théorie des quanta. (French) [Physical principles of quantum theory]. Gauthier-Villars, Paris, France, 1957. x + 124 + 2 pp. Préface de Louis de Broglie.

Heilbron:1974:HGJ

[Hei74] J. L. Heilbron. H. G. J. Moseley: The life and letters of an English physicist, 1887–1915. University of California Press, Berkeley, CA, USA, 1974. ISBN 0-520-02375-7. xiii + 312 pp. LCCN QC16.M68 H44.

Hestenes:1990:ZIQ

[Hes90] David Hestenes. The Zitterbewegung interpretation of quantum mechanics. Foundations of Physics, 20(10):1213–1232, October 1990. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01889466.

Hoyle:1965:MG

[HG65] Fred Hoyle and J. Gillis. Of men and galaxies. Physics Today, 18 (6):51, June 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3047483.

Hillers:1928:ENB

[Hil28] W. Hillers. Der experimentelle Nachweis der de Broglie-Wellen durch C. Davisson. (German) [Experimental proof of de Broglie waves by C. Davisson]. Z. f. math. Unterricht, 59(??):250–263, 1928.

Homberger:1970:CMN

[HJS70] Eric Homberger, William Janeway, and Simon Schama, editors. The Cambridge mind: ninety years of the Cambridge Review, 1879–1969. Jonathan Cape, London, UK, 1970. ISBN 0-224-61914-4. 315 + 8 pp. LCCN AC5 .H64.

Heitmann:1993:SQD

[HK93] Detlef Heitmann and Jörg P. Kotthaus. The spectroscopy of quantum dot arrays. Physics Today, 46(6):56–63, June 1, 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881355.

Haug:1994:QTO

[HKK94] Hartmut Haug, Stephen W. Koch, and Leonid V. Keldysh. Quantum theory of the optical and electronic properties of semiconductors. Physics Today, 47(2):106–108, February 1, 1994.

CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2808410.

Hurst:1994:RIS

[HL94] G. Samuel Hurst and Vladilen S. Letokhov. Resonance ionization spectroscopy. Physics Today, 47(10):38–45, October 1, 1994. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881420.

Holden:1966:NS

[HM66] Alan Holden and L. Muldawer. The nature of solids. Physics Today, August 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3048414.

Hudson:1979:BU

[HM79] Orlo K. Hudson and Isa Masrieh. Birzeit university. Physics Today, 32(8):13, August 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2995659.

Hoffmann:1984:RR

[HM84] B. Hoffmann and Robert March. Relativity and its roots. Physics Today, 37(5):86–89, May 1, 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2916251.

Hansen:1988:TSL

[HMH88] J.-P. P. Hansen, I. R. Mcdonald, and Douglas Henderson. Theory of simple liquids. Physics Today, 41(10):89–90, October 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2811600.

Hawkins:1966:PAM

[HÖ66] Gerald S. Hawkins and E. J. Öpik. The physics and astronomy of meteors, comets and meteorites. Physics Today, April 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3048195.

Hoffmann:1979:EC

[Hof79] Banesh Hoffmann. Einstein the catalyst. Physics Today, 32(3):36– 40, March 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2995457.

Hofstadter:1984:FB

[Hof84] Robert Hofstadter. Felix Bloch. Physics Today, 37(3):115–116, March 1, 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.2916128.

Hofmann:2017:IUM

[Hof17] Ralf Hofmann. The isolated, uniformly moving electron: Selfintersecting SU(2) Yang–Mills center vortex loop and Louis de Broglie’s hidden thermodynamics. arxiv.org, ??(??):1–5, June 27,

2017. URL https://arxiv.org/abs/1707.02858.

Holmes:1929:PRB

[Hol29] F. T. Holmes. Penetrating radiation and de Broglie waves. Nature, 123:943, 1929. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Holland:1987:GDB

[Hol87] P. R. Holland. Geometry of dislocated de Broglie waves. Foundations of Physics, 17(4):345–363, April 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF00733373.

Holland:1992:DEB

[Hol92] P. R. Holland. The Dirac equation in the de Broglie–Bohm theory of motion. Foundations of Physics, 22(10):1287–1301, October 1992. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF01889714.

Holland:1993:QTM

[Hol93] Peter R. Holland. The quantum theory of motion: an account of the de Broglie–Bohm causal interpretation of quantum mechanics. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-35404-8. xix + 598 pp. LCCN QC133 .H65 1993. URL http://

www.loc.gov/catdir/description/cam041/92020205.html;
http://www.loc.gov/catdir/toc/cam041/92020205.html.

Holland:1995:QTM

[Hol95] Peter R. Holland. The quantum theory of motion. An account of the de Broglie–Bohm causal interpretation of quantum mechanics. Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-48543-6 (paperback); 0-521-35404-8 (hardcover). xx + 618 pp.

Houston:2016:BWS

[Hou16] Louis M. Houston. A de Broglie wave solution to the Maxwell– Schrödinger–Einstein equations. Theor. Math. Appl., 6(3):1–13, 2016. ISSN 1792-9687 (print), 1792-9709 (electronic).

Himpsel:1985:PS

[HS85] Franz J. Himpsel and Neville V. Smith. Photoelectron spectroscopy. Physics Today, December 1, 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.880997.

Harvey:2005:SP

[HS05] Alex Harvey and Engelbert Schucking. A small puzzle from 1905. Physics Today, 58(3):34–36, March 1, 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.1897562.

Heller:1993:PQM

[HT93] Eric J. Heller and Steven Tomsovic. Postmodern quantum mechanics. Physics Today, 46(7):38–46, July 1, 1993. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881358.

Hromada:2014:BWF

[HTH^{+}14] Ivan Hromada, Raisa Trubko, William F. Holmgren, Maxwell D. Gregoire, and Alexander D. Cronin. de Broglie wave-front curvature induced by electric-field gradients and its effect on precision measurements with an atom interferometer. Physical Review A (Atomic, Molecular, and Optical Physics), 89(??):033612, ???? 2014. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Humbert:1936:PP

[Hum36] Pierre Humbert. Potentiels et prepotentiels. (French) [Potential and prepotential], volume 15 of Cahiers scientifiques. Gauthier-Villars, Paris, France, 1936. viii + 80 pp. With a preface by Louis de Broglie.

Humbert:1951:AON

[Hum51] Pierre Humbert. Analyses d’ouvrages: Nobel, Savants et découvertes, (Collection ^{≪}Les Savants et le Monde ^{≫}) par Louis de Broglie; André Georges. Revue d’Histoire des Sciences et de Leurs Applications, 4(3–4):369–370, July/December 1951. CO-DEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic).

URL http://www.jstor.org/stable/23905033.

Hund:1966:PQT

[Hun66] Friedrich Hund. Paths to quantum theory historically viewed. Physics Today, 19(8):23–29, August 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3048403.

Hutten:1954:RIP

[Hut54a] E. H. Hutten. Review: Indeterminism: La Physique quantique restera-t-elle indéterministe? Exposé du problème, suivi de la reproduction de certains documents et d’une contribution de M Jean-Pierre Vigier by Louis de Broglie. British Journal for the Philosophy of Science, 5(18):159–164, 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://

www.jstor.org/stable/685176.

Hutten:1954:RMA

[Hut54b] E. H. Hutten. Review: Methodological Aspects of Physics and Geometry: Méthodologie: Notions géometriques by Jean-Louis Destouches. British Journal for the Philosophy of Science, 5(19): 256–261, 1954. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://www.jstor.org/stable/685721.

Holland:1991:CET

[HV91] P. R. Holland and J. P. Vigier. Comment on “Experimental test of the de Broglie guided-wave theory for photons”. Physical Review Letters, 67(??):402, ???? 1991. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. See [WZM91].

Hurley:2013:NGF

[HV13] D. J. Hurley and M. A. Vandyck. A new geometrical framework for the de Broglie–Bohm quantum theory. Int. J. Geom. Methods Mod. Phys., 10(3):25, 2013. ISSN 0219-8878 (print), 1793-6977 (electronic).

Hoffmann:2004:GPS

[HW04] Dieter Hoffmann and Mark Walker. The German Physical Society under National Socialism. Physics Today, December 1, 2004. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/1.1878335.

Iona:1993:DAE

[IB93] Mario Iona and Philip Best. Different angles on errors in textbooks. Physics Today, 46(4):90–91, April 1, 1993. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2808889.

Imboden:2014:TN

[IB14] Matthias Imboden and David Bishop. Top-down nanomanufacturing. Physics Today, 67(12):45–50, December 1, 2014. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/PT.3.2621.

Izaz:1988:GUU

[II88] Mujaddid A. Izaz and M. Mansoor Ijaz. God as the ultimate unifying force. Physics Today, 41(1):129, January 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2811307.

Infeld:1931:ICE

[IS31a] Leopold Infeld and S. Szczeniowski. The influence of a cloud of electrons on the structure of de Broglie waves. Bull. Acad. Polon. Sci., Ser. A, ??(??):482–488, ???? 1931.

Infeld:1931:ISC

[IS31b] Leopold Infeld and S. Szczeniowski. The influence of space charge on the structure of de Broglie waves. Acta Physica Polonica, 1 (??):37–46, ???? 1931. CODEN APPOAK. ISSN 0001-673X.

Jackson:1987:ISR

[Jac87] J. David Jackson. The impact of special relativity on theoretical physics. Physics Today, 40(5):34–42, May 1, 1987.

CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881108.

Jacob:2001:LLR

[Jac01] Maurice Jacob. Louis Leprince-Ringuet. Physics Today, July 1, 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/1.1397405.

Jahnert:2010:QTC

[Jäh10] Martin Jähnert. Quantum theory at the crossroads: Reconsidering the 1927 Solvay Conference. Physics Today, October 1, 2010. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/1.3502549.

Jacobson:1995:PBW

[JBCY95] Joseph Jacobson, Gunnar Björk, Isaac Chuang, and Yoshihisa Yamamoto. Photonic de Broglie waves. Physical Review Letters, 74(??):4835, ???? 1995. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Jahn:1991:QMM

[JKA91] Robert G. Jahn, Paul Kolodner, and Philip W. Anderson. A question of mind over measurement. Physics Today, 44(10):13– 146, October 1, 1991. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2810271.

Jauch:1973:QR

[JL73] J. M. Jauch and R. Bruce Lindsay. Are quanta real? Physics Today, 26(8):55–56, August 1, 1973. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3128184.

Jammer:1967:CDQ

[JM67] Max Jammer and Eugen Merzbacher. The conceptual development of quantum mechanics. Physics Today, 20(3):102–106, March 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3034186.

John:2013:CSM

[JM13] Moncy V. John and Kiran Mathew. Coherent states and modified de Broglie–Bohm complex quantum trajectories. Foundations of Physics, 43(7):859–871, 2013. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Juffmann:2012:NPB

[JNA^{+}12] Thomas Juffmann, Stefan Nimmrichter, Markus Arndt, Herbert Gleiter, and Klaus Hornberger. New prospects for de Broglie interferometry. Foundations of Physics, 42(1):98–110, January 2012. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10701-010-9520-5.

Johnson:1977:NB

[Joh77] Laurence W. Johnson. Newton and de Broglie. Physics Today, 30 (5):100–101, May 1, 1977. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3037563.

John:1991:LL

[Joh91] Sajeev John. Localization of light. Physics Today, 44(5):32–40, May 1, 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.881300.

John:2002:MBB

[Joh02] Moncy V. John. Modified de Broglie–Bohm approach to quantum mechanics. Foundations of Physics Letters, 15(4):329–343, August 2002. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Jones:2008:QTS

[Jon08] Sheilla Jones. The Quantum Ten: a Story of Passion, Tragedy, Ambition and Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-536909-2. xii + 323 + 8 pp. LCCN QC174.12 .J66 2008.

Jewkes:1971:SI

[JSSS71] J. Jewkes, D. Sawyers, R. Stillerman, and Roland Schmitt. The sources of invention. Physics Today, 24(11):55–57, November 1, 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3022438.

K:1952:NSF

[K.52] W. K. Notes on science: First award of Kalinga Prize [to Louis de Broglie] — best light for the eyes. New York Times, ??(??): E9, June 15, 1952. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/112267262/.

Kaempffert:1931:HEU

[Kae31] Waldemar Kaempffert. How to explain the universe? Science in a quandary: One after another the theories put out by the scientists have been exploded and now science in its uncertainty has been forced to become idealistic and to drop the idea of a mechanical universe. New York Times, ???(??):120, January 11, 1931. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/99177884/13F390B7A1F783DA798.

Kaempffert:1948:RRB

[Kae48] Waldemar Kaempffert. The revolution that radium began: Fifty years after the Curies’ great discovery, nuclear physics is still a realm unbounded. New York Times, ??(??):SM13, SM25, SM27, December 26, 1948. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/108348269/.

Kaempffert:1953:BRP

[Kae53] Waldemar Kaempffert. Book review: In packages, not waves: The Revolution in Physics: A Non-Mathematical Survey of Quanta. By Louis de Broglie. Translated from the French by Ralph W. Niemeyer. 310 pp., New York: The Noonday Press. $4.50. New York Times, ??(??):BR10, July 19, 1953. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.

proquest.com/hnpnewyorktimes/docview/112668675.

Karapetoff:1929:RTO

[Kar29] V. Karapetoff. Relativity transformation of an oscillation into a traveling wave, and de Broglie’s postulate in terms of velocity angle. J. Opt. Soc. Amer., 19:253–265, 1929. ISSN 0030-3941.

Karplus:1971:WMP

[Kar71] Robert Karplus. What is modern physics? Physics Today, 24 (3):13–14, March 1, 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3022613.

Karl:2005:BRE

[Kar05] Gabriel Karl. Book reviews: Einstein Defiant: Genius Versus Genius in the Quantum Revolution, Edmund Blair Bolles, Joseph Henry Press, Washington, DC, 2004. $27.95 (348 pp.). ISBN 0-309-08998-0. Einstein’s Cosmos: How Albert Einstein’s Vision Transformed Our Understanding of Space and Time, Michio Kaku, Atlas Books / W. W. Norton, New York, 2004. $22.95 (251 pp.). ISBN 0-393-05165-X. Physics Today, 58(4):59–60, April 1, 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/1.1955481.

Kastner:2013:BWB

[Kas13] Ruth E. Kastner. de Broglie waves as the “bridge of becoming” between quantum theory and relativity. Foundations of Science, 18(1):1–9, March 2013. CODEN FOSCFI. ISSN 1233-1821 (print), 1572-8471 (electronic). URL http://link.springer.

com/article/10.1007/s10699-011-9273-4.

Kwal:1951:QCT

[KdB51] Bernard Kwal and Louis de Broglie. Quelques considérations sur les transformations de jauge et la définition des tenseurs de Hertz en théorie du corpuscule maxwellien de spin 1. (French) [Some considerations on gauge transformations and the definition of Hertz tensors in the theory of spin-1 Maxwellian particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 232(??):2056–2058, ???? 1951. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Ketterle:1999:ESB

[Ket99] Wolfgang Ketterle. Experimental studies of Bose–Einstein condensation. Physics Today, 52(12):30–35, December 1, 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.882898.

Khan:2005:CEH

[Kha05] Sameen Ahmed Khan. CERN’s early history revisited. Physics Today, 58(4):87–89, April 1, 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.4796963.

Khan:2020:TMN

[Kha20] Sieeka Khan. Top 10 most notable science anniversaries in 2020. Web site., January 8, 2020. URL https://www.sciencetimes.

com/articles/24639/20200107/science.htm.

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Khurana:1988:BET

[Khu88] Anil Khurana. Ballistic electron transport through a narrow channel is quantized. Physics Today, 41(11):21–23, November 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2811617.

Kiessling:2006:QAM

[Kie06] Michael K.-H. Kiessling. Quantum Abraham models with de Broglie–Bohm laws of electron motion. In Quantum mechanics. Are there quantum jumps? Trieste, Italy, 5 September 2005 and on the present status of quantum mechanics, BLošinj, Croatia, 7–9 September 2005. Proceedings. In honor of Giancarlo Ghirardi’s 70th birthday, pages 206–227. American Institute of Physics, Woodbury, NY, USA, 2006. ISBN 0-7354-0337-6 (hardcover).

Kiessling:2010:MSA

[Kie10] Michael K.-H. Kiessling. Misleading signposts along the de Broglie–Bohm road to quantum mechanics. Foundations of Physics, 40(4):418–429, 2010. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Kimball:2015:PCO

[Kim15] John Kimball. Physics curiosities, oddities, and novelties. CRC Press, Taylor and Francis Group, Boca Raton, FL, USA, 2015. ISBN 1-4665-7635-9 (paperback), 1-4665-7636-7 (e-book). xvii + 340 pp. LCCN QC24.5 .K56 2015eb. URL http://www.

crcnetbase.com/isbn/9781466576353.

Kirilyuk:1998:UCC

[Kir98] Andrei P. Kirilyuk. Universal concept of complexity by the dynamic redundance paradigm: Causal randomness, complete wave mechanics, and the ultimate unification of knowledge. arxiv.org, ??(??):1–37, May 30, 1998. URL https://arxiv.org/abs/

physics/9806002.

Kirilyuk:1999:YMW

[Kir99] Andrei P. Kirilyuk. 75 years of matter wave: Louis de Broglie and renaissance of the causally complete knowledge. arxiv.org, ?? (??):1–28, November 25, 1999. URL https://arxiv.org/abs/ quant-ph/9911107. Revised 25 November 2004.

Krachmalnicoff:2010:SFW

[KJB^{+}10] V. Krachmalnicoff, J.-C. Jaskula, M. Bonneau, V. Leung, G. B. Partridge, D. Boiron, C. I. Westbrook, P. Deuar, P. Ziń, M. Trippenbach, and K. V. Kheruntsyan. Spontaneous four-wave mixing of de Broglie waves: Beyond optics. Physical Review Letters, 104(??):150402, ???? 2010. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Kenmoku:1998:BBI

[KKTY98] Masakatsu Kenmoku, Hiroto Kubotani, Eiichi Takasugi, and Yuki Yamazaki. de Broglie–Bohm interpretation for the wave function of quantum black holes. Physical Review D (Particles and Fields), 57(??):4925, ???? 1998. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Kenmoku:2000:BBIb

[KKTY00a] Masakatsu Kenmoku, Hiroto Kubotani, Eiichi Takasugi, and Yuki Yamazaki. de Broglie–Bohm interpretation for wave function of Reissner–Nordström–de Sitter black hole. International Journal of Modern Physics. A, Particles and Fields, Gravitation, Cosmology, 15(14):2059–2075, 2000. CODEN IMPAEF. ISSN 0217-751X (print), 1793-656X (electronic).

Kenmoku:2000:BBIa

[KKTY00b] Masakatsu Kenmoku, Hiroto Kubotani, Eiichi Takasugi, and Yuki Yamazaki. de Broglie–Bohm interpretation of the wave function of a Reissner–Nordström–de Sitter black hole. Gravitation & Cosmology, 6(2(22)):116–122, 2000. ISSN 0202-2893 (print), 1995-0721 (electronic).

Kenmoku:2001:BBI

[KKTY01] Masakatsu Kenmoku, Hiroto Kubotani, Eiichi Takasugi, and Yuki Yamazaki. de Broglie–Bohm interpretation for analytic solutions of the Wheeler–De Witt equation in spherically symmetric spacetime. Progress of Theoretical Physics, 105(6):897–914, 2001. CO-DEN PTPKAV. ISSN 0033-068X (print), 1347-4081 (electronic).

Klein:1927:AMB

[Kle27] O. Klein. Sur l’article de M. L. de Broglie: “L’univers à cinq dimensions et la mécanique ondulatoire”. (French) [On the article by M. L. de Broglie: “The five-dimensional universe and wave mechanics”]. Radium (6), 8:242–243, 1927. ISSN 0368-3842.

Kleinman:1979:MP

[Kle79] David Kleinman. To modern physics. Physics Today, 32(10):15, October 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2995233.

Kleppner:1996:MSC

[Kle96] Daniel Kleppner. Memoirs of Schrödinger’s cat. Physics Today, 49(11):11–13, November 1, 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.881601.

Kleppner:1997:BGA

[Kle97] Daniel Kleppner. A beginner’s guide to the atom laser. Physics Today, August 1, 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.881877.

Kleppner:1999:YYH

[Kle99] Daniel Kleppner. The yin and yang of hydrogen. Physics Today, 52(4):11–13, April 1, 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.882619.

Kleppner:2014:BRE

[Kle14] Daniel Kleppner. Book review: Einstein and the Quantum: The Quest of the Valiant Swabian, A. Douglas Stone, Princeton U. Press, 2013. $29.95 (332 pp.). ISBN 978-0-691-13968-5. Physics Today, 67(4):48, April 1, 2014. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/PT.3.2349.

Klotz:1985:EV

[Klo85] Irving M. Klotz. Einstein and the Vatican. Physics Today, 38(7): 13–15, July 1, 1985. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2814628.

Kapteyn:2005:ENO

[KMC05] Henry C. Kapteyn, Margaret M. Murnane, and Ivan P. Christov. Extreme nonlinear optics: Coherent X-rays from lasers. Physics Today, 58(3):39–46, March 1, 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.1897563.

Kermisch:1974:MMW

[KMM74] Dorian Kermisch, George L. Murphy, and Heinrich A. Medicus. More on matter waves. Physics Today, August 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3128844.

Kenmoku:2002:CQS

[KMSU02] M. Kenmoku, T. Matsuyama, R. Sato, and S. Uchida. Classical and quantum solutions of (2+1)-dimensional gravity under the de Broglie–Bohm interpretation. Classical and quantum gravity, 19 (4):799–809, 2002. CODEN CQGRDG. ISSN 0264-9381 (print), 1361-6382 (electronic).

Kobe:2007:QPB

[Kob07] Donald H. Kobe. Quantum power in de Broglie–Bohm theory. Journal of Physics A: Mathematical and Theoretical, 40(19):5155– 5162, 2007. CODEN JPAMB5. ISSN 1751-8113 (print), 1751-8121 (electronic).

Kojima:1996:NPPa

[Koj96a] Chieko Kojima. Notes prises par Louis de Broglie lors des cours de Paul Langevin au collège de France sur la théorie des quanta. I. (French) [Notes taken by Louis de Broglie during Paul Langevin’s lectures at the College of France on quantum theory. I]. Annales de la Fondation Louis de Broglie, 21(1):81–89, 1996. ISSN 0182-4295 (print), 2108-6397 (electronic).

Kojima:1996:NPPb

[Koj96b] Chieko Kojima. Notes prises par Louis de Broglie lors des cours de Paul Langevin au collège de France sur la théorie des quanta. II. (French) [Notes taken by Louis de Broglie during Paul Langevin’s lectures at the College of France on quantum theory. II]. Annales de la Fondation Louis de Broglie, 21(3):335–343, 1996. ISSN 0182-4295 (print), 2108-6397 (electronic).

Kojima:2001:QPV

[Koj01] Chieko Kojima. Que pensait vraiment Langevin de la thèse de Louis de Broglie?. (French) [What did Langevin really think of Louis de Broglie’s thesis?]. Annales de la Fondation Louis de Broglie, 26(??):241–248, 2001. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://aflb.ensmp.fr/

AFLB-26j/aflb26jp241.htm.

Kojima:2004:PFA

[Koj04] Chieko Kojima. La physique française avant Louis de Broglie. (French) [French physics before Louis de Broglie]. Annales de la Fondation Louis de Broglie, 29(??):767–784, 2004. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://

aflb.ensmp.fr/AFLB-295/aflb295m312.htm.

Korn:1930:BWM

[Kor30] A. Korn. Die de Broglie-Wellen in mechanistischer Vorstellung und eine erweiterte Zustandsgleichung für Gase. (German) [de Broglie waves in mechanistic picture and an extended state equation for gases]. Physikalische Zeitschrift, 31:955–957, 1930. CO-DEN PHZTAO. ISSN 0369-982X.

Korff:1963:CRM

[Kor63] Serge A. Korff. Cosmic rays or machines? Physics Today, 16(9): 46–52, September 1, 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3051140.

King:1962:T

[KR62] Allen L. King and Stuart A. Rice. Thermophysics. Physics Today, 15(12):72–74, December 1, 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3057930.

Kwon:2010:OPB

[KRK10] Osung Kwon, Young-Sik Ra, and Yoon-Ho Kim. Observing photonic de Broglie waves without the maximally-path-entangled |N, 0 > +|0, N > state. Physical Review A (Atomic, Molecular, and Optical Physics), 81(??):063801, ???? 2010. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-

4519. URL http://www.w3.org/1998/Math/MathML.

Kopal:1962:AAA

[KS62] Zdenek Kopal and Otto Struve. Advances in astronomy and astrophysics, volume 1. Physics Today, 15(12):74, December 1, 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3057931.

King:1966:NBM

[KT66] Terry Johnson King and Dietrich E. Thomsen. The neutron beam murder. Physics Today, 19(1):128–129, January 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3047928.

Kubli:1970:LBE

[Kub70] Fritz Kubli. Louis de Broglie und die Entdeckung der Materiewellen. (German) [Louis de Broglie and the discovery of matter waves]. Archive for History of Exact Sciences, 7(1):26– 68, January 1970. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0003-9519&volume=7&
issue=1&spage=26.

Kulhanek:1969:EBW

[Kul69] J. Kulhanek. Equations of the de Broglie wavefield in the case of spherical symmetry. Journal of Mathematical Physics, 10 (7):1152–1156, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i7/p1152_s1.

Kulhanek:1970:EBW

[Kul70] J. Kulhánek. Equations of the de Broglie wave field and their relationship to Riemann’s curvature tensor. Journal of Mathematical Physics, 11:3199–3200, 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic).

Kulhanek:1971:STB

[Kul71] J. Kulhánek. Space–time of the de Broglie wave field. Journal of Mathematical Physics, 12(8):1535–1538, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i8/
p1535_s1.

L:1958:RLB

[L.58] A.-L. L. Review: Louis de Broglie, physicien et penseur by Albert Einstein, Louis de Broglie, and 44 others. Revue Philosophique de la France et de l’Étranger, 148:102, 1958. ISSN 0035-3833 (print),

2104-385X (electronic). URL http://www.jstor.org/stable/
41089668.

Laboccetta:1935:LBF

[Lab35] Letterio Laboccetta. L’onda di de Broglie e la frequenza di Compton nella meccanica classica. (Italian) [The de Broglie wave and the Compton frequency in classical mechanics]. Ricerca Scientifica ed il Progresso Tecnico, 2(??):107–108, 1935. CODEN RSPTB6.

Ladbury:1996:MSL

[Lad96] Ray Ladbury. Model sheds light on a tragedy and a new type of eruption. Physics Today, 49(5):20–22, May 1, 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2807610.

Loeb:1966:ECT

[LB66] Leonard B. Loeb and Sanborn C. Brown. Electrical coronas: Their basic physical mechanisms. Physics Today, 19(1):109–111, January 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.3047912.

Lande:1968:DD

[LBB68] Alfred Landé, Max Born, and Walter Biem. Dialog on dualism. Physics Today, 21(8):55–56, August 1, 1968. CODEN PHTOAD.

ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3035103.

Liou:1993:RCP

[LD93] K. N. Liou and Roger Davies. Radiation and cloud processes in the atmosphere. Physics Today, 46(9):66–67, September 1, 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2809044.

Langevin:1912:TDR

[LdB12] Paul Langevin and Louis de Broglie, editors. La théorie du rayonnement et les quanta: rapports et discussions de la réunion tenue à Bruxelles du 30 octobre au 3 novembre 1911 / sous les auspices de m. E. Solvay. (French) [Radiation theory and quanta: reports and discussions of the meeting held in Brussels from 30 October to 3 November 1911 / under the auspices of Mr. E. Solvay]. Gauthier-Villars, Paris, France, 1912. URL https://archive.

org/details/lathoriedurayo00inst.

Lee:1979:PEP

[Lee79] James B. Lee. Physicists and energy problem. Physics Today, 32 (12):78, December 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2995333.

Lee:1996:ABW

[Lee96] Chang Jae Lee. Atomic de Broglie waves in multiple optical standing waves. Physical Review A (Atomic, Molecular, and Optical Physics), 53(??):4238, ???? 1996. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Lenzen:1948:BRL

[Len48] V. F. Lenzen. Book review: L’évolution de la notion de phénomène physique, des primitifs à Bohr et Louis de Broglie by Jean Pelseneer. Isis, 39(3):194–196, August 1948. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/226331.

Lepore:1992:HDB

[Lep92] Vito Luigi Lepore. Homodyne detection of de Broglie waves. Foundations of Physics Letters, 5(5):469–478, October 1992. CO-DEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Lester:1979:CC

[Les79] William A. Lester, Jr. Chemistry computation. Physics Today, 32(8):11, August 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2995656.

Levi:1988:PMO

[Lev88] Barbara Goss Levi. Positron microscopes offer a new view of things. Physics Today, 41(12):17–19, December 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2811659.

Levi:1991:ANW

[Lev91] Barbara Goss Levi. Atoms are the new wave in interferometers. Physics Today, 44(7):17–20, July 1, 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2810166.

Levi:1993:ACA

[Lev93] Barbara Goss Levi. AIP Corporate Associates meet at Hewlett– Packard’s Palo Alto Lab. Physics Today, 46(2):77–78, February 1, 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2808811.

Levi:1995:AIP

[Lev95] Barbara Goss Levi. Atom interferometers prove their worth in atomic measurements. Physics Today, July 1, 1995. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2808088.

Levi:1996:RVA

[Lev96] Barbara Goss Levi. Researchers vie to achieve a quantumdot laser. Physics Today, 49(5):22–24, May 1, 1996. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2807611.

Levi:1997:BCC

[Lev97a] Barbara Goss Levi. Bose condensates are coherent inside and outside an atom trap. Physics Today, 50(3):17–18, March 1, 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.881708.

Levi:1997:THC

[Lev97b] Barbara Goss Levi. Trap holds condensates of two different spin states at once. Physics Today, 50(3):18–19, March 1, 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881726.

Levi:1998:LLB

[Lev98] Barbara Goss Levi. At long last, a Bose–Einstein condensate is formed in hydrogen. Physics Today, 51(10):17–19, October 1, 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.882393.

Levi:2001:CKW

[Lev01a] Barbara G. Levi. Cornell, Ketterle, and Wieman share Nobel Prize for Bose–Einstein condensates. Physics Today, 54(12):14– 16, December 1, 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/full/10.1063/1.1445529.

Levi:2001:MFN

[Lev01b] Barbara Goss Levi. Magnetic forces need not apply: Bose–einstein condensates can be made in an optical trap. Physics Today, 54(7): 20–22, July 1, 2001. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/1.1397387.

Levi:2004:UGF

[Lev04] Barbara Goss Levi. Ultracold gases of fermionic atoms offer another path to atom interferometry. Physics Today, 57(8):25–26, August 1, 2004. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/1.1801857.

Levin:2017:TCS

[Lev17] Frank S. Levin. On theory and consensus in science: a journey in time. Web blog., October 17, 2017. URL https://blog.oup.

com/2017/10/theory-consensus-in-science/.

Leydolt:2005:BWW

[Ley05] Hans J. Leydolt. de Broglie’s wavefunction and wave-particle dualism. In Foundations of probability and physics — 3. Proceedings of the international conference, Växjö, Sweden, June 7–12, 2004, pages 236–253. American Institute of Physics, Woodbury, NY, USA, 2005. ISBN 0-7354-0235-3 (hardcover).

Leute:1989:AIP

[LH89] Ulrich Leute and R. E. M. Hedges. Archaeometry: An introduction to physical methods in archaeology and the history of art. Physics Today, 42(11):93–94, November 1, 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2811218.

Liboff:1984:CPR

[Lib84] Richard L. Liboff. The correspondence principle revisited. Physics Today, 37(2):50–55, February 1, 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2916084.

Lindsay:1964:BRL

[Lin64] R. B. Lindsay. Book review: Louis de Broglie, Introduction to the Vigier Theory of Elementary Particles. Physics Today, 17(10):57, October 1, 1964. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3051176.

Lindenbaum:1965:LCC

[Lin65a] S. J. Lindenbaum. The on-line computer counter and digitalized spark-chamber technique. Physics Today, 18(4):19–28, April 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3047330.

Lindsay:1965:BRLb

[Lin65b] R. B. Lindsay. Book review: Louis de Broglie, La Thermodynamique de la Particule isolée (ou Thermodynamique cachée des Particules). Physics Today, 18(12):62, December 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3047037.

Lindsay:1965:BRLa

[Lin65c] R. B. Lindsay. Book review: Louis de Broglie, The Current Interpretation of Wave Mechanics. Physics Today, 18(3):67–68, March 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3047278.

Lindsay:1967:BRL

[Lin67] R. Bruce Lindsay. Book review: Louis de Broglie, Wave Mechanics and Molecular Biology. Physics Today, 20(6):77, 79, June 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3034372.

Lindsay:1969:BRO

[Lin69] R. Bruce Lindsay. Book review: Ondes Electromagnetiques et Photons, by Louis de Broglie, 99 pp. Gauthier-Villars, Paris, 1968. Physics Today, 22(5):97–98, May 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3035601.

Lipkin:2000:WOT

[Lip00] Harry J. Lipkin. Who ordered theorists? Physics Today, 53(7): 15–74, July 1, 2000. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.1292467.

Lochak:1996:ILB

[LKF96] G. Lochak, M. Karatchentzeff, and D. Fargue. Introduction to Louis de Broglie’s Diverses questions de mécanique et de thermodynamique classiques et relativistes. Annales de la Fondation Louis de Broglie, 21(2):117–130, 1996. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

Lorenz:1995:EC

[LM95] Edward N. Lorenz and Paul Martin. The essence of chaos. Physics Today, February 1, 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2807913.

Liu:2010:SPB

[LO10] B. Liu and Z. Y. Ou. Six-photon de Broglie wavelength and phasemeasurement sensitivity improving upon the standard quantum limit. Physical Review A (Atomic, Molecular, and Optical Physics), 81(??):023823, ???? 2010. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Lochak:1982:EIL

[Loc82] Georges Lochak. The evolution of the ideas of Louis de Broglie on the interpretation of wave mechanics. Foundations of Physics, 12 (10):931–953, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01889267. See comment [vdM82].

Lochak:1987:CDB

[Loc87a] Georges Lochak. Convergence and divergence between the ideas of de Broglie and Schrödinger in wave mechanics. Foundations of Physics, 17(12):1189–1203, December 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF01889593.

Lochak:1987:LB

[Loc87b] Georges Lochak. Louis de Broglie (1892–1987). Foundations of Physics, 17(10):967–970, October 1987. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.

springer.com/article/10.1007/BF00938006.

Lochak:1993:LBC

[Loc93] Georges Lochak. Louis de Broglie’s conception of physics. Foundations of Physics, 23(1):123–131, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF01883990.

Lochak:1995:PSP

[Loc95] G. Lochak. Sur la présence d’un second photon dans la théorie de la lumière de de Broglie. (French) [On the presence of a second photon in de Broglie’s theory of light]. Annales de la Fondation Louis de Broglie, 20(1):111–114, 1995. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

Lochak:2004:LBS

[Loc04a] G. Lochak. Louis de Broglie: sa conception du monde physique. (French) [Louis de Broglie: his conception of the physical world]. Annales de la Fondation Louis de Broglie, 29(??):749–766, 2004. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

URL http://aflb.ensmp.fr/AFLB-295/aflb295m311.htm.

Lochak:2004:PEP

[Loc04b] G. Lochak. “photons électriques” et “photons magnétiques” dans la théorie du photon de de Broglie. (French) [“Electric photons” and “magnetic photons” in de Broglie’s photon theory]. Annales de la Fondation Louis de Broglie, 29(1–2):297–316, 2004. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http:

//aflb.ensmp.fr/AFLB-291/aflb291p297.htm.

Lochak:2007:COL

[Loc07] Georges Lochak. A complementary opposition: Louis de Broglie and Werner Heisenberg. In Evans and Thorndike [ET07], pages 73–80. ISBN 3-540-32663-4 (hardcover), 3-540-32665-0 (e-book). ISSN 1612-3018. LCCN QC174.12 .Q346 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0824/2006934045-b.
html;   http://www.loc.gov/catdir/enhancements/fy0824/
2006934045-d.html;   http://www.loc.gov/catdir/toc/fy0708/
2006934045.html.

Lander:1985:NSS

[LP85] Gerard H. Lander and David L. Price. Neutron scattering with spallation sources. Physics Today, 38(1):38–45, January 1, 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881009.

Leibfried:1998:SMR

[LPM98] Dietrich Leibfried, Tilman Pfau, and Christopher Monroe. Shadows and mirrors: Reconstructing quantum states of atom motion. Physics Today, April 1, 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.882256.

Lande:1969:SR

[LS69a] Alfred Landé and Mendel Sachs. Subjective relativity. Physics Today, 22(11):11–17, November 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3035247.

Lawden:1969:MPQ

[LS69b] Derek F. Lawden and Garrison Sposito. The mathematical principles of quantum mechanics. Physics Today, 22(5):82–85, May 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3035593.

Ludwig:1969:WM

[LS69c] Gunther Ludwig and Garrison Sposito. Wave mechanics. Physics Today, 22(4):107–109, April 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3035501.

Lumsden:1974:SB

[LS74] Charles J. Lumsden and Melvin Silverman. Superlattices in biology. Physics Today, May 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3128574.

Liu:2008:DTP

[LSG^{+}08] B. H. Liu, F. W. Sun, Y. X. Gong, Y. F. Huang, Z. Y. Ou, and G. C. Guo. Demonstration of the three-photon de Broglie wavelength by projection measurement. Physical Review A (Atomic, Molecular, and Optical Physics), 77(??):023815, ???? 2008. CO-DEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Lipkin:1966:LGP

[LT66] Harry J. Lipkin and John G. Taylor. Lie groups for pedestrians. Physics Today, 19(1):111–113, January 1, 1966. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3047915.

Lubkin:1980:SSP

[Lub80] Gloria B. Lubkin. Solid state physics in China. Physics Today, 33 (6):20–21, June 1, 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2914113.

Lubkin:1984:SDG

[Lub84] Gloria B. Lubkin. SSC design goes to DOE: ICFA discusses CERN hadron collider. Physics Today, 37(6):17–19, June 1, 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2916267.

Lubkin:1996:ECH

[Lub96a] Gloria B. Lubkin. Experimenters cool helium below single-photon recoil limit in three dimensions. Physics Today, 49(1):22–24, January 1, 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.2807458.

Lubkin:1996:ORB

[Lub96b] Gloria B. Lubkin. Oak ridge builds facility for radioactive ion beams. Physics Today, 49(1):24, January 1, 1996. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2807459.

Lubkin:1999:NAL

[Lub99] Gloria B. Lubkin. New atom lasers eject atoms or run CW. Physics Today, 52(4):17–18, April 1, 1999. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.882646.

Lupasco:1949:RCD

[Lup49] Stéphane Lupasco. Review: Connaissance du monde physique. Éditions Albin Michel by Ulysse Filippi; Louis de Broglie; André George. Revue Philosophique de la France et de l’Étranger, 139: 243–244, 1949. ISSN 0035-3833 (print), 2104-385X (electronic).

URL http://www.jstor.org/stable/41087238.

Lagendijk:2009:FYA

[LvTW09] Ad Lagendijk, Bart van Tiggelen, and Diederik S. Wiersma. Fifty years of Anderson localization. Physics Today, August 1, 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/1.3206091.

Lynn:2002:BRN

[Lyn02] Jeffrey W. Lynn. Book review: Neutron Interferometry: Lessons in Experimental Quantum Mechanics, Helmut Rauch and Samuel A. Werner, Oxford U. Press, New York, 2000. $120.00 (393 pp.) ISBN 0-19-850027-0. Physics Today, 55(6):56, June 1, 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/1.1496381.

Mackinnon:1978:NBW

[Mac78] L. Mackinnon. A nondispersive de Broglie wave packet. Foundations of Physics, 8(3–4):157–176, April 1978. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://

link.springer.com/article/10.1007/BF00715205.

Malisoff:1939:BRN

[Mal39] William Marias Malisoff. Book review: On the new physics: Matter and Light. The New Physics, by Louis de Broglie. Translated by W. H. Johnston. 300 pp. New York: W. W. Norton & Co., Inc. New York Times, ??(??):106, October 1, 1939. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://

search.proquest.com/hnpnewyorktimes/docview/103076152.

Malamud:1977:WC

[Mal77] Herbert Malamud. Welfare costs. Physics Today, 30(5):101, May 1, 1977. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3037564.

Marton:1964:ADE

[Mar64] L. Marton. APS Division of Electron Physics the first 20 years. Physics Today, 17(10):44–50, October 1, 1964. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3051174.

Marquardt:1979:LA

[Mar79] Charles L. Marquardt. Laser annealing. Physics Today, 32(5): 13–15, May 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2995544.

Margaritondo:1988:YP

[Mar88] Giorgio Margaritondo. 100 years of photoemission. Physics Today, 41(4):66–72, April 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.881150.

Maroney:2005:DMB

[Mar05] O. J. E. Maroney. The density matrix in the de Broglie–Bohm approach. Foundations of Physics, 35(3):493–510, 2005. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic).

Matzkin:2017:TDS

[Mat17] A. Matzkin. The theory of the Double Solution: Dynamical issues in quantum systems in the semiclassical regime. Annales de la Fondation Louis de Broglie, 42(1):213–221, 2017. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://

aflb.ensmp.fr/AFLB-421/aflb421m9.htm.

May:1946:REI

[May46] Ed. May. Review: Die Elementarteilchen. Individualität und Wechselwirkung: Ergebnisse der Neuen Physik. II. Teil by Louis de Broglie. Zeitschrift für philosophische Forschung, 1(1):155– 156, 1946. ISSN 0044-3301 (print), 1439-2615 (electronic). URL

http://www.jstor.org/stable/20480077.

McKelvey:1979:SP

[MB79] John P. McKelvey and Lewis M. Branscomb. Science policy. Physics Today, 32(12):78–79, December 1, 1979. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2995334.

McDonald:1979:NWD

[McD79] Richard J. McDonald. Nuclear weapons disposal. Physics Today, 32(8):13, August 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2995658.

McWane:1977:PT

[McW77] John W. McWane. Physics of technology. Physics Today, 30(5): 100, May 1, 1977. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3037562.

Mehra:1976:SCP

[ME76] J. Mehra and Paul P. Ewald. The Solvay Conferences on Physics: Aspects of the development of physics since 1911. Physics Today, 29(7):47–49, July 1, 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3023564.

Medicus:1974:FYM

[Med74] Heinrich A. Medicus. Fifty years of matter waves. Physics Today, February 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3128444.

Meijer:1977:IA

[Mei77] Paul H. E. Meijer. International agreement. Physics Today, 30(5): 100, May 1, 1977. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3037561.

Mercier:1952:LPE

[Mer52] A. Mercier. Leçons sur les principes de l’électro-dynamique classique. (French) [Lessons on the principles of classical electrodynamics]. Number 23 in Bibliothèque scientifique. Éditions du Griffon, Neuchatel, Switzerland, 1952. 74 pp. Préface de Louis de Broglie.

Merzbacher:2002:EHQ

[Mer02] Eugen Merzbacher. The early history of quantum tunneling. Physics Today, 55(8):44–49, August 1, 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.1510281.

Mermin:2009:WBA

[Mer09] N. David Mermin. What’s bad about this habit. Physics Today, 62 (5):8–9, May 1, 2009. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/1.3141952.

Metadier:1938:TAC

[Met38] Jacques Metadier. La théorie de l’agitation chaotique. Mouvement brownien, agitation moléculaire, dispersion, floculation, etc. . . .. (French) [The theory of chaotic agitation. Brownian motion, molecular agitation, dispersion, flocculation, etc. …]. Hermann, Paris, France, 1938. 71 pp. Preface de Louis de Broglie.

Meyerson:1933:RDP

[Mey33] Emile Meyerson. Réel et déterminisme dans la physique quantique, volume 68 of Exposés de philosophie des sciences,; Actualités scientifiques et industrielles. Hermann, Paris, France, 1933. 49 pp. LCCN Q111 .A3 no. 68. Preface by Louis de Broglie (pages 3–5).

Maffei:1979:BM

[MH79] P. Maffei and Richard C. Hart. Beyond the moon. Physics Today, 32(4):54–55, April 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2995493.

Millet:1958:RTM

[Mil58a] Louis Millet. Review: La théorie de la mesure en mécanique ondulatoire (interprétation usuelle et interprétation causale) by Louis de Broglie. Les Études philosophiques, 13(1):66, 1958. ISSN 0012-2017 (print), 1746-8361 (electronic). URL http://www.jstor.

org/stable/20842581.

Millet:1958:RPP

[Mil58b] Louis Millet. Review: Les principes physiques de la théorie des quanta by W. Heisenberg; Louis de Broglie. Les Études philosophiques, 13(1):75, January/March 1958. ISSN 0012-2017 (print), 1746-8361 (electronic). URL http://www.jstor.org/

stable/20842599.

Miller:1966:IIP

[Mil66] Donald G. Miller. Ignored intellect: Pierre Duhem. Physics Today, 19(12):47–53, December 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3047853.

Millar:1967:RWM

[Mil67] David B. Millar. Review: Wave Mechanics and Molecular Biology by Louis de Broglie. Quarterly Review of Biology, 42(4):566– 567, 1967. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718

(electronic). URL http://www.jstor.org/stable/2819895.

Miller:1985:WHB

[Mil85] Arthur I. Miller. Werner Heisenberg and the beginning of nuclear physics. Physics Today, 38(11):60–68, November 1, 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.880993.

Miller:2013:UMR

[Mil13] Johanna L. Miller. Ultrasound measurements reveal a long-sought phase transition in superconducting cuprates. Physics Today, 66 (8):12–15, August 1, 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/full/10.1063/PT.3.2066.

Miller:2015:ETI

[Mil15] Johanna L. Miller. Efimov trimers imaged for the first time. Physics Today, 68(7):10–12, July 1, 2015. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/PT.3.2834.

Morse:1969:TA

[MIS69] Philip M. Morse, K. Uno Ingard, and R. S. Shankland. Theoretical acoustics. Physics Today, 22(5):98–99, May 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3035602.

Mullish:1967:ICP

[MLG67] Henry Mullish, John A. N. Lee, and Joseph Gillis. An introduction to computer programming and numerical analysis for computers. Physics Today, June 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3034367.

Morrey:2000:TPM

[MLM^{+}00] W. C. Morrey, Hoi-Kwong Lo, Pantazis Mouroulis, Charles K. Scharnberger, Gary Powell, Philip E. Kaldon, Phil Baringer, Moorad Alexanian, and Mano Singham. Teaching, propaganda, and the middle ground. Physics Today, 53(11):14–81, November 1, 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.1333277.

Marto:2001:BBF

[MM01] J. Marto and P. Vargas Moniz. de Broglie–Bohm FRW universes in quantum string cosmology. Physical Review D (Particles and Fields), 65(??):023516, ???? 2001. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Mallove:1989:SH

[MMJ89] Eugene Mallove, Gregory Matloff, and Eric M. Jones. The starflight handbook. Physics Today, 42(11):94–96, November 1, 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2811219.

Mohler:1979:NP

[Moh79] Sailor H. Mohler. Neutrino paradox. Physics Today, 32(12):78, December 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2995332.

Molski:1995:DDB

[Mol95] M. Molski. Does a dual de Broglie wave exist? Annales de la Fondation Louis de Broglie, 20(2):181–186, 1995. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

Morrison:1954:BRG

[Mor54] P. Morrison. Book review: Geometrical Mechanics and de Broglie Waves. J. L. Synge. Part of Cambridge Monographs on Mechanics and Applied Mathematics. G. K. Batchelor and H. Bondi, Eds. Cambridge Univ. Press, New York, 1954. vi + 167 pp. Illus. $4.75. Science, 120(3120):600, October 15, 1954. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Motz:1979:MFE

[Mot79] Hans Motz. More on free-electron lasers. Physics Today, 32(12): 15–78, December 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2995308.

McKinsey:1954:RST

[MS54] J. C. C. McKinsey and Patrick Suppes. Review: La Structure des Théories Physiques. by Paulette Destouches-Février; Louis de Broglie. Journal of Symbolic Logic, 19(1):52–55, 1954. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL

http://www.jstor.org/stable/2267651.

Mendelssohn:1967:QAZ

[MS67] K. Mendelssohn and Martin E. Straumanis. The quest for absolute zero: The meaning of low temperature physics. Physics Today, 20 (11):101, November 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3033999.

Meyer:1937:FTL

[MSB^{+}37] Stefan Meyer, A. Norman Shaw, Niels Bohr, George Hevesy, le Duc de Broglie, Johannes Stark, Otto Hahn, Enrico Fermi, L. Wertenstein, and Peter Kapitza. Further tributes to the late Lord Rutherford. Energia elettrica. Tribute to Lord Rutherford. Nature, 140(3555):1047–1054, December 18, 1937. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http:

//www.nature.com/nature/journal/v140/n3555/index.html.

Mucci:1974:MFB

[Muc74] Joseph F. Mucci. On the mathematical form of de Broglie’s cyclical action integral. Foundations of Physics, 4(1):91–95, March 1974. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF00708558.

Muchnik:1979:RE

[Muc79] Boris J. Muchnik. Right to emigrate. Physics Today, 32(10):15, October 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2995234.

Mumm:1965:FDC

[Mum65] T. Mumm. Fusion de deux corpuscules de spin 1/2 dans le cas des solutions singulières et théorème de guidage de M. Louis de Broglie dans le cas des particules de spin 1. (French) [Fusion of two spin-1/2 particles in the case of singular solutions and guidance theorem of M. Louis de Broglie in the case of spin-1 particles]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 260(??):1103–1106, ???? 1965. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Mills:1979:CPS

[MW79] C. B. Mills and Lester C. Welch. Computer program support. Physics Today, 32(5):13, May 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2995543.

Narayanamurti:1984:CSH

[Nar84] Venkatesh Narayanamurti. Crystalline semiconductor heterostructures. Physics Today, 37(10):24–32, October 1, 1984.

CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2915912.

Nicolle:1954:SIA

[Nic54] Jacques Nicolle. Die Symmetrie und ihre Anwendungen. (German) [Symmetry and its applications]. Deutscher Verlag der Wissenschaften, Berlin, West Germany, 1954. 172 pp. Foreword by Louis de Broglie.

Niaz:2012:RWP

[NM12] Mansoor Niaz and Cecilia Marcano. Reconstruction of Waveparticle Duality and Its Implications for General Chemistry Textbooks. Springer briefs in education. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-4395-5 (paperback). viii + 46 pp. LCCN QC476.W38

N53 2012. URL http://www.springerlink.com/content/978-
94-007-4395-3.

Naaman-Marom:2012:PMB

[NMEV12] Gillie Naaman-Marom, Noam Erez, and Lev Vaidman. Position measurements in the de Broglie–Bohm interpretation of quantum mechanics. Annals of Physics, 327(10):2522–2542, 2012. CODEN APNYA6. ISSN 0003-4916 (print), 1096-035x (electronic).

Norsen:2022:PWA

[Nor22] Travis Norsen. A pilot-wave approach to the many-body problem: Beyond the small entanglement approximation. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://link.

springer.com/article/10.1007/s10701-022-00621-6.

Nugent:2001:PO

[NPG01] Keith A. Nugent, David Paganin, and Tim E. Gureyev. A phase odyssey. Physics Today, August 1, 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.1404846.

Nye:1997:ACP

[Nye97] Mary Jo Nye. Aristocratic culture and the pursuit of science: The de Broglies in modern France. Isis, 88(3):397–421, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236150.

Olshanii:2000:BWF

[ODHP00] M. Olshanii, N. Dekker, C. Herzog, and M. Prentiss. de Broglie wave-front engineering. Physical Review A (Atomic, Molecular, and Optical Physics), 62(??):033612, ???? 2000. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Oliphant:1966:TE

[Oli66] Mark L. Oliphant. The two Ernests. I. Physics Today, 19(9): 35–49, September 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3048466.

Osche:2001:ECR

[Osc01] G. R. Osche. Electron channeling resonance and de Broglie’s internal clock. Annales de la Fondation Louis de Broglie, 36 (??):61–71, ???? 2001. CODEN ???? ISSN 0182-4295 (print),

2108-6397 (electronic). URL http://aflb.ensmp.fr/AFLB-361/
aflb361m718.pdf.

Osche:2011:ECR

[Osc11] G. R. Osche. Electron channeling resonance and de Broglie’s internal clock. Annales de la Fondation Louis de Broglie, 36(1):61–62, 2011. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (elec-

tronic). URL http://aflb.ensmp.fr/AFLB-361/aflb361m718.
htm.

Palevsky:1963:ISN

[Pal63] H. Palevsky. Inelastic scattering of neutrons in solids and liquids. Physics Today, 16(1):41–42, January 1, 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3050713.

Patri:1945:RPF

[Pat45] A. Patri. Review: Principes fondamentaux de Physique théorique. (T. I, II et III) by D. L. Detouches; Louis de Broglie. Les Études philosophiques, 20:50–54, 1945. ISSN 0014-2166 (print), 2101-0056

(electronic). URL http://www.jstor.org/stable/20840942.

Pauli:1953:RPP

[Pau53] W. Pauli. Remarques sur le problème des paramètres cachés dans la mécanique quantique et sur la théorie de l’onde pilote. (French)

[Remarks on the problem of hidden variables in quantum mechanics and on the pilot-wave theory]. In Louis de Broglie, physicien et penseur, pages 33–42. Éditions Albin Michel, Paris, 1953.

Pauling:1979:PHR

[Pau79] Linus Pauling. Pauling on human rights. Physics Today, 32(5): 11–13, May 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2995541.

Price:1973:WMF

[PCR73] William Charles Price, Seymour S. Chissick, and Tom Ravensdale, editors. Wave mechanics: the first fifty years. A tribute to Professor Louis de Broglie, Nobel Laureate, on the 50th anniversary of the discovery of the wave nature of the electron. Wiley, New York, NY, USA, 1973. ISBN 0-470-69731-8. 435 pp. LCCN QC174.22 .W38 1973.

Patton:2006:SPA

[PDEK06] Forest S. Patton, Daniel P. Deponte, Greg S. Elliott, and Stephen D. Kevan. Speckle patterns with atomic and molecular de Broglie waves. Physical Review Letters, 97(??):013202, ???? 2006. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Peierls:1985:BPR

[Pei85] Sir Rudolf E. (Rudolf Ernst) Peierls. Bird of Passage: Recollections of a Physicist. Princeton legacy library. Princeton University Press, Princeton, NJ, USA, 1985. ISBN 0-691-08390-8, 0-691-02416-2 (paperback), 0-691-60220-4, 1-4008-5461-X (e-book). xii + 350 + 12 pp. LCCN QC16.P375 A32

1985. URL http://site.ebrary.com/id/10897402;   https:
//www.jstor.org/stable/j.ctt7ztn4b.

Pelseneer:1947:LNP

[Pel47a] Jean Pelseneer. L’évolution de la notion de phénomène physique des primitifs à Bohr et Louis de Broglie. Leçons sur l’histoire de la pensée scientifique professées à l’Université libre de Bruxelles. (French) [The evolution of the notion of the physical phenomenon of primitives of Bohr and Louis de Broglie. Lessons on the history of scientific thought given at the Free University of Brussels]. Office international de librairie, Bruxelles, Belgique, 1947. 176 pp. LCCN ????

Pelseneer:1947:LND

[Pel47b] Jean Pelseneer. L’évolution de la notion du phénomène physique des primitifs à Bohr et Louis de Broglie. (French) [Evolution of the concept of the physical phenomenon of primitives in Bohr and Louis de Broglie]. Office des cours du Cercle des Sciences de l’Université Libre de Bruxelles, Bruxelles, Belgium, 1947.

Perillan:2018:QNP

[Per18] José G. Perillán. Quantum narratives and the power of rhetorical omission: An early history of the pilot wave interpretation of quantum theory. Historical Studies in the Natural Sciences, 48 (1):24–55, February 2018. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic).

Pestre:1988:BRL

[Pes88] Dominique Pestre. Book review: Louis de Broglie: Un itinéraire scientifique by Louis de Broglie; Georges Lochak. Isis, 79(4):740– 741, December 1988. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
234813.

Pike:1977:NLJ

[Pik77] E. R. Pike. New letters journals again. Physics Today, 30(5):15– 100, May 1, 1977. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3037538.

Papenfuss:2002:YWH

[PLS02] Dietrich Papenfuss, D. (Dieter) Lüst, and Wolfgang Schleich, editors. 100 Years Werner Heisenberg: works and impact. Wiley-VCH, Berlin, Germany, Weinheim, Germany, and New York, NY, USA, 2002. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). ix + 299 pp. LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Pinto-Neto:2013:QCB

[PNF13] N. Pinto-Neto and J. C. Fabris. Quantum cosmology from the de Broglie–Bohm perspective. Classical and quantum gravity, 30(14):

57, 2013. CODEN CQGRDG. ISSN 0264-9381 (print), 1361-6382 (electronic).

Pinto-Neto:2012:QCT

[PNSS12] Nelson Pinto-Neto, Grasiele Santos, and Ward Struyve. Quantumto-classical transition of primordial cosmological perturbations in de Broglie–Bohm quantum theory. Physical Review D (Particles and Fields), 85(??):083506, ???? 2012. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Pinto-Neto:2014:QCT

[PNSS14] Nelson Pinto-Neto, Grasiele Santos, and Ward Struyve. Quantumto-classical transition of primordial cosmological perturbations in de Broglie–Bohm quantum theory: The bouncing scenario. Physical Review D (Particles and Fields), 89(??):023517, ???? 2014. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Poincare:1925:PMS

[Poi25] L. Poincaré. La physique moderne. Son évolution. (French) [Modern physics. Its evolution]. Bibliothèque de philosophie scientifique. E. Flammarion, Paris, France, second edition, 1925. iv + 288 pp. Followed by 3 complementary chapters by M. de Broglie.

Poincare:1954:O

[Poi54] Henri Poincaré. Œuvres. (French) [Works]. Gauthier-Villars, Paris, France, 1954. xvi + 704 pp. Publiées sous les auspices de l’Académie des Sciences par la Section de Géométrie. Tome IX. Publié avec la collaboration de Gérard Petiau. Préface de Louis de Broglie.

Poincare:1974:CIT

[Poi74] Henri Poincaré. Izbrannye trudy v trekh tomakh. Tom III: Matematika. Teoreticheskaya fizika. Analiz matematicheskikh i estestvennonauchnykh rabot Anri Puankare. (Russian) [Selected works in three volumes. Volume III: Mathematics. Theoretical physics. Analysis of mathematical and natural sciences of Henri Poincaré]. Seriya “Klassiki Nauki”. Izdat. “Nauka”, Moscow, USSR, 1974. 771 pp.

Pondrom:1981:BR

[PPF^{+}81] Lee G. Pondrom, Peter Pesch, Gordon R. Freeman, Kenneth Sassen, and N. David Mermin. The boojum returns. Physics Today, 34(9):11–15, September 1, 1981. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2914739.

Putterman:1971:QNS

[PR71] Seth J. Putterman and Isadore Rudnick. Quantum nature of superfluid helium. Physics Today, August 1, 1971. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3022882.

Proca:1938:ESGa

[Pro38] A. Proca. Sur une équation symbolique groupant les équations du mésoton (électron lourd), celles de Kemmer, de Klein–Gordon et les équations du photon de L. de Broglie. (French) [On a symbolic equation grouping the equations of the mesoton (heavy electron), those of Kemmer, Klein–Gordon and the photon equations of L. of Broglie]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 207(??):1182–1184, ???? 1938. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Pu:2002:AOB

[PSZM02] Han Pu, Chris Search, Weiping Zhang, and Pierre Meystre. Atom optics from de Broglie waves to Heisenberg ferromagnets. In Papenfuss et al. [PLS02], pages 235–240. ISBN 3-527-40392-2 (paperback), 3-527-61085-5 (e-book). LCCN QC16.H35 A15 2002. EUR 119.00. URL http://www.

loc.gov/catdir/bios/wiley046/2003268171.html;   http:
//www.loc.gov/catdir/description/wiley0310/2003268171.
html;   http://www.loc.gov/catdir/toc/wiley031/2003268171.
html.

Papini:1995:GSB

[PW95] G. Papini and W. R. Wood. A geometric solution to the de Broglie variable mass problem. Physics Letters A, 202(1):46–49, 1995. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Rabi:2006:SED

[RC06] Isidor Isaac Rabi and R. Fraser Code. Stories from the early days of quantum mechanics. Physics Today, 59(8):36–41, August 1, 2006. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/1.2349731.

Reichenbach:1944:DSE

[Rei44a] Hans Reichenbach. Derivation of Schrödinger’s equation from de Broglie’s principle. In Philosophic Foundations of Quantum Mechanics [Rei44b], chapter 13, pages 66–71. LCCN QC174.1 .R4.

Reichenbach:1944:PFQ

[Rei44b] Hans Reichenbach. Philosophic Foundations of Quantum Mechanics. University of California Press, Berkeley, CA, USA, 1944. x + 182 pp. LCCN QC174.1 .R4.

Reichenbach:1998:DSE

[Rei98a] Hans Reichenbach. Derivation of Schrödinger’s equation from de Broglie’s principle. In Philosophic foundations of quantum mechanics [Rei98b], chapter 13, pages 66–71. ISBN 0-486-40459-5 (paperback). LCCN QC174.12 .R45 1998. URL http://www.

loc.gov/catdir/description/dover032/98040052.html.

Reichenbach:1998:PFQ

[Rei98b] Hans Reichenbach. Philosophic foundations of quantum mechanics. Dover Publications, Inc., New York, NY, USA, 1998. ISBN 0-486-40459-5 (paperback). x + 182 pp. LCCN QC174.12

.R45 1998. URL http://www.loc.gov/catdir/description/
dover032/98040052.html.

Raman:1969:WWI

[RF69] V. V. Raman and Paul Forman. Why was it Schrödinger who developed de Broglie’s ideas? Historical Studies in the Physical Sciences, 1(??):291–314, ???? 1969. CODEN HSPSAS. ISSN

0073-2672. URL http://www.jstor.org/stable/27757299.

Rechenberg:1991:BBL

[RFS^{+}91] H. Rechenberg, A. Förster, W. Stamm, D. Stauffer, H. Rauch, W. Weppner, H. W. Schrötter, and W. Zinth. Buchbesprechung: Broglie, Lochak: Heisenberg’s Uncertainties and the Probabilistic Interpretation of Wave Mechanics/Spatschek: Theoretische Plasmaphysik Eine Einführung/Gaisser: Cosmic Rays and Particle Physics/Hüfner: Nur ein Störfall /Weisbuch: Complex Systems Dynamics/Audretsch u. Mainzer: Wieviele Leben hat Schrödingers Katze?/Herder Lexikon Physik/Kudo u. Fueki: Solid State Ionics/Kovetz: The Principles of Electromagnetic Theory/Brückner, Feller u. Grummt: Applications of Time-Resolved Optical Spectroscopy. Physikalische Blätter, 47(9):857– 859, September 1991. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://onlinelibrary.

wiley.com/doi/10.1002/phbl.19910470915/abstract.

Rollet:1966:CMC

[RG66] J. S. Rollet and J. Gillis. Computing methods in crystallography. Physics Today, 19(9):101–102, September 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3048436.

Rowe:1967:NLE

[RH67] Joseph E. Rowe and H. J. Hagger. Non-linear electron-wave interaction phenomena. Physics Today, June 1, 1967. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3034371.

Risley:1978:IPU

[RM78] John S. Risley and Eugen Merzbacher. International position of US physics. Physics Today, 31(7):9, July 1, 1978. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2995132.

Rousseau:1974:PO

[RMW74] M. Rousseau, J. P. Mathieu, and Dudley Williams. Problems in optics. Physics Today, April 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3128550.

Rubin:1993:WME

[RO93] Harry Rubin and Peter Olson. Walter M. Elsasser. Physics Today, 46(2):98–99, February 1, 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2808822.

Robert:2017:BWP

[Rob17] J. Robert. L. de Broglie waves: a particular solution. Annales de la Fondation Louis de Broglie, 42(1):197–212, 2017. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http:

//aflb.ensmp.fr/AFLB-421/aflb421m8.htm.

Rodgers:2019:TAS

[Rod19] Glen E. Rodgers. Travelling with the Atom: a Scientific Guide to Europe and Beyond. Royal Society of Chemistry, Cambridge, UK, 2019. ISBN 1-78801-528-2 (paperback), 1-78801-702-1 (e-book). xxxii + 551 pp. LCCN QC171.2 .R63 2020.

Rosser:1948:RPF

[Ros48] J. Barkley Rosser. Review: Principes Fondamentaux de Physique Théorique. by Jean-Louis Destouches; Louis de Broglie . Journal of Symbolic Logic, 13(4):213, 1948. CODEN JSYLA6. ISSN 0022-4812 (print), 1943-5886 (electronic). URL http://www.jstor.

org/stable/2267138.

Rosenfeld:1958:RTI

[Ros58] L. Rosenfeld. Review: Une Tentative d’interprétation causale et non linéaire de la mécanique ondulatoire (la théorie de la double solution) by Louis de Broglie. British Journal for the Philosophy of Science, 9(34):167–168, 1958. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-3537 (electronic). URL http://www.jstor.

org/stable/685156.

Rosen:1966:MF

[Ros66] Louis Rosen. Meson factories. Physics Today, 19(12):21–36, December 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3047851.

Rosaler:2015:BBT

[Ros15] Joshua Rosaler. Is de Broglie–Bohm theory specially equipped to recover classical behavior? Philosophy of Science, 82(5):1175– 1187, December 2015. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic). URL http://www.jstor.org/stable/
10.1086/683468.

Rose:1953:EFB

[RP53] M. E. Rose and C. L. Perry. The effect of the finite de Broglie wavelength in the theory of beta-decay. Physical Review (2), 90 (3):479–482, May 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Ridley:1997:EPS

[RP97] Brian K. Ridley and Peter J. Price. Electrons and phonons in semiconductor multilayers. Physics Today, 50(12):72–74, December 1, 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.2806751.

Rhodin:1975:SAS

[RT75] Thor N. Rhodin and David S. Y. Tong. Structure analysis of solid surfaces. Physics Today, October 1, 1975. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2998917.

Ruark:1969:WDW

[Rua69] Arthur E. Ruark. Where do we go from here? Physics Today, September 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3035779.

Rubinowicz:1950:PCT

[Rub50] A. Rubinowicz. Propagation of the cut-off train of de Broglie waves. Acta Physica Polonica, 10:79–88, 1950. CODEN AP-POAK. ISSN 0001-673X.

Reiter:2013:ESY

[RY13] Wolfgang Reiter and Jakob Yngvason, editors. Erwin Schrödinger — 50 Years After. European Mathematical Society Publishing House, Zürich, Switzerland, 2013. ISBN 3-03719-121-X (paperback), 3-03719-621-1 (e-book). viii + 185 pp. LCCN QC16.S265 I58 2011.

Rossi:1997:MA

[RZB97] Harald H. Rossi, Marco Zaider, and Wesley Bolch. Microdosimetry and its applications. Physics Today, 50(12):70–71, December 1, 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.882031.

S:1952:RSD

[S.52] P. M. S. Review: Savants et découvertes by Louis de Broglie. Revue Philosophique de la France et de l’Étranger, 142:249, 1952. ISSN 0035-3833 (print), 2104-385X (electronic). URL http://

www.jstor.org/stable/41087365.

Snow:1965:TCS

[SA65] C. P. Snow and C. G. Amstutz. The two cultures: And a second look. Physics Today, 18(3):68–70, March 1, 1965.

CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3047279.

Smith:1990:N

[SC90] Henry I. Smith and Harold G. Craighead. Nanofabrication. Physics Today, 43(2):24–30, February 1, 1990. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881222.

Sjoqvist:1997:GPQ

[SC97] Erik Sjöqvist and Henrik Carlsen. Geometric phase, quantum measurements, and the de Broglie–Bohm model. Physical Review A (Atomic, Molecular, and Optical Physics), 56(??):1638, ???? 1997. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Sedov:1965:TDP

[SCC^{+}65] L. I. Sedov, C. K. Chu, H. Cohen, B. Seckler, and J. Gillis. Two dimensional problems in hydrodynamics and aerodynamics. Physics Today, December 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3047038.

Schilpp:1949:AEPa

[Sch49a] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 1 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951, 1969, and 1982.

Schilpp:1949:AEPb

[Sch49b] Paul Arthur Schilpp, editor. Albert Einstein: Philosopher-Scientist, volume 2 of The Library of Living Philosophers. Cambridge University Press, Cambridge, UK, 1949. ISBN 0-87548-286-4. ISSN 0075-9139. xviii + 781 pp. LCCN QC16.E5 S3 1970. Reprinted 1951 and 1970.

Schrodinger:1955:BWG

[Sch55a] Erwin Schrödinger. Die Bedeutung der Wellenmechanik. (German) [The meaning of wave mechanics]. In Louis de Broglie und die Physiker. (German) [Louis de Broglie and the Physicist], pages 18–25. Claassen, Hamburg, Germany, 1955.

Schrodinger:1955:SMO

[Sch55b] Erwin Schrödinger. La signification de la mécanique ondulatoire. (French) [The meaning of wave mechanics]. In Louis de Broglie, Physicien et Penseur. (French) [Louis de Broglie, Physicist and Thinker], pages 16–32. Albin Michel, Paris, France, 1955.

Schwarzschild:1980:LLI

[Sch80] Bertram M. Schwarzschild. Laue–Langevin Institute expands. Physics Today, 33(6):21–22, June 1, 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2914114.

Schwarzschild:1982:SSN

[Sch82] Bertram M. Schwarzschild. Spallation sources for neutron scattering. Physics Today, 35(11):19–21, November 1, 1982. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2914843.

Schwarzschild:1984:NSW

[Sch84] Bertram M. Schwarzschild. Nuclear spin waves seen in dilute polarized gases. Physics Today, 37(6):19–21, June 1, 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2916268.

Schopf:1985:BRB

[Sch85a] H.-G. Schöpf. Book review: Barut, A. O. /van der Merwe, A. /Vigier, J.-P. (eds.), Quantum, Space and Time — The Quest Continues. Studies and Essays in Honour of Louis de Broglie, Paul Dirac and Eugene Wigner. Cambridge et al., Cambridge University Press 1984. Vii, 662 S., £25.00 A P/b. US $49.50. ISBN 0-521-31911-0. (Cambridge Monographs on Physics). Zeitschrift für Angewandte Mathematik und Mechanik, 65(10), 1985. CO-DEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (elec-

tronic). URL http://onlinelibrary.wiley.com/doi/10.1002/
zamm.19850651014/abstract.

Schwarzschild:1985:ABS

[Sch85b] Bertram Schwarzschild. Atomic beams stopped by laser light, and trapped. Physics Today, 38(6):17–21, June 1, 1985. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2814588.

Schwarzschild:1985:SRI

[Sch85c] Bertram Schwarzschild. Storage ring intensifies Los Alamos neutron source. Physics Today, 38(6):21, June 1, 1985. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2814589.

Schwarzschild:1986:LBF

[Sch86a] Bertram Schwarzschild. Laser beam focus forms optical trap for neutral atoms. Physics Today, 39(9):17–19, September 1, 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2815137.

Schwarzschild:1986:NTE

[Sch86b] Bertram Schwarzschild. Now they’re even trapping antiprotons. Physics Today, 39(9):19, September 1, 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2815138.

Schewe:1987:PN

[Sch87a] Phillip F. Schewe. Physics news in 1986. Physics Today, January 1, 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2815287.

Schwarzschild:1987:PNP

[Sch87b] Bertram Schwarzschild. Physics Nobel Prize Awarded for microscopies old and new. Physics Today, 40(1):17–21, January 1, 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.2815291.

Schewe:1988:SSP

[Sch88a] Phillip F. Schewe. Special supplement: Physics news in 1987. Physics Today, January 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2811267.

Schwarzschild:1988:UEC

[Sch88b] Bertram Schwarzschild. Are the ultra-energetic cosmic gammas really photons? Physics Today, 41(11):17–21, November 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2811616.

Schewe:1999:PU

[Sch99] Phillip F. Schewe. Physics update. Physics Today, 52(12):9, December 1, 1999. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2802844.

Schwarzschild:2002:UNE

[Sch02] Bertram Schwarzschild. Ultracold neutrons exhibit quantum states in the Earth’s gravitational field. Physics Today, 55(3): 20–, 22–23, March 1, 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/full/10.1063/1.1472382.

Schewe:2003:PBW

[Sch03] Phil F. Schewe. The photonic de Broglie wavelength. Physics Today, 56(1):9, January 1, 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.4796891.

Schaerf:2005:PGR

[Sch05] Carlo Schaerf. Polarized gamma-ray beams. Physics Today, 58 (8):44–50, August 1, 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/full/10.1063/1.2062918.

Schewe:2007:GAL

[Sch07a] Phillip F. Schewe. Guided atom laser. Physics Today, 60(2):22–23, February 1, 2007. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/1.4796328.

Schucking:2007:BRE

[Sch07b] Engelbert L. Schucking. Book review: Einstein: His Life and Universe, Walter Isaacson, Simon & Schuster, New York, 2007. $32.00 (675 pp.). ISBN 978-0-7432-6473-0. Physics Today, 60(11):59–61, November 1, 2007. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/1.2812127.

Schwarzschild:2013:EHA

[Sch13] Bertram M. Schwarzschild. Englert and Higgs are awarded the Nobel Prize in Physics. Physics Today, 66(12):10–12, December 1, 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/PT.3.2196.

Smoot:1994:WTT

[SDBP94] George Smoot, Keay Davidson, Marcia Bartusiak, and Joel R. Primack. Wrinkles in time and through a universe darkly. Physics Today, 47(9):90–91, September 1, 1994. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2808636.

Selleri:1993:FQQ

[Sel93] Franco Selleri, editor. Fundamental questions in quantum physics and relativity. Collected papers in honor of Louis de Broglie. Hadronic Press, Inc.1, Palm Harbor, FL, USA, 1993. ISBN 0-911767-63-0 (paperback). 184 pp.

Sergescu:1958:TDP

[Ser58] P. Sergescu. Review: L’évolution de la notion de phénomène physique des primitifs à Bohr et Louis de Broglie by J. Pelseneer. Revue philosophique de la France et de l’Étranger, 148:102–103, 1958. ISSN 0035-3833 (print), 2104-385X (electronic). URL http:

//www.jstor.org/stable/41089669.

Schlegel:1978:WDM

[SG78] Richard Schlegel and Richard K. Gehrenbeck. Who discovered matter waves? Physics Today, 31(7):9–13, July 1, 1978. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2995133.

Sakharov:1969:PCI

[SH69] Andrei D. Sakharov and Jack M. Hollander. Progress, coexistence and intellectual freedom. Physics Today, May 1, 1969.

CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3035592.

Sharma:1974:SNB

[Sha74] Jagadish Sharma. Satyendra Nath Bose. Physics Today, April 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3128568.

Shimony:1985:BRH

[Shi85] Abner Shimony. Book review: Henry Folse, The Philosophy of Niels Bohr: The Framework of Complementarity. Physics Today, 38(10):108–109, October 1, 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2814739.

Shostak:2000:JH

[Sho00] Seth Shostak. Jean Heidmann. Physics Today, December 1, 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.1341931.

Szczeniowski:1931:ICEa

[SI31] S. Szczeniowski and L. Infeld. The influence of a cloud of electrons on the structure of de Broglie waves. Bulletin International de l’Académie Polonaise des Sciences et des Lettres. Classe des sciences mathématiques et naturelles. Série A: Sciences mathématiques, 1931(6):482–488, 1931. CODEN BIAPAC. ISSN 0366-0303.

Szczeniowski:1932:WCE

[SI32] S. E. Szczeniowski and Leopold Infeld. O wpływie chmury elektronowej na strukture fali de Broglie’a. (Polish). [On the effect of the electron cloud on de Broglie’s wave structure]. Acta Physica Polonica, 1(1–2):37–46, ???? 1932. CODEN APPOAK. ISSN 0001-673X.

Sievers:1997:LBQ

[Sie97] Henning Sievers. Louis de Broglie und die Quantenmechanik. (German) [Louis de Broglie and quantum mechanics]. arxiv.org, ??(??):1–87, December 17, 1997. URL https://arxiv.org/abs/

physics/9807012.

Sivashinsky:2011:BSL

[Siv11] Gregory I. Sivashinsky. The de Broglie soliton as a localized excitation of the action function. Physica. D, Nonlinear phenomena, 240(4–5):406–409, 2011. CODEN PDNPDT. ISSN 0167-2789 (print), 1872-8022 (electronic).

Slater:1974:QCC

[SJ74] John C. Slater and Keith H. Johnson. Quantum chemistry and catalysis. Physics Today, 27(10):34–41, October 1, 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3128915.

Sachdev:2011:QC

[SK11] Subir Sachdev and Bernhard Keimer. Quantum criticality. Physics Today, 64(2):29–35, February 1, 2011. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://physicstoday.scitation.org/doi/full/10.1063/1.
3554314.

Sarkar:2015:ACS

[SKFS15] Resham Sarkar, May E. Kim, Renpeng Fang, and Selim M. Shahriar. N-atom collective-state atomic interferometer with ultrahigh Compton frequency and ultrashort de Broglie wavelength, with \sqrt{N} reduction in fringe width. Physical Review A (Atomic, Molecular, and Optical Physics), 92(??):063612, ???? 2015. CO-DEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446,

1538-4519. URL http://www.w3.org/1998/Math/MathML.

Sakurai:1965:IPE

[SL65] J. J. Sakurai and D. B. Lichtenberg. Invariance principles and elementary particles. Physics Today, March 1, 1965. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3047277.

Stanley:1969:LSE

[SL69] R. C. Stanley and Robert Lindsay. Light and sound for engineers. Physics Today, 22(12):75, December 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3035316.

Snoke:2010:PC

[SL10] David Snoke and Peter Littlewood. Polariton condensates. Physics Today, August 1, 2010. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.3480075.

Slater:1968:QPA

[Sla68] John C. Slater. Quantum physics in America between the wars. Physics Today, 21(1):43–51, January 1, 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3034740.

Seymour:1969:WCS

[SLF69] S. Frederick Seymour, Morton Leeds, and M. W. Friedlander. Washington Colloquium on Science and Society (First Series) and Washington Colloquium on Science and Society (Second Series). Physics Today, May 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3035600.

Sun:2006:OFP

[SLH^{+}06] F. W. Sun, B. H. Liu, Y. F. Huang, Z. Y. Ou, and G. C. Guo. Observation of the four-photon de Broglie wavelength by stateprojection measurement. Physical Review A (Atomic, Molecular, and Optical Physics), 74(??):033812, ???? 2006. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Slichter:1976:FWL

[Sli76] Charles P. Slichter. F. Wheeler Loomis. Physics Today, 29(5): 83–85, May 1, 1976. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3023484.

Shapley:1969:BO

[SM69] Harlow Shapley and Martin F. McCarthy. Beyond the observatory. Physics Today, April 1, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3035499.

Smart:2010:QET

[Sma10] Ashley G. Smart. Quantum effects turn pore diffusion on its head. Physics Today, October 1, 2010. CODEN PHTOAD.

ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.4797248.

Smolin:2019:CRB

[Smo19a] Lee Smolin. The challenge of realism: de Broglie and Einstein. In Einstein’s Unfinished Revolution: the Search for What Lies Beyond the Quantum [Smo19b], chapter 7, pages 97–106. ISBN 1-59420-619-8 (hardcover), 0-698-16913-1 (ebook). LCCN QC174.13 .S6545 2019. URL https://www.

penguinrandomhouse.com/books/316818/einsteins-unfinished-
revolution-by-lee-smolin/9781594206191. Illustrations by

Kaca Bradonjic.

Smolin:2019:EUR

[Smo19b] Lee Smolin. Einstein’s Unfinished Revolution: the Search for What Lies Beyond the Quantum. Penguin Press, New York, NY, USA, 2019. ISBN 1-59420-619-8 (hardcover), 0-698-16913-1 (e-book). xxix + 322 pp. LCCN QC174.13 .S6545

2019. URL https://www.penguinrandomhouse.com/books/
316818/einsteins-unfinished-revolution-by-lee-smolin/

9781594206191. Illustrations by Kaca Bradonjic.

Smyk:2015:SSL

[Smy15] A. Smyk. Scientific school of Louis de Broglie and its role in the development of theoretical physics. Annales de la Fondation Louis de Broglie, 40(1):139–148, 2015. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://aflb.ensmp.

fr/AFLB-401/aflb401m819.htm.

Shafrir:1965:CET

[SNG65] Uri Shafrir, Morris Neiburger, and J. Gillis. Collision efficiencies of two spheres falling in a viscous medium for Reynolds numbers up to 19.2. Physics Today, 18(2):62, February 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3047195.

Snow:2013:EPS

[Sno13] W. Michael Snow. Exotic physics with slow neutrons. Physics Today, 66(3):50–55, March 1, 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/PT.3.1918.

Sun:2006:PMM

[SOG06] F. W. Sun, Z. Y. Ou, and G. C. Guo. Projection measurement of the maximally entangled N-photon state for a demonstration of the N-photon de Broglie wavelength. Physical Review A (Atomic, Molecular, and Optical Physics), 73(??):023808, ???? 2006. CO-DEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446,

1538-4519. URL http://www.w3.org/1998/Math/MathML.

Sommerfeld:1911:PRC

[Som11] Arnold Sommerfeld. Projet de rapport concernant les discussions du congrès de Bruxelles de M. de Broglie au Professeur Langevin. Lettre du 23 décembre 1911. (French) [draft report on the discussions of the Brussels from M. De Broglie to Professor Langevin. Letter of 23 December 1911]. Letter., 1911.

Sommerfeld:1934:LEM

[Som34] Arnold Sommerfeld. Literaturberichte: L’électron magnétique. (German) [Literature report: the magnetic electron]. Monatshefte für Mathematik und Physik, 41(1):A24, December 1934. CODEN

MMPHA8. ISSN 1812-8076. URL http://link.springer.com/

article/10.1007/BF01697933. L. de Broglie, (Théorie de Dirac). 315 S. Hermann et Cie., Paris 1934. Preis Frs. 100.

Sato:2012:SHI

[SP12] Yuki Sato and Richard Packard. Superfluid helium interferometers. Physics Today, 65(10):31–36, October 1, 2012. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/PT.3.1749.

Speziali:1968:AOL

[Spe68] Pierre Speziali. Analyses d’ouvrages: Louis de Broglie et la mécanique ondulatoire, (Savants du monde entier) par Marie-Antoinette Tonnelat. Revue d’Histoire des Sciences et de Leurs Applications, 21(2):194–196, avril–juin 1968. CODEN RHSAAM. ISSN 0048-7996 (print), 1969-6582 (electronic). URL http://

www.jstor.org/stable/23904489.

Squires:1992:ACB

[Squ92] Euan J. Squires. An apparent conflict between the de Broglie– Bohm model and orthodoxy in quantum cosmology. Foundations of Physics Letters, 5(1):71–75, February 1992. CODEN FPLEET. ISSN 0894-9875 (print), 1572-9524 (electronic).

Siong:2013:BBI

[SR13] Ch’ng Han Siong and Shahidan Radiman. de Broglie–Bohm interpretation of canonical quantum cosmology implies early stages of the universe dominated by radiation. Commun. Theor. Phys., 60 (6):691–694, 2013. ISSN 0253-6102 (print), 1572-9494 (electronic).

Sebera:1966:ESC

[SS66] Donald K. Sebera and M. E. Straumanis. Electronic structure and chemical bonding. Physics Today, 19(4):89–92, April 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3048194.

Searl:1992:JEE

[SS92] Milton Searl and Chauncey Starr. Japan: Not an energy efficiency model. Physics Today, 45(2):95–97, February 1, 1992. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2809538.

Stein:1995:PU

[SS95] Benjamin P. Stein and Phillip F. Schewe. Physics update. Physics Today, 48(8):9, August 1, 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2808115.

Shojai:2000:BBQ

[SS00] A. Shojai and F. Shojai. de Broglie–Bohm quantum theory and perihelion precession. Annales de la Fondation Louis de Broglie, 25(3–4):293–302, 2000. CODEN ???? ISSN 0182-4295 (print),

2108-6397 (electronic). URL http://aflb.ensmp.fr/AFLB-253/
aflb253p293.htm.

Shojai:2001:ASP

[SS01] Ali Shojai and Fatimah Shojai. About some problems raised by the relativistic form of de-Broglie–Bohm theory of pilot wave. Physica Scripta, 64(5):413–416, 2001. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

Schewe:1996:PU

[SSB96] Phillip F. Schewe, Benjamin P. Stein, and Stephen G. Benka. Physics update. Physics Today, 49(7):9, July 1, 1996. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2807678.

Steane:1995:PMA

[SSDD95] A. Steane, P. Szriftgiser, P. Desbiolles, and J. Dalibard. Phase modulation of atomic de Broglie waves. Physical Review Letters, 74(??):4972, ???? 1995. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Schulman:1998:TAQ

[SSL98] Lawrence S. Schulman, Steven F. Savitt, and Joel L. Lebowitz. Time’s arrows and quantum measurement and time’s arrows today: Recent physical and philosophical work on the direction of time. Physics Today, 51(10):79–81, October 1, 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.882412.

Schleich:1991:JSW

[SSP^{+}91] Wolfgang P. Schleich, Georg Süssmann, John Philpott, Malvin C. Teich, and Bahaa E. A. Saleh. A jump shot at the Wigner distribution. Physics Today, October 1, 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2810308.

Steck:1935:MHB

[Ste35] Max Steck. Die Massenvariabilität aus der Hypothese von de Broglie. (German) [Mass variability from the hypothesis of de Broglie]. Zeitschrift für Physik, 93:634–635, 1935. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831 (electronic).

Stern:1997:RPP

[Ste97] David P. Stern. Reprint pact proposed for saving time, effort when reusing figures. Physics Today, 50(2):11–13, February 1, 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.881667.

Steuernagel:2002:BWR

[Ste02] O. Steuernagel. de Broglie wavelength reduction for a multiphoton wave packet. Physical Review A (Atomic, Molecular, and Optical Physics), 65(??):033820, ???? 2002. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Stein:2003:SID

[Ste03] Benjamin P. Stein. A strongly interacting degenerate Fermi gas. Physics Today, 56(3):9, March 1, 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.2409951.

Stone:2005:EUI

[Sto05] A. Douglas Stone. Einstein’s unknown insight and the problem of quantizing chaos. Physics Today, 58(8):37–43, August 1, 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/1.2062917.

Straumanis:1965:BRH

[Str65] M. E. Straumanis. Book review: H. Hartmann, Die chemische Bindung, 2nd edition, 109 pp. Springer–Verlag, Berlin, 1964. (German) [Chemical binding]. Physics Today, 18(2):62–63, February 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3047196.

Straumanis:1966:BRK

[Str66] M. E. Straumanis. Book review: Klaus Schäfer, Physikalische Chemie: Ein Vorlesungskurs, second edition. 432 pp. Springer– Verlag, Berlin and New York, 1965. Physics Today, 19(1):128, January 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3047927.

Stumpf:2003:SPP

[Stu03] H. Stumpf. Symmetry properties of photon eigenstates of generalized de Broglie–Bargmann–Wigner equations. Annales de la Fondation Louis de Broglie, 28(1):65–76, 2003. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://aflb.

ensmp.fr/AFLB-281/aflb281p065.htm.

Stumpf:2004:GBB

[Stu04] H. Stumpf. Generalized de Broglie–Bargmann–Wigner equations, a modern formulation of de Broglie’s fusion theory. Annales de la Fondation Louis de Broglie, 29(??):785–806, 2004. CODEN ????

ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://

aflb.ensmp.fr/AFLB-295/aflb295m307.htm.

Summers:1988:BRG

[Sum88] Stephen J. Summers. Book review: Günther Ludwig, An Axiomatic Basis for Quantum Mechanics Volume 1: Derivation of Hilbert Space Structure and Volume 2: Quantum Mechanics and Macrosystems. Physics Today, 41(8):72–74, August 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2811530.

Surdin:1982:BWR

[Sur82] M. Surdin. De Broglie’s wave revisited. Foundations of Physics, 12 (9):873–888, September 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://link.springer.

com/article/10.1007/BF01884998.

Sutherland:1997:PSG

[Sut97] Roderick I. Sutherland. Phase space generalization of the de Broglie–Bohm model. Foundations of Physics, 27(6):845–863, June 1997. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF02550344.

Suvorov:1967:VBV

[Suv67] S. G. Suvorov. Vklad de Brojlja v razvitie fiziki (k publikacii perevodov statej de Brojlja) (k 75-letiju lui de brojlja) (Russian) [de Broglie’s contribution to the development of physics (publication of translations of de Broglie’s articles) (on the 75th birthday of Louis de Broglie)]. Physics-Uspekhi, 93(??):177, September 1967. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-

4780 (electronic). URL http://ufn.ru/ru/authors/11960/de-
broglie-louis/.

Smart:1966:EFT

[SV66] J. Samuel Smart and J. H. Van Vleck. Effective field theories of magnetism. Physics Today, 19(8):77–78, August 1, 1966. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3048415.

Struyve:2009:BBG

[SV09] Ward Struyve and Antony Valentini. de Broglie–Bohm guidance equations for arbitrary Hamiltonians. Journal of Physics A: Mathematical and Theoretical, 42(3):18, 2009. CODEN JPAMB5. ISSN 1751-8113 (print), 1751-8121 (electronic).

Schapira:1979:MHR

[SW79] Morey Schapira and Lincoln Wolfenstein. More on human rights. Physics Today, 32(8):13–15, August 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2995660.

Swann:1939:RML

[Swa39] W. F. G. Swann. Review: Matter and Light by Louis de Broglie. Science, 90(2338):373, October 20, 1939. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1666482.

Synge:1954:GMB

[Syn54] J. L. Synge. Geometrical mechanics and de Broglie waves. Cambridge monographs on mechanics and applied mathematics. Cambridge University Press, Cambridge, UK, 1954. viii + 167 + 54 pp.

Synge:2010:GMB

[Syn10] J. L. Synge. Geometrical mechanics and de Broglie waves. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-15688-2 (paperback). vii + 167 pp. Reprint of the 1954 hardback edition.

Takeuehi:1927:TSR

[Tak27] T. Takeuehi. Time in special relativity and de Broglie’s phase velocity. Proc. Phys.-Math. Soc. Japan, III. Ser., 9:81, 1927. ISSN 0370-1239 (print), 2185-2707 (electronic).

Tappert:1938:EIB

[Tap38] J. G. Tappert. An experimental investigation of de Broglie’s equation. Physical Review (2), 54(??):1085–1088, ???? 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Tappenden:2022:PWT

[Tap22] Paul Tappenden. Pilot-wave theory without nonlocality. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4.

ISSN 0015-9018 (print), 1572-9516 (electronic). URL https://

link.springer.com/article/10.1007/s10701-022-00627-0.

Tarter:2000:REB

[Tar00] C. Bruce Tarter. Roger Elwood Batzel. Physics Today, December 1, 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.1341930.

Taylor:1991:EWH

[Tay91] Kelly Taylor. Ease the way to hiring foreigners. Physics Today, 44(10):146, October 1, 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2810307.

Tilbi:2007:KCP

[TBMC07] A. Tilbi, T. Boudjedaa, M. Merad, and L. Chetouani. The Kanai– Caldirola propagator in the de Broglie–Bohm theory. Physica Scripta, 75(4):474–479, 2007. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

Terzian:2005:AP

[TBMW05] Joseph E. Terzian, Charles H. Bennett, Ady Mann, and William K. Wootters. Asher Peres. Physics Today, 58(8):65– 66, August 1, 2005. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/full/10.1063/1.2062925.

Terletsky:1971:PMS

[TdB71] Jacob Terletsky and Louis de Broglie. Sur la possibilité de mesurer simultanément les coordonnées et l’impulsion d’une particule pour un événement passé. (French) [On the possibility of simultaneously measuring the coordinates and the momentum of a particle for a past event]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 272(??):1161–1163, May

17, 1971. ISSN 0366-6077. URL http://gallica.bnf.fr/ark:/
12148/bpt6k480293k/f1185/.

Thackray:1993:HST

[Tha93] Barbara M. Thackray. High school teachers need college contact. Physics Today, 46(4):91–92, April 1, 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2808890.

Thomson:1967:SE

[Tho67] Sir George Thomson. The septuagenarian electron. Physics Today, May 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3034306.

Thomas:2010:NPF

[Tho10] John E. Thomas. The nearly perfect Fermi gas. Physics Today, 63(5):34–37, May 1, 2010. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/full/10.1063/1.3431329.

Tisza:1965:BRT

[Tis65] Laszlo Tisza. Book review: La Thermodynamique de la particule isolée (ou Thermodynamique cachée des particules). Louis de Broglie. Gauthier-Villars, Paris, 1964. vi + 125 pp. Illus. Paper, $7.50. Science, 148(3668):362, April 16, 1965. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.jstor.org/stable/1715087.

Turner:1989:TUC

[TJLB89] Leaf Turner, Robert Jones, Yale Jay Lubkin, and Saul Birnbaum. Thoughts unbottled by cold fusion. Physics Today, 42(9):140–144, September 1, 1989. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2811168.

Taylor:1968:IGS

[TM68] A. M. Taylor and L. Marton. Imagination and the growth of science. Physics Today, June 1, 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3035027.

Tonolo:1956:IEE

[Ton56] Angelo Tonolo. Sur les intégrales des équations de l’électromagnétisme de M. L. de Broglie. I, II. (French) [On the integrals of the electromagnetism equations of M. L. de Broglie. I, II]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 242(??): 2626–2629, 2699–2702, ???? 1956. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Tonomura:1990:EHN

[Ton90] Akira Tonomura. Electron holography: A new view of the microscopic. Physics Today, 43(4):22–29, April 1, 1990. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881230.

Torretti:1987:BRA

[Tor87] Roberto Torretti. Book review: Asim O. Barut, Alwyn van der Merwe, and Jean-Pierre Vigier, editors, Quantum, Space, and Time — The Quest Continues, Studies and Essays in Honour of Louis de Broglie, Paul Dirac, and Eugene Wigner. Noûs, 21 (3):442–444, September 1987. CODEN ???? ISSN 0029-4624 (print), 1468-0068 (electronic). URL http://www.jstor.org/

stable/2215196.

Tellez-Plasencia:1943:GDP

[TP43] H. Tellez-Plasencia. Généralisation du principe de focalisation rotatoire de Bragg–de Broglie pour la spectrographie Röntgen. Application aux spectrographes à deux cristaux. Focalisation et structure mosaique. (French) [Generalization of the Bragg–de Broglie rotatory focusing principle for Röntgen spectroscopy. Application to two-crystal spectrographs. Focal and mosaic structure]. Arch. Sci. Phys. etc., 25(??):35–54, 1943.

Taylor:1976:GTW

[TQ76] J. C. Taylor and Helen R. Quinn. Gauge theories of weak interactions. Physics Today, 29(11):76–78, November 1, 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3024523.

Tsong:2006:FYS

[Tso06] Tien T. Tsong. Fifty years of seeing atoms. Physics Today, March 1, 2006. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
full/10.1063/1.2195313.

Turlapov:2005:TLE

[TTS05] Andrey Turlapov, Alexei Tonyushkin, and Tycho Sleator. Talbot– Lau effect for atomic de Broglie waves manipulated with light. Physical Review A (Atomic, Molecular, and Optical Physics), 71 (??):043612, ???? 2005. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Topinka:2003:IEF

[TWH03] Mark A. Topinka, Robert M. Westervelt, and Eric J. Heller. Imaging electron flow. Physics Today, December 1, 2003. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/1.1650228.

Utsuro:2002:ASB

[Uts02] M. Utsuro. Analytical solution for the de Broglie wave packet description of the neutron in subcritical transmission through a mirror. Physics Letters A, 292(4–5):222–232, 2002. CODEN PY-LAAG. ISSN 0375-9601 (print), 1873-2429 (electronic).

Vallee:1990:LBC

[Val90] Robert Vallée. Louis de Broglie and cybernetics. Kybernetes, 19 (2):32–33, 1990. ISSN 0368-492X.

Valentini::1992:PWT

[Val92] Antony Valentini. On the Pilot-Wave Theory of Classical, Quantum and Subquantum Physics. Ph.D. thesis, ISAS, Trieste, Italy, 1992.

VanVleck:1964:APC

[Van64] J. H. Van Vleck. American physics comes of age. Physics Today, 17(6):21–26, June 1, 1964. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.3051643.

VanHorn:1979:PWD

[Van79] Hugh M. Van Horn. The physics of white dwarfs. Physics Today, 32(1):23–32, January 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2995367.

Vandyck:1994:DHO

[Van94] M. A. Vandyck. On the damped harmonic oscillator in the de Broglie–Bohm hidden-variable theory. Journal of Physics A (Mathematical and General), 27(5):1743–1750, 1994. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Vankov:2005:BWN

[Van05] A. Vankov. On de Broglie wave nature. Annales de la Fondation Louis de Broglie, 30(1):15–34, 2005. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic). URL http://aflb.ensmp.

fr/AFLB-301/aflb301m176.htm.

Varlamov:2002:AAF

[Var02] V. V. Varlamov. About algebraic foundations of Majorana– Oppenheimer quantum electrodynamics and de Broglie–Jordan neutrino theory of light. Annales de la Fondation Louis de Broglie, 27(2–3):273–286, 2002. CODEN ???? ISSN 0182-4295 (print),

2108-6397 (electronic). URL http://aflb.ensmp.fr/AFLB-272/
aflb272p273.htm.

vanderMerwe:1982:EPE

[vdM82] Alwyn van der Merwe. Editorial postscript to “The evolution of the ideas of Louis de Broglie on the interpretation of wave mechanics”. Foundations of Physics, 12(10):955–962, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/

BF01889268. See [Loc82].

vonFoerster:1982:NMD

[vF82] Thomas von Foerster. New method for determining the phases of diffracted x rays. Physics Today, November 1, 1982. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2914842.

Vigano:1970:RNP

[Vig70] Mario Viganò. Review: Nuove prospettive in microfisica. (Collana di Filosofi Contemporanei) by Louis de Broglie; Filippo Selvaggi. Gregorianum, 51(2):393–394, 1970. ISSN 0017-4114. URL http:

//www.jstor.org/stable/23574695.

Vigier:1971:MDP

[Vig71] Jean-Pierre Vigier. Masse du photon et expériences de Kunz sur l’effet photoélectrique du plomb supraconducteur. (French) [Mass of the photon and Kunz’s experiments on the photoelectric effect in superconducting lead]. Comptes Rendus des Séances de l’Académie des Sciences, Série B: Sciences Physiques, 273(23):

993–996, December 8, 1971. ISSN 0366-6077. URL http://

gallica.bnf.fr/ark:/12148/bpt6k56192021/f87/. Presented

by Louis de Broglie.

Vigier:1982:LBP

[Vig82] Jean-Pierre Vigier. Louis de Broglie — physicist and thinker. Foundations of Physics, 12(10):923–930, October 1982. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF01889266.

Vigier:1991:EMC

[Vig91] Jean-Pierre Vigier. Explicit mathematical construction of relativistic nonlinear de Broglie waves described by three-dimensional (wave and electromagnetic) solitons “piloted” (controlled) by corresponding solutions of associated linear Klein–Gordon and Schrödinger equations. Foundations of Physics, 21(2):125–148, February 1991. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF01889529.

Vigier:1993:DNE

[Vig93] J. P. Vigier. From Descartes and Newton to Einstein and de Broglie. Foundations of Physics, 23(1):1–4, January 1993. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF01883986.

Vigier:1994:PTR

[Vig94] J. P. Vigier. Possible test of the reality of superluminal phase waves and particle phase space motions in the Einstein–de Broglie–Bohm causal stochastic interpretation of quantum mechanics. Foundations of Physics, 24(1):61–83, January 1994. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (elec-

tronic). URL http://link.springer.com/article/10.1007/
BF02053908.

Vigier:1995:DIF

[Vig95] Jean-Pierre Vigier. Derivation of inertial forces from the Einstein– de Broglie–Bohm (E.d.B.B.) causal stochastic interpretation of quantum mechanics. Foundations of Physics, 25(10):1461–1494, October 1995. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL http://link.springer.com/article/
10.1007/BF02057462.

Vink:2022:SCE

[Vin22] Jeroen C. Vink. Spin and contextuality in extended de Broglie– Bohm–Bell quantum mechanics. Foundations of Physics, 52(5):??, October 2022. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-

9516 (electronic). URL https://link.springer.com/article/
10.1007/s10701-022-00584-8.

Visconti:1951:TDC

[Vis51] Antoine Visconti. Sur la théorie du champ soustractif de Louis de Broglie et la fonction d’interaction de R. P. Feynman. (French) [On the subtractive field theory of Louis de Broglie and the interaction function of R. P. Feynman]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 232(??):217– 219, ???? 1951. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

vonMeyenn:2001:WP

[vMS01] Karl von Meyenn and Engelbert Schucking. Wolfgang Pauli. Physics Today, 54(2):43–48, February 1, 2001. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://physicstoday.scitation.org/doi/full/10.1063/1.
1359709.

Vrkljan:1951:ETP

[Vrk51] V. S. Vrkljan. Existe-t-il dans la théorie de la particule à spin 1 de M. de Broglie une grandeur analogue au vecteur de Hertz?. (French) [Does there exist in the theory of the spin 1 particle of M. De Broglie a magnitude analogous to the vector of Hertz?]. Comptes Rendus Hebdomadaires des Séances de l’Académie des Sciences, 232(??):1908–1910, ???? 1951. CODEN COREAF. ISSN 0001-4036 (print), 2419-6304 (electronic).

Vesely:2003:LBE

[VV03] S. L. Vesely and A. A. Vesely. Louis De Broglie’s experiment. arxiv.org, ??(??):1–30, March 8, 2003. URL https://arxiv.org/

abs/physics/0303033.

Vila-Valls:2014:LBD

[VV14] Adrien Vila-Valls. Louis de Broglie et la diffusion de la mécanique quantique en France (1925–1960). (French) [Louis de Broglie and the spread of quantum mechanics in France (1925–1960)]. docteur en Epistémologie et Histoire des Sciences et des Techniques, Université Claude Bernard, Lyon I, France, May 19, 2014. 297

pp. URL https://tel.archives-ouvertes.fr/tel-00993036/
document.

Vysin:1971:SMT

[Vys71] Vratislav Vysin. Still more on tachyons. Physics Today, 24(3):14– 65, March 1, 1971. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.3022615.

Waite:1995:SVL

[Wai95] T. Waite. Smooth vortex-like de Broglie particle/waves. Annales de la Fondation Louis de Broglie, 20(4):427–471, 1995. CODEN ???? ISSN 0182-4295 (print), 2108-6397 (electronic).

Wali:2006:MBB

[Wal06] Kameshwar Wali. The man behind Bose statistics. Physics Today, 59(10):46–52, October 1, 2006. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/1.2387088.

Wang:2009:BBI

[Wan09] Bo-Bo Wang. de Broglie–Bohm interpretation for wave function of a toroidal black hole. Gen. Relativ. Gravitation, 41(5):1181–1193, 2009. ISSN 0001-7701 (print), 1572-9532 (electronic).

Wigner:1930:BBL

[WBBE30] E. Wigner, J. Bartels, A. Betz, and S. Erk. Besprechungen: de Broglie, Louis, Einführung in die Wellenmechanik (German) [Book reviews: de Broglie, Louis, Introduction to Wave Mechanics]. Die Naturwissenschaften, 18(14):307–310, ???? 1930. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/m5071k1380584921/.

Woolard:1967:SA

[WCÖ67] Edgar W. Woolard, Gerald M. Clemence, and Ernst J. Öpik. Spherical astronomy. Physics Today, 20(6):73–74, June 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3034368.

Weinberg:2006:VEM

[Wei06a] Steven Weinberg. The value of Einstein’s mistakes. Physics Today, 59(4):15–16, April 1, 2006. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/full/10.1063/1.4797363.

Weinberger:2006:RLB

[Wei06b] P. Weinberger. Revisiting Louis de Broglie’s famous 1924 paper in the Philosophical Magazine. Philosophical Magazine Letters, 86 (7):405–410, 2006. CODEN PMLEEG. ISSN 0950-0839 (print), 1362-3036 (electronic).

Weinberg:2011:PPR

[Wei11] Steven Weinberg. Particle physics, from Rutherford to the LHC. Physics Today, 64(8):29–32, August 1, 2011. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/PT.3.1216.

Weinberger:2014:NBD

[Wei14] P. Weinberger. Niels Bohr and the dawn of quantum theory. Philosophical Magazine, 94(27):3072–3087, August 2014. CODEN PH-MAA4. ISSN 0031-8086.

Wertheim:1965:BRA

[Wer65] G. K. Wertheim. Book review: A. Abragam, L’Effet Mössbauer et ses Applications à l’Étude des Champs internes. (French)

[The] Mössbauer effect and its applications to the study of internal fields. Physics Today, 18(2):60–62, February 1, 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.3047193.

Werner:1980:NI

[Wer80] Samuel A. Werner. Neutron interferometry. Physics Today, 33(12):24–30, December 1, 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2913855.

Ward:1967:RAA

[WG67] W. Dixon Ward and S. A. Goudsmit. Reviewer and author anonymity. Physics Today, January 1, 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3034118.

Wilde:1979:ME

[WG79] S. E. Wilde and Robert H. Good. More on ERA. Physics Today, 32(10):13–15, October 1, 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.2995232.

Wheeler:1985:NBM

[Whe85] John Archibald Wheeler. Niels Bohr, the man. Physics Today, 38 (10):66–72, October 1, 1985. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.881017.

White:1962:RNP

[Whi62] Frederick A. White. Review: -New Perspectives in Physics (Where Does Physical Theory Stand Today?) by Louis de Broglie; A. J. Pomerans. American Scientist, 50(4):422A–424A, 1962. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL

http://www.jstor.org/stable/27838556.

Whitaker:1995:BBB

[Whi95] M. A. B. Whitaker. Bethe, Bell and de Broglie. American Journal of Physics, 63(10):873, 1995. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Wignall:1985:BWN

[Wig85] J. W. G. Wignall. De Broglie waves and the nature of mass. Foundations of Physics, 15(2):207–227, February 1985. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL

http://link.springer.com/article/10.1007/BF00735293.

Wignall:2007:MDM

[Wig07] J. W. G. Wignall. A microscopic definition of mass. Physics Today, 60(4):12, April 1, 2007. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/full/10.1063/1.2731956.

Williamson:1948:NMB

[Wil48] Marjorie Williamson. A note on M. de Broglie’s theory of the photon. Philosophical Magazine Series 7 (1926–1955), 39:314– 324, 1948. CODEN PHMAA4. ISSN 0031-8086.

Wilson:1961:MDB

[Wil61] William Wilson. Maurice, Le Duc de Broglie. 1875–1960. Biographical Memoirs of Fellows of the Royal Society, 7:31–36, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://www.jstor.org/stable/
769394.

Wilson:2011:ERB

[Wil11] R. Mark Wilson. Experiments reveal a Bose–Einstein condensate of photons. Physics Today, 64(2):10–11, February 1, 2011. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/1.3554306.

Winter:1931:IEM

[Win31] M. Winter. Introduction à l’étude de la mécanique ondulatoire par Louis de Broglie. (French) [Introduction to the study of wave mechanics by Louis de Broglie]. Revue de Métaphys. et de Morale, 38(??):125–131, 1931.

Winter:1934:ETD

[Win34] Jacques Winter. Études théoriques sur la diffusion des ondes de de Broglie. (French) [Theoretical studies on the diffusion of the waves of de Broglie]. Ann. Physique, XI. Sér., 11(2): 455–560, 1934. ISSN 0003-4169 (print), 1286-4838 (electronic).

URL https://www.annphys.org/articles/anphys/abs/1934/
02/anphys19341102p455/anphys19341102p455.html.

Witte:1998:ESR

[Wit98] N. S. Witte. Exact solution for the reflection and diffraction of atomic de Broglie waves by a travelling evanescent laser wave. Journal of Physics A (Mathematical and General), 31(2):807–832, 1998. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Wheaton:1984:TSE

[WN84] Bruce R. Wheaton and Pierre Noyes. The tiger and the shark: Empirical roots of wave–particle dualism. Physics Today, 37(10): 95, October 1, 1984. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.2915925.

Weidner:1962:EMP

[WSAS62] Richard T. Weidner, Robert L. Sells, and Fay Ajzenberg-Selove. Elementary modern physics. Physics Today, 15(12):74, December 1, 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.scitation.org/doi/
abs/10.1063/1.3057932.

Webb:1988:QIF

[WW88] Richard A. Webb and Sean Washburn. Quantum interference fluctuations in disordered metals. Physics Today, 41(12):46–53, December 1, 1988. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.scitation.
org/doi/abs/10.1063/1.881140.

Weissman:1997:ASH

[WW97] Samuel I. Weissman and Michael Weissman. Alan Sokal’s hoax and A. Lunn’s theory of quantum mechanics. Physics Today, 50 (6):15–114, June 1, 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.881789.

Wang:1991:ETB

[WZM91] L. J. Wang, X. Y. Zou, and L. Mandel. Experimental test of the de Broglie guided-wave theory for photons. Physical Review Letters, 66(??):1111, ???? 1991. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. See comments [HV91, CGLM92].

Yang:2012:QNC

[Yan12] Chen Ning Yang. Quantum numbers, Chern classes, and a bodhisattva. Physics Today, 65(1):33–34, January 1, 2012. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/full/
10.1063/PT.3.1398.

Yaes:1997:GSS

[YF97] Robert Yaes and Kenneth Fox. Gatekeepers in science stifle individual creativity, those in legal system should study science. Physics Today, 50(2):11, February 1, 1997. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.881956.

Yourgrau:1976:AL

[You76] Wolfgang Yourgrau. Alfred landé. Physics Today, 29(5):82–83, May 1, 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.scitation.org/
doi/abs/10.1063/1.3023483.

Yang:1999:SNT

[YY99] Wenxiong Yang and Changjun Yang. Super nonlinear total energy of a particle and the theory of de Broglie wave. Applied Mathematics and Mechanics = Ying yung shu hsueh ho li hsueh, 20 (5):527–531, 1999. CODEN AMMEEQ. ISSN 0253-4827 (print), 1573-2754 (electronic).

Zare:1980:SSR

[ZB80] Richard N. Zare and Richard B. Bernstein. State-to-state reaction dynamics. Physics Today, 33(11):43–50, November 1, 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.2913823.

Zhu:2015:FEM

[ZD15] Yimei Zhu and Hermann Dürr. The future of electron microscopy. Physics Today, 68(4):32–38, April 1, 2015. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/full/10.1063/PT.3.2747.

Zou:1992:CEB

[ZGWM92] X. Y. Zou, T. Grayson, L. J. Wang, and L. Mandel. Can an “empty” de Broglie pilot wave induce coherence? Physical Review Letters, 68(??):3667, ???? 1992. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Zucker:1963:MF

[ZS63] Alexander Zucker and Arthur H. Snell. Meson factories? Physics Today, 16(4):19–24, April 1, 1963. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.scitation.org/doi/abs/10.1063/1.3050872.

Zurek:1991:DTQ

[Zur91] Wojciech H. Zurek. Decoherence and the transition from quantum to classical. Physics Today, 44(10):36–44, October 1, 1991.

CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.scitation.org/doi/abs/
10.1063/1.881293.

Zhang:1992:QSA

[ZW92] Weiping Zhang and Daniel F. Walls. Quantum superpositions of atomic de Broglie waves by an atomic mirror. Physical Review Letters, 68(??):3287, ???? 1992. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.